A Complete Bibliography of IEEE Transactions on Computers (1970–1979)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 December 2023 Version 1.21

Title word cross-reference

#14 [Fre74b].

(d, k) [Sto70d]. (d, r) [Spi73]. −2 [Yue75b, KO73, Zoh73d]. 1 [Dav78]. 16 [Ath75, Erc73]. 2 [Agr77, Bak75a, BR71, Chi77, GBMC72, Liu75, SB77a, SB79, TL73b]. 2^{n} × 2^{n} [Ari79]. 3 [Ano76t, Ano77u, Ano78-42, Dav78, SK78a]. 4 [Ano78-43, CSY75]. 40 [GSB^{+}76]. ^{2} [McC79]. A = (r − 1)p^{2} [Rao78]. B [Hwa72]. C [Ben71]. d [PHG77]. GF (2^{m}) [LR71b, Sar78]. H [MS73b, MA71, RRNA74]. j [PHG77]. k [AM79c, FN75, Hwa72, TOS71, Wyn71]. L^{2} [Pav75b]. log_{2} x [KP75]. LU [SK74]. M [AH76, AM73, AM79c, MF78]. N [Boy74, HN71, LSB77, RH75b, Sin72, TM71, TI72a, WM71, AM73, Goo70, MF78, Pea77, RW74, RH74, Wyn71]. n ≥ 1 [Dru71a]. O(n) [CW79b]. P [RW74, Kar74, Kri77, Swa74a]. Q [SB74b, HC78]. QR [SK77a, SK77b]. R [Kun75, SV74, Ulm70, PHG77]. r = 2^{n} [Dru71a]. T [HK77, HK78, RH75b, PHG77]. X [Ath75]. x^{P} + x^{k} + 1 [AM79c]. Y [Ath75]. y = Ay^{n} [Max77]. Z [Cor77].

-2 [Zoh76b]. -adic [Kri77]. -Algorithm [SK74]. -ary [HC78]. -Asummability [SB77a, SB79]. -Class [Ben71]. -Cube [Pea77, Wyn71]. -Dimensional [Goo70]. -Fail-Safe [TM71, TI72a]. -Gate [HK77, HK78]. -Grams [RH74]. -Graphs [HN71]. -Input [SK78a]. -level [Liu75]. -Multiple [Kar74]. -Nearest [FN75].

-Ordered [Swa74a]. -Out-of- [AM73, Dav78, MF78, RW74]. -out-of-n [TOS71]. -Output [SK78a]. -Sequences [AM79c]. -Space [LSB77]. -Stable [AH76]. -Summability [GBMC72, Hwa72]. -Transform [Ulm70]. -Universal [SB74b]. -Valued [SV74, Sin72]. -Vector [Ath75].

1 [Har70a, Har70b, HA70, Ric74a, Sil73, Su75, Wex71]. 10 [Arb70, Gel73a, Rem72, Ros74a]. 11 [Fai73, Kin72, Poe74, Yam70]. 12 [Bib74, Rot70, Su73]. 13 [Fre74b, Win70a]. 14 [Cut70, Jos74]. 15 [Bar74c, She70b]. 16 [Lud74, Skl70]. 17 [Fri70, Lew74, Rip72]. 17-Bit [Pez71a, Pez71b]. 18 [Bar74b, Goo70, Jos72a]. 19 [Ano70a, Cox72, Gil74a, New70]. 1969 [Shi70]. 1972 [Ano72-40]. 1973 [Ano73f]. 1976 [Ano76v, Ano76z, Ano76y, Ano76w]. 1977 [Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, Ano78t].

2 [And72, And70, Bas71a, Bon74, DLL^{+}73, Fos73, Gor75, Hop70, Tab70b]. 20 [Ano71a, Den74, Luc72, She70c]. 21 [Ano72a, Arn70, Pav74, Wan72]. 22 [Ano73a, Lom74, Swa72b, Tur70]. 220 [SB71]. 23 [Ano74a, Har74a, Kam70a, McL72]. 24 [Ano75a, Dee72, Rob70b, Vic74]. 25 [Ano76b, Ben70, Coo72, Mil74a, SB71]. 254 [Ano79b]. 26 [Ano77a, Duk74, Flo72, Yoe70a]. 27 [Ano78b, Ber70a, Gag72, Wul74]. 28 [Ano79a, Ham74, Pre70]. 29 [Jum70, Lei74]. 2nd [Jos72a, Luc72].

3 [Aar70, Cha74c, Fly72a, Pag70a, Pav75a, Rux71, Uhr73]. 30 [Fu74, Joh71, Su70]. 31 [Kuo74]. 32 [Voi74, Wag70]. 33 [Bar74a, Paz70a]. 34 [Bar70, Fre74a]. 35 [Kra74, Tsi70]. 36 [Hed70, Mah74]. 360 [Cam72, SB71, Wal74]. 360/25 [SB71]. 37 [Mui74, She70a]. 370 [Flo72]. 38 [BS74, Lev70]. 39 [Kam70b, Wal74].

4 [Bre71c, Cam72, Geo70, Jos72b, Pla73, Sch74b, Sto70a, Voi75]. 40 [Ake70, Sie74]. 40-ns [Pez71a]. 40ns [Pez71b]. 41 [Cha70, Rei74b]. 42 [Dav70a, Rei74a]. 43 [Kim74, Yeh70]. 44 [Brz70, Gri74]. 45 [Paz70b, Ris74]. 46 [Mic74].

5 [Bre73a, Gil74b, Gra70, de 70]. 5th

[Ano79c, Ano79d].

6 [Bas71b, Hob73, Jor74b, Kar70, Tab70b].

7 [Gil70b, Jor74a, Roo73]. 70 [Wak70c]. 74 [Bar74a, Duk74, Fre74a, Fu74, Ham74, Kuo74, Lei74, Voi74, Wal74, Wul74]. 74-26 [Duk74]. 74-27 [Wul74]. 74-28 [Ham74]. 74-29 [Lei74]. 74-30 [Fu74]. 74-31 [Kuo74]. 74-32 [Voi74]. 74-33 [Bar74a]. 74-34 [Fre74a]. 74-39 [Wal74]. 79 [Ano79m, Ano79n, Ano79o, Ano79p, Ano79q].

8 [Els74, Ker73, You70].

9 [CDO78, Feu72, Law70, Ram73a, Swa74b]. 9-V [CDO78]. 900 [Guf71].

A/D [Mor75, Zoh73d]. Abbreviated [CM79c]. ABE [AT70]. Abelian [Cai71, Kar77b]. Abort [FD74a]. Abstract [BT73, Kim79, Poe74, de 70]. Abstracts [Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano73e]. Academic [Tab70b]. Accelerated [Cha71]. Acceptors [Har74b]. Access [AU74, Bat77, DI71, Hay75, Kau73, KH77b, KH77c, Law75, Len77, NTA78, Nai79, Nel74, SD79, Sri78, Sti78a, Sto72, Sto75, Wil78]. Accessing [MVW79]. Accuracy [Boh77, Kul78, Mar73c]. Accurate [AS74]. ACDL [BT74]. Achieve [BCH^{+}73]. Achieving [AN74]. Acting [Man71]. Activity [KR75]. Adaptation [Dee72, Mak71]. Adaptive [BI72, BR76b, Bur74, GW77, Ios75, Koh74, Mil77, Mor70a, TI72b]. Addendum [LE79]. Adder [Gad70, She77, Svo70]. Adders [BK71a, LM79, LHSM74, Pra74, Rao72, WL74, WM71]. Addition [AR78, AO79, Ban74b, Bre70a, DS72b, Gar78, HC73, KO73, SW71a, SW73]. Additional [CF70b, Den76, Gou77, WA73]. Additive [EGG72]. Address [HPP77, Sch78c, Sie77]. Addressable [LC79]. adic [Kri77]. Adjacencies [Myk75, Van73a]. Adjacency [Cur77]. Adjunct [Har74b]. Admissibility [GW77]. Adoption [PB71]. Advance [Ano78c, Ano78f, Ano78d, Ano78e, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j]. Advanced [BCA^{+}75]. Advantage [CS70]. Advantages [Feu73]. Aerospace [Mal72]. AFA [GG73]. Affecting [GM76]. Affine [DN77]. AGC [Mor75]. AI [EFLR76]. Aid [Ber70b, KR71a, Kar77a, Lan78]. Aided [EZS78, Lar71, Mar78, SN71]. Air [PS78]. Aircraft [DBM77]. Algebra [DM75, DM76, Eps73b, Lee76, SAPR70, SS72, Svo73, Tis71, VLS70]. Algebraic [BF75a, CSJ74, Den76, Gou77, GM76, HC78, Pra71a, RS71a, Rud74, SG74, UB74, Whi71c]. Algebras [Eps73a, NSR74, Yeh70]. Algol [Fre74a]. Algorithm [AS74, AA77, All75a, All75b, Ami77, Ath75, BC71a, BW73, BI72, Ber72, Bla74, BE70a, BG79, Bro77, Bur74, But71, Car71a, CM74, CDO78, Cha74a, CL77, Cor77, Das71c, Den76, DH79, Eks73, EO78, FT76, FD77a, Fre79a, FC77, FBS75, FV79, FK70b, FO71a, FN75, Ful72, GM72a, GL70, God71, Gou77, Gri78a, HPP79, HO79, Hoe76, HW75, IR76, Jac73, JB76a, JLH73, Kau74a, Ken74, KH77b, KH77c, KS73, KF72a, KNF75, Kro74, LNL72, LH76, Lee78b, Len78, LK75, Lig77, Lin70, Lit74, MT74a, MM73b, MM78b, MF76d, MC70, NK74, Nai79, NF77, Nec79, NW79, OS76, Pap78, Par79, PM77b, Pic74, Pre78b, PR71, Ram75, Ram77a, RT79, Rub74, Rub75a, Rub76a, Rub75c, SG74, SK77a, SK77b, SCK73b, SB77a, SV75, SK74, SM70a]. Algorithm [SMW71, Sil79, Sil77, SCL70, Smi76, SO79b, Sto74b, SKK79, Sur74, SH79, Tab70a, Tan78, TI72b, Tom70, Tom73, Tru71, TE76, Van77, Wak70b, Wan75a, WG73, Weg72, Wet78, Wil70, Wil77b, WL77, YT76, YT71c, Yue77d, Bel76, FT74]. Algorithmic [Bre71d]. Algorithms [Agr75a, AOR74, ABK76, Avi73, Bak75b, BB72, Bar77, BS72, BS78, BF78, BO79, BHP^{+}71, BG78, BV79, Che73, CF78b, Des79, DTI78, Dow73, DW73b, Far73, Fer77, FL75, GHJ77, GC77, Gel73b, Geo70, GL74, GF78, Han75, KKK74, Kri71a, LT73, Liu75, Mah72, Mar72, Mor70b, MM70, Muk79, NTA78, PBC74, RT73b, Ros74a, SCK73a, SR72a, Smi79e, Sta74, Tab74, Tom74, TD70, Tri76, TE77, Tri77a, Yan74a, Yan75a, Bak75a]. Aligned [SP72a]. Alignment [Law75]. Allocation [BE70a, BE70b, GSB^{+}76, RC74, Ung77a]. Allocator [Von75]. Almost [Hal74, Kau74a]. Almost-Optimal [Kau74a]. Alpha [Gri76]. Alphanumeric [Ern72]. Alternate [She78a]. Alternate-Data [She78a]. Alternating [RM78]. Alternative [BCB74, But71, O’C75]. Alternatives [Ano77b, Ano77c]. Ambiguity [BEGO71, MS77a]. Amenable [Buz73]. Among [New70]. Amplitude [SP72b, SP72c]. Analog [Aar70, All75b, BS77, Gil70b, Gra70, KES72, KR71a, Kon71, KV76, Lew77, LC72, Mar78, New75b, RS72a, RC71c, TP72, WK73, Wil77b, You70, Yue78]. Analog-Computer [KES72]. Analog-to-Digital [Aar70]. Analog-to-Gray [Yue78]. Analog/Digital [Gil70b]. Analog/Hybrid [KV76]. Analysis [AD75, AA75, AKT75, AC70, Ano78h, Ano78i, Ano78j, Ano78-37, Ano78-38, Ano79k, Ano79l, ABK76, BW76a, Bha75a, BO70, BCL75, BH71a, Bre71a, BG78, CR72d, CM73b, CA77, CS79, Cle73, CDL78, Cov74, Dav72b, DB74, FB74, FT79, GS79, Geo70, GGMM71, Gil70b, HND75, HO74, HE73, HF71, HI72, KES72, KF79, Kim70, KKA71, KF72b, KF72c, KNF76, Kor79, KS79, Kru71, KM75b, KPS77, LB74, Led72, LF72, Low73, LF77, Mal72, MAB78, Mat71c, Mey76, MMB77, Mor70a, NA74, OH79, PM75a, Pav74, PM78b, RH75a, Rat76, Rip72, RT71, RL72, Sam70a, SF71, Sch79b, Shr71, Sie77, SR71, SR77, Tho74, Tho77, UM77, WKP75, Wal76, WN76, Wel78, Whi72, Whi71c, FT74, NA76b]. Analytic [FS74, Smi79a]. Analytical [Hoe74, Var79]. Analytically [RS71b]. Analyzed [Met73]. Analyzer [DEA73, MN70, MB75]. and-Exclusive-or [Fis74]. Angle [RJ73, RW75]. Angles [Dav77b]. Announcement [Ano76-28, Ano78g, Ano78c, Ano78d, Ano78e, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j]. Announcements [Ano78f]. Announces [Ano78w, Ano78v, Ano78x]. Announcing [Ano76c, Ano76d, Ano77b, Ano77c, Ano78h, Ano78i, Ano78j, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79k, Ano79l]. Anomalies [YRT75]. Anomalous [CM73a, FD79, Pec76]. Answering [BCSW76]. Anticipatory [WY73]. Any [Bar71, CS70, Dru71a, Len77, MF78]. Aperture [Wei75]. Apes [WBRG76]. APL [Har71, TM70, Wal74]. APL/360 [Wal74]. Apparent [Tun70a]. Application [ARS71a, Arm73, Avi71, BM77b, BB74a, CH72a, CH73b, CC73, CM79d, Dav72b, DW72a, DVS74, DV78b, Edw75, Ern72, Fel79, FK70a, GL70, HHJ^{+}76, HK77, HK78, HA70, Hou72, Hun73, KT72, LF72, Mag70, Mam77, MM73b, RS71b, Sar71, Sch73, SF72, Sho74, SB72, Sil79, Su70, TM76, TD73, The75, Tis71, TE76, WG73, Wei75, Yan75b, Yue76, ZHR77a, ZHR77b]. Application-Oriented [Fel79]. Applications [Ale71, Ano72n, BR76a, Bha78, Brz70, CS75, Duk74, HLD70, Har74a, HSS78, HP72, Lem70, LJ79, RS71a, Rem72, RH73, SNOS79, Tri76, Wal74, YT70]. Applicative [FW78]. Applied [Moh71]. Applying [BT73]. Approach [AM79a, Agr79c, Agr75b, AKT75, BK76, BO70, CM77, CCB70, Coo72, Das71b, DK72, Das73, Eks74, Gel73b, Hu75, JB76b, Kam75, KS78a, KNF76, KSW74, KV73, LD75, Lec70b, Man77, MA71, Mil74a, MS75b, MF76c, Mon74, Mur70, PW79b, RM71, Ram76b, Reu75, RV73, RH71, Rud74, SB79, SD74, SR72b, SB77b, Sto77, SC72b, Swa72b, TBY78]. Approaches [AO79, Bre73a, Gil74b, SM71b]. Approximate [BM70, Fos73, HLD70, Mam77, Mar72]. Approximating [CY76]. Approximation [ADMR78, Kar70, KP75, LT70, Pav73, Pav75b, PR74a, Tom74, Van75, Vil74, Yue75e]. Approximations [BB78c, Bre71b, Chu74, Hun71, Pav77]. Arbiter [CM73a, PFL75]. Arbiters [Plu72]. Arbitrarily [Sut74]. Arbitrary [DEA73, Han72, Hen71, KN70, LH76, MS77b, OW72, OY70, Pap77b, PB76a, Sik70, Smi71a, SP72b, SP72c, YT71b]. Architectural [FSLC78, HS75, SNEL79]. Architecture [AAK74, BV78, BS75, Chu79, Col77, Feu73, KK79a, RA78, Sri76]. Architectures [EFLR76, MM78b]. Arcs [CY76, SKK79]. Arctan [SK71]. Area [LH79, Stu75]. Argument [Hoe74]. Arising [CB72, GHJ77]. Arithmetic [Agr75a, Agr77, Agr79a, Ano77g, Atk70, AG73, AT70, Avi71, Avi73, Ban74c, Bar77, BM78, BM77b, BKM73, CR72c, CH72b, Cod73, Dev75, Gar76, Gri78b, HF72, HL86, KKK74, KS77c, Kri77, KPS77, Kul77, LW78, Maj76, Man72b, Meo75, NR75, OW73a, PA78, RPL73, Rao70, RT70, Rao78, Red79, SCK73a, Shi70, Spi73, SM71c, Sut74, SCNS83, Swa90, Tri77b, Tri78, Tun70b, Wak75b, Wil77a, YC76, Yoh73, Yue75b, NA76b]. Arithmetical [TG72a]. Arithmetics [MM73a]. Array [Ake72, BBC72a, BBC72b, BW73, Bla74, CH73a, DS70, Dia76, FL77b, Gar74, HC73, Hua71, HK74, Hwa79, KSA74, KS78c, Kro74, Law75, LR71b, LC79, Lip77, MK71a, MK71b, Maj72a, Maj73, Nut77, Ogu74, PM77a, Pea77, Pez71a, Red79, Rot78, SC74, Tom75, Tri76, Tri77a, Woo79, Pez71b]. Arrays [Agr79c, Agr79a, AB78, BM77a, Ber77, Cur79, Deu70, Dev75, Dir72, Fle79, FM72b, Goo70, GT78a, Ham71, Hou72, HI72, IH72, IC78, Jul78, JF72, Jum74, Kam79, KAN79, LY71, Lar71, Maj76, Man77, MF71, Nic71, OH79, Pap72, Pra78d, PG75, SR74a, Smi79c, Stu75, Sun76, YO70]. Arrivals [Chu70, MN78]. Art [Fre74b, Hob73, Kra74, Lom74, Mui74, Pla73]. Artificial [Arb76, BW76b, Coo72, FL77a, Hol76, Rei74b]. ary [HC78]. Asbestos [PS78]. Aspect [SC74]. Aspects [BR76b, Fre79b, FW78, TL73a, Yip76, Yoe70a]. Assembler [Hus73]. Assembles [EUY^{+}72]. Assembly [Kau74a, WG75]. Assessing [Mal73]. Assessment [Kla71]. Assignment [FG71c, GLL79, HA74, HB70, KR78a, MT70, MT71, MSJ72, Noe73, NR74, NR76b, RSH79, Sau72b, SHR72, TB74]. Assignments [Cur73a, FK78a, Kin71c, MN78, NT78, NT79, Par72a, Pra78a, RN77, Smi74, Tan71]. Associated [KR73]. Associative [CCF79, DB74, FL77b, FS71, Kin71b, Koh74, Lan77, LC79, RTW78, RS71a, SNN70, SH77, SNOS79]. Associative-Search [LC79].

Asummability [SB77a, SB79, Sur74]. Asymmetric [Hu75]. Asymptotes [Win70b]. Asymptotic [BB78c, CF70b, KF72b]. Asymptotically [CW79c]. Asynchronous [Bat70, BH71b, BH73, Bre75, BA71, BW70, CD73, Chu76, Dav77a, Fri70, FG71b, FG71c, FG76, HD71, HC71b, HA76, Jum74, KM73, Kin70, Kin71c, KR78a, LF74, Mag71, MT70, MT71, MS73a, MS74, Mar77, MT74b, NT78, NT79, NM77, PM74, PFL75, Plu72, PR73, Pra78a, Pra78b, PR71, RV74, Sau72a, Sau72b, SM74, Saw74, SM75a, SM75b, Ser73, She70a, SY75, SS71, Sin71, SR71, Smi74, Tan71, Ung71, Wak75a, WM74, Yen79, BT74]. Asynchronously [FM71]. Atomic [PC78]. Atrubin [Goy76]. ATS [SP72a]. Attainable [Bre73b]. Auditing [Ano76u]. Augmented [CH73a, Gar74, KM73, SK74]. August [Ano76v, Ano77w]. Author [Agr79b, All77, Ano70n, Ano72o, Ano72p, Ano72q, Ano72r, Ano72s, Ano72t, Bab74, BH74a, CM79a, Che72, CBD70, Chu72, CD75, Coh70, Das72, DM76, Edw76, Ekl73, En79, FCC72, FO71b, GRR74, GR73a, GM72b, HC74, HF70, KAA77, Koh72a, Koo78, LM71a, Mar75, Mar76, MB79, Men70a, PR75, Pra71b, Rao78, RS73a, RS73b, SM79, SS75b, To76, Yue77a, Yue77b, Zoh71]. Authors [Ano70-38, Ano70-39, Ano70-40, Ano70-41, Ano70-42, Ano70-43, Ano70-44, Ano70-45, Ano70-46, Ano70-47, Ano70-48, Ano71-37, Ano71-38, Ano71-39, Ano71-40, Ano71-41, Ano71-42, Ano71-43, Ano71-44, Ano71-45, Ano71-46, Ano71-47, Ano71-48, Ano72-65, Ano72-66, Ano72-67, Ano72-68, Ano72-69, Ano72-70, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano74-27, Ano75-27, Ano78-44, Ano79-51, HW78]. Autocorrelation [Mar73b, NM76b, NM76a]. Automata [Arb70, Ban74a, Brz70, Cho74, Cut70, EGG71, Gaj75, Hed70, Hop70, Kam70a, Kam70b, KW70, LF78, Moo71, Pag70b, Par72b, Paz70a, PR74a, Poe74, Pu72, RS71a, SIF76, Tsi70, Wag70, Yan74a, Yan75b, YM78, de 70, de 71]. Automated [CRB72, HND75, KJ79, MOW76, Mor76b, SP72a]. Automatic [Abe72, AA72a, Bec71, Dee72, OS76, PS78, RS74, RJ72, Rei76, Ros75, SW71b, SH72, WN76]. Automation [Cha74c, NR76a]. Automaton [Arn70, Bar71, HN78, MF76d, RS72a]. Automorphism [TV79]. Autonomous [IH72, Kri73, MR72, Wan71]. Autoregressive [Tjo77]. Autotesting [Cio78]. Auxiliary [Dru71a]. Availability [CLL78]. Available [Ano72-50, Ano73h, Ano79-53]. Average [Bri73, Yue74a]. Averaged [Dun75]. Avoidance [HD71]. Axiomatization [Eps73a]. Axioms [San76].

B [Bar74a, Duk74, Fre74a, Fu74, Ham74, Kuo74, Lei74, Voi74, Wal74, Wul74]. B70 [Aar70, And70, Har70a, Sto70a, Tab70b, de 70]. B70-1 [Har70a]. B70-2 [And70]. B70-3 [Aar70]. B70-4 [Sto70a]. B70-5 [de 70]. B70-6 [Tab70b]. B71 [Bas71a, Bas71b, Bre71c, Rux71, Wex71]. B71-1 [Wex71]. B71-2 [Bas71a]. B71-3 [Rux71]. B71-4 [Bre71c]. B71-6 [Bas71b]. B715 [Har71]. B72 [And72, Cam72, Coo72, Cox72, Dee72, Feu72, Flo72, Fly72a, Gag72, Jos72a, Jos72b, Kin72, Luc72, McL72, Rem72, Rip72, Swa72b, Wan72]. B72-10 [Rem72]. B72-11 [Kin72]. B72-17 [Rip72]. B72-18 [Jos72a]. B72-19 [Cox72]. B72-2 [And72]. B72-20 [Luc72]. B72-21 [Wan72]. B72-22 [Swa72b]. B72-23 [McL72]. B72-24 [Dee72]. B72-25 [Coo72]. B72-26 [Flo72]. B72-27 [Gag72]. B72-3 [Fly72a]. B72-4 [Cam72, Jos72b]. B72-9 [Feu72]. B73 [Bre73a, Fai73, Fos73, Gel73a, Hob73, Ker73, Pla73, Ram73a, Roo73, Sil73, Su73, Uhr73]. B73-1 [Sil73]. B73-10 [Gel73a]. B73-11 [Fai73]. B73-12 [Su73]. B73-2 [Fos73]. B73-3 [Uhr73]. B73-4 [Pla73]. B73-5 [Bre73a]. B73-6 [Hob73]. B73-7 [Roo73]. B73-8 [Ker73]. B73-9 [Ram73a]. B74 [Bar74b, Bar74c, BS74, Bib74, Bon74, Cha74c, Den74, Els74, Fre74b, Gil74b, Gil74a, Gri74, Har74a, Jor74a, Jor74b, Jos74, Kim74, Kra74, Lew74, Lom74, Lud74, Mah74, Mic74, Mil74a, Mui74, Pav74, Poe74, Rei74a, Rei74b, Ric74a, Ris74, Ros74a, Sch74b, Sie74, Swa74b, Vic74]. B74-1 [Ric74a]. B74-10 [Ros74a]. B74-11 [Poe74]. B74-12 [Bib74]. B74-13 [Fre74b]. B74-14 [Jos74]. B74-15 [Bar74c]. B74-16 [Lud74]. B74-17 [Lew74]. B74-18 [Bar74b]. B74-19 [Gil74a]. B74-2 [Bon74]. B74-20 [Den74]. B74-21 [Pav74]. B74-22 [Lom74]. B74-23 [Har74a]. B74-24 [Vic74]. B74-25 [Mil74a]. B74-3 [Cha74c]. B74-35 [Kra74]. B74-36 [Mah74]. B74-37 [Mui74]. B74-38 [BS74]. B74-4 [Sch74b]. B74-40 [Sie74]. B74-41 [Rei74b]. B74-42 [Rei74a]. B74-43 [Kim74]. B74-44 [Gri74]. B74-45 [Ris74]. B74-46 [Mic74]. B74-5 [Gil74b]. B74-6 [Jor74b]. B74-7 [Jor74a]. B74-8 [Els74]. B74-9 [Swa74b]. B75 [Gor75, Pav75a, Su75, Voi75]. B75-1 [Su75]. B75-2 [Gor75]. B75-3 [Pav75a]. B75-4 [Voi75]. Balakrishnam [Tab70b]. Balanced [FL75, KV73, LP76, LP78]. Balancing [Cha74d, CK79]. Band [GSJ^{+}76, Sok72]. Bandpass [Wai70]. Bandwidth [CKL77, Rau79a, Rav72]. Bandwidths [CBS71]. Base [Agr75a, Agr77, Cha74a, CR72d, Chi77, Jos74, KO73, LJ79, Mat70, Mui74, OW73a, O’K75, OW72, SCK73a, SCK73b, WM71, Yue75b]. Base- [Agr77, Chi77]. Based [AKT^{+}77, Ber71, Che71b, Che73, DM78b, Gel76, JP73, KH77a, Mey76, MS76, OTK76, PW79a, Per78, Rub79, Sah74, SR72a, TG79, YC76, RT73a]. Bases [GS79]. Basic [Jor74a]. Basis [HA77]. Batch [Ano72-39, MN78, Par71]. Batch-Fabricated [Ano72-39, Par71]. Baye [Tou75]. Bayes [Dev74, GY77]. Bayesian [Hun77]. BCD [Ben77, BR71, Maj73]. Be [SO73, Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, Kin71a, LF74]. Beam [Wan77]. Behavior [BF72, CM73a, Chu70, CW75, FD79, Hei79, HI72, Rau79b, She77]. Benes [Len78]. Berkeley [DS70]. Bernoulli [EGG72]. Beta [Gri76]. Better [Ano75c, Ano75d, Ano75e, Ano75f]. Between [DM75, DM76, Eps73b, Goo71, Kun79, Lan76, MR72, Moo71, RHR76a, RHR76b, Rob70a, Rus72, Stu75, SS72, Tak76, Tou72e, Ull72, Van73a]. Bias [DF78]. Bidirectional [NT70]. BIFORE [ARS71b]. Big [SO73]. Bilateral [GT78a, Maz71]. Bilevel [JR76]. Bilinear [Lam79a]. Binary [Agr76, AR78, Ake78, AO79, BBC72b, Ban78, Bar77, Ben77, BF75a, Boy74, Bre70a, BP72, BL75, Byr71, CH73a, CS70, CMS79, Cha74d, Che71b, Dav71, Dor72, EI77, FL75, Gad70, Gar72b, GD72, Gra71, GP74, HLD70, Har73, HW79, Hor71a, IC78, JB76a, Klu77, Kri73, LM79, LF74, LW78, Lyn72, MK71b, Maj72a, Maj73, Mar72, Met73, MBM77, NR77b, Obe74, Pak71, PW79a, PM77b, Pea77, PC72, RC71b, RT70, Rat76, Rhy70, RE71, RH74, Rom78, Sag77, SV74, SB72, Sit74, Sok72, Ste72, Sur70a, Tom75, Toy71b, WL74, Wan75c, WC74, Bak76]. Binary-Coded-Decimal [LW78]. Binary-to-Decimal [Hor71a, Rhy70]. Biresidue [Rao70]. Bit [CW79b, EF79, Lin70, LHSM74, Pez71a, PBC74, Rob79a, Pez71b]. Bit-Sequential [CW79b]. Bitonic [NS79]. Bits [JW73]. Black [Mor74]. Black-White [Mor74]. Blake [Bha78]. Blobs [AL71, SCL76]. Block [LR71a, Sto70c]. Blocking [RM71]. Blocks [BT72, MC78, O’K73, Sch78b]. Blurred [Jai77b]. BN [CR72c]. Board [SO73]. Boards [AA72a]. Bodies [SF77, SF78]. Book [Ano72u, Ano72v, Ano72x, Ano72w, Ano72y, Ano72z, Ano73b, Ano73c, Ano73d, Shi75]. Books [Ano77d, Ano78k]. Boolean [Bis73, CM79d, DDSM76, DM75, Eps73b, MST^{+}75, Yue76, AB78, Bar70, Bau72, Bis70, Bis71, BG78, Bro70b, CRB72, Chi77, Cur76, DS75, DM76, Edw75, GGMM71, GBMC72, GC73, GC71b, HG78a, Hay71b, Hig73, HO72, HC71b, Hwa74, Kli74, KM75b, Lee76, Mar71, MPS78, Pap77a, Pic74, PD79, PC78, PM71, Pre71a, PM76b, PMB77, RS71a, Rud74, Rud76, SB77a, She70c, Spi73, SS72, Sur74, Svo73, TD73, Ulr71, Win70a]. Booth [Rub75c]. Bound [Du72, FN75, Han72, Hol72a, KNF75, NF77, TB75, Van72, Vil70, Wet78]. Boundaries [Nev76]. Boundary [ACJ75, FD73, FC77, Hu73, NJ77, PCM70, SB72, Tho77, Vic71, WK73]. Boundary-Value [WK73]. Bounded [NA72, SF78]. Bounds [BCKT79, Bau71, CK75, CK77, Chu71, Chu74, Der73, Dev74, FS70, FB73a, Fre79b, Moo71, Red73b, Rub75b, Tou71, Tou72b, Tou75, Wei72, Yue72]. Bragg [WKP75]. Bragg-Diffraction [WKP75]. Brain [Rei74b]. Branch [FN75, KNF75, NF77, Wet78]. Bridge [Ios78, WS76]. Bridging [Mei74]. Bruijn [Lem70]. Bubble [BD71, Bha75b, BCH^{+}73, CW79a, Che78, Gar72a, KSM76, LC79, LJ79, MBSS75, Min75, WC76]. Bubble-Memory [BCH^{+}73]. Bubbles [Whi71a, Wil77a]. Buddy [Von75]. Buffer [Chu70, Hei79, Hil79, MH70a]. Building [AT70, O’K73, ZT75]. Buried [YRT75]. Burroughs [SB71]. Bus [FT79, Nut77]. Business [Luc72]. Bypass [Kau70a]. Byte [Gri78b, NR75, Red78a]. Byte-Organized [Gri78b, NR75]. Byte-per-Card [Red78a].

C

[Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano76b, Ano77a, Ano78b, Ano79a]. C-19 [Ano70a]. C-20 [Ano71a]. C-21 [Ano72a]. C-22 [Ano73a]. C-23 [Ano74a]. C-24 [Ano75a]. C-25 [Ano76b]. C-26 [Ano77a]. C-27 [Ano78b]. C-28 [Ano79a]. Cache [BCB74, Cho76]. Calculation [Fos73, Van75]. Calculus [Lee76, San76, TD73]. Call [Ano77e, Ano78l, Ano78m, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano79r, Ano79s, Ano79t, Ano79u, Ano79v, Ano79w, Ano79x, Ano79z, Ano79-27, Ano79-28, Ano79y]. can [LF74]. Canonic [PP75, SR75, SO79a]. Canonical [KS75, KS77b]. Capabilities [PRD71, RM78]. Capability [HPP77]. Capable [Gra73, MKO70]. Capacity [Bit77, Cho76]. Card [Red78a]. Cardiac [CC73]. Cards [AB79b]. CARE’S [Vin78]. Carlo [AA76, SF74]. Carry [Bri73, Chi77, Pra74, She77, Yue74a]. Carry-Save [Pra74]. Cascade [BPC75, Cut70, DSBC74, Mat71b, MBSS75, Pap78, Pap76, SK78a, SK79b]. Cascaded [GC71a]. Cascades [But75, HE79, Kau70a, Kol72, Pap77b, Yoe70b]. Case [Bri73, Cha73b, CDL78, FL77a, MH70b, Rub75b, Yue74a]. CASSM [SNEL79]. Catastrophic [Sri71b]. Categorization [TG72b]. Cause [BH71a, LB74]. Cause-Effect [BH71a, LB74]. Cayley [Vil73a]. CC [Cur71a, HS70, VVA78]. CC-Tables [Cur71a, VVA78]. CCD [FM78]. Cell [BM77a, BB73, Nel74, Ogu74, SM73a, Sri71a]. Cells [But78b, CK74, CW79c, Coy79, Gec77, GT78a, Led72, Liu77a, MH73, PG75]. Cellular [Agr79c, BBC72a, BM77a, But76, CL72, Gar72a, HN78, Hu72, Hua71, HK74, Hwa79, Kau70a, KS78c, Kol72, KC73, LR71b, MK71a, Maj72a, Maj73, Man77, NK75, Pap72, RTW78, RW74, SR74a, Skl70, SCL76, SC74, Sun76, TM70, Tom75, Yam70, YO70, Yoe70b]. Cellular-Array [LR71b]. Certain [BB74a, Erc73, Hen78, Hol72b, Lam79a, Pag72, SK75, Sik70]. Certifying [Lee75]. Chain [FD77a, God73].

Chain-Coded [FD77a]. Chaining [TL73b]. Chains [Wyn71]. Chamber [BD71]. Chaney [Wor79]. Change [Ano73e, BH73, Bre75, CD73, Lil72]. Changes [Chu76, Ung71]. Channel [Gil70a, GLL79, Ric72, Smi79a]. Channel-Assignment [GLL79]. Character [ARS71a, Dos77, Dut74, Gra72, HTD72, Koz71, Nay71, Wil70, Win75]. Character-Recognition [Dos77]. Characteristic [Has74]. Characteristics [BV78, BB73, Cod73, LS71, Pea76]. Characterization [BPC75, DM78b, HA74, Kin71c, Wan75c]. Characterizing [Dir72, TB75]. Characters [DT70, TG72b, TD70]. Charge [PM76a]. Charge-Coupled [PM76a]. Chart [Lan78]. Charts [BM70]. Chebyshev [VD76b]. Check [AM73, BT70b, KM73]. Checked [OW73b]. Checker [Dav78, MF78, Usa75]. Checkers [AR77, Red74]. Checking [AM73, AR77, BCH^{+}73, Bou74a, Bou74b, BG79, CDJ71, CDS73, CHSS73, CSST73, Dav78, DAA79, Dor72, Fon71, FM73, Hol72a, Hsi71b, MF78, Red74, She78b, TB75, Usa75, Wak74, WY75]. Chest [BS76]. Chief [Ano71-49]. Chip [Woo79]. Choice [Dos77, Dui76, Sok72]. Choosing [MG71b, Tou72a]. Chromosome [KKA71]. Chronology [HH76]. Cineangiography [CC73]. CINTEL [DWT70]. Cipher [Rub79]. Circle [MB75]. Circuit [BW76a, BT74, CDJ71, CDJ73, CK77, DTG79, DW73a, FS70, Fri74, GC71b, HH77, Hou72, Kau70b, Mut76, NT70, Obe72, Ogu75, PM78a, Rus72, Saw74, SR73, SM76, SO73, TW77, WR77]. Circuit-to-Pin [Rus72]. Circuitry [BT70b]. Circuits [AB79a, AM73, Ano75-28, Arm72, ATN70, BC71b, Bet71, BF75a, BHP^{+}71, BH73, Bre70b, Bre71d, Bre73c, BW70, CM74, CM77, Cer78, CDO78, CM73a, Cha70, CE75, Cio78, Cur71b, Dan73, Dav77a, Dia75, DAA79, Du71, DW73b, EGG72, EI77, Fai73, FM74, Fis74, FD79, Fri70, Fri74, FG71b, FG71c, FG76, GR72a, GR73b, HD71, Hay76c, Hu73, HA76, Iba74, JM70, Jum70, Kai70, Kau70b, Kin70, KSM76, Kje72, KS71, LR72, Mag71, Mag73, MT70, MT71, Mar79, Mat71b, McC79, MJ77, MJ78, MT74b, MS70, NA72, NA76a, Ols70, PM75a, PR71, RHR76a, RHR76b, Red72a, Red77b, Rip72, Riv77, Rot70, Roy74, SK78a, SK78b, Saw75, SM72b, Ser73, Ser75, She70a, Sin71, Smi78, SHR72, SC72b, Su73, TM76, Tis71, To73, Toh74, TV76, Ung71, Ung77d, Ung77b, Ung77c]. Circuits [Voi77a, Voi77b, Wak74, WM74, WA73, Wyn71, YT71c, YY75, Yua74, ZKK79, Spi73]. Circuits-General [GR72a]. Circular [Red73b, SKK79]. Circulating [Myk75, Van73a]. Class [AOR74, BS72, Ben71, BV79, BR74, Che71a, Cor71a, Dev74, Ehr72, Gel73b, GT74, Har74b, Has73, Has74, HP72, IH72, KS73, Mar73a, Mat71a, PCM70, Red78a, RS71b, Tou72c, Tou75]. Classes [Bau72, Bro70a, DH79, Har73, HI72, JM71, Sag77, Tsi70, Ull72, YT76]. Classification [BC71a, BT70a, BB78b, CH72a, CH73b, DTI78, Edw75, Fri77, FK71, FO71c, FF75, Glu71, Hus72, Hus74, Kit78, LT70, LF72, MM73b, MS75b, NFS77, Oza78, PR78, PF72, Pav74, Sam70a, SH72, Yue76]. Classifier [Sto74a]. Classifiers [Cha74b, Kul78, Tou72d]. Classify [Uhr72]. Classifying [GD72]. Clause [DK72, SD74, Sic76]. Clause-Column [DK72]. Clipped [Sok72]. Clisp [Tei76]. Clock [BH71b, HA76]. Clocked [Sin77]. Closed [DBC71, Kau70b, VN78, Yan74b, ZR72]. Closed-Loop [VN78]. Closure [DW72a, RB76, RK75a, Tsi70, Yan73, YB75, YT76]. Cloud [HEWB72, SP72a]. Cluster [BCL75, KF72c, KNF76]. Clustering [Ake70, FS78, GL70, JP73, KF72b, KNF75, PS71, RL72, Wet78, WL77]. Clusters [SV75, Zah71]. Coaxial [Ano72-39, Par71].

Cobol [Sil73]. Code

[Agr76, Ano70o, Dev79, En77, FGR71, FR72, Gre75, IC78, Kau70c, MF78, Obe74, Paa79, PB76b, TOS71, Yue78]. Coded [Avi73, FD77a, FL75, LW78, Wei75]. Codes [AM73, AR77, Avi71, BM74, BM78, Bha78, BCC78, CR72c, CD78, FG71a, Gri78b, HP72, KM73, LW78, NR75, PA78, Rao70, Rao78, Red78a, Sri71a, Sri71b, Sun79, TL73b, Wak75b]. Coding [LR72, RC70, Sti78a]. Coherence [CF78a]. Coherent [FHTS78, LH75, Sri76, Win75]. Cold [Eas78]. Cold-Start [Eas78]. Collection [Tru76]. Collinearity [OC76]. Color [SM72a]. Colored [Jai77a]. Coloring [Ake74]. Column [DK72]. Columns [CW79a]. Combinational [AM77, AA75, BK76, BC71b, BK73, Bos79, BH71a, BCS76, CM77, Cer78, CDO78, CK77, Coy79, DB76, Dia75, DAA79, Dor72, DW73b, ES77, EZS78, FM74, FD74b, FD77b, Fri74, GRR72, GS74, GM71, GT78a, GT78b, Hay71a, Hay76c, Kau70b, KS78b, KS71, KK72, KB75, KK77a, KK77b, KC73, LB74, LD74, Mar79, MC71, NA76a, Ogu75, Pap77a, PM75b, PG75, Red77a, Riv77, Rot70, Roy74, RK71, SM72b, SD74, SM75c, Smi71b, Smi78, Smi79b, SS77, Sto77, SC72b, SC74, To73, Tun70b, TV76, TBY78, Ung77d, Wan75a, Wan75b, Whi71c, YT71c, YY75, ZKK79]. Combinatorial [Abr75, Gol76, Par79, Tom73]. Combinatoric [Jac73]. Combined [IC78]. Comment [Ano70o, Bas73, Bel76, Bro70a, BW70, Car71a, Hen70, Hor71a, Lec70a, Paw71, Ree70, SW70a, Sch79a, She77, Swa79a, Wad71, Whi72, Yue74a]. Comments [Ada78, Agr77, Agr79c, Ale71, Bar78, Bau77, BT70a, Bis73, Bla79, Bla74, Bos79, Bre75, Car71b, CH73b, Cer76, Cha79, Chi72, Col77, CM79c, CM79d, Das71c, DDSM76, Das78, DTG79, Dev79, Du71, DM75, DM76, Eps73b, FHTS78, Flo75, FM72b, Gar74, GS74, Gou77, Gri78b, HO79, Hla78, HL86, Iba74, Kan73a, KS72, Kro74, Kru71, LB74, LD72a, MH78, Maj72b, Maj76, MS73a, Mas78, MMB78, MST^{+}75, Nai79, New75a, NA76a, NRS74, Paa79, RB76, Rao78, Red77a, Ric72, RS73b, Sag77, SM75a, Sch73, Sha75, Shi75, Smi79b, Sur70a, Sur73, Sur78, Tab74, Tou72a, Tou72b, TV72, Tou72c, Tou74, Tou75, Tru71, Van73b, Van77, Wak75a, WY75, WK71, Wel74, Wet78, WZ72, Wil77b, Wor79, YB75, YM78, Yue76, Zoh74]. Committee [Osb77]. Common [Gri78a, JM73, Sch79b]. Communicating [SN79]. Communication [Hei79, LTE77, Mah74, MF76d, Ris74, Sah77]. Communications [Kle76]. Commutative [Kar77a]. Compact [But72, Los78, SKG77]. Compaction [AN74]. Comparative [WM76]. Compared [CS70]. Comparison [Bar75a, CCES74, Che71a, FD75, GG75, GF78, Mil73, MG71b, Sie79, Tou72a, Tou72c]. Comparisons [Dig73]. Compatibility [Bro70a]. Compatible [Rub75b, Tom72, YT76]. Compatibles [Han75, RS72b, RS73b, Sto74b, Sur73, Yan75a]. Compcon [Ano79b, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano77f]. Compensating [Sie75]. Compensation [ASS72, KES72]. Compilation [RPL73]. Compiler [And72, Feu72]. Compiling [Lew74]. Complement [BW73, Bla74, DM75, DM76, Eps73b, GG75, Hwa79, JW76, Kro74, MG75, SS72]. Complementary [CCB70, Kri71a]. Complementation [HO72]. Complemented [Vin78]. Complete [Ano76t, Ano77u, Ano78-42, Ano79-47, Ano79-48, Ano79-49, Ano79-50, DDSM76, Fon71, IS75, Mar71, MST^{+}75, Pu72, Red73a, Rob79a, Tom70, YT71c, Ano78-43]. Completely [LH76, Set70a]. Completeness [KF78]. Completion [Ung77d]. Complex [ARS71b, Bau77, CHSS73, Hig73, YT77, Yue75c].

Complexities [Liu77a]. Complexity

[Arn70, AO79, BV79, CJ74, FK78b, Ham71, Han72, Hay78b, Lam79a, Mor70b, Ric74a, SM73b, Smi71b, VD76a, Wak70a]. Component [AO79, CK77]. Components [BF75a, FP75]. Composed [Ish77a]. Composer [BL75]. Compound [Hus74]. Computable [Ehr76]. Computation [Ano73w, Arb76, Ban74c, Boh77, Buz73, Cha73b, Dru71a, Eas78, God71, God73, Goo70, Hen70, Hun72, KV76, Kun75, Lec70a, Lit73, Mey76, Mil73, Muk79, Pra72, SF71, SR73, Spi73, Ulm70, VD76b, WS76]. Computation-Based [Mey76]. Computational [BV79, CY76, CDL78, DW73b, ES77, FK78b, Hel72, Hel74, Mas78, Moo76, Mor70b, PBC74, PD77, SG74, TL73a, VD76a, Wel74]. Computations [CA74, Der73, Erc77, Kri77, Ric74a, Roo73, Vee78]. Compute [Gri78a, Spi71]. Computer [AAK74, AD75, AB76, Agr79c, All75a, AK75, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano72-50, Ano72-57, Ano72-58, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74j, Ano74p, Ano74k, Ano74q, Ano74l, Ano74r, Ano74m, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74n, Ano74y, Ano74o, Ano74z, Ano75o, Ano75p, Ano75q, Ano75i, Ano75r, Ano75j, Ano75s, Ano75k, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75l, Ano75y, Ano75m, Ano75z, Ano75n]. Computer [Ano76c, Ano76d, Ano76h, Ano76k, Ano76l, Ano76i, Ano76m, Ano76j, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano76t, Ano76u, Ano77b, Ano77c, Ano77d, Ano77g, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano77u, Ano78k, Ano78y, Ano78z, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78-32, Ano78-33, Ano78v, Ano78-34, Ano78w, Ano78-35, Ano78x, Ano78-36, Ano78-42, Ano78-43, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42, Ano79-43, Ano79-44, Ano79-47, Ano79-48, Ano79-49, Ano79-50, Ano79-53, Ath75, AG73, AGM^{+}71, BHK79, Bar74b, Bas71a, BS75, Bha75b, Bha78, Bib74, BS77, BO70, Bon74, BE70b, BB74a]. Computer [BR77, BW77, Bre73a, CG74, CS73, Cas73, CM74, CR72a, Cha73b, Chi78, CA77, Col77, CSY75, DRS75, Ekl72, EZS78, Els74, Erc77, EFLR76, FD74a, Flo72, Fly72a, Fly72b, Fre77, Gar76, Gar78, GGMM71, GM72a, GSB^{+}76, Gil74b, GD72, Guf71, Hal74, Ham74, HC70, Hei79, Hob73, HP72, Hun73, HB75b, JP76, KES72, Kar70, KR71a, Kar77a, KC71, Kim74, Kje72, KKA71, Koc71, Kon71, KP71, Kor76, KK70, Kra74, KF73, Kul77, Kuo74, Lam79b, Lar71, Lew77, Lin70, Lom74, Lou70, Lud74, Mai71, Maj72b, Man77, Mar78, MB75, MA77, MJTW75, MR74, MG71a, Mil75, MM70, Mui74, NR77a, NS79, New75a, NS75, NK75, PM78b, Pla73, Ple77, PGS70, RM71, Ram73a, Ram76a, Rao70, Rau79a, RA78, RF78, Ric74a, RC71c, Rip72, Ris74, Ros74a, Ros71]. Computer [RH73, ROW71, SF74, SW70a, SW70b, SK75, Sch73, SF78, Shi70, Sho74, Sta75b, Sto70b, SN71, SC72c, SSR72, Swa74b, Swa90, TI72b, TM70, TK74, Tri77b, Tri78, Ung77a, Van77, Vic74, Voi75, WBRG76, Whi70, WG75, RT73a, DWT70]. Computer-Aided [EZS78, Lar71, Mar78, SN71]. Computer-Communication [Ris74]. Computer-Controlled [Ros71, WBRG76]. Computer-Drawn [Lou70]. Computer-Maintained [Agr79c, Man77].

Computer-Oriented [Bre73a, Gil74b].

Computers

[Ano75b, Ano76e, Bar75a, Bas71b, BS78, BH76, Bla79, BCJ^{+}71, Cor71a, DBC73, Des74, ET75, Feu72, FM72a, Gra70, GM70, HG78b, Hec73, HHJ^{+}76, Hou72, Jos72a, Jos72b, Luc72, Mon74, Mor79, Orc76, RC71a, RS71b, Sar71, SGR78, Wex71, Ano70a, Ano70-29, Ano70-30, Ano70-31, Ano70-32, Ano70-33, Ano70-34, Ano70-35, Ano70-36, Ano71a, Ano71z, Ano71-27, Ano71-28, Ano71-29, Ano71-30, Ano71-31, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano72a, Ano72-59, Ano72-60, Ano72-61, Ano72-62, Ano72-63, Ano72-64, Ano73a, Ano73m, Ano73n, Ano74a, Ano75a, Ano76b, Ano77a, Ano78b, Ano79a]. Computerss [Ano70-37]. Computing [Agr78a, Ano73h, Bak79, Bea78, Ber71, Car73, CH75, CM79c, FN75, Har70b, Hay76a, Hob73, Hun71, HH76, Ken74, Kim75, Lei74, LR75, MN70, Met73, MR76, Ram71, Red78b, RT79, Sch74a, SM71b, SK74, Stu75, Tan78, TM78a, Yue75a, Ano72-50, Tab70b]. Concave [VD76b]. Concavities [SCL76]. Concavity [Skl72]. Concentrators [TWB78]. Concept [Arn73, Hop71]. Concepts [UM77, Vai71]. Concerning [Maj76, YM78]. Concurrency [TF73]. Concurrent [RH75a]. Conditional [FG71a, KO73, RF72]. Conditional-Sum [KO73]. Conditions [Hu73, Smi79d]. Cone [Uhr72]. Conference [Ano79c, Ano79d, Ano79-54, Shi70]. Confidentiality [Ple77]. Configuration [BW76a, HB75a, Sch79b, Sti78b, Wel78]. Conflict [Kur74, Nut77]. Conflicts [DI71, FT79]. Congruences [Reu72]. Congruential [Gal78, NM76b]. Connected [ADL72, NS79, RS72a]. Connection [FK78a, HA74, Rub74, Rub76a, Tho78, Tom70]. Connections [Abe72, HO74, KJ79]. Connectives [Set70b]. Consecutive [WG74]. Consensus [CM79d, Mag70].

Consequences [RSL76]. Conservative [SK78a, SK78b, SK79a]. Considerations [DR74, HB75b, Kle70]. Constant [Chu70, MH70a, TL73b]. Constant-Weight [TL73b]. Constrained [BI72, Bro75a, BW70, CK74, Hun73, Ram77a, Whi71c]. Constrained-Input [Bro75a]. Constraint [PMB77]. Constraints [DT70, SN71]. Construct [PR76]. Constructed [MF76b]. Constructing [BF78, PM77b]. Construction [Ban78, Ben70, Feu72, HC78, Oja79, Sto70d]. Constructions [JW76]. Content [LC79, Red73b]. Content-Addressable [LC79]. Contentions [FT79]. Context [Ano70z, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46, Gre75, Har70b, Sel74, Tho75]. Context-Free [Gre75, Har70b, Tho75]. Contextual [Chu71, DT70, Dos77, EK75, RE71, RH74, Tou72b]. Continuation [WK73]. Continued [BB74a, BB74b, Tri77b, Tri78, VV74, RT73a]. Continuity [Tun70a]. Continuous [AOR74, Gou71, JR76, Lew77, SKM78, Tom74]. Continuous-Parameter [SKM78]. Contour [DT70, TD70]. Contour-Traced [DT70, TD70]. Contradiction [Hwa72]. Contributions [Gar76]. Contributor [Ano73f]. Contributors [Ano70p, Ano70q, Ano70r, Ano70s, Ano70t, Ano70u, Ano70v, Ano70w, Ano70x, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r, Ano71s, Ano71t, Ano71u, Ano71v, Ano71w, Ano72-40, Ano72-27, Ano72-28, Ano72-29, Ano72-30, Ano72-31, Ano72-32, Ano72-33, Ano72-34, Ano72-35, Ano72-36, Ano72-37, Ano72-38]. Control [Ast78, BK79, BT70b, BW76b, BA71, Cam72, CHSS73, CSST73, Cov74, DBC73, HG78b, Har74a, JB76b, JM73, Jum74, KP71, Kor76, LS76a, LNL72, LT73, Len78, Les71, LM78, Man72a, MG71a, NM77, PR72, Red78a, RS75, RS78, Svo70, Swa79b, Tab70a, Tab74, Toy71a, WY73].

Controllability [Cer78, Gal78]. Controlled [Ath75, Bre75, CD73, Ros71, WBRG76, WG75]. Controllers [Smi79a]. Controls [Gil70b]. Conventional [Ish77a]. Convergence [GW77, GH72, RF75]. Convergent [MC70]. Converging [NW79]. Conversational [Tei76]. Conversion [Aar70, Ben77, Hor71a, IC78, Maj73, Mat70, Mor75, Nic71, Paw71, Rhy70, San76, Wad71, Yue78, Zoh70, Zoh73d]. Conversions [BR71, Kri71a]. Converter [Mar74, TS75, TP72]. Convertor [WS76]. Convex [Sch78a]. Convolution [DV78b, PH73, Rem74, Zal71]. Convolutions [Rad72, Rad73]. Cooperation [Dos77]. Cooperative [Arb76]. Coordinate [BF78, EO78]. Copy [DLL^{+}73, GS79]. Copyright [Ano79-45, Ano79-46]. CORDIC [Des74]. Core [Mcd72, RB78]. Coreless [Pin71]. Corner [FD77a]. Corner-Finding [FD77a]. Corners [Dun75]. Correct [Cha74a, Kin71a, Max77, Ogu75, Pat76]. Correcting [BM73, BCC78, Dev79, Gri78b, KR71a, Kau70c, LW78, LF77, LF78, NR75, Rao70, TF78, WL74]. Correction [Ano70y, Ano72-39, Ano72-40, Ano73f, Ano78t, Ano78s, BM74, Bei71, BK71a, Cat77, CH72b, En77, HK78, Hla70, LL75, MF76a, Man72b, MJ78, OTK76, Paa79, PB76b, Rad73, Ram77b, Rao72, RH74, Rub76a, SK77a, Saw77, Ser73, Ser75, She78a, Sri71a, Sri71b, Tho75, Tho76, WSB79, YM78, YL73, ZHR77b]. Corrections [MH78, RHR76a, RL78, Tri78]. Correctly [Lam79b]. Correctness [Hus72, RS74]. Correlated [MM73a]. Correlation [Koh72b, RF75]. Correlations [Ken74, Yue75d]. Correlator [Win75]. Correspondence [Hwa72, Rob70a]. Corresponding [San76]. COS [MJ78, MJ77]. COS/MOS [MJ78, MJ77]. Cosine [ANR74, Har76, Sha75, SP77, TM78a, Bak76].

Cost [Avi71, CF73, CY76, CDL78, KAN79, LL79, Mas78, ML76, PA78, SA74, She71, Wak75b, WM71]. Cotes [Hun71]. Count [Hay76c, Hay78a]. Counter [Ano70o, Hol72a, Toy71b]. Counters [Byr71, CM79b, Dor72, HK77, HK78, KO78, MF76b, O’K73, RC70, Swa73a, VV74]. Counting [BO79, Byr71, Dev79, FS71, Kau70c, Myk75, PS78, Red77b, Zoh73a, Zoh73b, Zoh74]. Coupled [Den76, Gou77, PM76a, SG74]. Coupling [MMO74]. Course [Fre74a, Wal74]. Cover [Ano78t, Das71b]. Coverability [Sil79]. Coverage [AM79a, AM79b, Arn73, Sti78b]. Covering [DW72a, Du72, Reu75, Sch75, Tom73]. Covering-Closure [DW72a]. Creases [Huf76]. Creating [Mil75]. Criteria [Che71a, MH70c, Tou72c]. Criterion [Deu71, FE78, FK70b, VV71]. Critical [BH71b, Koh75]. Criticalities [Wan75b]. Cross [MMO74, Wan77]. Cross-Coupling [MMO74]. Cross-Section [Wan77]. Crossing [NS75]. Cube [Pea77, Wyn71]. Cubic [All75b, Wil77b]. Current [Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano77v, Cat77]. Currents [Lei74]. Cursive [EK75, Far79]. Curvature [BD75a, Huf76]. Curve [But71, Can71, Dan70, DR77, RT71, RTL72, Sha78b, Skl78, ZHR77a, ZHR77b]. Curved [AB76, Gou71, MA77]. Curves [Dub79, FHTS78, FD77a, JLH73, LH75, MF75, PH74, Ram76c, RJ73, RW75, Rub76b, ZR72, Bel76]. Curvilinear [CA77]. Customized [Lar71]. Cut [Cur79, Das73]. Cut-Sets [Das73]. Cutpoint [Pap72, YO70]. Cybernetics [Rei74b].

Cycle [HO79, Sil77, Smi71a, WR77]. Cycles [Hol72a, Per70, Van73a]. Cyclic [RRNA74, SAPR70, SF77, TL73b, YC76, Yue75d, Zal71]. Cyclic-Order [SF77]. Cyclically [WY73].

D [Hil79, Mar74, Moo76, Mor75, NR76b, Zoh73d]. D-Script [Moo76]. D/A [Mar74]. Damage [Smi75]. Damage- [Smi75]. Data [Anz71, Bas71b, BH76, BT70b, BCL75, CR72d, Chu79, DR75, DB74, Dow73, Dru71a, DB73, Far75, Fen74, FK70b, FO71a, FO71c, GS79, GSB^{+}76, Gol72, HND75, HEWB72, Hig76, How75, HTD72, JM71, Jos74, KR73, Kru71, Law75, Len78, LJ79, LG74, McL72, Mui74, NW79, OF73, Pat72, Ram73a, Ram77b, RNR75, Rum77, Sch78b, Sch74b, Sha78b, She78a, Sto70c, Sto72, SSR72, Tru71, Wan72, Whi72, WL77, FT74]. Data-Storage [Anz71]. Database [BHK79, BH76, De 79, FSLC78, Lan79]. Databases [BHK79, SNOS79]. DBC [BHK79]. DDA [MMB78, HO74, HW76, Max77, MB75]. Deadlock [LL79]. December [Ano76w, Ano78t]. Decentralized [GR72b]. Decimal [Hor71a, LW78, Rhy70, SW71a, Yue77c]. Decimal-to-Binary [Hor71a, Rhy70]. Decimations [AM79c]. Decision [Ake78, BS76, Bit77, CMS79, Chi71, Ehr76, Fri77, KS78a, KKA71, MM73b, PCM70, PS71, PM77b, WL77]. Decision-Directed [PCM70, WL77]. Decisions [EF79]. Declustering [FE78]. Decoder [CDJ71, Wym73]. Decoders [GT78b, Voi77a]. Decoding [FF75, Lin70, SM72a, Sun78]. Decomposability [Kol72, New70]. Decomposable [DF75]. Decomposition [Cur76, Cut70, DF70, FP75, GB71a, HSP78, Hig73, Kan76, Kin70, Kje72, Lan78, Mar70, Paz71, Per70, Ram73b, RS72b, RS73b, Sch78a, SM70a, SMW71, Sur73, Vic71, Wil75, Yoe70a]. Decompositions [Paz70b, Rud76, Tur70]. Deconvolution [Dep70]. Decrypting [Rub79]. Deductive [Arm72, BCSW76, CCES74, MC78, Rei76]. Defects [Sri71a]. Defined [Dav77a, Smi71b]. Definite [Brz70, Pre70]. Definition [But72, Car71b, Dav72a, Ern76, KS72, LM71b, O’C75, Pag70b, Van73b]. Definition-Driven [Ern76]. Definitions [DM75, DM76, Eps73b, SS72]. Degraded [Jai77a]. Degrading [BF75b]. Degree [AN74]. Degrees [AC71]. Delay [Ake70, Bre74, BW70, Dav70b, Hla70, PM71, Sch79b, Ser73, She70a, Vil70]. Delay-Free [BW70]. Delayed [LV75, MBM77, Sin71, Sur78]. Delayed-Input [Sin71]. Delays [Bat70, Bre74, BW70, SS71]. Delegates [Ano72n]. Deletion [Gel76]. Deletion-Directed [Gel76]. Delta [Lee78a]. Demand [Yan74a]. Denominator [Gri78a]. Density [AB79b, Dui76, FS74, SM71b, Woo79]. Dependence [BCL75, Das71a, HG78a, TV72, Vil72, Vil73b]. Dependent [CB72, Hus74, RC71b]. Depth [PMB77]. Derivation [Ben71, Bet71, Dan73, DDSM76, Hil79, Mar71, MST^{+}75]. Derivatives [Tha78]. Deriving [BG79, Cur79, Cur77, YT76]. Describe [Uhr72]. Describing [Bar75a, SN72a, Zah71]. Description [AB76, BEGW72, Ber77, GL74, JK77, SF78, Sta70b]. Descriptions [Moo76]. Descriptor [Ano70z, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46]. Descriptor-in-Context [Ano70z, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46]. Descriptors [Tan78, ZR72]. Design [AM73, And74, Arm73, AM79c, Ast78, Atk70, Aus70, AT70, Avi71, AGM^{+}71, Bat70, BC71b, BT74, Ber70b, Bha77, BT76, CJW^{+}71, CR72b, CDS73, Cha74c, CD78, CS79, CF70a, CSST73, Cor71a, CLL78, Coy79, DR74, Dav77a, DAA79, EZS78, Elm76, Els74, FM72b, FK74, GHJ77, GD72, Gon70, Gra70, GT78b, HSS78, HP75, Hol79, KH74, Kam79, Kin71b, Kin72, LY71, Lar71, Lec70b, LD72b, Lem70, Lud74, Maj76, MA71, MH70c, MF78, MPS78, McC79, MJ77, MJ78, MI72, Nay71, New76, NM77, O’K73, PR74b, Pra78a, Pre71a, RTW78, Red72a, RSL76, Saw75, SM71a, SF72, Sho74, SY75, SG71, SM73a, Sie74, Sto74a, Sto77, Su75, ST76, TG78, TM70, Toh74, TW77, TK74, TS79, Toy71a, Toy71b, TG79, Usa75, VV74, Voi74, YC76, KD79]. design [RT73a]. Designed [SM74, Wak75a]. Designer [FS70]. Designing [BF75a, Far73, Hu75, KS71]. Designs [Abd71, IS76]. Detect [Dun75]. Detecting [BDR77, DS71, Gad70, GL70, MF75, SR75, Zah71]. Detection [AB79a, BM74, BK73, Bos79, BH71a, BHP^{+}71, Bre71d, DF75, DB76, DR75, DV78a, DW73b, Ehr77, FM74, Far73, FD75, FC77, GRR72, GS74, Gon70, GT78a, Gri73, HF74, Hay75, HC71b, HPP77, IS75, Ios78, KP74, Kar77a, KS71, KK72, KR78b, LL75, LB74, Lil72, MF71, NR77a, OS76, PA78, Pra78d, PG75, RH75a, Ram76b, RM78, RT71, RTL72, RJ73, RW75, Ros75, SM74, Sha78b, SV74, SCL76, Skl78, Smi79c, SS77, Su73, SKM78, TF70, Toy71a, Toy71b, TM78b, TBY78, Usa75, Wak75a, Wak75b, WR77, YY75, Zis76, ZKK79]. Detector [DF78]. Detectors [RJ72, Sur70a]. Determination [Bro70a, Bub72, Cho76, DBC71, IR76, Maj72c, MT74a, MBM77, Mor70c, NRS74, Sar71, Sik70, Sto74b]. Determine [Set70b]. Determining [BK72, Bor72, HEWB72, Han75, PC72, RS72b, RS73b, SV75, SO79b, Sur73, SB74a, Tom70, Tom72]. Deterministic [Dav72b, Kri73, Moo71, San76, SCK73b]. Developing [Ben77]. Development [Pat76]. Developments [Cox72]. Devices [Bak79, Bar74a, DS72a, Ern72, Fai73, Flo75, HH76, Kar77a, LC79, PM76a, Stu75]. Diagnosability [AKT75, BGM76, Hay74, RK75a, RK75b]. Diagnosable [CS79, FM72b, FK74, FK78a, FK78c, HA74, Jos78, LY71, MH76, MH78, MM78a, Pai73, SM71a]. Diagnosing [Mar79]. Diagnosis [Ake74, Cha70, CRB72, Fri74, FK78b, GST79, Hay71a, Kam75, KF79, Kim70, Kim79, KB75, KK77a, KK77b, MH76, MH78, MS75a, MM78b, MKO70, RM71, RC71a, Rot70, Roy74, RK71, RK75a, RK75b, SK77c, Smi79e, ST76, Toi76, YO70]. Diagnostic [BK72, MK79, Rob70b, Sel74]. Diagnostics [HF71]. Diagram [MS73b, MA71, RRNA74, WR77]. Diagrams [Ake78]. Dialogues [Els74]. Diameter [TS79]. Dielectric [YRT75]. Difference [CRB72, HC71b, KM75b, Wes78]. Differences [DDSM76, Mar71, MST^{+}75]. Different [DM75, DM76, Eps73b, Ges70, LHSM74, SS72]. Differential [All75b, Bak79, BW76a, BM77b, DEA73, Fra78, Kar70, MB75, MN70, New75b, RS71b, SF74, TD73, Uhr73, Wil77b, Wit70, Wor76]. Diffraction [WKP75]. Digit [Tun70b]. Digital [Aar70, Agr75b, Ake70, AP76b, AP76a, ACJ75, Ano78v, Ano78w, Ano78x, Ath75, Avi71, Bar75a, BS72, BGM76, BV78, BCA^{+}75, BB74a, BdB70, Bre70b, Bre73c, But79, Cas73, CS75, CR78, CDO78, Cha74c, DBC73, DTG79, Dav71, DEA73, DRS75, EH78, Erc77, Fai73, Feu72, Fre79a, GSB^{+}76, GSJ^{+}76, Gil70b, GM70, HND75, HLD70, HO79, HC70, Hun72, Hun73, HB75b, Hun77, Jai77a, JA78, KR71a, Kar77a, KF79, Kim74, Kim70, Kim79, KV76, LL75, LC72, Lit74, Maj72b, Mar73c, Mat71c, MMO74, MN70, MK79, Mon74, Mor76a, ND73, NA74, PGS70, RHR76a, RHR76b, Red78a, RC71c, RS71b, Rob70a, RJ73, Ros74b, RW75, RH71, SF74, SAPR70, Sar71, SM72b, SM71b, SM76, SC77, SG71, Sie74, Sil77, SP72a, Sta75b, SP77, SM70b, Stu75, Su73, Swa74b]. Digital [ST76, TL73a, TP72, TW77, Tri77b, Tri78, TE76, Var79, Voi74, Wai70, Wex71, Whi70, Yue77d, Zoh73a, Zoh73c, Zoh76a, NA76b, MB75]. Digital-to-Analog [Aar70, TP72]. Digital/Optical [Cas73]. Digitally [CB72]. Digitized [Ami77, AN74, SCH72]. Digits [Sok72, Sut74]. Digrams [RE71]. Digraphs [DS73, PC78]. Dimensional [Ano72-39, BDR77, CC73, DR75, DM78b, FO71c, Goo70, HND75, IH72, Jai77c, Kun79, MS78, NA74, NW79, Par71, Roe78, SR74a, Sar71, SM72a, UM77, Vil73a, Wei75, NA76b]. Dimensionality [CA74, Ehr76, FO71a, JA74, SV75, Tru71, Tru76]. Dimensions [Gra71]. Direct [AK75, BH71c, FHTS78, LH75, Les71, LR75, Sho74, SB77b, Whi71b, De 79]. Directed [CR72d, Gel76, PCM70, Tho75, WL77]. Directly [Bae74, Hoe74]. Discrete [AP76b, BLZ75, Boo70, CF79, DF70, Dot70, DV78b, Duk74, Kun79, Law70, Rad72, Rad73, RT79, Rei74a, Sha75, Sto74a, TL73a, TD77, Tou72d, Van75, VR79, WL77, ANR74, Har76, Hun71, TM78a]. Discrete-Time [DF70, Dot70]. Discrete-Valued [Tou72d]. Discretization [HO74]. Discriminant [Cha73a, FS75, Kit77, Kit78, Sam70b, VD76b]. Discrimination [Vil72, ZRD75]. Discussion [Han75, Hel74]. Disease [SH72]. Disjoint [Eps73a]. Disjunctive [But75, Hig73, SM70a, SMW71]. Disk [FS74]. Dispatching [FR72]. Displacement [Ano77v, Cat77]. Display [FO71c, HSS78, HB75b, JR76, KF73, LH79, LG74, New76]. Displaying [JR76]. Dissent [HP73]. Distance [Dev79, Kau70c, KF72a, LR76, TL73b, Tou72e, VV71, ZT75]. Distance- [TL73b]. Distinct [Ban78, RN77]. Distinguishing [Bou74a]. Distributed [Coo77, RSH79, Svo70]. Distribution [ABK76, BT72, JW73, Obe72, Rom78, SP72b, SP72c, Wei75].

Distribution-Free [ABK76, BT72]. Distributions [BD75b, Hal74]. Diversified [RS75]. Diversity [WL75]. Divide [Cha74a, She71]. Divide-and-Correct [Cha74a]. Divided [Wes78]. Divider [Ste72]. Dividers [Boy74]. Division [AKT^{+}77, BJ75, BKM73, CH73a, Cha74a, Chi72, Fly70, Gar74, Jac73, KKK73, Kri70a, Kri70b, Kri71b, Lig77, Rhy71, Rob70a, SCK73b, SR72a, Sit74, TE77, Tun70b]. Divisors [Jac73]. Does [Ano75c, Ano75d, Ano75e, Ano75f]. Doesn’t [Ano76t, Ano77u, Ano78-42, Ano78-43, Ano79-47, Ano79-48, Ano79-49, Ano79-50]. Domain [KH74]. Domains [Gar72a]. DON’T [Vin78]. Dotted [FM72a]. Double [Cur73b, GH78, HC70, MF76a, Myk75, PB76b, Van73a]. Double-Precision [HC70]. Double-Rank [Cur73b, GH78]. Down [Har74b]. Drawing [RL72]. Drawings [EUY^{+}72]. Drawn [Lou70]. Driven [Ern76]. Drivers [Wym73]. Drum [Ful72]. Dual [FS74]. Dual-Density [FS74]. Dualize [HPP79]. Due [CB72]. Dyadic [GC77, Pea75, RRNA74]. Dynamic [AU74, BBLT73, BE70a, BH74b, CJW^{+}71, CK79, Cod73, CF70a, DI71, GSB^{+}76, Hla78, KK79a, LT73, Len77, MVW79, RA78, RS78, RH73, Si78, Sto72, Sto75, Tab74, Uhr73, Vic74]. Dynamical [HN78]. Dynamically [BR76b, OW73b]. Dynostat [GM72a].

E-Nets [NN73]. E2 [Sil79]. Easier [Ano75c, Ano75d, Ano75e, Ano75f]. Easily [Coy79, Fri73, FNSK75, Jos78, Kod74, Red72b, RW74, RH75b, SR74a]. Easy [Lio76]. Economic [SK71]. Economical [Kri71b]. ed [Jos72a, Luc72]. Edge [DF78, FD75, Gri73, Nie75, PR78, RT71, RTL72, SR71]. Edge-Sensitive [SR71]. Edges [ADL72, DR75, Dun75, Ehr77, OC76, Sto77]. Editing [Wil70]. Editor [Ano70-27, Ano71x, Ano71y, Ano71-49, Ano72-47, Ano72-48, Ano72-49, Ano73g, Ano74g, Ano74h, Ano74i, Ano75g, Ano75h, Ano76f, Ano77i, Ano77j, Ano77h, Ano78u, Ano79-29, Ano79-30, Ano79-31, Ano79-32, Shi75]. Editor-in-Chief [Ano71-49]. Editorial [Ano73e]. Editors [Ano76h]. Eds [Tab70b]. Education [Ram76a]. Effect [Arn73, BH71a, BCHL73, Gra70, LB74, Mar77]. Effective [CKL77, JA78]. Effectiveness [Avi71, Fly72b]. Effects [Bre74, FLBK76, Hun71, Sie77]. Effects-of [FLBK76]. Efficiency [Bit77, FLR78, Kan75, KR71a, Los78]. Efficient [DH79, EO78, God73, Gri78a, Har76, Hur74, KS73, Los76, MF78, MM78b, NTA78, NS75, PGS70, PB76a, PM76b, RC71b, Sha78a, STG71, Sil79, SO79b, Weg72, YT71c, Bak75a]. EFL [Elm76]. Eigenvalue [Has74, The75]. Eigenvectors [BC71a]. Eight [MTB70]. Ek [Hil79]. Ek/D/1 [Hil79]. Eklundh [Ram75, TE76]. Electrocardiogram [Ges71]. Electrochemical [Nel74]. Electronic [CDS73, CHSS73, KES72]. Element [AT70, HK77, HK78, KT72, Nel74, Tak76]. Elementary [Bak75b, Erc73, Mar72, Meo75, Oja79, Bak75a]. Elements [Ama72, Dav70b, DVS74, GC73, Gra73, Gri74, Hen71, Hla70, HS75, Ish77a, JK77, Lyo70, OY70, SK78b, SK79a, Ser73, She70b, SR72b, Win70a]. Eliminate [BH71b]. Eliminating [BH74b, Hla78]. Elimination [HD71, Jol76, Kan73b]. Ellipses [TM78b]. Emerging [FLBK76]. Emitter [ET75]. Emitter-Function [ET75]. Empirical [Gri76]. Employing [HA76, NA76b]. Emulation [FLBK76, Hoe74, Les71]. Emulator [SB71]. Encoding [KM73, KKA71, SM72a, STG71, SB72]. Encryption [Ple77, KD79]. End [She77]. End-Around-Carry [She77]. Engineering [Ano76t, Ano77u, Ano78-42, Ano78-43, Ano79-47, Ano79-48, Ano79-49, Ano79-50, Boe76, Jor74b, Lom74, Ram76a, Ano78-39, Ano78-40, Ano78-41]. English [Hub76]. Enhance [FR72]. Enhanced [Sto72]. Enhancement [Hal74, HB75a, KCLB71, KS77a, ZHR77a, ZHR77b]. Enhancing [WA73]. Ensemble [WL75]. Entirely [MF76b]. Entropy [CF73, Mas78, WD77]. Entry [Ano73f]. Enumeration [AA72b, MTB70]. Enviromnent [Sch79b]. Environment [AI78, BD75a, BCHL73, KM78, Tun70a]. Environments [LD75, Li77, NC78, Nev76]. Equal [KP75, Per70]. Equation [BM77b, BLZ75, Kar70, MR75, Sch73]. Equational [Bro74, Bro75b, Cer76, Eps73a]. Equations [AH76, Bak79, BB74b, Bro70b, Bro71, CW77, Den76, Fra78, Gou77, Has73, Hwa72, Kar77a, KS73, MS73a, Max77, MN70, MC70, New75b, Pic74, RS71b, Rud74, SF74, SG74, Sau72a, She70c, Sur70b, Uhr73, Ulr71, Weh74, Wit70, Wor76]. Equivalence [Bau72, Kun79, Mat71a, MC71, Moo71, MS70, NSR74, RS74, Roy74, Yan75a]. Equivalent [Bar71, BT72]. Equivalents [Gaj75, GB71b]. Erasure [Sun78, WSB79]. Erlang [MH70a]. Errico [Ano73z]. Error [Avi71, Avi73, BM73, BM74, BCC78, CB72, CH72b, Chu71, Chu74, DTG79, Dev79, DTI78, En77, FK71, Gad70, GY77, Gri78b, HO74, KR71a, Kar77a, Kau70c, LW78, LF77, LF78, MF76a, Man72b, Man72a, MM73a, Nad71, NR75, NA74, NA76b, Paa79, PA78, Pea76, PR72, PB76b, RH75a, Rao70, Rao72, Red78a, RH74, SM76, She78a, Sri71a, Sri71b, SR77, Sun78, TF78, Tou71, Tou72b, Toy71a, Toy71b, WSB79, WL74, YL73, Yue75a]. Error-Coded [Avi73]. Error-Control [PR72]. Error-Correcting [Gri78b, KR71a, LW78, LF77, LF78, NR75, Rao70, TF78]. Error-Correction [MF76a, PB76b]. Error-Detecting [Gad70]. Errors [HC71a, KES72, KP75, Lit73, PA78, Sri71b, Tho75, Usa75, Wak75b]. Essential [Hla70, Maj72c]. Estimates [MR75].

Estimating [BD75b]. Estimation [AOR74, Boo70, BR74, Dav72b, DM78b, FK71, GY77, Hun73, KS77a, Maj72b, MS76, NJ77, PCM70, TL73a, Tru76, Whi70]. Estimators [Dui76]. Evaluating [BW77, FV79]. Evaluation [AS74, BB74a, BG78, BKM73, Erc73, Erc77, FD75, Gay78, Gel73b, Kim74, Koh75, Kor79, Mar72, Mar73d, Moh71, Nec79, PD79, PM76b, SB77b, Tou72e, Vil73b, Win76, Yue74b]. Events [Paz70b, Paz70c, Paz70a, Rey70, Whi71a, Yam70]. Every [LF74]. Exact [Kri77]. Exactly [RW74]. Examination [CSJ74, KAN79]. Example [MR74, REFN76]. Examples [Fly72a]. Exchange [Ano72n, Che73, Lan76, LS76a]. Excitation [Bre75, CD73]. Exclusive [Fis74, MS70, RR74]. Exclusive-or [RR74]. Executable [Bae74, KMC72]. Executed [Hoe74]. Executes [Lam79b]. Execution [Cha72, FR72, GR72b, LM78, Tho74, TF70]. Exhibiting [Smi78]. Existence [FK78c, RS72a, Smi78]. EXOR [Pra78d]. Expandable [Kau73]. Expanded [Ano72-50, Ano73h]. Expanding [Jos72b, RB72]. Expansion [FK70a, Ges71, God71, Gri78a, KT72, NFS77, SO79a, TM76, VV74]. Expansions [AP76a, Swa72a]. Experience [MM70]. Experimental [CM76, CW75, Dut74, Smi79a, WG75]. Experiments [AB79b, Bou74a, Bou74b, BG79, DF75, EK75, Far73, FT76, Gon70, Hol72a, Hsi71b, MOW76, RTL72, TB75, WKP75, WY75]. Explanation [Obe74]. Exploration [Gri76, SF72]. exploratory [FT74]. Exposure [RK75b]. Expression [Hwa74]. Expressions [BEGO71, BKM73, GLR71, PM78a, Pat71, PC78, PM76b, RPL73, SB77b]. Expressive [Set70b]. Extended [MN70, Mcd72, Roc70]. Extending [Ano79b]. Extension [LG74, OW73a, O’K75, SD74, Sli70, TE76].

Extensions [She70a]. External [Dru71a, Pat73]. Externally [DS74]. Extra [FK74, FNSK75]. Extract [Ges71]. Extraction [ADL72, Bab72, Cal70, DW72a, FE78, GP74, HKD^{+}71, Hey71, KF72a, Koz71, MK71a, Maj72a, Mar78, Pre71b, The75, Tjo77, TG72b, Tou74]. Extractors [GY77, You71].

Fabricated [Ano72-39, Par71]. Face [SF77]. Facility [MCRD72]. Factor [FLR78]. Factoring [HW75]. Factorization [Dru71b, Jai77b]. Factors [PB76a, RK71, Yue73a]. Fail [Chu76, CD78, DGC74, DAA79, GCP^{+}77, HG79, MT74b, PM74, SM75b, TM71, TI72a, TOS71, Toh74]. Fail-Safe [Chu76, CD78, DGC74, DAA79, GCP^{+}77, HG79, MT74b, PM74, SM75b, TOS71, Toh74]. Failsafe [TKH71]. Failure [Koc71, LR72, RC70, TA71]. Failure-Tolerant [LR72, RC70, TA71]. Failures [BCC78, CLL78, Rot70, Sie75, SS77, Wak75c]. Fall [Ano79b, Ano77f]. Fan [Ast78, CK74, FM74, KS78b, NA72, PMB77, SN71, Wan77]. Fan-Beam [Wan77]. Fan-In [CK74, PMB77, SN71]. Fan-Out [Ast78, FM74, KS78b, NA72, SN71]. Fanout [BB78c, BK73, Mar79, SK79b]. Fanout-Free [BB78c, BK73, Mar79, SK79b]. Fast [ARS71b, BS72, BF78, BM70, BB74a, Bro77, Buz73, Car71a, Cor77, Des79, DV78b, Ekl72, FA76, FC77, GM72a, GLMR71, GD73, HW70, Hab77, Hen70, Jai77a, Lec70a, Len77, Lig77, Maj72b, Nec79, NW79, Nus78, O’K75, PBC74, PD77, RTW78, RT79, RNMP77, Sar78, Sch73, SMW71, Sto75, Sur75, Whi70, Zal71, Zoh73b, Zoh74, Bak76, Ber72, Cai71, Cor71a, Cor71b, CSY75, CL72, Dru71a, Gla70, Goo71, GW70, Kar77b, MS73b, PB76a, SR77, WK71]. Fast-Converging [NW79]. Fast-Division [Lig77]. Fault [AB79a, AM77, Ake74, Ake76, AGM^{+}71, Avi76, BC71b, BCA^{+}75, BK73, Bos79, BH71a, BCJ^{+}71, Bre71d, Bre72a, Car73, CM76, Cer78, CCES74, CRB72, DF75, DB76, DTG79, Dia75, DV78a, DW73b, FM74, Far73, FM72a, FD74b, FD77b, GRR72, GS74, Gon70, Gra73, GM76, GT78a, GCP^{+}77, GF78, Hay71a, HF74, Hay76a, Hec73, Hop71, HS75, HC71b, HPP77, IS75, Ios75, Kim75, Kim79, KS71, KB75, KK77b, KM75b, LB74, MH76, MH78, MS74, Mat71c, MC71, MF71, MC78, Mey71, MR76, MM78b, MKO70, NA76a, Ogu74, OW73b, OH79, PR73, PR74b, Pra74, Pra78a, Pra78b, Pra78d, PG75, Ram71, Red72a, Red77a, Red78b, RM78, Roy74, RK71, RK75a, RK75b, SR75, SM74, SM75a, SM71a, Sch74a, SM76, SV74, SG71, Smi75, SM78, Smi79d, Sri78, Sti78b, SC72b, Su73]. Fault [To73, TBY78, Wak75a, Wei72, Whi71c, YO70, YY75]. Fault-Detection [DF75, Far73, HPP77, KS71]. Fault-Locatable [Red72a]. Fault-Secure [Pra78a]. Fault-Tolerance [Ogu74]. Fault-Tolerant [AGM^{+}71, Avi76, BCJ^{+}71, Car73, CM76, FM72a, GM76, GCP^{+}77, Hay76a, Hec73, Hop71, HS75, Ios75, Kim75, MS74, Mat71c, MR76, OW73b, PR73, PR74b, Pra74, Pra78b, Ram71, Red78b, SM75a, Sch74a, SG71, Smi75, Sti78b]. Faults [Arm72, BHP^{+}71, Bou74b, Bre73c, Bre74, BCS76, CSJ74, CLL78, Fri74, GM71, GST79, Hay75, Hu75, Ios78, KP74, Kam75, Kau74b, KH77c, KK72, KK77a, KS79, KR78b, MM78a, Mar79, Mei74, MS75a, Nai79, PM75a, SM72b, SK77c, Smi79c, SKM78, Sun79, Var79, Wan75b]. Faulty [Cha70, Cle73, Dir72]. FDP [GLMR71]. Feasibility [Far75]. Feature [ARS71a, And71, Cal70, Che71a, Chi71, Das71a, Deu71, FE78, FP75, FK70a, GY77, GP74, HKD^{+}71, Hey71, KCLB71, Kit77, KKA71, KF72a, Koz71, LD75, Moh71, NF77, OC76, Pre71b, RJ72, The75, Tjo77, TG72b, TV72, Tou72c, Tou72e, Vil73b, WL75, You71].

Features [Bab72, Bec71, BB78a, Chi77, Ges71, Ram76b, STG71, SNEL79, Tou72d, Tou74]. Featuring [CLL78]. Feed [Liu77a, Liu77b]. Feed-Forward [Liu77a, Liu77b]. Feedback [Brz71, Fri70, HC72, LNL72, Lem70, LF74, Mar73a, Mar73b, MAB78, Mor75, Per70, Riv77, Zei71]. Feedback-Control [LNL72]. Feedforward [Ols70]. Ferroelectric [Kau73]. FET [Pai73]. Fewer [Bar71]. FFT [HL86, But79, Dru71b, SCNS83, TJM79]. Fibers [PS78]. Field [Red79, RTKH77, Sar71]. Figural [VV71]. Figure [Hec73]. File [Ste77, WG74]. Files [Anz71, CW79a, RC71b]. Filling [But71, Fre77]. Filter [MS78, Zoh73a, Zoh76a]. Filtered [CH72a, CH73b]. Filtering [BV78, But79, Dav71, Far75, Gol72, Hab77, HLD70, Hun72, Jai77c, NA74, Pra72, SM72a, Yue77d]. Filters [HC70, KH74, LL75, Lit74, SF72, Zoh73c, NA76b]. Final [de 71]. Find [Du72]. Finding [Cha74b, Das73, FD77a, FBS75, FO71a, OC76, PHG77, RB78, Tru71, Whi71a, Yan75a]. Fingerprint [MJTW75, MF76c, RB78]. Finite [BV78, BF72, Cai71, Cha72, CL75, DV78b, Fre79b, Gaj75, Gil70a, Goo70, HN78, KYO70, KW70, Kar77a, Kar77b, Kri73, LF74, MM73a, MB77, Moo71, Paz70a, Pfl73, Pu72, Red79, RTKH77, RS72a, Ric72, Vai72, Vil70, Yan74a, Yan75b, YM78, Yeh70]. Finite-Memory [KYO70]. Finite-Precision [MM73a]. Finite-State [BF72, Cha72, CL75, Gil70a, Goo70, Kri73, LF74, MB77, Pfl73, Ric72, Vai72, Yeh70]. FIR [MS78]. Firing [Wak70c]. Firmware [Pag72, Pat76]. First [She70a, Weg76, Shi70]. Fit [Lio76]. Fitting [Can71]. Fixed [Jac73, LK75, SR77]. Fixed-Increment [LK75]. Fixed-Integer [Jac73]. Fixed-Point [SR77]. Flexibility [MH75]. Flexible [CK74, KKA71, MH73, Obe72]. Flight [FD74a]. Flip [BH73, Bre75, Bro71, Brz71, CD73, Cur70, Cur71b, EW77, FD79, ISN76, MF76b, NR76b, Rub79, Sin77, SR71, Sur70b, Toh74, TB74]. Flip-Flop [Bro71, Cur70, Cur71b, EW77, NR76b, Sin77, Sur70b]. Flip-Flops [BH73, Bre75, Brz71, CD73, FD79, ISN76, MF76b, Rub79, SR71, Toh74, TB74]. Flipping [Chi77]. Floating [Bau77, Boh77, Bre73b, Cod73, HW76, Hig76, Ken77a, KKK74, KPS77, Mat70, MMB78, Mor71, Pak71, Yoh73, Yue75c, NA76b]. Floating-Point [Boh77, Bre73b, Cod73, HW76, Hig76, Ken77a, KKK74, KPS77, Mat70, MMB78, Mor71, Pak71, Yoh73, NA76b]. Flop [Bro71, Cur70, Cur71b, EW77, NR76b, Sin77, Sur70b]. Flops [BH73, Bre75, Brz71, CD73, FD79, ISN76, MF76b, Rub79, SR71, Toh74, TB74]. Flow [CCF79, Rej72, Rum77]. Fluctuating [BG75]. Focus [LE79, Swa79a, LE77]. Folding [Lee78a, NA76a, To73]. Following [SB72]. Forests [KLCC77]. Foreword [Ano76g, Ano77g]. Forgeries [NR77a]. Form [AM79a, Ano79-45, Ano79-46, Kun75, MT74a, Max77, Mor70c, Ulm70]. Formal [TF78]. Formalism [GL74]. Formalization [Mat70]. Format [Anz71, Hig76]. Formed [RF75]. Forms [CM79c, HG78a, Hen78, KS75, KS77b, Lam79a, PP75, Sri75]. Formula [SA74]. Formulas [Ben70, Pin71]. Formulating [Nie75]. FORTRAN [BCKT79, KMC72, Roc70, Swa72b, Fre74a, LG74, Rej72]. FORTRAN-Like [BCKT79, KMC72]. Forward [Liu77a, Liu77b]. Foundation [Fu74, Kul77]. Four [CYNM79, CM79b, Dir72, NK75]. Four-Phase [Dir72]. Four-State [NK75]. Four-Valued [CM79b]. Four-Variable [CYNM79]. Fourier [Ber72, Cai71, Cor71a, Cor71b, CSY75, CL72, Dru71a, FHTS78, Gla70, Goo71, GW75, GW70, Hen70, Hun71, Kar77b, LH75, Lec70a, PB76a, SR77, Zoh74, CF79, Des74, Des79, DV78b, Gra72, GW75, Kun79, RT79, Rej72, Sar78, Tan78, Yue75a, ZR72, Zoh73b]. Fourth [Gag72]. Fraction [VV74]. Fractions [BB74a, BB74b, God71, Tri77b, Tri78, RT73a]. Frame [SH77]. Free [ABK76, BT72, BB78c, BK73, BW70, FM74, Gre75, Har70b, HK77, HK78, Mar79, SK79b, Tho75, WN76]. Freedom [AC71]. Freehand [NR77a]. Frequency [FV79, GF78, KH74, New70, RC71b, Sok72, TS75]. Frequency-Dependent [RC71b]. Frequently [Len78]. Frills [JW76]. Front [Ano78t]. Full [LHSM74]. Fully [Mor79]. Function [AB78, ASS72, Bar70, Bor72, BT76, CCB70, DW73a, Edw75, EH78, ET75, GC71b, HD71, HG78a, HPP79, Han72, Hen71, HA76, KAN79, LH76, Lee78a, Lyn72, Mar73b, ML76, OY70, Pap77b, RB72, Van75, YT71b, Yue75e, Yue76, ZT75]. Function-Unit [BT76]. Functional [AM79a, AK75, Ber70b, Bre70b, Fly70, Gar71, Kar70, KK70, LS71, MC78, Pav73, Rud76, SF71, VW72]. Functions [AA72b, AA71, And70, Bak75b, Bau71, Bau72, BCSM72, BB78c, BR76a, BR78, Bis70, Bis71, Bis73, Bit77, Boh77, BB74a, BG78, BV79, Bro75b, Bro75c, But72, But78a, But75, Can71, Car71b, CR78, Cha73a, CL77, CM73b, CYNM79, Cur76, CM79c, CM79d, Dah73, Das71b, DS71, DK72, Dav72a, DS75, Dui76, Eks73, Erc73, Erc77, Fen74, FD73, FB74, GBMC72, GC73, God71, GM76, HF72, Har70a, Hay71b, Hig73, HE79, HC73, Hol72b, HO72, HA77, Iba74, Kan73a, Kan73c, Kan76, KT72, KS72, Kit78, Kod74, KS77b, KS78c, Kol72, KN70, LM71b, LL72, Lew77, Mag70, Mag73, Mar72, Mat71b, MS77a, MTB70, Nec79, Orc76, OW72, Pav75b, Pin71, PD79, PH73, PP75, Pra78c, PM71, Pre71a, Red72b, Red73a, RW74, Red73b, Red71b, RS71a, RNMP77, RH75b, Rud76]. Functions [SA70, SB77a, SK79b, Sch75, Set70a, SM70a, SMW71, Sin72, SC72a, Smi71b, Spi71, SP77, SC72c, Sur74, Sur75, TG72a, TD77, Tom74, VD76b, Van73b, Vin78, VW72, Wes78, Win70a, Wri73, YT71a, Yeh70, Yue75d, Bak75a, Spi73]. Fundamental [BR78]. Fundamentals [Swa74b]. Further [Cur71a, RB76, RTL72]. Future [Cox72, TP73]. Fuzzy [BD75b, Cha72, GL70, Kan73a, Kan73c, Kan76, Mam77, Paz70a, SC72a, Sri75, Su70].

G101 [PRD71]. Galois

[BR76a, BR78, Pra78c, Wes78]. Gap [Nel74]. Garbled [OTK76]. Gate [BY79, CC72, HK77, HK78, Iba71, Ken77b, NTK72, Nak79, Obe76, RM71, RH75b, Rot70]. Gate-Interconnection [Iba71]. Gate-Minimum [Nak79]. Gates [BCSM72, CYNM79, Ham73, Hu75, Iba71, IM71, LM79, LHSM74, NT70, TG79]. Gaussian [Mar73c]. General [Bar70, Brz70, CH72a, CH73b, Dru71a, Erc77, GR72a, HP72, Hoo77, HA76, KKK73, KS73, Mur70, PM75b, Rey70, Smi71a, Sri76, Win70a, Zis76]. Generalization [CM79d, FC77, Gla70, Ram75]. Generalized [AA77, AC70, CL79, CF79, FS78, Hsi71a, KSA74, ML76, Pra72, SK74, Swa72a, Tho78, VD76a, Wak70c]. Generate [Pap77b]. Generated [CL71, Rao78]. Generating [Agr75b, AM79c, Ber70a, CL72, Das71c, DK72, Fre79a, Hwa74, MAB78, MS77b, Mur70, Myk75, SCL70, Sok72, YT71c]. Generation [AA72a, AA75, Ake76, Bau72, BCC78, BH71b, BH71c, Bre71d, Bre72a, Bre74, BH74b, CSJ74, CDO78, Che78, Dan70, Dav72a, FP75, HLD70, HO79, Hay78a, Hla78, How75, Hub76, Hur74, JLH73, KN70, Lee78b, LE71, Lew77, MB75, Mut76, OH79, Pap77a, Pat71, Qua74, RJ72, Reu75, Ros71, Rub76b, Sil77, Smi74, SKK79, Ung77d, Van73b, Wan75a, WC74, WC76, Yan74b, Bel76, Gag72]. Generator [And72, Arm73, HPP77, MH70b, MH70c, Mar73c, PW79a, SLC75, Bak76]. Generators [NM76b, NM76a, Yue77e]. Geometric [Ban74a, BO79, Mor75, Ram77b]. Geometry [Fer77, Wan77]. Gestalt [Zah71, ZT75]. Getting [Ano74-28, Ano75-29, Ano75-30]. Global [BD71, Ehr77, Hwa79, MN78]. GNAF [LR75]. Goal [LS76b]. Goes [Ano74b, Ano74c, Ano74e, Ano74f, Ano74d]. Golay [Pre71b, PC72]. Golomb [Rob79b]. Good [Hur74, She70a]. Gracefully [BF75b]. Gradients [Ros75]. Grammars [MB77, Pat71, Tho76]. Grammatical [Gai79]. Grams [RH74]. Graph [Ake74, Das73, Hay76a, KNF76, Law70, Lem70, MR74, Rej72, Rub75a, Sau72b, Yan75b, YM78, Zah71]. Graph-Theoretic [KNF76, Law70]. Graph-Theoretical [Zah71]. Graphic [Arm73, New76, PRD71, SA70]. Graphical [DLL^{+}73, Fon71, MA71]. Graphics [Bib74, KC71, LH79, MB75, New75a, SAFV71]. Graphs [BEGO71, Dav77a, Har70a, HN71, LR71a, Reu72, ROW71, Sic76, Sto70d, Tur70, Vil70]. Gray [Ano70o, Mor74, PR78, Yue78]. Gray-Scale [Mor74]. Great [WBRG76]. Greatest [Gri78a]. Grey [Ami77]. Grey-Level [Ami77]. Grid [Fre79a, LR76]. Group [Dun75, Kol72, Lev70]. Grouping [FK70b, ZT75]. Groups [Cai71, Kar77a, Kar77b, YC76]. Guidance [LS76b]. Guide [Har71]. Guided [Rei76]. Guidelines [Ano76h].

Haar [RNR75, Rej72]. Hadamard [Car71a, MS73b, WK71, AK72, CL77, FA76, GC77, Hen78, Kun75, Kun79, RNR75, RRNA74, SO79b, Ulm70]. Halftone [KC71]. Halftones [Ros71]. Halting [Nie70b]. Hamilton [Vil73a]. Hand [Gra72]. Handling [Ram77b]. Handprinted [DT70, HTD72, TD70]. Handwritten [Dut74, TG72b]. Hard [DLL^{+}73]. Hard-Copy [DLL^{+}73]. Hardware [BW77, CS75, Chu79, CLL78, Des79, Erc77, Gad70, GSJ^{+}76, GL74, Ham74, Mar72, Muk79, Pag72, Rot77, Sie74, SW76, SP72c, SP77, Von75, Zoh73b, Zoh73c, Zoh74, RT73a]. Hardware-Oriented [Erc77]. Hardware-Shared [Gad70]. Harmonics [RB72]. Harrison [Nam70]. Hartley [Hil79]. Having [Ges70, God71, Sti78b]. Hazard [HK77, HK78, Hla70, Kan73b, Ser73, Ser75, Zis76]. Hazardless [Fra72]. Hazards [BH74b, HD71, Hla78, ZKK79]. Head [Kam70a]. Hearsay [FL77a, REFN76]. Hearsay-I [REFN76]. Height [LP76, LP78]. Height-Balanced [LP76]. Heterogeneous [CK79]. Heuristic [AAK74, BD75b, Col77, Coo72, Gri76, PR71, RL72, ROW71]. Heuristics [AI78]. Hexagonal [Deu70, LR76, Pre71b]. Hidden [Lou70, Wri73]. Hidden-Line [Lou70]. Hierarchical [Kul78, Nad71, WH78]. Hierarchy [Bae74, Cho76, Wel78]. High [Agr79a, AN74, BCH^{+}73, CH73a, CL75, Chi78, CH72b, Cor71b, Gar74, Glu71, Hou72, JM70, KES72, KG72, Ken77b, Kri70a, Kuc70, LE71, Lin70, Lit74, LM70, Loo70, Mar73c, NW79, Pla73, Qua74, RGK72, RT74, SW71a, SW76, Sri71a, Sri71b, Ste72, SKG77, Sto70c, SKK79, Tom75, Woo79]. High-Dimensional [NW79]. High-Level [RT74]. High-Order [Glu71]. High-Speed [Agr79a, CH73a, Chi78, CH72b, Cor71b, Hou72, JM70, KES72, KG72, Ken77b, Kri70a, Kuc70, Lin70, Lit74, LM70, Loo70, Qua74, RGK72, Sri71a, Sri71b, Ste72, SKG77, Sto70c, SKK79, Tom75]. Higher [Atk70, DV78a, EH76]. Highlights [Ken77a]. Highly [Agr79c, GLW74, Los76, Mai71, Man77]. Hilbert [But71]. Hole [LS76b]. Holographic [BL75, GW75]. Holography [SNN70]. Homomorphic [Zei71]. Homomorphism [Lem70].

Homomorphisms [Kam70b]. Horizontal [RT74, TG76]. Horn [Hen76b]. Hough [Skl78, TM78b]. Huffman [Par79]. Human [FD75]. Hybrid [AK75, CS73, Cas73, CM70, Deu71, GM72a, Gil70b, IS76, KK70, KV76, Lit73, MJTW75, MG71a, Ogu74, RT79, RH73, SF71, SM73a, SM73b, TL73a, Wit70]. Hybrid-Computer [KK70]. Hyperresolution [Win76].

I/O [VN78]. IBM [Joh71, SB71]. IC [Gra70]. Ideal [Hoe74]. Identification [Ada78, Bar78, Bis70, Bis73, Bou74a, BCS76, Cor77, DT76, Das78, Dep70, DBM77, Gar72b, GH72, Kel71, MJTW75, Moh71, WZ72, YT71a, You70]. Identified [Bab72, Tou74]. Identifier [Ano70-28, Ano72-51, Ano72-52, Ano72-53, Ano72-54, Ano72-55, Ano72-56]. Identifying [Eks73]. IEEE [AG73, Shi70, Ano70a, Ano70-29, Ano70-30, Ano70-31, Ano70-32, Ano70-33, Ano70-37, Ano70-34, Ano70-35, Ano70-36, Ano71a, Ano71z, Ano71-27, Ano71-28, Ano71-29, Ano71-30, Ano71-31, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano72n, Ano72-50, Ano72-57, Ano72-58, Ano72a, Ano72-59, Ano72-60, Ano72-61, Ano72-62, Ano72-63, Ano72-64, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano73a, Ano73m, Ano73n, Ano74j, Ano74p, Ano74k, Ano74q, Ano74l, Ano74r, Ano74m, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74n, Ano74y, Ano74o, Ano74z, Ano74a, Ano75o, Ano75p, Ano75q, Ano75i, Ano75r, Ano75j, Ano75s, Ano75k, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75l, Ano75y, Ano75m, Ano75z, Ano75n, Ano75a, Ano76c, Ano76d, Ano76h, Ano76k, Ano76l]. IEEE [Ano76i, Ano76m, Ano76j, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano76b, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano77a, Ano78h, Ano78i, Ano78j, Ano78y, Ano78z, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78-32, Ano78-33, Ano78v, Ano78-34, Ano78w, Ano78-35, Ano78x, Ano78-36, Ano78b, Ano78-37, Ano78-38, Ano78-39, Ano78-40, Ano78-41, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79k, Ano79l, Ano79-45, Ano79-46, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42, Ano79-43, Ano79-44, Ano79a, Ano79-53]. IEEE/Popov [Ano72n]. If [Ano76t, Ano77u, Ano78-42, Ano79-47, Ano79-48, Ano79-49, Ano79-50, Ano78-43]. II [Cur70, FL77a, Lew74, Loo70, SP72c, Sri71b, WN76]. III [Atk70]. ILLIAC [Atk70, AA72a, Dow73, Orc76]. Image [AP76a, ACJ75, Ano78v, Ano78w, Ano78x, BS72, CDL78, DM78b, FR76, Hal74, HA77, Hun73, Hun77, JA74, KF73, MMB77, NJ77, PR78, PD77, PC72, RNR75, RTKH77, Roe78, Sha78b, SC77, Sta75a, TE76, VM78, WD77]. Image-Processing [SC77]. Imagery [ASS72]. Images [AD75, Ami77, AP76b, CB72, CA77, Ehr77, GSJ^{+}76, Gra71, HKD^{+}71, HE73, HB75b, Jai77a, Jai77b, Jai77c, JR76, KC71, KS77a, Ram77b, ST79, Sta75b]. Imaging [Eks73, Saw77, WKP75, Wei75, YRT75]. Impact [Chi78, FD74a]. Imperfect [SF78, She78b]. Imperfectly [Bab72, TG75, Tou74]. Implement [Har76]. Implementation [Ban74b, Ber72, CM76, Des79, EI77, Gad70, GM72a, GL74, GC71b, HL86, Kar74, MS78, Orc76, Red71b, SP72c, SNEL79, SCNS83, TJM79, Win76]. Implementations [Alm78, Fen74]. Implemented [BB75, Dev79, Kau70c]. Implementing [CM79b]. Implicant [BM70, HW75, IR76, Mor70c]. Implicants [BH71c, Bub72, Das71c, DK72, Hwa74, Maj72c, NRS74, Reu75, SCL70, Voi77b]. Implication [VVA78]. Implications [EFLR76, Gar71, PM76a]. Implicit [SM70b]. Importance [Mal73]. Important [Ano74-28, Ano75-29, Ano75-30]. Improve [Hay74, IR76]. Improved [BK72, Ben71, BG79, Hol72a, JLH73, MSJ72, RHR76a, RHR76b, RW75, Rub75a, Sto70d, TH79, TB75, Van72, Vil70, Bel76]. Improvement [CF70b, Eks74]. Impulse [BV78]. Incompatibility [Das73]. Incompatible [SB74a]. Incomplete [Ehr72, Kit78, RB76, Yan73, YB75]. Incompletely [Bis74, DH79, GD73, Hig73, Kel70, Pfl73, RB75, Wil75]. Incorporating [Boy74]. Increasing [KR71a, SM71b]. Increment [LK75]. Incremental [BR77, Dan70]. Independent [CJ74, Cio78, Kel74, TF70, Tou72d]. Indeterminate [She77]. Index [Ano70n, Ano70z, Ano70-28, Ano70a, Ano71a, Ano71-51, Ano72o, Ano72p, Ano72q, Ano72r, Ano72s, Ano72t, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46, Ano72a, Ano72-51, Ano72-52, Ano72-53, Ano72-54, Ano72-55, Ano72-56, Ano73a, Ano74a, Ano75a, Ano76b, Ano77a, Ano78b, Ano79a, BM78, IA76, Man71, RT70]. Indirect [Pea77]. Industrial [NR76a, Per78, YT77]. Inequalities [Cha71, Men70b, NK74, Tou72e, WG73]. Inertial [Ser73]. Inessential [HC71a]. Inexact [Pea76]. Inference [MB77, Vee78]. Inferencer [Gai79]. Inferred [SF71]. Information [And70, Ano70-38, Ano70-39, Ano70-40, Ano70-41, Ano70-42, Ano70-43, Ano70-44, Ano70-45, Ano70-46, Ano70-47, Ano70-48, Ano71-37, Ano71-38, Ano71-39, Ano71-40, Ano71-41, Ano71-42, Ano71-43, Ano71-44, Ano71-45, Ano71-46, Ano71-47, Ano71-48, Ano72-65, Ano72-66, Ano72-67, Ano72-68, Ano72-69, Ano72-70, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano74-27, Ano75-27, Ano78-44, Ano79-51, Anz71, BR74, Dot70, Gri73, Hop71, Key77, Low73, Mat71a, SM72a, Sri76, Ste77, TA71, TW76, Vil74, Wei75].

Information-Lossless [Dot70].

Informational [Mat71b]. Infotech [Fre74b]. Inhibition [RF72]. Initial [Kar70, MR75]. Initialization [Kri73]. Initializing [Gra73]. Injection [DMR77]. Ink [Ern72]. Inner [Bla79, SGR78, Wak70b]. Input [AH76, Ano75-28, BH73, Bre75, Bro75a, CW79b, CD73, Chu76, DDSM76, Fre79b, FG71b, FG71c, GSB^{+}76, Hig71, Hu75, KN72, Mar71, MH70a, MST^{+}75, PM75a, Pec76, SK78a, SK78b, Sin71, Smi75, Sur70b, TB75, Ung71, Vai71, Vai72, Vin78]. Input-Memory [Fre79b, Vai71, Vai72]. Input-Output [TB75]. Input/Output [GSB^{+}76, Smi75]. Inputs [FNSK75, Ges70, Mar77]. Instruction [KV73]. Instructions [Ken77a, TF70]. Instrumentation [WBRG76]. Integer [Jac73, LR75, MI72, Sik70, Wil77a]. Integral [Lee78a]. Integrals [Fos73]. Integrated [Agr79c, BT70b, DMR77, EI77, Man77, Sri71a, WA73]. Integration [Ben70, Max77, RH71]. Intelligence [Arb76, BW76b, Coo72, FL77a, Hol76, Rei74b, Ano78h, Ano78i, Ano78j, Ano78-37, Ano78-38, Ano79k, Ano79l]. Intelligent [EUY^{+}72, LJ79, SAFV71]. Intensity [Ern72, Saw77]. Intensity-Modulated [Ern72]. Interactive [AKM^{+}71, Bib74, BO70, CJW^{+}71, Nay71, New75a, PS71, Sam70a, WN76]. Intercell [MF76d]. Intercommunication [Tho78]. Interconnected [Hur74, Pat73]. Interconnecting [Mor79]. Interconnection [CW79c, Gec77, GC71b, Ham71, Iba71, JM70, Sie77, Sie79, TW77]. Interconnections [Ano72-39, DEA73, Lan76, Par71, Swa74a]. Interconnectivity [Sic76]. Interdisciplinary [Ham74]. Interface [DWT70, NS75, Stu75]. Interfaces [Tro75]. Interference [Bha75a, Hoo77, Rav72, SD79]. Interleaved [CBS71, KR75, Rau79a, Rau79b, Rav72].

Interlock [Cha79, Wor77, Wor79]. Intermittent [Bre73c, KP74, Kam75, KK77a, KS79, MM78a, SKM78, Var79]. Internal [HB70, Smi74, Toh74]. International [Ano79c, Ano79d, Hob73, Kra74, Lom74, Mui74, Pla73]. Interpolation [AP76b]. Interpolator [Kor76]. Interpretation [FG71a, SW71b]. Interpreter [Chu79]. Interrelationships [AM77]. Interruption [Mcd72]. Intersections [BO79]. Interval [NS75, She78b]. Intervals [KP75]. Interwoven [Abr75]. Intrinsic [CA74, FO71a, Tru71, Tru76]. Introduction [Ano73w, Ano76g, AG73, Bas71a, Car73, CH75, Den74, Fen77, Fle79, FR71, Ham74, Hen76a, Kim75, Kin72, Lan79, Les71, Luc72, MR76, Ram71, Ram73a, Red78b, Rei74a, Rei74b, Sch74a, Sil73, Vra77]. Invariant [Eks73, Pea75]. Invariants [DBM77, Mat71a, WGS78]. Inverse [AA77]. Inverses [Ols70, SK74, Yua74]. Inversion [Che73, NC78]. Inversions [Wak70a]. Inverted [Ste77]. Inverters [BS77]. Invertible [SSK72]. Investigation [AGM^{+}71, BCB74, Hen71]. investigations [RT73a]. Invited [Ano72n]. Irredundant [AA75, Bub72, CM79c, DBC71, HG78a, Hay76b, Mor70c, NA76a, NRS74, To73]. Irrelevant [BB78a]. Isobaricity [GBMC72]. Isolation [BCA^{+}75]. Isomorphism [Yan75b, YM78]. Issue [Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, Ano78t, CH75]. Issues [Bha77, TW76]. Iteration [Fly70]. Iterations [Des74]. Iterative [ARS71a, BBC72b, Bar74c, Bor72, CH73a, CM70, CW77, Coy79, Dev75, Dia76, Eks74, FM72b, Fri73, Gar74, GR72a, GR73b, Goo70, Goy76, Ham71, HI72, IH72, Kon71, Kri70a, Kri71b, LY71, Mag70, MK71b, MF71, Nic71, Ogu74, RGK72, SM73a, Sri79, Ung77d, Ung77c]. Iterative-Improvement [Eks74].

Iteratively [ATN70, Hu73, RF75]. IV [AA72a, Dow73, Orc76, Swa72b]. IV-Type [Orc76].

J [Wor79, EW77, MF76b, Obe76, Rub79, TB74]. James [Ano73z]. January [Ano73f, Ano76x]. Jets [Ern72]. Job [Cam72, Mal72, SW76]. Job-Mix [Mal72]. Join [Tha78]. Joining [Sta74]. July [Ano77x]. Jumps [RF72]. June [Shi70, Ano72-40].

Karhunen [FK70a, Jai77a]. Keyhole [Oza78]. Keys [Sel74]. Kind [She70a]. Knee [ADL72]. Knowledge [PS71]. Kronecker [Dru71b].

L [Tab70b, McC79]. Labeling

[DR77, ZHR77a, ZHR77b]. Ladder [BS76]. Ladder-Structured [BS76]. Landmarks [HEWB72]. Language [BT74, Fel79, HP75, KS77c, RT74, RC71c, Tho76, Cam72]. Languages [Arn70, Bar75a, BW76b, BT73, Gre75, Har74b, Har70b, Hoe74, Pag70a, Pla73, Ros74a, Sta70b, TF78, Tho75, Weg76]. Laplace [MR75]. Large [Ari79, BHK79, Ben77, BM70, CM70, GC71b, KO78, New70, RC71b, Ram75, Ram77b, Smi74, WA73]. Large-Scale [CM70, GC71b, WA73]. Latency [DTG79, FM78, SM76]. Latin [HB75a]. Lattice [CW79a, Yeh70]. Layer [Tak76]. Layered [Uhr72]. Layout [Lar71]. Learning [Ama72, CF70b, Dee72, DTI78, Fu74, LD75, Mic74, MS75b, MS76, Ull72, UC75]. Least [All75b, Gol72, HA77, Hun73, Wil77b]. Lee [Rub76a, Hoe76, Rub74]. Length [Bak79, BC71b, But79, Hil79, Hol72a, LE71, MH70a, MBM77, PW79a, Per70, Qua74, Smi71a, Wei72]. Level [Ami77, BK76, BS75, CC72, CCB70, EZS78, ET75, Far79, HP75, Ios78, KV73, MK79, Nak79, OY70, PR78, Pla73, PR74b, RT74, RR74, RH75b, RV73, Liu75]. Levenshtein [OTK76]. Library [Ano76t, Ano77u, Ano78-42, Ano78-43, Ano79-47, Ano79-48, Ano79-49, Ano79-50]. Like [BCKT79, Cur73a, KMC72, PP75]. Limitations [Sha78a]. Limited [Mak71, RSH79]. Limits [Ano77b, Ano77c]. Line [Ada78, Agr75b, All75b, Ano70y, Bar78, BLZ75, BW70, BL75, DT76, Das78, Fre79a, HPP77, Kau74a, Lou70, Lud74, Mac70, Meo75, MS75a, RL72, Ros74b, SW70a, SW70b, Sta74, SSR72, TE77, Wei75, Wil70, Wil77b, Wri73, You70, ZHR77a, ZHR77b]. Line-Drawing [RL72]. Linear [Bre72a, Can71, Cha71, Cha73a, Che74, CK75, CW77, Das71a, DS72a, Dav70a, DM78b, Dep70, Dun75, EGG72, FL77b, Gal78, GR71a, GR72a, GR73b, HO74, Has73, HI72, IH72, Kar77a, KCLB71, KP75, Kri77, Mar70, Mar73a, MAB78, Men70b, MBM77, NK74, Ols70, Pat71, PR74a, Pea76, Per70, Red78a, SR73, SK77c, SIF76, Tom74, TV72, VD76b, Van75, WG73, You71, Yua74, RS71b]. Linearity [GCC74, KD73]. Linearly [Hur74]. Lines [CY76, Hor71b, Hor72, MMO74, SKK79]. Linguistic [Mam77, WBRG76]. Linkages [CR72b]. Lisp [Tei76]. List [Ano76a, Ano78s, Ano78a, Ano79-52, Hol79, Mor70b, MM70, Weg72]. Literature [Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m]. Little [Yue77d]. Load [CK79]. Loaded [SW70a, SW70b]. Loading [Hou72, Wym73]. Local [Gra71, RS75, TS79]. Localization [FS78]. Localized [SD79]. Locatability [GM71]. Locatable [Red72a]. Locating [Nev76].

Location [FM74, Sri78, SC72b]. Loeve [Jai77a, FK70a]. Logarithm [HL86, SK71, SA75, SCNS83]. Logarithmic [DG72]. Logarithms [BB75, HLD70, Mar72]. Logic [Abr75, AM77, AA72a, AA75, AA76, Agr77, Ake72, Ake73, Ake76, Ale71, Alm78, AB78, Arm72, Bar77, Bar70, BW76a, BCA^{+}75, Ber70b, BF75a, BHP^{+}71, Bre72a, Bre72b, Bro74, CJW^{+}71, CS70, CM74, Cer76, CE75, Chi77, Cle73, CM79b, Dan73, DR74, DMR77, Dia75, Dia76, Dir72, Dor72, DVS74, Edw75, EZS78, ET75, Elm76, EH76, FS70, Fis74, Fle79, FM72a, GW77, Gar72a, GC73, GLR71, GT74, GT78b, Hay71a, Hay74, Hay76b, Hay76c, Hay78b, HC73, Hu75, HK74, Ios75, Ios78, ISN76, IC78, JM70, Kam79, KH77a, KSM76, Kli71, KAN79, Kod74, KS75, KS78c, KS71, KK72, LR71a, LV75, Lec70b, LC79, LM70, Loo70, Lyo70, Mam77, MA71, MC71, McC79, MR74, MF71, MBSS75, Min75, Mor70a, ML76, Ops72, Osb77, OW73b, OH79, Pap78]. Logic [PM75a, PW79b, Pat73, PR72, Pre72, RT70, Red72b, Red73a, RW74, Red77b, Ree70, RM78, Rot78, ROW71, Rus72, SR74a, SR74b, SK78a, SK78b, SK79a, Si78, Sin72, Skl70, Sli70, Smi78, SM78, Smi79c, Sri79, Sto77, Sto70b, Su70, Su75, SB74b, Sur78, TM76, TD77, Tis71, TKH71, Tom75, TK74, Toy71a, TG79, Voi77a, Voi77b, Vra77, Wan75a, WR77, Wea75, Weh74, WA73, Win70a, Win70b, Woo79, YI74, YT70, YT71c, YY75, Yue75b, Yue76]. Logic-in-Memory [ET75, Sto70b]. Logical [Aus70, Bet71, CF73, GL74, KF78, KT72, MH73, MH75, Mas78, MS70, TM71, Wak74]. Logics [Mor76b]. Long [Sut74, Wil70]. Lookaside [CDJ73]. Lookup [MS77b]. Loop [Gal78, PB76a, VN78]. Loopfree [RS74]. Loops [BCKT79, Gil70b, Kau70b, Lit73, Pag72]. Lossless [Dot70]. Low [PA78, She71, Wak75b]. Low-Cost [PA78, She71, Wak75b]. Lower [Du72, FS70, Han72, Van72]. LR [Jol76]. LSI [Chi78, HB75a, Rus72, Toy71a]. LX [HA70]. LX-1 [HA70].

M. [Ano73z]. Machine [Abd71, Alm78, Ano78h, Ano78i, Ano78j, Ano78-37, Ano78-38, Ano79k, Ano79l, Arb76, Ber71, Cur73b, Fel79, God71, Gra73, Ham71, HC72, Har70b, JK77, Kel71, Koz71, Kru73, LV75, LF74, Osb77, PM74, PHG77, RN77, Sie77, SM71c, Sur78, Tom72, TK74, WZ72, YT77, Yan74b, Yoe70a]. Machines [BF72, Bis74, Bou74b, BH71b, Bre75, Brz70, Cha72, CL75, Che74, CD73, Chu76, CD78, Cur70, Cur73a, DF75, Dav70a, Dav70b, DSBC74, DS73, DGC74, DAA79, DW72b, DH79, ES77, Ehr72, Far73, Fre79b, FM73, FK74, FNSK75, GH78, Gel71, GR71a, Ges70, Gil70a, Goo70, GB71a, GB71b, GD73, GCC74, GCP^{+}77, HSP78, HG79, HN73, HC71a, Hol72a, HC71b, Hsi71b, HI72, Hu72, Hwa73, IH72, Jos78, KYO70, Kan75, Kar74, KM73, Kel70, KN72, Kri73, KR78a, Lan79, MS73a, MS74, Man72a, Mar70, Mey71, MR72, MKO70, NT79, New70, Nie70a, PM77a, Par72a, Par72b, Paz70c, Pfl73, Pre70, RB75, RB76, RV74, Rey70, Ric72, Roo73, RT73b, Sau72a, Sau72b, SM74, SM75b, She70b, SV74, SG71, Sie79, Sil79, Smi74, SB74a, TI72a, Tan71, TA71, TOS71]. Machines [TB74, Vai71, Vai72, Wak75a, WY75, Wan71, Wil78, Wil75, Yan73, YB75, Yeh70, Yen79, Vee78, Lyn72]. Macro [Aus70, NN73]. Made [Lio76]. Magnetic [Bha75b, CW79a, Che78, Gar72a, Hoa76, KSM76, LJ79, MBSS75, Min75, PB76b, Wil77a, WC76]. Magnitude [BM78, MMO74, RT70]. Main [JM73, Tho74]. Maintained [Agr79c, CLL78, Man77]. Maintenance [CHSS73, Gay78]. Majority [HP73, RR74, Sie75]. Make [Lam79b]. Making [KKA71, PS71]. Maloperation [Cha79, Wor77, Wor79]. Man [Arb76, Bon74, Els74]. Man-Computer [Els74]. Management [Ano76u, DB74, De 79, Dow73, LJ79, Mui74, Wan72]. Management-International [Mui74]. Manipulating [Fen74, Wil70]. Manned [WKD70]. Many [Dig73, DVS74, DV78a, KT72, VLS70]. Many-to-Many [Dig73]. Many-Valued [DVS74, DV78a, KT72, VLS70]. Mapping [EH78, Kru71, LSB77, NW79, ROW71, Sik70, Whi72]. Mappings [MR74]. Maps [ACJ75]. Maritsas [Hil79]. Markets [Jos72b]. Marking [PC72]. Markov [Hab77, SKM78]. Maryanski [Gai79]. Maryland [AG73]. Masking [Dia75, RK75b]. Masks [Sie77]. Mass [PA78]. Master [FG76]. Master-Slave [FG76]. Match [And74]. Matching [Cho76, DW78, DN77, FE73, Oza78, VR77, WH78]. Mathematical [Bre71c, Dir72, Duk74, FS70, Kul77]. Mathematics [Har74a]. Matings [And76]. Matrices [All75a, Ari79, Ban78, DS74, Har73, Hen78, Oja79, Ram75, Sag77, SK77a, SK77b, SSK72, TF73, Van77]. Matrix [ARS71b, AC71, Ast78, Cap76, Car71a, Che73, CL79, Den76, Dru71b, Ekl72, FA76, God73, Gou77, Hwa72, Koh72b, KR73, Kri77, NC78, Ram77a, Rat76, RF75, SG74, SR73, Sch73, SK74, Tom70, Tom72, TE76, VR79]. Max [MMB77]. Max-Min [MMB77]. Maximal [Ada78, ADMR78, Bar78, DT76, Das78, HP72, LE71, MBM77, PW79a, Qua74, Rub75b, Yan75a]. Maximal-Length [PW79a]. Maximin [Hed70]. Maximization [Ram77a]. Maximum [Boh77, Bro70a, Han75, KP75, RS72b, RS73b, Sto74b, Sur73, TH79, WD77]. Maxwell [Cat77, Ano77v]. May [AG73, Ano77y]. Mean [All75b, AOR74, Ano75c, Ano75d, Ano75e, Ano75f, Kul78, Maj72b, Whi70, Wil77b]. Means [AN74, Pag72]. Measure [Das71a, Hel72, Hel74, JP73, KF79, Kor79, Mas78, MMB77, TV72, Wel74]. Measured [Ung77a]. Measurement [BD75a, BCC78, JP76, Kim74, NS75, Whi71b]. Measurements [CJ74, Glu71, You70]. Measures [Bea78, BT73, FGR71, SH72, Tou72e, Vil73b]. Measuring [FLR78, Skl72]. Mechanical [HF71, WG75]. Mechanism [MF75, SN72a]. MECRA [Mai71]. Media [BG75]. Medium [Red71a]. Meet [Tha78]. Megasamples [GSB^{+}76]. Megasamples/s [GSB^{+}76]. Members [Ano72-50, Ano73h]. Membership [Ano72-57, Ano72-58, Ano73i, Ano73j, Ano73k, Ano73l, Ano74j, Ano74k, Ano74l, Ano74m, Ano74n, Ano74o, Ano75i, Ano75j, Ano75k, Ano75l, Ano75m, Ano75n, Ano76i, Ano76j]. Memories [AU74, Bae74, Bha75b, BD77, BK71b, CKL77, Che78, CBS71, DB74, FM78, GSB^{+}76, GM70, Hay75, JM73, Klu77, KH77b, KH77c, Koh72b, Len77, LJ79, Mon74, MVW79, NTA78, Nai79, PA78, Pra78b, Rau79b, RF75, Sha78a, SY75, Sri71a, Sti78a, Sto72, Sto75, Sun78, Sun79, Swa74a, Wil71, WC76]. Memory [BK79, Bat77, Bha75a, BCH^{+}73, BE70a, BE70b, BB75, CDJ73, CM76, Cov74, Cur70, Cur71b, DI71, DBC73, Dig73, Dru71a, ET75, FS71, Fre79b, FT79, Gar71, GSB^{+}76, GLW74, Gra73, HK77, HK78, Hoo77, Hou72, HB75a, KYO70, Kar76, Kau73, Kin71b, KR75, Koh74, Kuc70, Kur74, Lan76, LC79, LM78, MJ77, MJ78, Nel74, Nut77, Pea76, RTW78, RSH79, Rau79a, Rav72, SNN70, SD79, Sho74, SW76, Sri78, Sri71a, Sri71b, Sto70b, Sto70c, Tho74, TI72b, Toh74, Vai71, Vai72, VM78, Von75, WSB79, Wel78, Wil78, WY73, Yua74]. Memoryless [CM74]. Merging [Gol76, Hol79, Lam79a, Par79]. Merit [Hec73]. Mersenne [Rad73, Rad72]. MESFET [TG79]. Mesh [NS79]. Mesh-Connected [NS79]. Message [Sah74]. Message-Based [Sah74].

Metaphorical [Rei74b]. Metastable [CW75]. Method [AA75, And74, ACJ75, Ano75-28, AB78, Arm72, Ban74b, Ben71, BF75a, BCL75, BM70, BB74b, Bub72, Cha71, Che73, Cur79, Cur77, DS74, DB73, Ekl72, Erc77, Fon71, FG71c, GLR71, Gon70, HB70, Hwa72, Hwa74, Jai77b, Kit77, Kla71, KO78, Koh75, KP71, KK70, LSB77, Lee78a, LR75, Maj72b, Mar78, MR75, MT74b, Nie75, NRS74, OTK76, Oza78, Pap72, Pap77b, Pav77, RB76, RW75, RS72b, RS73b, SA70, Sch73, Set70b, Sha78b, SS75a, SW71b, SN72b, Sur73, SSR72, The75, Tom72, Vic71, WK73, Whi70, Whi71b, Yan73, YB75, Zis76]. Methodology [CM77, How75]. Methods [AK75, AKT^{+}77, Atk70, AT70, Bar70, Bar74c, CSJ74, CCES74, EH78, GBMC72, HEWB72, Hun77, Kar70, KM78, KPS77, Mag71, Mil75, MS77b, Mor76b, NC78, Pav75a, RGK72, RS75, Rob70a, Tab70b, TH79, UB74, Wes78, WS76, Wor76, Wul74, You70, YRT75, Zah71]. Metric [OTK76]. Microcellular [Ber77]. Microcode [Rob79a]. Microcomputer [Kar76, Les71]. Microdiagnostics [Guf71, Joh71, RC72]. Microelectronic [Ano72-39, Par71]. Microfilm [Ros71]. Microinstructions [JB76b]. Micromultiprocessing [RV73]. Microprocessor [BBLT73, CF70a, FT79, HA70, Pea77, Roe78, TW77, VN78]. Microprocessors [Ano79b, JM73, Lip77]. Microprogram [Age76, CSST73, Pat76, Sta70a]. Microprogrammed [AAK74, BK79, BT70b, CDS73, CHSS73, Col77, DBC73, GM70, HG78b, JF72, KSW74, Mon74, Red71a, SAFV71, Swa79b]. Microprogramming [AR74, Ber70b, FR71, FLBK76, Gar71, Hig78, JM74, RT74, RA78, Rux71, Shr71]. Microprograms [Ada78, Bar78, DT76, Das78, KR71b, Lee75, RS74, Tho74, TG76]. Microwave [YRT75]. millions [Ano76e]. MIMD [PM78b]. Min [MMB77].

Minicomputers [Kra74, Sch74b, Tro75]. Minicomputers-International [Kra74]. Minimal [AB78, BF78, Bor72, Bre71b, CMS79, Cur77, Dan73, DDSM76, Du72, Fra72, FD74b, FD77b, HG78a, Hay71b, Hig71, KLCC77, Kau70b, KS77b, Lee78b, MT74a, Mar71, MST^{+}75, NTK72, Nie70a, Red77a, Reu72, Riv77, TM76, Vai72, Vin78, Yua74]. Minimal-State [Nie70a]. Minimalization [Arb70]. Minimally [SR74b]. Minimization [BK79, BI72, BS77, Bis71, Bis74, BV79, Che74, CE75, CM79d, DBC73, DH79, Ehr72, GR73b, GM70, GD73, HG78b, Iba71, KW70, Kan73a, Kan73c, Kel70, Koh71, LD72a, Meo75, MS70, ML76, Pap79, RB75, RNMP77, SO79a, SC72a, SC72c, Sur75, Voi77b, YI74, Yue75a]. Minimize [Ake70]. Minimizes [JB76b]. Minimizing [FM78, Hun72, RB76, Yan73, YB75]. Minimum [BC71b, Bet71, CC72, CDJ73, DTI78, IM71, LM79, LL79, Nak79, Nie75, Pat72, SK78b, SCH72, Voi77a]. Minimum-Perimeter [SCH72]. Minneapolis [Shi70]. Minnesota [Shi70]. Minor [Gri78a]. Miss [Eas78]. Mission [FD74a]. Mix [Mal72]. Mixed [Yen79]. Mixed-Operation [Yen79]. Mixture [BD75b]. MLP [Guf71]. MLP-900 [Guf71]. Modal [JW76]. Mode [Byr71, MS76, Ung77b, Yen79]. Model [Arn73, Ban74a, BW76b, BF75b, BP72, BY79, Chi71, FS74, Gel71, Hay71a, Hay76a, Hoo77, IS76, Jai77c, Joh71, KP74, Kim70, Kim79, Law70, Low73, Mat71b, MK79, Mut76, Per78, Rau79a, Sie79, SN79, SKM78, Ung77a, VN78, Yam70]. Model-Based [Per78]. Modeling [BCSW76, BCJ^{+}71, BCH^{+}73, BT76, Mal72, Mat71c, RC72, Sie75, ST76, Var79, Yan74a]. Modelling [JA74]. Models [CK79, CLL78, Dir72, Fre77, MH76, MH78, Mil73, Pea75, Reg73]. Modem [McL72, Swa72b]. Modern [SM71b].

Modification [NR74]. Modified

[Agr76, DS72a, Das73, Gra73, Hil79, Kit78, LK75, Rub75c, Smi76, Tom70, TM78b]. Modifying [Hay74]. Modular [AS74, AC71, Cer78, Dav77a, DM78a, DW72b, GR71a, GST79, Hwa79, KK78, KS79, NM77, OW73a, SM75c, Sut74, TG78, TK74, Toy71a, Wak75c]. Modularity [O’K73]. Modulated [Ern72]. Modulation [Kon71]. Module [Ake70, BK76, KH77a, PFL75, Sie75, Voi77a, Voi77b]. Module-Level [BK76]. Modules [Ale71, Alm78, BEGW72, FM71, GT74, HK74, Kel74, Lan76, LV75, NA72, Ops72, OW72, Pat73, Pre72, SN79, SB74b, Sur78, TW77, Wak75c, YT70]. Moduli [O’K75]. Modulo [Boy74, CR72c, Pap79, YC76]. Modulo- [Boy74]. Modulo-2 [Pap79]. Molecular [Fre77]. Moment [DBM77]. Moments [SW71b]. Monitored [de 71]. Monotone [Iba74, Lam79a, Mag73, Riv77]. Monotonic [BG78, Dan73, Mat71b]. Monotonicity [Fon71]. Monte [AA76, SF74]. Morphisms [Yan75b]. MOS [MJ78, Dir72, EZS78, Lar71, Liu75, Liu77a, Liu77b, MJ77]. Mosaic [Skl72]. Mosaics [DS75]. Motions [HEWB72]. Movement [CW79a]. Moving [AD75, CA77]. MSI [PGS70]. MU5 [LTE77]. Muller [KS75, KS77b, PP75, SR75, SO79a, Swa72a]. Multi [Kam70a, Nak79, SD74]. Multi-Head [Kam70a]. Multi-Output [Nak79, SD74]. Multi-Tape [Kam70a]. Multicache [CF78a]. Multicategory [Glu71]. Multicell [SNEL79]. Multichannel [Gil70a, Ric72, Win75]. Multiclass [Tou71]. Multicode [KR78a, NT79]. Multicomputer [KK79a, LTE77, She70d]. Multiconnected [TWB78]. Multidimensional [And71, Bat77, CF79, GR72a, GR73b, Kun79]. Multifont [Sch78c]. Multiform [KD73]. MultiFunction [Ham73]. Multigate [Win70a]. Multihypothesis [Dev74, Tou75].

Multilevel [Glu71, Liu77b, Ram76b, SN71, TG72b]. Multilingual [LS71]. Multioperand [AO79]. Multiple [AR78, And74, Bos79, BH71a, BT76, BH73, Bre75, BCS76, CSJ74, CK79, Chu70, CD73, Chu76, CBS71, DS75, Dos77, DW73b, FM74, Fos73, FT76, FT79, GRR72, GS79, Ges70, GLR71, God71, GS74, HC73, ISN76, JM73, KH77a, Kar74, KK72, KM75b, Lan77, LB74, Lin70, Mag70, Maj72c, MM78b, Mil74b, Mor79, Pra78d, PG75, RT73b, SM71a, SW73, Smi78, Smi79a, Smi79d, SN71, TG79, UC75, VVA78, Vra77, Wak75b, YY75]. Multiple-Bit [Lin70]. Multiple-Choice [Dos77]. Multiple-Fault [FM74]. Multiple-Input [BH73, Bre75, CD73, Chu76]. Multiple-Microprocessor [FT79]. Multiple-Output [DS75, Mag70, SN71]. Multiple-Response [Lan77]. Multiple-Stream [Mil74b]. Multiple-Valued [GLR71, ISN76, KH77a, TG79, Vra77]. Multiple-Word [CBS71]. Multiplexed [HW76, MMB78]. Multiplexer [TM76]. Multiplexers [Lan78]. Multiplexing [LC72]. Multiplication [AC71, BBC72b, BW73, Bla74, BH71c, BB75, Che71b, CF78b, DS72b, JB76a, KG72, Kro74, Lev70, Lin70, MK71b, RGK72, Rub75c, SR72a, SW73, SKG77, Tom75, TE77]. Multiplicative [Bak75b, Bar77]. Multiplier [CW79b, Goy76, Ken77b, LR71b, MG75, MS78, Obe72, Pez71a, Rob70a, Swa73b, Pez71b]. Multipliers [CL72, GG75, HW70, Hwa79, NR77b, Nus78, Sik70]. Multiplies [HO79, Sil77]. Multiply [She71]. Multiplying [Mar74, TP72]. Multipoint [CR72b]. Multipole [Ges71]. Multiprocess [Lam79b]. Multiprocessing [Bae76, BCHL73, KV73, Reg73, RV73]. Multiprocessor [Bha77, Cov74, De 79, EFLR76, FB73a, HS75, Koh75, Kur74, Lam79b, Mal72, RCG72, Rau79a, Rum77, SK75, Sch79b, SD79]. Multiprocessors [Bha75a, Hoo77]. Multiprogrammed [RC74]. Multiprogramming [AI78, BCHL73]. Multipurpose [GHJ77]. Multiserver [Reg73]. Multispectral [Deu71]. Multistep [Ben70]. Multitalker [KM78]. Multithreshold [DVS74, DV78a, GC71a, GC73, Ish77a, Ish77b, Set70a, SR72b]. Multivalued [BF75a, DMR77, DM75, DM76, Eps73b, HE79, Ish77a, KS75, KS78c, Kol72, LL72, Mar78, McC79, NSR74, PP75, Sin77, SS77, Sri79, SC72c, SS72, VW72, Yoe70b]. Multivariate [BCL75, FO71c]. Multiwire [MMO74]. Munson [HTD72]. Mutual [BR74, Vil74]. Myopic [BB78b].

n [TOS71, KS79, Rat76]. N-Modular [KS79]. N-Tuple [Rat76]. N.Y [Tab70b]. NAND [CCB70, Cur79, Hay71a, LM79, LD72b, Pap78, SN71]. Natural [BDR77, GC77, Mur70]. Near [AB78, Bou74b, Hay71b, JP73, Pat73]. Near-Minimal [Hay71b]. Near-Optimal [Bou74b, Pat73]. Nearest [Cha74b, FBS75, FN75, GY77]. Nearly [Cur79, Yue74b]. Necessary [Smi79d, Spi71]. Necessity [Kau70b, Riv77]. Needs [Jos72b]. Negabinary [Agr78a]. Negative [Agr75a, Cha74a, Iba71, IM71, Kri71a, NTK72, Nak79, Paw71, SCK73a, SCK73b, Wad71, Zoh70]. Neighbor [Cha74b, GY77]. Neighborhood [O’C75]. Neighborhoods [ADMR78]. Neighbors [FBS75, FN75, JP73]. Nested [Hop70]. Nets [Ama72, BB73, Mor79, NN73, TKH71, Tun70b]. Network [Bos79, CG74, CCB70, CR72b, CCF79, CF73, EFLR76, Hen71, Hol79, KO78, Lan76, LS76a, LD72b, Lee78b, Len78, Mar77, Mas78, MPS78, OY70, SA74, Smi75, Tom70]. Networks [AS74, Abr75, AM77, AM79a, AA75, AA76, Ake70, Ake73, BK76, BCSM72, Bei70, Bei71, Ben71, BK73, BH71a, Bre72a, BCS76, BA71, BY79, But75, But76, But78b, CK74, CW79c, Cle73, CYNM79, DR74, DB76, Dav74, DW72b, DVS74, DV78a, EZS78, Fra72, FD74b, FD77b, GRR72, Gec77, GC71a, GS74, GM71, Har72, Hay71a, Hay74, Hay76b, Hay78b, HW79, Iba71, IM71, Ios75, IA76, Ish77a, Ish77b, KH77a, KK78, Kau74b, Kin71c, Kle76, KS78b, Koh71, KK72, KB75, KK77a, KK77b, Lam79a, LB74, LD72a, LD74, Liu75, Liu77a, Liu77b, Mah74, MH73, MH75, Maz71, MC71, Meo75, MI72, ML76, NTK72, Nak79, Pai73, Pap77a, PM75b, PR73, PR74b, Pra78b, PMB77, Ram73b, Red77a, Ree70, Ris74, RR74, RH75b, RK71, Sah74, Sah77, SR74b, SR75]. Networks [SM75a, SM71a, SD74, SK77c, She78b, SY75, Si78, Sie77, Sie79, SR71, SM78, Smi79b, SN79, SS77, SN71, SH79, Tak76, TWB78, Tho78, TS79, TBY78, Uhr72, Van72, Vin78, Wan75a, Wan75b, Wea75, Whi71c, YI74, KD79]. Networls [SM75c]. Neumann [Yam70]. Neustadt [Tab70b]. Newton [Hun71, Pav77]. Next [AH76, Ano75b, Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, MS73a, Sau72a]. Next-State [AH76, MS73a, Sau72a]. Nine [Mut76]. Nine-Valued [Mut76]. Noise [Boy74, CY76, EGG72, Jai77a, Mar78, MMO74, Sok72]. Noise-Perturbed [CY76]. Noisy [DR75, KS77a, LF77, Tru76, You70]. Non [Kar77b, MH70b, MH70c, RS71b]. Non-Abelian [Kar77b]. Non-linear [RS71b]. Non-Poisson [MH70b, MH70c]. Nonclassical [Mor76b]. Nondeterministic [DW72b, KW70, Moo71]. Nonfundamental [Ung77b]. Nonlinear [Bar74c, CA74, Du71, Far75, FB74, GH72, Hor72, Hun77, Kai70, KF72a, Kru71, LT73, Max77, ND73, NW79, OF73, Tab70a, Tab74, TL73a, WK73, Whi72]. Nonlinearities [Saw77]. Nonlinearly [AA71]. Nonnumeric [Muk79]. Nonorthogonal [Cal70]. Nonparametric [BT70a, DTI78, Fri77, GY77, JLH73, KF72b, KF72c, KNF76, MS76, Ram76c, Rub76b, Whi71b, Bel76]. Nonrecursive [Dav71, Zoh73c]. Nonredundant [Wei75]. Nonresetting [JM71]. Nonrestoring [BJ75, Chi72, MK71a, Rhy71]. Nonsquare [All75a, Van77]. Nonsupervised [CF70b]. NOR [BCSM72, LM79, Wan71]. Normal [BD75b, CM79c, HG78a, Mor70c]. Normalization [KCLB71]. Notation [Ber77]. Note [Agr77, Ano77h, Bre72b, Cha79, Ehr72, Gon79, Goy76, Gra73, Har70b, KO73, Kan73b, Kod74, KS77b, Mcd72, O’K75, Ols70, Pre70, Qua74, Red74, Red77b, Reu72, Ros75, SR72a, SIF76, SB74b, Ulr71, Wor77, Wor79, Wyn71, Yue75b, Zal71]. Notes [CR72c, Pag72]. Notice [Ano70-27, Ano71x, Ano71y, Ano72-47, Ano72-48, Ano72-49, Ano73g, Ano74g, Ano74h, Ano74i, Ano75g, Ano75h, Ano76f, Ano77i, Ano77j, Ano78u, Ano79-29, Ano79-30, Ano79-31, Ano79-32]. Novel [Agr78a, Ban74b, NT70, Toy71b]. November [Ano76y, Ano77z]. NP [Rob79a]. NP-Complete [Rob79a]. NPN [Bau72]. ns [Pez71a]. Number [Ano75-28, BB72, Ban74b, BM73, BM74, BM78, Bau71, BB78c, Bor72, Bre73b, But75, DS72b, FB73a, Har73, Hay71b, HE79, IM71, Jul78, KKK73, KKK74, KMC72, LE77, LE79, NM76b, NM76a, New70, Nus78, O’K75, PW79a, Par72a, PB71, RN77, Sag77, SK79b, SB77b, SA75, Swa79a, Tom75, TJM79, YL73, Yue77e]. Numbers [AR78, BBC72b, Bau77, Bre70a, MK71b, Maj72a, Mar73c, MS77b, Yue75c, Yue77c]. Numerals [Yue73b]. Numeric [KS77c, Mat70]. Numerical [Cod73, Eks74, Eks73, Fer77, Har74a, Kor76, MR75, Pav75a, Pic74, PH73, RH73, Shr71, TM76, Wor76].

O [VN78]. Object [Nev76]. Objective [HEWB72]. Objects [AB76, BG75, CC73, EUY^{+}72, MA77]. Observability [Cer78, MR72]. Obtained [Dav71, HTD72, SC77]. Obtaining [SK71, SS75a]. Occurrence [MR72]. Octal [PB71, Yue73b]. October [Ano76z, Ano77-27]. Office [HSS78]. On-Line [All75b, HPP77, Lud74, MS75a, SSR72, TE77, Wei75, Wil70, Wil77b, You70]. On-Set [DGC74]. One [Bar71, BDR77, DR75, Geo70, IH72, Kun79, LHSM74, Pav75b, PB76a, Tom74]. One-Bit [LHSM74]. One-Dimensional [BDR77, DR75, IH72, Kun79]. One-Way [Bar71]. Ones [KO78]. Only [BB75, Dav70b, Mon74, Pra78b, SY75, GM70]. Op [FGR71]. Op-Code [FGR71]. Open [Gon79]. Operand [AR78, SW73]. Operands [Avi73]. Operating [Bas73, BS74, BR76b, CW75, Fre74b, FM71, Mil74a, WKD70]. Operation [FG71a, Hig78, Ung77b, Yen79]. Operations [Ano76u, Cha73b, Deu70, FL75, Jul78, KMC72, SC77, Wak74]. Operator [Jai77b]. Operators [KR73, Mor70c, The75]. Optical [Cas73, CH75, CS75, CR78, Far75, FHTS78, HND75, Koc75, Koz71, MJTW75, Sta75b, Stu75, Win75, GW75, LH75]. Optical-Digital [HND75, Sta75b]. Optical/Digital [CS75, CR78]. Optimal [BS78, BCSM72, BD75b, Bou74a, Bou74b, Can71, CG74, CW79c, CS79, CYNM79, Cur79, FB73a, FS75, Ful72, GLL79, Hay78a, Ios78, Kau74a, KH77c, Kri70a, LT73, LHSM74, MS77a, MG71a, Mon74, MI72, Nai79, NFS77, Noe73, NR74, Pap78, Pap77a, Pat73, Pav75b, PM77b, Pea75, RCG72, Sam70b, SH77, SF72, Tab74, Tan78, TW77, WG73]. Optimization [Age76, AK75, Bar74c, Bur74, DS75, JM71, KR71b, KK70, LNL72, MM73b, Rob79a, Rot78, Saw74, Tab70b, TWB78, TG76]. Optimum [GM76, NR76b, Ops72, PC72, Rob79b, Sok72, SHR72]. Order [Car71a, EH76, GC77, Glu71, RS71b, SF77, WK71]. Ordered [Bro77, GY77, Kun75, RTW78, Swa74a, Ulm70]. Ordering [Abe72, But78a, CL77, FK70a, KJ79, TF73]. Orders [Pea76]. Ordinary [Fra78, FL75, New75b, RS71b, Wor76]. Organization [Bar74b, Bas71a, BCH^{+}73, BD77, BK71b, De 79, Flo72, GLW74, HC70, Hsi71a, PM78b, RC71b, Ram73a, Sie74, Sto70c, Tho74, Voi75, WG74]. Organizations [BCB74, Cov74, EFLR76, Fly72b, SK75, VN78]. Organized [Cha73b, Gri78b, NR75, Red78a]. Organizing [Ama72, JB76b]. Orientation [DF78]. Oriented [AM77, BBLT73, Bre73a, Erc77, Fel79, Fis76, Gil74b, RA78]. Orthogonal [And70, CC73, CL79, HB75a, KT72, NFS77]. Other [Jul78]. our [Ano79b, DM76]. Outcome [CF70b]. Outer [AP76a]. Outlook [BG75]. Output [Ano70y, Bou74b, CW79b, DS75, Fre79b, GSB^{+}76, Kri73, Mac70, Mag70, Maj72c, MR72, Nak79, Ogu75, PRD71, PM78a, SK78a, SD74, Smi75, SN71, TB75, Vai71, Vai72]. Output-Memory [Fre79b, Vai71, Vai72]. Outputs [Chu70, Dav70a, FK74, Obe72, SM70b]. Overflow [BK71a, CF78b, LL75]. Overflows [Gar78]. Overlapped [Hig78]. Overlapping [AL71]. Overview [AG73, Ram71, Vra77].

Package [BCC78]. Packaging

[Chi78, Hou72]. Packet [CG74, MN78]. Packet-Switched [CG74]. Page [GF78]. Paged [TI72b]. Paged-Memory [TI72b]. Pages [Pat72]. Paging [ABK76, GF78, Smi76, Tri77a, Tun70a, Yan74a]. Pair [Pag70b, Yeh70]. Pairs [PHG77, SB74a, Tom72, YT76]. Paper [Ano71-52, Huf76]. Papers [Ano73e, Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77e, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, Ano78l, Ano78m, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano78v, Ano78w, Ano78x, Ano79r, Ano79s, Ano79t, Ano79u, Ano79v, Ano79w, Ano79x, Ano79z, Ano79-27, Ano79-28, Ano79y]. Parallel [Agr75b, Ano73w, AM79c, Bak75b, BCKT79, BS78, BW73, Ber72, Bla74, Bre71a, BKM73, BD77, BK71b, Buz73, Cha73b, CCES74, Cha74d, CKL77, CK75, CW79b, CW77, CSY75, CM79b, DF75, DEA73, DS72b, Deu70, Fen74, Fen77, FR72, Fra78, FW78, GG75, GR71b, GR72b, IC78, KO78, Koc75, KS73, Kro74, Kru73, LM79, LT73, Len78, Les71, Li77, Man71, Mar73d, MF75, Mil73, NS79, NN73, PM76b, Pre78a, RC71a, RPL73, RFL76, RL78, SK77a, SK77b, Sha78a, SN72a, SCL76, SB77b, Sta74, Sta75a, Sta75b, Ste72, SKG77, Sto70c, Sto71, Svo73, Swa74a, Swa73a, Tab74, Tho78, TP73, TF70, Toy71b, WC74, Wor76]. Parallel-Acting [Man71]. Parallel-Decomposable [DF75]. Parallel-Pipelined [BD77]. Parallel-Sorting [Pre78a]. Parallel-to-Serial [IC78]. Parallel/Pipelined [Li77]. Parallelism [Ada78, Bar78, DT76, Das78, FL77a, RF72]. Parameter [Boo70, Bro71, LNL72, SKM78, SSR72]. Parameters [BPC75, BD75b, Dui76]. Parametric [New75b]. Parity [Hen71, KM73, OY70, Toy71b]. Parsers [Jol76, TF78]. Parsing [Lew74]. Part [Cur70, LM70, Loo70, SP72b, SP72c, Sri71a, Sri71b]. Partial [BM77b, CRB72, Cov74, DS71, God71, Jol76, Kar70, KD73, MA77, SF74, Uhr73, Wit70]. Partial-Fractions [God71]. Partially [Sch75, Smi71b, Wak74]. Partition [GC73, Pag70b, PHG77, RB76, Yan73, YB75]. Partitionable [GB71b]. Partitioned [Mor70b, MM70, Pat72]. Partitioning [ARS71b, BT70a, Bre70b, DM78b, Fri77, GHJ77, MM73b, PGS70, RH73, ROW71, Sto77]. Partitions [BD75b, LR71a, MR74, Yan74b]. Parts [Per78, YT77]. Parzen [Dui76]. PAS-II [WN76]. Past [TA71]. Patch [RC71c]. Patches [Mah72]. Patching [RC71c]. Patern [Pav74]. Path [CRB72, Hay78b, Koh75, Rub74, Rub76a]. Paths [KK79b]. Pattern [AA77, Ama72, Ano78h, Ano78i, Ano78j, Ano78-37, Ano78-38, Ano79k, Ano79l, Ano79c, Ano79d, BC71a, Bab72, BD71, BT70a, BT72, BB78a, Bre73a, BP72, BR74, Cal70, CDO78, Cha73a, Che71a, Chi71, CR72d, Chu74, Dev74, DN77, FE78, FP75, FB73b, FR76, FF75, Gil74b, Gil74a, Hay75, HF71, Hus72, Kit78, LT70, LNL72, LS72, LF72, LF78, MM73b, Mic74, MF76c, MG71b, NC78, Nad71, PM76a, Pre71b, Rem74, Sam70a, SLC75, Sel74, STG71, SN72b, Sto74a, Tak76, Tou71, Tou72a, Tou72c, Tou74, Tou75, Vil72, VV71, WG73, Wea75, WL75]. Pattern-Recognition [Tak76]. Pattern-Sensitive [Hay75]. Patterns [Ama72, BDR77, BL75, Glu71, GD72, Hus74, KKA71, LF77, NFS77, Rat76, TG75, Wan75c]. Paull [Pra71a]. PC [AB79b]. PCN [Mat71a]. Peaks [NM76b]. Pen [BdB70]. Penalties [Low73]. People [Ano74-28, Ano75-29, Ano75-30]. Percolation [FR72]. Perfect [Sto71]. Performance [Bea78, Bha75b, Bha77, CM73b, CBS71, FD75, Gri76, GF78, Kla71, Rau79b, Rus72, SK75, Tri77a]. Performance-Related [Bea78]. Performing [Wak74]. Perimeter [SCH72]. Periodic [Gaj75, GB71a, GB71b, Hwa73, Law70, RS72a, SW70a, SW70b, Set70a, Usa75]. Peripheral [Bar74a, Flo75, MF76a]. Peripherals [Ano73x, Ano73y]. Permanent [KS79]. Permutable [WY73]. Permutation [CW79a, Har72, Orc76, Ram73b, KD79]. Permutations [Che78, Len78, WC76]. Permuter [BBC72a]. Perspective [AR74, Kim75]. Perturbed [CY76]. Phase [Dir72, MAB78]. Phase-Shift [MAB78]. Photo [SW71b]. Photographs [BD71]. Photomosaicking [Mil77]. Photomosaics [Mil75]. Phrases [KM78]. Physical [Key77]. Pi [IR76]. Pi-Algorithm [IR76]. Pictorial [Ano70y, FE73, Mac70, Ram76b]. Picture [AL71, KS78a, KM75a, Kru73, Mor74, ND73, OC76, SB72, ZRD75]. Pictures [ADMR78, Hun72, MF75, SC77, SP72a]. Piecewise [ADMR78, Can71, Cha73a, KP75, Lee78a, Pav75b, PH73, Tom74, VD76b, Van75]. Piecewise-Linear [Tom74]. Pin [LR71a, Rus72]. Pipeline [Coo77, Dev75, GW70, KSA74, Maj76]. Pipelined [BD77, Li77, Mor79]. Pipelines [Roe78]. Pipelining [HF72, JA78, KV73]. Pixel [PR78]. PL [Jor74b, Rem72, Wex71]. PL/I [Jor74b, Rem72, Wex71]. PLA [OH79]. Placement [Bae74, HF74]. Plan [Ano76c, Ano76d, EUY^{+}72]. Planar [Ano72-39, CA77, Par71]. Planarity [Rub75a]. Plane [CY76, NW79, PH74, Sam70b, ZR72]. Plateaus [BDR77]. Plotter [Ath75]. Plotting [Wri73]. PLZT [BL75]. Pneumoconiosis [JP76]. Point [Bau77, Boh77, Bre73b, Cod73, HW76, HF74, Hig76, Ken77a, KKK74, KPS77, Mat70, MMB78, Mor71, O’C75, Pak71, RB78, SV75, SR77, Vic71, Wei75, Yoh73, Yue75c, NA76b]. Points [FD73, LSB77, OC76, Sta74, WA73]. Poisson [Buz73, Chu70, MH70b, MH70c, MN78, Sch73]. Polarized [MT74a]. Poles [God71]. Policies [BB78b]. Policy [Ano73e]. Polish [Hus73]. POLLY [Bec71]. Polygonal [AD75, Pav77]. Polygons [FP75, Sch78a, SCH72]. Polyhedra [Lou70]. Polylinear [Cur71b, HSP78]. Polynomial [BB74b, Gri78a, Lee78a, MT74a, MC70, Pav75b, PH73, Rem74, Tou72d]. Polynomially [IS75]. Polynomials [Has74, Mar73d, VD76b]. Popov [Ano72n]. Positive [HG79]. Post [Eps73a, NSR74]. Postal [Sch78c]. Postcorrection [BJ75, Rhy71]. Postcorrections [Chi72]. Posteriori [TH79]. Postfix [Ber77]. Postprocessing [Dos77, RH74]. Potential [RF72]. Power [Set70b, WGS78, YI74, Yue74b]. pp. [Tab70b]. Practical [TBY78]. Practice [AGM^{+}71, Bre71c]. Practices [Rux71]. Pragmatic [Mil74a]. Precisely [SP72a]. Precision [Bre73b, HC70, MM73a]. Predicate [JW76, San76]. Predicate-Calculus [San76]. Prediction [Boo70, Toy71b]. Preemptive [Gon79, VN78]. Preface [Ano71-50, Ano72-71, Ano76-27, AG73, Shi70]. Preliminary [Dub79, Koh75, Lar71, Pin71]. Prepaging [Tri76]. Preprocess [Uhr72]. Preprocessing [Gra72, HKD^{+}71, Kuc70]. Presence [Ast78, CM79c, Mar78]. Preserved [WL74]. Preserving [Yan75b]. Preset [FT76]. Press [Tab70b]. Previous [Hen71]. Prime [Ben71, BM70, BH71c, Bub72, Das71c, DK72, DBC71, HW75, Hwa74, IR76, Maj72c, Mor70c, NRS74, Pap77b, Reu75, SCL70]. Primer [Huf76]. Primitives [Kli71, MBSS75]. Principally [BF75a]. Principle [Koh74]. Principles [Bib74, New75a, Rux71]. Print [Gra72]. Printed [AKM^{+}71, AN74, SO73]. Printer [Ano70y, Mac70]. Printing [Ern72, Yue73b]. Priori [Gri73]. Privacy [TW76]. Probabilistic [AA75, Hu75, Kri73, MH76, MH78, MB77, PM75b, Paz70a, Tho76, Vil73b, Yoe70a, de 71]. Probabilistically [FK78a, FK78c]. Probabilities [MPS78, PM75a, PM78a, SD79]. Probability [BT73, Dui76, Hus74, Ogu75, Tou71]. Problem [Ake74, Bro75a, CH72a, CH73b, DTI78, Du72, FL77a, Fis76, FS78, GHJ77, Geo70, GLL79, Har70a, Kau74a, Kur74, Lou70, Nam70, Nie70b, PS71, PR74a, PHG77, RA78, Reu75, RC74, Sch75, She70a, Sri76, SIF76, VD76a, Wak70c, Wex71, Wri73]. Problem-Oriented [Fis76, RA78]. Problems [Bar74c, CF78a, CCF79, Gar72b, Has74, IS75, Kar70, Lam79a, Lee75, MOW76, MG71a, Par72b, Sto74a, Tab70b, Tom73, Vic71, WK73, YI74]. Procedure [ARS71a, BT70a, Chu71, DS72a, Dut74, EH78, Fis76, Gel76, HTD72, KP74, Kor79, LF72, MT71, Noe73, NR74, Red72a, Rob70b, ROW71, TG78, Tou72b]. Procedures [BK72, BH74b, Cur70, FG76, Gol72, Hla78, RC72, Rob70a, SKM78]. Proceedings [Ano79b, Ano79-54]. Process [Mai71, REFN76, Tab70a]. Processable [RPL73, RFL76, RL78]. Processed [CB72]. Processes [MVW79, SK71]. Processing [AP76a, Ano78v, Ano78w, Ano78x, AL71, BG75, CCF79, CDL78, Cor71b, DRS75, Far75, Fen77, FW78, FR76, GSJ^{+}76, GR71b, HF71, Hop71, HA70, HB75b, Koc75, KSW74, KF73, Kru73, LT73, Li77, MJTW75, ND73, PC72, RNR75, RTKH77, Roe78, Sch78b, Sch74b, SC77, Sta75a, Sta75b, Sto71, Svo73, SH79, Tab74, TP73, Tun70a, TE76, VM78]. Processor [Aus70, BCKT79, Cap76, CS75, CR78, CDS73, CK75, CK79, CW77, CL72, DS70, FL77b, GLMR71, Guf71, HND75, HSS78, Law75, LS71, MM78b, Mil74b, Nut77, RSH79, Red71a, Red79, RSL76, SH77, SNOS79, Sie77, SN79, Ste77, Tho78]. Processors [AT70, BK79, BT76, BD77, Fen74, Fen77, FB73a, GC77, Gri78b, Kri77, Lan76, McL72, NR75, PR72, Red71a, Ros74a, SB77b, Swa79b]. Product [AP76a, BM74, Bla79, Dru71b, Mar73a, SGR78, Yue75a]. Productions [Jol76]. Products [AB78, God73, HLD70, Oja79, Pap79, Wak70b]. Professionals [Ano77d, Ano78k]. Program [Ano78c, Ano78f, Ano78d, Ano78e, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Bae74, Cov74, EH76, FLR78, GR71b, How75, Kje72, LM78, Rau79b, Rej72, Sch78b, Wul74]. Programmable [AB78, Fle79, GSB^{+}76, GLMR71, HND75, Kam79, KAN79, NR76a, OH79, PM77a, PW79b, Smi79c, Woo79]. Programmed [Fre74a, Rot78]. Programming [BS77, BW76b, Bre71c, Coo72, Den74, Fel79, FW78, Geo70, Gri74, Jor74a, Jor74b, Ker73, MI72, Pap77a, Rem72, RH73, Sil73, Uhr73, Voi75, Weg76, Wex71]. Programs [Bas71b, BBP75, CR72a, Cha72, Kar77a, Kin71a, KMC72, Lam79b, MOW76, PB76a, San76]. Progress [SG71]. Progressive [ST79]. Projection [AM79b, SC77, The75, FT74]. Projections [Cal70, CC73, KM75a, SF78, Wan75c]. Projectors [Oja79]. Proof [Gel76, Pra71a, Rub75c, Vil73a]. Proofs [Moo71]. Propagation [MM73a, RHR76a, RHR76b, SW70a, SW70b]. Properties [BM73, Bre71b, CL71, CF79, Dav71, Fis74, Gal78, GR72a, Gra71, GM76, HN78, Hay76b, KK79b, Kri73, Mar73a, MG71b, Pra78a, RGK72, SC77, Sun79, TD77, The75, Tou72a, Wan75b, Wea75]. Property [DF70, HN71, Hwa72, SF77, She78b, SB72, WG74]. Proposed [KK70]. Protection [FSLC78]. Protocol [WN76]. Prototype [EUY^{+}72]. Prototypes [Cha74b]. Prover [Ern76]. Provided [CF70b]. Proving [Abd71, Fis76, Kin71a, MOW76, Mor76b, Rei76]. Pseudo [Kli74, PW79a, Pap77a]. Pseudo-Boolean [Kli74, Pap77a]. Pseudo-Random [PW79a]. Pseudodivision [SK71]. Pseudoinverse [PD77]. Pseudonoise [Hur74, SP72b, SP72c]. Pseudorandom [Dav71, MH70b, MH70c, Mar73b, Mar73c, MAB78, NM76b, NM76a]. Pseudothreshold [Bau71]. Publications [Ano72-57, Ano72-58, Ano73i, Ano73j, Ano73k, Ano73l, Ano74j, Ano74p, Ano74k, Ano74q, Ano74l, Ano74r, Ano74m, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74n, Ano74y, Ano74o, Ano74z, Ano75o, Ano75p, Ano75q, Ano75i, Ano75r, Ano75j, Ano75s, Ano75k, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75l, Ano75y, Ano75m, Ano75z, Ano75n, Ano76k, Ano76l, Ano76i, Ano76m, Ano76j, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano78y, Ano78z, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78-32, Ano78-33, Ano78-34, Ano78-35, Ano78-36, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42]. Publications [Ano79-43, Ano79-44]. Published [Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z]. Pulmonary [SH72]. Pulse [BH71b, FG71b, FG71c, GD72, Hor71b, Kon71, Obe72, RHR76a, RHR76b, SW70a, SW70b]. Pulse-Ratio [Kon71]. Pulses [MH70b, MH70c]. Pure [Myk75]. Purging [Los76]. Purpose [DRS75, GGMM71, HC70, SF74]. Purposes [Ben70]. pursuit [FT74]. Push [Har74b]. Push-Down [Har74b]. Pushdown [Arn70, Kam70a].

Quadratic [CY76]. Quadric [Mah72, SF78]. Qualified [Win76]. Quality [FLR78]. Quantitative [DB74, DF78, FD75]. Quantities [Mat71a]. Quantization [CB72, CJ74]. Quantized [BD75a]. Quarter [JB76a]. Quarterly [Ano78h, Ano78i, Ano78j, Ano79k, Ano79l]. Quasi [MG75, Swa73b]. Quasi-Serial [MG75, Swa73b]. Question [BCSW76]. Question-Answering [BCSW76]. Queue [Sch79b]. Queueing [Reg73]. Quinary [PW79a]. Quotients [HLD70].

R [AH76, Wak70c]. R-70-31 [Wak70c]. R70 [Ake70, Arb70, Arn70, Bar70, Ben70, Ber70a, Brz70, Cha70, Cut70, Dav70a, Fri70, Geo70, Gil70b, Goo70, Gra70, Har70b, Hed70, Hop70, Jum70, Kam70a, Kam70b, Kar70, Law70, Lev70, New70, Pag70a, Paz70b, Paz70a, Pre70, Rob70b, Rot70, She70a, She70b, She70c, Skl70, Su70, Tsi70, Tur70, Wag70, Win70a, Yam70, Yeh70, Yoe70a, You70]. R70-1 [Har70b]. R70-10 [Arb70]. R70-11 [Yam70]. R70-12 [Rot70]. R70-13 [Win70a]. R70-14 [Cut70]. R70-15 [She70b]. R70-16 [Skl70]. R70-17 [Fri70]. R70-18 [Goo70]. R70-19 [New70]. R70-2 [Hop70]. R70-20 [She70c]. R70-21 [Arn70]. R70-22 [Tur70]. R70-23 [Kam70a]. R70-24 [Rob70b]. R70-25 [Ben70]. R70-26 [Yoe70a]. R70-27 [Ber70a]. R70-28 [Pre70]. R70-29 [Jum70]. R70-3 [Pag70a]. R70-30 [Su70]. R70-32 [Wag70]. R70-33 [Paz70a]. R70-34 [Bar70]. R70-35 [Tsi70]. R70-36 [Hed70]. R70-37 [She70a]. R70-38 [Lev70]. R70-39 [Kam70b]. R70-4 [Geo70]. R70-40 [Ake70]. R70-41 [Cha70]. R70-42 [Dav70a]. R70-43 [Yeh70]. R70-44 [Brz70]. R70-45 [Paz70b]. R70-5 [Gra70]. R70-6 [Kar70]. R70-7 [Gil70b]. R70-8 [You70]. R70-9 [Law70]. Races [BH71b, Bre74]. Rademacher [Yue76, Edw75]. Radiation [BG75]. Radiographic [HKD^{+}71, HE73]. Radiographs [ADL72, BS76]. Radix [Atk70, BR71, CSY75, Dru71a, DV78a, Erc73, Kri71a, Nic71, Paw71, Wad71, Zoh70, Zoh73d, Zoh76b, Bak75a]. Radix- [BR71, CSY75, Erc73, Bak75a]. Rado [Lyn72]. Rahman [Bha78]. RAM [Zoh76a]. Random [AA75, Agr78b, AU74, BI72, Bre71d, DB76, EF76, En77, GD72, Hay75, Kau73, KH77b, KH77c, Len77, Los78, MS77b, MVW79, Mur70, NTA78, Nai79, Paa79, PW79a, Sch79a, SLC75, Sok72, Sri78, Sti78a, Sto75, Wil78, Yue77e]. Random-Access [Hay75, NTA78, Sri78, Sti78a].

Random-Error-Correction [En77, Paa79]. Random-Pulse [GD72]. Randomized [LS72]. Randomly [BG75]. Range [Kri70b, Kri71b, MH75, Sik70]. Range-Transformation [Kri70b, Kri71b]. Rank [Cur73b, GH78, SK74]. Rank-Augmented [SK74]. RAP.2 [SNOS79]. Rapid [AU74, MC70, Roc70]. Raster [HSS78, ST79]. Rate [CL75, MH70a, MR72, NR77b, Obe72, Sok72]. Rates [Nad71]. Ratio [Hus74, Kon71, Rus72]. Rational [SB77b, VV74]. Ratios [Eas78]. Ray [Wei75]. Reach [Ano79b]. Read [BB75, Mon74, Pra78b, SY75, GM70]. Read-Only [BB75, Mon74, Pra78b, SY75, GM70]. Readily [Dev79, Ehr76, Kau70c]. Reading [AKM^{+}71, JM74, Sch78c]. Readings [Fly72a]. Real [AK72, Dep70, Dow73, Gay78, GSJ^{+}76, Goo70, Goy76, HO79, Hen78, HA70, Koc75, Mar78, Pol71, Saw77, Sil77, WKD70]. Real-Time [AK72, Dep70, Dow73, Gay78, GSJ^{+}76, Goo70, Goy76, HA70, Koc75, Mar78, Pol71, Saw77, WKD70]. Realizability [BPC75]. Realizable [Dav70b, Gel71, HE79, Mat71b]. Realization [AA71, CL75, Cur71b, DSBC74, DGC74, DVS74, EGG71, FT76, GH78, GC73, GCC74, GCP^{+}77, HG79, Har70a, HC70, Hen71, Hwa73, KT72, Mag71, MT74b, OY70, PC78, SK78a, SK78b, She70b, Sin72, Sri75, Sri79, TOS71, Zoh76a]. Realizations [Bro75b, Bro75c, CC72, GR71a, GM76, HC72, Hay71b, Kod74, Lan78, MBSS75, OW73b, PM77a, Red72b, RW74, RT73b, Saw74, Ung77c, Wil78, Zei71, Zoh73c]. Realize [PM71]. Realized [ATN70, But75, Hu73, LF74, Paz70a]. Realizers [Han72]. Realizing [Cur70, OW72]. Reasoning [Mam77]. Rebalancing [LP78]. Rebuttal [DM76].

Received [BG75]. Reciprocal [Ric74b]. Recoding [Rob70a]. Recognition [ARS71a, AA77, Ano79c, Ano79d, BD71, BB78a, Bre73a, BP72, BR74, Cal70, Cha73a, Che71a, Chi71, CR72d, Chu71, Chu74, Dev74, DT70, Dos77, Dut74, EK75, Far79, FE78, FP75, FB73b, FR76, Gil74b, Gil74a, Gra72, Har70b, HF71, Hus72, HTD72, JK77, KM78, Koz71, LF77, LF78, Mak71, Mat71b, MA77, Mic74, MF76c, MG71b, NC78, Nay71, PM76a, Pol71, RPL73, REFN76, Rem74, RE71, Sch78c, Sel74, STG71, SN72b, Sto74a, Tak76, TG75, Tjo77, TG72b, Tou71, Tou72a, Tou72b, Tou72c, Tou75, Uhr72, Ull72, Vil72, VV71, WG73, Wea75, Win75]. Recognizer [Nad71]. Recognizing [BS76, CY76, TD70]. Reconfigurable [Mai71]. Reconfiguration [GT74]. Reconstruction [CC73, DB73, KM75a, Wan77, WD77, YRT75]. Reconstructors [JM71]. Record [Hsi71a]. Recorder [Ros71]. Recording [BL75, Hoa76]. Recovery [CJW^{+}71, CR72a, LL79]. Rectangular [Ake72, Skl72]. Recurrence [CK75, KS73, RB72]. Recurrent [Kar77a]. Recursion [TV76]. Recursive [AM79b, AOR74, Bro77, Fri77, Jai77c, Maj72b, Mor70c, NC78, Whi70, Zoh73a, NA76b]. Redefinition [RA78]. Reduced [Bro70b]. Reducing [Ano75-28, HS70]. Reduction [Ben71, Cur71a, JA74, KLCC77, Pfl73, Pin71, PMB77, VVA78]. Redundancy [AS74, AB79a, Atk70, DR74, DM78a, FM72a, Hey71, IS76, Kla71, KS79, LR72, LD74, Los76, Ogu74, SM73a, SM73b, Sie75, Smi78, Smi79b, Wak75c, YT71a]. Redundant [Abr75, BM73, Dah73, DW78, FD74a, FD77b, Lyo70, NA76a, Si78, To73, Toi76, YL73]. Reed [KS75, KS77b, PP75, SR75, SO79a, Swa72a]. Referees [Ano71-52, Ano78s, Ano78a, Ano79-52]. Refinement [ST79]. Refinements [WM76].

Reflected [Agr76, Obe74]. Reflection [Dav72b, Hor71b]. Refutations [And76]. Region [CW75]. Register [Bar75a, BEGW72, HC78, HP75, JW73, Kar76, KK78, Klu77, LF74, Mar73a, Mar73b, MAB78, Max77, Myk75, NR77b, Per70, RT73b, Smi71a, Van73a]. Register-Transfer [BEGW72]. Registering [HEWB72]. Registerless [Mil74b]. Registers [AM79c, Hur74, Lem70, Man71, O’K73]. Registration [ACJ75, BS72]. Regular [Ber70a, HPP79, PM78a, Pat71, Paz70b, Paz70c]. Reject [Nad71]. Related [Bea78]. Relational [De 79, EF76, Kam70b]. Relations [Kli74]. Relationship [DM75, DM76, Eps73b, Goo71, LR71a, MR72, SS72, Tak76]. Relaxation [Cha71, DR77, ZHR77a, ZHR77b]. Reliability [AS74, Abr75, Arn73, Bea78, BF75b, BCJ^{+}71, BCH^{+}73, CLL78, FD74a, HB75a, IS76, IA76, Kor79, KS79, Mal73, Mar77, Mat71c, Mey76, Par72b, RH75a, Sie75, Sti78b, You71]. Reliable [Ano79-54, Hu75, Mai71, Saw75, WSB79]. Remark [Gre75, Vai71]. Remarks [Noe73, OW73a]. Remote [Arm73, PRD71]. Removal [MH70a]. Repair [CS79, GT74, RK75a, RK75b, YO70]. Repairable [Arn73]. Repairing [AGM^{+}71]. Repeatable [MH70b, MH70c]. Repeated [Bor72]. Repetitive [Brz70, Rey70]. Rephrased [Nie70b]. Replacement [EF79, Gel73b, SM75c, TI72b]. Replication [Cov74, LR72]. Reply [Agr79b, All77, Ano73z, Bab74, Bar75b, Bib75, BH74a, CM79a, Che72, CBD70, Chu72, CD75, Coh70, Das72, DM76, Edw76, Ekl73, En79, FCC72, FO71b, GRR74, GR73a, GM72b, HC74, HW78, HF70, KAA77, Koh72a, Koo78, LM71a, Mar75, Mar76, MB79, Men70a, PR75, Pra71b, Rao78, RS73a, RS73b, SM79, SS75b, To76, Yue77a, Yue77b, Zoh71]. Report [Fre74b, Hob73, Kra74, Lom74, Mui74, Pla73]. Reporting [BO79]. Representation [Bau77, Das71b, EF76, FE73, HC73, KR73, KN70, KS77c, Lee78a, Mor74, Mor71, NN73, OF73, SM72b, Sel74, SB74a, TL73a, TF73, Tou72d, Yam70, Yue75c, Yue77c]. Representations [GB71a, GB71b, Pea76]. Representative [Bau72]. Represented [Glu71]. Required [Lev70, Nie70b, PM71]. Requirements [FLBK76]. Requiring [Hay71b]. Research [BBLT73]. Research-Oriented [BBLT73]. Resident [Gar72a]. Resident-Bubble [Gar72a]. Residue [BB72, Ban74b, Ban74c, BM73, BM74, BM78, HC73, Jul78, KKK73, KKK74, Man72b, O’K75, RT70, TJM79, YL73]. Resistor [Pak71]. Resolution [AM77, BK72, DI71, Gel76, Hen76a, Hen76b, Lan77, LS76b, MN70, Pic74, WM76, YK76]. Resolution-Based [Gel76]. Resolution-Oriented [AM77]. Resolver [MS77a]. Resolvers [And74]. Resource [Mil74b, PM78b, SN79, Ung77a]. Responders [FS71]. Response [BV78, FS74, Lan77, Pec76, Sah74, Sah77]. Restoration [HA77, Hun73, Hun77, JA74, Jai77a, Jai77b, PD77, TH79]. Restricted [But76, FM73, KS78b]. Restrictions [KN72]. Restructurable [RF78]. Result [DW72a]. Resulting [KMC72]. Results [ABK76, Bri73, CYNM79, Cut70, DW73b, EZS78, HTD72, Lyn72, SR77, Ung77a, YK76, Yue74a]. Retrieval [CM73b, RTW78, Ste77, WG74]. Retrieving [Wei75]. Retry [She78a]. Reversal [PBC74]. Review [Ano72u, Ano72v, Ano72w, Bas73, Flo75, New75a]. Reviewer [Bar75b, Bib75]. Reviewers [Ano76a]. Reviews [Ano72x, Ano72y, Ano72z, Ano73b, Ano73c, Ano73d, Ano73h, Shi75, Ano72-50]. Revision [Hil79]. Riccati [Gou77, Den76, SG74]. Right [Pat71, She71]. Right-Linear [Pat71]. Rings [DV78b]. Risk [Dev74, Tou75]. Robot [EUY^{+}72]. Robust [Yue75a]. Rollback [CR72a, She78b]. ROM [Jul78]. Root [SK71]. Rooting [RGK72]. Roots [MK71a, Maj72a, Rud76]. Rotary [JK77]. Rotate [NT70]. Rotating [FV79]. Rotation [EO78]. Rotations [And71]. Rounding [KPS77, Zoh76b]. Roundings [Yoh73]. Router [AB79b]. Routing [Abe72, AKT^{+}77, CG74, HW79, KJ79, Lam79a, Sah74, Sah77, SH79]. RST [Sur70b]. RTM [BEGW72]. Rule [Ehr76, Fri77, HS70, KK78]. Rulers [Rob79b]. Runs [Rom78].

s [GSB^{+}76, AH76]. Safe [Chu76, CD78, DGC74, DAA79, GCP^{+}77, HG79, MT74b, PM74, SM75b, TM71, TI72a, TOS71, Toh74, WM74]. sample [Vee78]. Sampled [All75b, ASS72, DB73, SSR72, Wil77b]. Sampled-Data [SSR72]. Samples [Bab72, BF72, PS78, Tou74]. Sampling [Sok72]. Satellite [HEWB72, MN78]. Satisfying [KK78]. Save [Pra74]. Scale [CM70, GC71b, Mor74, VV74, WA73]. Scaling [Jul78, SM71c]. Scan [AN74, HB75b]. Scan-Digitized [AN74]. Scanner [HND75]. Scanner-Processor [HND75]. Scanning [Bec71, EF79, Stu75, Wan77]. Scene [Pav74, RT71, WH78]. Scenes [Gri73]. Schedules [FB73a, Gon79]. Scheduling [AI78, Ful72, Kau74a, Koh75, Li77, RCG72, RFL76, RL78, Sto70a, VD76a]. Schemata [Tho75]. Scheme [AKT^{+}77, Che71b, Kla71, LM70, Loo70, Los76, MF76a, SKG77, Toy71b, WSB79, WL74]. Schemes [BT70b, DF78, FD75, Hus72, Kri70a, Kri71b, Ple77, Pre78a]. Science [Fre74a, Jor74b, Ram76a]. SCIRTSS [HH77]. SCR [WS76]. Script [EK75, Far79, Moo76]. Scroll [Wil70]. Sea [Whi71a]. Search [BI72, Bur74, Cha74d, Dig73, Fis76, Gel76, HH77, LNL72, LS72, LC79, Pag70b, Pat72, Sic76, TS79, WM76, WC74]. Searching [Gri76, RC71b]. Second [JM74, RS71b, AG73]. Second-Order [RS71b]. Section [Ano72-40, Fle79, Wan77]. Sectioning [Hun72]. Sections [Ano76-28]. Secure [Pra78a, SM78]. Security [FSLC78, TW76]. Seeking [KF72c]. Segment [Sta74]. Segmentation [DR77, PR78, PF72, Pav73, PH74, RC71a, Wil71, ZRD75]. Segmented [Sah77]. Segments [Ros74b]. Seismic [SF72]. Seismology [Dav72b]. Selcuk [Sah74, Sah77]. Selected [Ano78v, Ano78w, Ano78x]. Selecting [Chi71, Rob70b]. Selection [ARS71a, And71, Che71a, Das71a, Deu71, FK70a, Kit77, LD75, MT70, MSJ72, NF77, SW76, TV72, Tou72c, WNR74, Whi71b]. Selector [Hol72b]. Self [Ama72, AM73, AR77, BCH^{+}73, CDJ71, CDS73, CHSS73, CS79, CSST73, Dav78, DAA79, Dor72, HA76, Los76, Mai71, MF78, OW73b, Red74, RV74, She78b, Ung77b, Usa75, Wak74, AGM^{+}71]. Self-Checked [OW73b]. Self-Checking [AM73, AR77, BCH^{+}73, CDS73, CHSS73, CSST73, Dav78, DAA79, Dor72, MF78, Red74, She78b, Usa75, Wak74]. Self-Diagnosable [CS79]. Self-Organizing [Ama72]. Self-Purging [Los76]. Self-Reconfigurable [Mai71]. Self-Synchronization [HA76]. Self-Synchronized [RV74]. Self-Synchronizing [Ung77b]. Self-Testing [CDJ71, AGM^{+}71]. Semantic [BCSW76, Hen76b, Wal76]. Semantically [Rei76]. Semantics [JW76, Ros74a]. Semi [Sar78]. Semi-Fast [Sar78]. Semicausal [Jai77c]. Semiconductor [BD77, NTA78, Sri78, Sun78]. Seniority [Osb77]. Sensitive [BH73, Hay75, SR71].

Sensitization [CRB72]. Sensitized [KK79b]. Sentences [Hub76]. Separable [AA71, AR77, CD78]. Separating [Sto77]. September [Ano77-28]. Sequence [Kri73, Sur70a]. Sequences [Ama72, AM79c, CL71, DDSM76, Dav71, FM73, Gal78, Hab77, HC78, HH77, JW73, LE71, Mar71, Mar73a, MH70b, MH70c, MAB78, MST^{+}75, MBM77, PW79a, Qua74, Smi71a, Sut74, TB75]. Sequency [Bro77, CL77, GC77, KH74, Red73b, SO79b]. Sequency-Ordered [Bro77]. Sequential [Abd71, AB79a, Agr75b, AU74, Alm78, Ano75-28, ATN70, Bei70, Bei71, BB78b, Ben71, Bis74, Bou74b, BH71b, Bre71d, Che74, CW79b, Chi71, CD78, Cio78, Cur70, Cur71b, Cur73b, Cur73a, DF75, DTG79, Dav70a, Dav70b, DSBC74, DGC74, DAA79, DW72b, Du71, DH79, Ehr72, EGG72, Far73, Fel79, Fre79b, Fri70, FM73, FG71b, FG71c, FK74, FNSK75, GH78, Gel71, Ges70, GB71a, GB71b, GD73, GCC74, GCP^{+}77, HD71, HSP78, HG79, Har72, HC71a, HC72, HH77, Hol72a, HC71b, Hsi71b, Hu72, Hu73, HA76, Hus72, Hus74, Hwa73, Iba74, Jos78, Jum70, Kai70, KYO70, Kar74, KM73, Kel70, Kel71, KS77a, KN72, Kin70, Kin71c, KK77b, KR78a, KC73, LR72, LV75, LSB77, Len77, Mag71, Mag73, MT70, MT71, MS73a, MS74, Man72a, Mar70, Mar77, Mey71, MVW79]. Sequential [MJ77, MJ78, MT74b, MKO70, NT79, NA72, Nie70a, Ols70, OS76, PM77a, Par72a, Par72b, PM78a, PM74, Paz70c, PHG77, PD79, Pre70, Ram76b, RB75, RB76, RV74, RN77, RT73b, Sau72a, Sau72b, SM74, Saw74, Saw75, SM75b, SR73, Ser73, Ser75, SM76, She77, She70b, She70c, SV74, SY75, Sil79, Sin71, SR71, Smi74, SN72b, Sto74b, Sto75, SHR72, Sun76, SB74a, Sur78, TI72a, Tan71, TA71, TOS71, Toh74, Tom72, TB74, Ulr71, Ung71, Vai71, Voi77a, Wak75a, Wak75c, WY75, Wan71, WM74, Wil78, WZ72, Wil75, WH78, Yan73, Yan74b, YB75, Yua74, Vee78].

Serial [BK71a, Hor71a, IC78, Lit73, MG75, Rhy70, Sit74, Swa73b, Tro75, Wil77a]. Series [Ric74b, Yue75e]. Set [AM79a, Bub72, DGC74, DW72a, FS75, GHJ77, GF78, Lig77, MBSS75, Moo71, Nie75, NRS74, RS71a, Set70b, SF78, Smi76, WGS78, Yan74b]. Sets [AM79b, Ake73, Bet71, Bou74a, Dan73, Das73, DDSM76, DBC71, GY77, GL70, Gre75, HG78a, Hen76b, KS71, Mar71, MST^{+}75, Pra78d, Red73a, SR75, Sch78a, SC77, Wan75a, YT71c, de 71]. Setting [OS76]. Seven [MG71b, Tou72a]. Shading [Gou71, LH79]. Shape [Dav77b, UM77]. Shaped [Oza78]. Share [PRD71]. Shared [Gad70, JP73, Kle70, Mil74b, OW73b, PM78b]. Shared-Resource [Mil74b]. Sharing [Gel73a]. Shift [CF78b, Dav74, Eks73, HC78, Hur74, JW73, Kar76, KK78, Klu77, Lem70, LF74, Mar73a, Mar73b, MAB78, Myk75, NR77b, NT70, Per70, RT73b, She71, Smi71a, TG72a, Van73a]. Shift-Invariant [Eks73]. Shift-Register [JW73, Kar76, KK78, Klu77, LF74, Mar73a, Mar73b, Smi71a]. Shifts [RRNA74]. Ship [SW71b]. Shop [Gon79]. Short [Bak79, Cur79, Fri74]. Short-Circuit [Fri74]. Short-Cut [Cur79]. Short-Word-Length [Bak79]. Shortened [Sun79]. Should [SO73]. SHR [Noe73, NR74]. SHR-Optimal [Noe73, NR74]. Shuffle [Lan76, LS76a, Sto71]. Shuffle-Exchange [Lan76, LS76a]. Sides [Dav77b]. Sieve [Mar78]. Sift [DM78a]. Sift-Out [DM78a]. Sigma [Lyn72]. Sign [Chi77, HL86, SA75, SCNS83]. Sign-Flipping [Chi77]. Sign/Logarithm [HL86, SA75, SCNS83]. Signal [CB72, Cor71b, DRS75, GLMR71, HF71, HA70, Kor79, KSW74, NA72, PM75a, Rej72, SH79, Tru76]. Signal-Dependent [CB72]. Signals [All75b, Kuo74, LC72, Mar78, Ung77d, Usa75, Wil77b, Yue74b]. Signature [HF71]. Signed [Agr76, AR78, BBC72b, Chi77, DS72b, MK71b, Tom75, Tun70b]. Signed-Carry [Chi77]. Signed-Digit [Tun70b]. Signed-Number [DS72b]. Significance [BM77b]. Significant [Ges71]. Silhouettes [SCH72, SN72a]. SIMD [Sie77, Sie79]. Similar [Liu77a]. Similarity [JP73, OS76]. Simple [BJ75, BV78, CDJ71, Gar72b, GBMC72, Gri73, HTD72, IR76, Jol76, Maj72c, MBM77, MC70, PC72, Rhy71, SN79, TB75, Von75]. Simpler [FP75]. Simplex [Sil79]. Simplification [KN72, VW72]. Simplified [Car71b, Cur76, HK74, KS72, LM71b, LS76a, Zis76]. Simplify [AB78, HF74]. Simplifying [Das71b, Pin71]. Simply [Has73, SSK72]. Simulated [LF74]. Simulating [Arm72]. Simulation [Ben70, Bre70b, Bre72b, CCES74, Fre77, Gor75, JM71, KP71, KV76, MCRD72, MC78, Mor76a, RHR76a, RHR76b, RC71c, SW70a, SW70b, Sch74b, She70d, SM70b, Stu75, ST76, Ung77a, Vic74, Wai70]. Simulations [RH73]. Simulative [MM73a]. Simulator [BW77, Gra73]. Simultaneous [Has73]. Simultaneously [KMC72]. Sine [SP77, Bak76]. sine/cosine [Bak76]. Single [Bro71, Dev79, FS74, Gil70a, Har70b, Kau70c, KT72, KR78a, Mar79, MS78, NT78, Ric72, Rot70, SS77, WSB79, WL74]. Single-and [FS74]. Single-Channel [Gil70a, Ric72]. Single-Error-Correcting [Dev79, Kau70c, WL74]. Single-Gate [Rot70]. Single-Parameter [Bro71]. Single-Tape [Har70b]. Single-Threshold [KT72]. Sinusoids [HO79, Sil77]. Size [Len77, Moo71]. Sized [Red71a]. Skip [Lig77]. Skip-and-Set [Lig77]. Skit [O’K73]. Slave [FG76, Wil71]. Small [MBSS75, RV73, SB77b, TS79]. Small-Diameter [TS79]. Smaller [KO78]. Smooth [CL71]. Smoothing [Boo70, Dui76, Gol72, KS78a]. Society [Ano72n, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano72-50, Ano72-57, Ano72-58, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74j, Ano74p, Ano74k, Ano74q, Ano74l, Ano74r, Ano74m, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74n, Ano74y, Ano74o, Ano74z, Ano75o, Ano75p, Ano75q, Ano75i, Ano75r, Ano75j, Ano75s, Ano75k, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75l, Ano75y, Ano75m, Ano75z, Ano75n, Ano76c, Ano76d, Ano76h, Ano76k, Ano76l, Ano76i, Ano76m, Ano76j, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano78y, Ano78z, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78-32, Ano78-33, Ano78v, Ano78-34]. Society [Ano78w, Ano78-35, Ano78x, Ano78-36, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42, Ano79-43, Ano79-44, Ano79-53]. Software [Ano78-39, Ano78-40, Ano78-41, Boe76, Gay78, Ham74, Lom74, Pag72, Ano79-54]. Solid [Stu75]. Solid-State [Stu75]. Solution [Abr75, Bak79, BM70, CF78a, CM70, CW77, DTI78, Du72, Eks74, Fra78, GLR71, GLL79, Kar70, Kli74, KS73, Lou70, Maz71, Men70b, MG71a, MR75, MC70, NK74, Nam70, New75b, SF74, Sch73, Ulr71, WG73, Wit70, Wor76, Wri73]. Solutions [Bro70b, Bro71, Hig71, WK73]. Solvable [Has73]. Solve [RS71b]. Solver [Buz73]. Solving [BB74b, FL77a, Max77, MN70, SG74, Sri76, Tom73, Wex71]. Some [ABK76, Bak75b, Bha77, Bha78, Bre71b, Bri73, BB73, Chi72, Chu74, Cut70, DF78, DH79, Fly72b, FK78c, Gar76, HN78, Hoe76, HI72, Hus72, JW76, KF78, Kin71c, Kri73, Lee75, LS76b, Maj76, Mil73, Nay71, New75b, Pag72, RGK72, RS71a, SC77, SR77, Tou71, Tou72e, Ung77a, Vin78, Wea75, WK71, YM78, Yip76, Yue74a, Bak75a]. Sonic [BdB70]. Sort [NS79]. Sorting [BS78, Lam79a, Mor79, Pre78a, Roc70, SM70b, Van72]. Source [Mar73b]. Sources [Boy74]. Soviet [SG71]. Space [But71, Der73, Fre77, Hec73, Hop71, LSB77, PR78, SIF76, Weg72, Wri73]. Space-Efficient [Weg72]. Space-Filling [But71, Fre77]. Spaceborne [WSB79]. Spaceflight [WKD70]. Spaces [BF78]. Spanning [BF78, KLCC77]. Sparing [BF75b]. Sparse [Gri78a]. Spatial [SM72a]. Speaker [Mak71, Moh71]. Special [Ano76-28, Bec71, CH75, DRS75, DH79, Ehr72, Fle79, GGMM71, HC70, SF74]. Special-Purpose [DRS75, GGMM71, HC70]. Specialized [Swa79b]. Specification [Dub79, Mon74]. Specifications [Ano79-54]. Specified [Bis74, DH79, GD73, Hig73, Kel70, Pfl73, RB75, Sch75, TG75, Wil75]. Spectral [AC70, CL71, Hun71]. Spectroscopy [GW75]. Spectrum [WGS78, Yue74b]. Speech [EFLR76, Mak71, REFN76, Wal76]. Speed [Agr79a, CH73a, Chi78, CH72b, Cio78, Cor71b, Gar74, Hou72, JM70, KES72, KG72, Kel74, Ken77b, Kri70a, Kuc70, LE71, Lin70, Lit74, LM70, Loo70, Mar73c, Qua74, RGK72, SW71a, SW76, Sri71a, Sri71b, Ste72, SKG77, Sto70c, SKK79, Tom75]. Speed-Independent [Cio78, Kel74]. Speeding [BBP75, Pag72]. Speedup [KMC72]. Spherical [RB72]. Spline [Dub79, HA77, Lio76]. Spoken [KM78, Pol71]. SPOOF [Cle73, SS75a]. Spread [Eks73, Wyn71]. Squad [Wak70c]. Square [HB75a, JB76a, MK71a, Maj72a, RGK72, Rud76, SK71]. Square-Rooting [RGK72]. Squares [Agr78a, All75b, Gol72, HA77, Hun73, Maj72a, Wil77b]. Squaring [Che71b]. SR [Brz71]. Stability [GW77, Gil70b, SIF76]. Stabilization [Tab70a]. Stable [AH76, HO79, Sil77]. Stack [Hop70]. Stage [Ast78, VR77]. Standard [GSB^{+}76, Hen78, Guf71]. Standby [BF75b]. Standby-Sparing [BF75b]. Standing [Ano76c, Ano76d]. STAR [AGM^{+}71]. STARAN [Bat77]. Stare [DLL^{+}73]. Start [Eas78]. Stastical [Tru76]. State [AH76, Ben71, BF72, Bis74, Bou74a, Cha72, CL75, CW79c, Cur73a, DH79, Ehr72, Fre74b, FG71c, Gec77, Gil70a, Goo70, GD73, Hob73, HB70, KW70, Kel70, Kin71c, Kra74, Kri73, KR78a, LF74, Lom74, MT70, MT71, MSJ72, MS73a, MB77, Moo71, Mui74, NT78, NT79, Nie70a, Noe73, NR74, NR76b, NK75, Par72a, Pfl73, Pla73, Pra78a, RN77, Ric72, Sau72a, Sau72b, Smi74, SHR72, Stu75, Tan71, TB74, Vai72, Wai70, WR77, Yan74b, Yeh70, de 71]. State-Set [Moo71]. State-Variable [Wai70]. States [Bar71, New70, Per70, Tom72, YT76]. Static [BE70b, BH74b, Cod73, HK77, HK78, Hla78, Zis76]. Static-Hazard-Free [HK77, HK78]. Statistical [BW76a, BCL75, EZS78, Mar73a, Moh71]. Statistically [BT72, Hur74, Hus74]. Status [AR74, Ano77b, Ano77c, RT73a]. Step [BH76, DR75, PB76a]. Stochastic [Ban74a, EGG71, FF75, Gel71, LT70, LF72, LF77, Nie70a, Pag70a, Pag70b, Par72b, Paz70c, Paz71, Pre70, Sil79, Tho75, VN78]. Storage [Anz71, Cho76, GG73, Har76, Hoa76, KLCC77, Mcd72, Nel74, PA78, RC74, SM71b]. Store [HG78b, HPP77, JB76b]. Stored [DEA73, DS74, Dru71a, IC78]. Straight [Ada78, Bar78, BLZ75, Byr71, CY76, DT76, Das78, Fre79a, Ros74b, Sta74, SKK79]. Straight-Line [Ada78, Bar78, DT76, Das78, Sta74]. Strategies [CR72a, FD74a, KR71b, RCG72, WM76]. Stream [KV73, Mil74b, Rub79]. Strings [Wil70]. Strongly [SM78]. Structural [RK71, Yan75b]. Structure [AC71, Bas71b, BR78, Cho74, DS72b, DW73a, Kru71, MN70, NM76a, Weg72, Whi72]. Structured [BS76, BS75, GG73, KD79, Pat76].

Structures [BM77a, Ben77, Ber71, BW76b, Duk74, ET75, Elm76, FE73, Fly72a, Klu77, KC73, Law70, Les71, MAB78, Pat72, Ram73a, Rei74a, RS71a, TJM79, Wan72]. Strum [Pat76]. STT [FG71c, NT79]. Stuck [BCS76, Ios78, KH77c, KR78b, Mei74, Nai79, SS77, Sun79, Wei72]. Stuck-At [Ios78, Mei74, KH77c, Nai79, Sun79]. Stuck-Fault [Wei72]. Stuck-Faults [KR78b]. Stuck-Type [BCS76, SS77]. Students [Fre74a]. Studies [Avi71, Fel79, Nay71, Ros74a, WBRG76]. Study [BCL75, Cha73b, CDL78, DF78, Far75, FL77a, HN73, Ken77a, MM73a, MH70b, Nel74, Red71a, Smi79a, Stu75, WM76]. Style [Gri74]. Stylus [BdB70]. Subcodes [Rao72]. Subdirect [Tur70]. Subfunctions [Reu75]. Subject [Ano71-51]. Submatrices [Ban78]. Suboptimal [Hab77]. Subresultants [Gri78a]. Subroutine [Roc70]. Subroutines [KR71a]. Subscribe [Ano76t, Ano77u, Ano78-42, Ano78-43, Ano79-47, Ano79-48, Ano79-49, Ano79-50]. Subscription [Ano76c, Ano76d]. Subset [NF77]. Subsets [Chi71, MG71b, Rub75b, Tou72a]. Subspace [The75]. substitution [KD79]. substitution-permutation [KD79]. Subsystem [Mal73]. Subtraction [Ban74b]. Successive [JW73]. Sufficient [Smi79d]. Suggested [Hen78]. Suggestion [Bak76, Ste72]. Suitability [GR71b]. Sum [AB78, KO73, Pap79]. Sum-of-Products [AB78]. Summability [GBMC72, Hwa72]. Summation [HK74]. Sums [Cur77, JW73]. Supervised [CF70b]. Support [FSLC78, KS77c, Wor79]. Supporting [De 79]. Surface [Ges71, Mah72]. Surfaces [Gou71, OF73, SF78]. Survey [Age76, Ano77g, Bas73, Gar76, HKD^{+}71]. Survivable [GLW74]. Swap [Kle70]. Swap-Time [Kle70]. Switch [Ast78, IS76, Ogu74, SM73a, SM73b].

Switched [CG74]. Switching

[AA71, Bei70, Bei71, BR76a, BR78, BV79, Bro75b, Bro75c, CM77, CDS73, CHSS73, CYNM79, Das71b, DS71, DK72, DW73a, DM75, DM76, Eps73b, HPP79, Han72, Hen71, HE79, Iba71, Kau70a, Kin70, Kin72, Kje72, KS77b, LH76, MN78, Mat71b, Maz71, MS70, MI72, OW72, OY70, Pap77b, Pra78c, Red72a, RNMP77, Sch75, SD74, SM75c, SM70a, SMW71, Smi71b, Spi71, Su70, SC72c, SS72, Sur75, TD73, TD77, Tha78, Tis71, Tom70, UB74, Ung71, VLS70, VW72, Wes78, YT71a, YT71b, ZKK79]. Symbiosis [Arb76]. Symbol [MR72]. Symbolic [Hol72b]. Symmetric [Bis70, Bis73, Bor72, Dah73, HS75, KKK73, KKK74, LL72, LH76, MM78b, RH75b, SK77a, SK77b, YT71b]. Symmetries [EH78]. Symmetry [DS71, KD73, TV79, YT71a]. Symposium [AG73]. Synchronization [Brz70, DW78, GS79, HA76, SH77]. Synchronized [RV74]. Synchronizer [CM73a, Cha79, FD79, Wor77, Wor79]. Synchronizers [CW75, EW77, Pec76]. Synchronizing [LM70, Loo70, Ung77b]. Synchronous [Ano75-28, Chu70, Cur70, Cur73a, DSBC74, FG76, HSP78, HG79, Hu72, Hwa73, LF74, MF76b, NA72, PM77a, Saw75, Ser75, SHR72]. Syntactic [FP75, FB73b, FF75, JK77, KM78, LF78, TG75, Wal76]. Syntax [CR72d, LF72, LF77, Tho75]. Syntax-Directed [CR72d, Tho75]. Synthesis [AAK74, Bar70, BF72, BBP75, BH73, Bre75, CM74, CMS79, CCB70, CK77, CD73, Col77, CYNM79, Dav70a, Edw75, EH78, Fra72, FG76, GGMM71, Ges70, GG75, GL74, Hu72, IM71, Ish77b, Jum70, KH77a, Kje72, Kol72, LV75, Liu75, Liu77a, Liu77b, Lyo70, Mam77, MS77a, Nak79, Pai73, Pap78, Pap72, SAPR70, SA70, SR72b, SR71, SN71, Sur70a, Sur78, Tho74, UB74, Win70a, Yen79, Yoe70b, Yue76]. Synthesizing [KO78]. System [AA72a, Ake76, AKT75, Anz71, Arn73, AKM^{+}71, Avi71, Bar74b, Bas71b, BCA^{+}75, Bha77, BR76b, BdB70, CM76, Cas73, CR78, CDS73, CHSS73, CK79, CW77, CF70a, Cor77, Den76, Dep70, Dos77, DLL^{+}73, Eps73a, EFLR76, Fer77, FSLC78, FK78b, FT79, GSB^{+}76, GSJ^{+}76, GR72b, Gou77, HW76, HP73, HH77, Joh71, KK79a, KF79, Kim70, Kim79, Kin72, KKK73, KKK74, Koc71, Kor76, KF73, Kuc70, LNL72, LTE77, LE77, LE79, LL79, LH79, Low73, Mak71, Mal72, MMB78, Min75, MF76c, O’K73, Pat76, Per78, PB71, RC71a, RCG72, RC72, RSH79, Rau79a, RF78, REFN76, Rei76, RH74, RK75a, RK75b, SNN70, SG74, Sch78c, She70d, Smi79e, Sri76, Sta70b, SA75, Swa79a, TM70, Toi76, TW77, TK74, TJM79, Von75, WN76, WKD70, WG75]. System [YT77, You70, Cam72, Flo72]. System/360 [Joh71, Cam72]. System/370 [Flo72]. Systematic [Cur70, Den74, Rao72]. Systems [Ano72-39, Avi76, Bae76, BB72, Ban74b, Bar75a, BM73, BM74, BGM76, BM78, Bas73, BS74, Bea78, Ber70a, BS75, Bha75b, Boo70, BF75b, BE70b, BW77, Bre71a, Bre73b, BP72, CF78a, Cha74c, CM70, CK75, CS79, CLL78, Coy79, DW78, De 79, Dee72, DF70, Dot70, Eks74, Eks73, FD74a, FS74, Fre74b, FB74, FM71, Fri73, FB73b, Fu74, FK78a, FK78c, Gay78, Gel73a, Gor75, GH72, HA74, Hay76a, Hei79, Hil79, Jos74, JA78, KG72, KO73, KF79, Ker73, Kim74, Kle70, Koh75, KP71, KS79, Kuo74, Kur74, Lan77, LT73, LJ79, Lud74, MH76, MH78, MF76a, MM78a, Mat71c, McL72, MK79, Mil74a, Mor76a, Nay71, NN73, O’K75, Par71, Paz71, Pea75, Pu72, RH75a, Rav72, Red78a, Reg73, SAPR70, Saw77, SM71b, SD79, SM73b, Sie74, Sri79]. Systems [Sri71a, Sri71b, Stu75, Su70, SSR72, Tab74, TM71, TI72b, Tis71, TW76, UB74, VM78, Var79, Vic74, VLS70, Wai70, Wak75c, YL73, Zoh76b]. Systems-Switching [Tis71].

T [GB71b]. T-Partitionable [GB71b]. T. [Wor79]. Table [Cur77, Das71b, DK72, Dia76, MS77b, SD74]. Table-Lookup [MS77b]. Tables [AM79c, Cur71a, DW72a, HS70, VVA78]. Tac [JW76]. Tagged [Feu73]. Tan [Cur73a]. Tan-Like [Cur73a]. Tandem [But78b]. Tangent [Pre78b]. TANT [LD72a, Fra72, Koh71, Lee78b, Vin78]. Tape [Arn70, Har70b, Kam70a, PB76b]. Tapered [Mor71]. Task [GR72b, WN76]. Task-Free [WN76]. Tasks [Koh75, RPL73, RFL76, RL78, RSH79, RV73]. Tausworthe [NM76a]. Teacher [MS76]. Teaching [WBRG76]. Technical [Ano71-52]. Technique [Agr78a, AC70, Bor72, Bre71d, DV78a, GL70, GH72, GP74, Hun72, IR76, JR76, Jol76, KS77a, Koc75, Koh71, KF72b, KF72c, Lan78, LD72a, Maj72c, MJTW75, MBM77, PF72, PH73, Red71b, RNMP77, RB72, SH77, Sic76, Skl78, Sto74a, Sur75, TM76, Toh74, TB74, WNR74]. Techniques [Aar70, CDJ73, CHSS73, Cha74c, CM70, DM78b, FM72a, Gil74a, HKD^{+}71, Kim74, Kri70b, LT70, Lil72, LM78, Mil77, Moh71, MG71b, New75b, PR72, PR76, Pra72, PD77, RPL73, RC70, SF71, Sau72b, STG71, Sie77, Sto70d, SNEL79, Tou72a, Wai70, Weh74, YK76]. Technological [Rem72]. Technology [Ano77b, Ano77c, Fre74a, FLBK76, HHJ^{+}76, Jos72b]. Telecommunications [Cox72]. Teleprocessing [CR72b, IA76, TWB78]. Television [DWT70]. Template [Oza78, VR77]. Temporary [Sri71b]. Ten [Sit74]. Tensor [WS76]. Terminal [Arm73, Har70a, HI72, KR78b, PRD71, SAFV71, Wei72]. Terminals [Hob73, Pat73, TWB78]. Termination [Hor72]. Ternary [AA72b, Bar77, BM77a, BY79, CS70, EI77, FL75, Gra73, HK77, HK78, MJ77, MJ78, SAPR70, SA70, SA74]. Tesselated [MF76d]. Tessellation [SC74, Wag70]. Test [AM79a, AM79b, AA72a, AA75, Ake73, Ake76, Bar70, Bet71, Bre71d, Bre74, BH74b, CSJ74, CDO78, Dan73, DDSM76, HF74, HH77, Hla78, How75, Hus74, Mar71, MST^{+}75, Mut76, OH79, Pap77a, Pra78d, Red73a, RS75, SR75, SLC75, SV74, TV79, Wan75a, Wei72, WA73, Wul74, YT71c]. Test-Input [DDSM76, Mar71, MST^{+}75]. Test-Pattern [SLC75]. Testability [DR74, RSL76, WA73]. Testable [Coy79, Fri73, FNSK75, Kod74, Red72b, RW74, SR74a, SR74b, Sun76]. Tested [RH75b]. Testing [AA76, Agr78b, BK76, Bre73c, CDJ71, GBMC72, GCC74, Hay76c, Kau74b, KH77b, KH77c, LD74, Los78, NTA78, Nai79, PR71, RC72, RS74, Red77b, RS78, Rub75a, SB77a, SB79, Sch79a, Si78, Smi79b, SC74, Sur74, TV76, Yue77e, AGM^{+}71]. Tests [BC71b, Bre72a, FD74b, FD77b, Hay71b, Hay78a, KS71, MSJ72, Red77a, Rob70b, Toi76, Wan75b]. Text [AKM^{+}71, AN74]. Textural [Tho77]. Texture [DM78b, MMB77, RJ72, Ros75, SH72, Tan78, ZRD75]. Texture-Based [DM78b]. Textured [Ehr77, Nev76, Sha78b]. th [AM79c]. Their [Ale71, AP76a, BR76a, BF72, Duk74, Fen74, Fly72b, Har70a, Hay74, HC73, Hwa72, Jos72a, KN70, KMC72, LJ79, Meo75, SK79a, Tha78, The75, Wak74, Wan75c, YT70, FD75, GL74, KT72]. Theorem [Ern76, Fis76, MOW76, Mor76b, Pra71a, Rei76, TL73a, Vil73a, Yue75d]. Theorem-Proving [MOW76]. Theorems [FK78c]. Theoretic [KNF76, Law70, Nus78]. Theoretical [Che71a, CW75, Mil73, Sha78a, Tou72c, Zah71]. Theories [de 70]. Theory [AGM^{+}71, BGM76, BS74, Bei70, Bei71, BA71, Cha74c, CM79d, Fu74, Gar78, Hou72, KS78a, Kar74, Kel74, Lei74, Lew74, LS76b, Moo76, Pra78c, Rei74b, RS71a, Roo73, SP72b, Sto70a, TD73, Tha78, UB74, YK76, Yoe70a]. There [CS70]. Thermoplastic [BL75]. Thought [Mic74]. Three [AA72b, Ano72-39, Ast78, BM77a, BCSM72, Bre72b, CCB70, CC73, Gec77, Koc71, Mor76a, O’K75, Par71, RH75b, Sar71, SF77, UM77, Wei75]. Three-Dimensional [Ano72-39, CC73, Par71, Sar71, UM77, Wei75]. Three-Face [SF77]. Three-Failure-Tolerant [Koc71]. Three-Level [CCB70, RH75b]. Three-Stage [Ast78]. Three-Valued [Bre72b, Mor76a]. Threshold [AA72b, Ama72, Bar70, BW76a, Bre71b, CC72, CM79b, DS72a, Edw75, FD73, GW77, GBMC72, GC73, Ham73, Hen71, HC73, Ish77a, Ken77b, KT72, Lyo70, Mor70a, MTB70, OS76, OY70, SA70, She70b, SR72b, Sli70, Tak76, Wea75, WNR74, Win70a, Win70b, Yue76]. Threshold-Element [Tak76]. Threshold-Logic [Lyo70]. Throughput [BCHL73, Sch79b]. Tic [JW76]. Tic-Tac-Toe [JW76]. Time [AK72, Arn70, AO79, BCKT79, CK75, CK77, Dep70, Der73, DF70, Dot70, Dow73, FB73a, FS74, Gay78, Gel73a, GSJ^{+}76, Goo70, Goy76, HO79, Har70b, HA70, Hun72, Kle70, Koc75, KR78a, Lev70, Mar78, NT78, Nec79, Nie70b, PRD71, Pea75, Pol71, Saw77, Sch79b, Sil77, Spi71, WKD70]. Time-Component [AO79]. Time-Invariant [Pea75]. Time-Shared [Kle70]. Times [Pec76, Spi73]. Timing [Nie70b]. Toe [JW76]. Toeplitz [Eks74]. Toggle [AM79c]. Toggle-Registers [AM79c]. Tolerance [BO70, Ogu74]. Tolerant [AGM^{+}71, Avi76, BCJ^{+}71, Car73, CM76, FM72a, GM76, GCP^{+}77, Hay76a, Hec73, Hop71, HS75, Ios75, Kim75, Koc71, LR72, MS74, Mat71c, Mey71, MR76, OW73b, PR73, PR74b, Pra74, Pra78b, Ram71, Red78b, RC70, SM75a, Sch74a, SG71, Smi75, Sti78b, TA71]. Tombs [Oza78]. Tone [JR76]. Tool [BW77]. TOPI [GGMM71]. Topoglyphs [Gol78]. Topological [Maz71, SV75, TG72b, UB74]. Total [DS71]. Totally [AM73, AR77, Bis70, Bis73, Bor72, Dav78, Usa75, YT71b]. Tours [DS73]. Trace [Ram77a]. Traced [DT70, TD70]. Traces [BBP75]. Tracey [HB70]. Tracing [Weg72]. Track [PB76b]. Tracking [SP72a, SSR72]. Tradeoff [Rus72]. Training [DS72a, LK75]. Transactions [Ano70a, Ano70-29, Ano70-30, Ano70-31, Ano70-32, Ano70-33, Ano70-37, Ano70-34, Ano70-35, Ano70-36, Ano71a, Ano71z, Ano71-27, Ano71-28, Ano71-29, Ano71-30, Ano71-31, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano72a, Ano72-59, Ano72-60, Ano72-61, Ano72-62, Ano72-63, Ano72-64, Ano73a, Ano73m, Ano73n, Ano74a, Ano75a, Ano76b, Ano77a, Ano78h, Ano78i, Ano78j, Ano78b, Ano78-37, Ano78-38, Ano78-39, Ano78-40, Ano78-41, Ano79k, Ano79l, Ano79a, Ano73e, Ano78h, Ano78i, Ano78j, Ano79k, Ano79l]. Transcoder [DG72]. Transfer [ASS72, Bar75a, BEGW72, BT70b, HP75, Low73, Sto70c]. Transferable [Roc70]. Transference [Ull72]. Transfers [Max77, Mcd72]. Transfom [ANR74]. Transform [ARS71b, Cap76, Car71a, CS73, CF79, Des74, Des79, DV78b, Edw75, FA76, GC77, Hen70, Jai77a, Kun75, Lec70a, Lec70b, LD72b, RNR75, RT79, Ree70, Rej72, RRNA74, Sha75, Sha78b, Ulm70, Yip76, Yue75a, Yue76, Yue77e, Ber72, Cai71, Cor71a, Cor71b, CSY75, CL72, Dru71a, Gla70, GW70, Har76, LH75, MS73b, PB76a, SR77, TM78a, WK71]. Transformation [AK72, CL79, Cor77, KF72a, Kri70b, Kri71b, LH76, ND73, Pap76, TM78b, Tur70, VW72, Weh74, YT71b, Zoh73b, Zoh74]. Transformations [Ami77, BD71, DN77, NSR74, PC72, Rat76, Sta70a, WGS78]. Transforms [Bro77, CH72a, CH73b, Dun75, FHTS78, Kun79, Nus78, Pre71b, Rad72, Rad73, RTKH77, Sar78, VR79, Yue72, Goo71, GW75, Hun71, Kar77b]. Transient [Wak75c]. Transistor [HHJ^{+}76].

Transition [BH73, Hay76c, Hay78a, KR78a, NT78, Red77b, Reu72, SR73, Sho74]. Transition-Sensitive [BH73]. Transition-Time [KR78a]. Transitive [Tsi70]. Translation [BB72, IC78, Lew74, Tho75]. Translators [BCH^{+}73, CDJ73]. Transmission [And70, Hor71b, Hor72, Mah74, MMO74, SW70a, SW70b]. Transparent [Sun79]. Transportation [Mah74]. Transposing [All75a, Ari79, DS74, Ekl72, Ram75, Sch73, Van77]. Transposition [TE76]. Trapezoidal [Max77]. Traversing [Klu77]. Treatment [FA76, Hol72b, PM75b, VR79]. Tree [AM79a, AA76, Arb70, BS76, Ber70a, Ber71, CK74, CR72b, Cha70, CE75, FB73b, Gri76, IA76, Ish77b, KH77a, Klu77, KB75, LF78, MH75, MF76c, Pat72, SK77c, Ung77c]. Tree-Searching [Gri76]. Tree-Type [CE75, KH77a]. Trees [BF78, CMS79, Cha74d, Cur79, EF76, GST79, KLCC77, Li77, LS76b, LP76, LP78, MM73b, PM77b, TG78, WC74]. Trends [Hol76, JM74, New76]. Triangular [Fre79a]. Triangulation [LSB77]. Tridiagonal [SK77a, SK77b]. Trigger [HN73]. Triple [AS74, Wak75c]. Truncation [Zoh76b]. Trunk [AKT^{+}77]. Trunk-Division [AKT^{+}77]. Truth [Dia76, Har70a, Pin71]. Truth-Table [Dia76]. Tumors [BS76]. Tuple [Rat76]. Turing [Har70b, Lyn72]. Tutorial [Ano78v, Ano78w, Ano78x, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Hen76a]. Twelve [Car71a, WK71]. Twiddling [Yue73a]. Two [AO79, Bar70, Bar71, BW73, Bla74, BR74, CC72, CC73, CW79c, DM78b, EZS78, ET75, FO71c, GG75, Goo71, Gra71, HND75, Han75, Har70a, HO79, Hwa79, Ios78, Jai77c, JM71, Kri71a, Kro74, LSB77, Mar73b, MG75, MK79, MS78, Moh71, Moo71, Nak79, NA74, OY70, PCM70, Pav75b, PR74b, RR74, Roe78, SR74a, SK78b, SM72a, SM71b, Sil77, Sta74, Tak76, Tom74, VR77, Vic71, Vil73a, Wri73, Yan75a, YRT75, NA76b]. Two-Class [BR74, PCM70]. Two-Dimensional [DM78b, FO71c, HND75, Jai77c, MS78, NA74, Roe78, SR74a, SM72a, Vil73a, NA76b]. Two-Input [SK78b]. Two-Layer [Tak76]. Two-Level [CC72, EZS78, ET75, Ios78, MK79, Nak79, OY70, PR74b, RR74]. Two-Multiplies-per-Cycle [HO79, Sil77]. Two-Point [Vic71]. Two-Space [LSB77, Wri73]. Two-Stage [VR77]. Two-State [CW79c]. Two-Terminal [Har70a]. Two-Way [Bar71, Kri71a, Moo71]. Type [AH76, BCS76, CE75, Cur73b, KH77a, Orc76, PHG77, Pre72, Sik70, SS77]. Types [KF78, LHSM74, Tak76]. Typicality [WL75].

ULM [Voi77b]. Ultimate [Brz70]. Ultrareliable [Mat71c]. Ultrasonic [GW75]. Unate [BPC75, Bau72, Bet71, DSBC74, Mat71b, SK79b, Skl70, TD77]. Unateness [Bar70, Fis74, Pra78a]. Uncertainty [Key77]. Understanding [Dav77b, EFLR76, REFN76, Wal76]. Undetectability [Smi79d]. Unformatted [AKM^{+}71]. Unger [Pra71a]. Unidirectional [Wak75b]. Unified [CM77, DAA79, DB73, FA76, Gel73b, KS77c, LD75, UB74, VR79]. Uniform [CF78b, Dav74, GR71a, Hal74, Hou72, Jum70, Obe72, Wil75]. Unifying [Reu75]. Unimodal [GL70]. Uniprocessor [RFL76, RL78]. Uniqueness [Nie70a]. Unit [Bat70, BT76, Dev79, Jol76, Kau70c, KS77c, Mor70a, SS71, Sri78, Wea75]. Unit-Distance [Dev79, Kau70c]. Unitary [WGS78]. Units [Atk70, GHJ77, Meo75, RS75, RS78, YC76]. Universal [Ake73, Ale71, Alm78, AT70, BB73, But78b, Cha70, CE75, FG71c, GT74, Hua71, Ios75, KH77a, Kan75, Kel74, Kli71, LV75, NT78, NT79, NA72, NK75, OW72, Pat73, Pra78d, Pre71a, Pre72, Smi79e, SB74b, Sur78, TG78, Voi77a, Voi77b, YT70]. Universality [SK79a]. University [AG73]. Unrestricted [MS75a, Ung71]. Unscramble [Swa74a]. Unscrambling [PB76a]. Unspecified [Dav70a]. Unstable [Vic71]. Unsupervised [LD75, MS75b]. Update [GS79]. Upper [Fre79b, Tou71, Vil70, Yue72]. Use [Agr78b, Atk70, Aus70, Ban74c, BEGW72, Ber77, BK71b, Cle73, DT70, EF79, EH76, FB74, Ges71, Hla70, Hun71, Jos72b, Kar70, Mal73, PS71, Sch79a, Sha78a, Tha78, Tri77b, Tri78, TB75, Wak74, RT73a]. Used [Bec71, Len78, O’K73]. User [Har71, Tho74]. Uses [AP76a, Jos72a]. Using [ADMR78, Alm78, ACJ75, Ano79b, BC71a, BPC75, Bar77, Bar70, BT72, BCL75, BH73, Bre75, BB75, Cha72, CD73, CD78, Cur70, Cur71b, DDSM76, DM78b, Dep70, Des74, Elm76, FM72a, Gar72a, GY77, Gel76, Gri73, Gri78a, HLD70, HA77, Hun72, HTD72, Iba71, IC78, JP73, JB76a, Jul78, KF79, KM78, KC71, Kit78, Kon71, KF72a, Kri77, LT70, Lan76, Lin70, Mam77, MS73b, Mar71, Mat71b, Meo75, MST^{+}75, MBSS75, MS78, Mor75, NT70, Oja79, PA78, PM75a, Pra78b, RM71, RGK72, RT70, RE71, RH74, Sau72b, SH77, SM72a, Ser73, SV74, SY75, SR71, SP72a, Tho76, Tjo77, TOS71, Toh74, TJM79, TB74, Tun70b, VD76b, VV74, VVA78, Wak75b, Wil78, WS76, Yen79, You70, BF75a, Rao72]. Utilities [BH76]. Utilization [Cha70, DB74, FGR71, SN79]. Utilizing [Das71b, FK74, SNN70].

V [Tab70b, CDO78]. Valley [KF72c]. Valley-Seeking [KF72c]. Value [Kar70, PR78, Vic71, WK73]. Valued [Bre72b, CM79b, DVS74, DV78a, GLR71, ISN76, KH77a, KT72, Mor76a, Mut76, SV74, Sin72, Tou72d, TG79, VLS70, Vra77]. Values [Wan75b]. Variable [Byr71, CYNM79, Ehr76, Sti78b, Sto74a, Tom74, Wai70]. Variable-Mode [Byr71]. Variables [AA72b, Ano75-28, BCSM72, Bor72, Dah73, MTB70, Mur70, Pav75b, Vin78, Wri73]. Variables-Weighted [Dah73]. Variational [RH71]. Variations [Hoe76]. Various [Bre73b, IS76, MM73a, Sie79]. Varistructured [Lip77]. Vector [Ath75, Eks74, FV79, Hig76, Kit77, Lee76]. Vectorcardiograms [KCLB71]. Vectors [BP72, FS75, Glu71, Kit78, SO79b, Swa74a]. Vehicles [Hop71]. Velocity [Dav72b, SF72]. Verification [BT72, Dia76, DW73a, EH76, HP75, Rot77, ST76]. Versatile [MH70b, TP72, YT77]. Version [WN76]. Versus [Ham71, LR71a, LR72, Pea76]. Vertex [Hwa72]. Vertices [Hwa72, SF77]. Very [BHK79, Des79, RV73]. via [GW75, Rad72, Rad73, RA78, WA73]. Video [GSJ^{+}76, KF73, MCRD72]. Viewed [BG75]. Viewpoint [Car73, FR71]. Views [MA77, UC75]. Visible [Mah72]. Vision [Per78, YT77]. Visual [RT71, UC75]. VLSI [PW79b]. Vol [Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano76b, Ano77a, Ano78b, Ano79a]. Voltage [TS75]. Volume [Cha74c, Lew74]. Voted [Sie75]. Voting [HP73].

W [Tab70b]. Walsh [Car71b, Hen70, KS72, Lec70a, Van73b, Yue76, AK72, Bro77, But72, But78a, CS73, Dav72a, Edw75, FA76, FB74, GC77, Kun79, LM71b, Red71b, RRNA74, Ric74b, TG72a, Yue72, Yue74b, Yue75a, Yue75e, Yue75d, Yue77e]. Washington [Ano74b, Ano74c, Ano74d, Ano74e, Ano74f]. Waveform [PF72, Pav73]. Waveforms [CL71, Dav71, EF76, Tjo77]. Way [Bar71, Du72, Har76, Kri71a, Moo71]. Week [Ano79m, Ano79n, Ano79o, Ano79p, Ano79q]. Weight [Hwa72, TL73b, WL74]. Weight-Preserved [WL74]. Weighted [AA77, Bro75c, Dah73, SLC75, VV74]. Weighting [Bub72, NRS74]. Weights [Mal73]. which [JB76b]. Whirl [Paz71]. White [Jai77a, Mor74, Sok72, Yue74b]. Wide [GSJ^{+}76]. Wide-Band [GSJ^{+}76]. Width [BK79, JB76b]. Wiener [FB74, Hab77, Pra72, PD77]. Window [Cha73b]. Winograd [Wak70b]. Wire [Abe72, AKT^{+}77]. Wire-Routing [AKT^{+}77]. Wiring [Kau74b]. with-Binary [EI77]. Within [WGS78]. Without [BKM73, CF78b, Hla70, MS76, RK75b]. Word [Bak79, CBS71, Far79, RE71, Sch78c]. Word-Level [Far79]. Words [JB76b, KM78, OTK76, Pol71, Rom78]. Work [ES77, Hel72, Hel74, Mas78, Wel74]. Working [GF78, Smi76]. Worst [Bri73, Rub75b, Yue74a]. Writable [LM78].

X [Wei75]. X-Ray [Wei75].

Years [Ano75b, Weg76]. Yield [HB75a].

Zadeh [Tab70b]. Zero [Geo70, NS75]. Zero-Crossing [NS75]. Zero-One [Geo70]. Zoom [Yip76].

References

Ali:1971:RNS

[AA71] M. Ali and R. Ahmed. Realization of nonlinearly separable switching functions. IEEE Transactions on Computers, C-20(6):695–698, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                                 tronic). URL http://ieeexplore.
                                                 ieee.org/stamp/stamp.jsp?tp=&
                                                 arnumber=1671923.

Agrawal:1972:ATG

[AA72a] Vishwani D. Agrawal and Prathima Agrawal. An automatic test generation system for Illiac IV logic boards. IEEE Transactions on Computers, C-21(9):1015– 1017, September 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009081.

Aibara:1972:ETT

[AA72b] Tsunehiro Aibara and Michihiro Akagi. Enumeration of ternary threshold functions of three variables. IEEE Transactions on Computers, C-21(4):402–407, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008986.

Agrawal:1975:PAR

[AA75] P. Agrawal and V. D. Agrawal.

Probabilistic analysis of random test generation method for irredundant combinational logic networks. IEEE Transactions on Computers, C-24(7):691–695, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672882.

Agrawal:1976:MCT

[AA76] P. Agrawal and V. D. Agrawal. On Monte Carlo testing of logic tree networks. IEEE Transactions on Computers, C-25(6):664–667, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674670.

Al-Alaoui:1977:NWG

[AA77] M. A. Al-Alaoui. A new weighted generalized inverse algorithm for pattern recognition. IEEE Transactions on Computers, C-26(10): 1009–1017, October 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674736.

Abd-Alla:1974:HSM

[AAK74] A. M. Abd-Alla and D. C. Karlgaard. Heuristic synthesis of microprogrammed computer architecture. IEEE Transactions on Computers, C-23(8):802–807, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672629. See comments

[Col77].

Aaron:1970:BAD

[Aar70] M. R. Aaron. B70-3 analog-to-digital and digital-to-analog conversion techniques. IEEE Transactions on Computers, C-19(8): 762–763, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671622.

Agin:1976:CDC

[AB76] Gerald J. Agin and Thomas O. Binford. Computer description of curved objects. IEEE Transactions on Computers, C-25(4):439–449, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674626.

Arevalo:1978:MSB

[AB78] Z. Arevalo and J. G. Bredeson.

A method to simplify a Boolean function into a near minimal sum-of-products for programmable logic arrays. IEEE Transactions on Computers, C-27(11): 1028–1039, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674992.

Abramovici:1979:RFD

[AB79a] M. Abramovici and M. A. Breuer.

On redundancy and fault detection in sequential circuits. IEEE Transactions on Computers, C-28(11):864–865, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675267.

Agrawal:1979:EDR

[AB79b] P. Agrawal and M. A. Breuer. Experiments with a density router for PC cards. IEEE Transactions on Computers, C-28(3): 262–267, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675333.

Abdali:1971:PSM

[Abd71] S. K. Abdali. On proving sequential machine designs. IEEE Transactions on Computers, C-20 (12):1563–1566, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671763.

Abel:1972:OCA

[Abe72] L. C. Abel. On the ordering of connections for automatic wire routing. IEEE Transactions on Computers, C-21(11):1227–1233, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672075.

Aven:1976:SRD

[ABK76] O. I. Aven, L. B. Boguslavsky, and Y. A. Kogan. Some results on distribution-free analysis of paging algorithms. IEEE Transactions on Computers, C-25(7):737–745, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674682.

Abraham:1975:CSR

[Abr75] J. A. Abraham. A combinatorial solution to the reliability of interwoven redundant logic networks. IEEE Transactions on Computers, C-24(5):578–584, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672858.

Andrews:1970:GTS

[AC70] Harry C. Andrews and Kenneth L. Caspari. A generalized technique for spectral analysis. IEEE Transactions on Computers, C-19(1):16–25, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008895.

Andrews:1971:DFM

[AC71] H. C. Andrews and K. L. Caspari. Degrees of freedom and modular structure in matrix multiplication. IEEE Transactions on Computers, C-20(2):133–141, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671795.

Andrus:1975:DIR

[ACJ75] J. F. Andrus, C. W. Campbell, and R. R. Jayroe. Digital image registration method using boundary maps. IEEE Transactions on Computers, C-24(9):935–940, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672934.

Aggarwal:1975:CAM

[AD75] J. K. Aggarwal and R. O. Duda.

Computer analysis of moving polygonal images. IEEE Transactions on Computers, C-24(10):

966–976, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672695.

Adams:1978:CIM

[Ada78] E. N. Adams. Comments on “The Identification of Maximal Parallelism in Straight-Line Microprograms”. IEEE Transactions on Computers, C-27(3):287, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675091. See [DT76,

Das78, Bar78].

Ausherman:1972:ECE

[ADL72] D. A. Ausherman, S. J. Dwyer, and G. S. Lodwick. Extraction of connected edges from knee radiographs. IEEE Transactions on Computers, C-21(7):753–758, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672171.

Ahuja:1978:PAP

[ADMR78] N. Ahuja, L. S. Davis, D. L. Milgram, and A. Rosenfeld. Piecewise approximation of pictures using maximal neighborhoods. IEEE Transactions on Computers, C-27(4):375–379, April 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675110.

Atkins:1973:PCA

[AG73] David E. Atkins, III and Harvey L. Garner. Preface: Computer arithmetic: An introduction and overview [Second IEEE Symposium on Computer Arithmetic, May 15–16, 1972, University of Maryland]. IEEE Transactions on Computers, C-22(6):549–551, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009104;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Atkins.
pdf.

Agerwala:1976:MOS

[Age76] T. Agerwala. Microprogram optimization: a survey. IEEE Transactions on Computers, C-25(10): 962–973, October 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674537.

Avizienis:1971:SST

[AGM^{+}71] A. Avizienis, G. C. Gilley, F. P.

Mathur, D. A. Rennels, J. A. Rohr, and D. K. Rubin. The STAR (Self-Testing And Repairing) computer: An investigation of the theory and practice of fault-tolerant computer design. IEEE Transactions on Computers, C-20(11):1312–1321, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671726.

Agrawal:1975:AAN

[Agr75a] D. P. Agrawal. Arithmetic algorithms in a negative base. IEEE Transactions on Computers, C-24(10):998–1000, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672702.

Agrawala:1975:GLP

[Agr75b] A. K. Agrawala. On generating a line “parallel” to a digital line — a sequential approach. IEEE Transactions on Computers, C-24(10):983–988, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672698.

Agrawal:1976:SMR

[Agr76] D. P. Agrawal. Signed modified reflected binary code. IEEE Transactions on Computers, C-25(5):549–552, May 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674648.

Agrawal:1977:CNB

[Agr77] D. P. Agrawal. Comments on “A Note on Base-2 Arithmetic Logic”. IEEE Transactions on Computers, C-26(5):511, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674869. See [Yue75b].

Agrawal:1978:NTC

[Agr78a] D. P. Agrawal. A novel technique for computing negabinary squares. IEEE Transactions on Computers, C-27(3):266–270, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675084.

Agrawal:1978:WUR

[Agr78b] V. D. Agrawal. When to use random testing. IEEE Transactions on Computers, C-27(11): 1054–1055, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674994. See comment

[Sch79a].

Agrawal:1979:HSA

[Agr79a] D. P. Agrawal. High-speed arithmetic arrays. IEEE Transactions on Computers, C-28(3): 215–224, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675322.

Agrawal:1979:AR

[Agr79b] V. D. Agrawal. Author’s reply. IEEE Transactions on Computers, C-28(8):581, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675416.

Agrawal:1979:CAH

[Agr79c] V. D. Agrawal. Comments on “An Approach to Highly Integrated Computer-Maintained Cellular Arrays”. IEEE Transactions on Computers, C-28(9):691–693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675440. See [Man77].

Acha:1976:INS

[AH76] J. I. Acha and J. L. Huertas. On input and next-state equations of the R–S type Mstable. IEEE Transactions on Computers, C-25(7):759–763, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674685.

Appelbe:1978:SHM

[AI78] W. F. Appelbe and M. R. Ito. Scheduling heuristics in a multiprogramming environment. IEEE Transactions on Computers, C-27(7):628–637, July 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675163.

Alexandridis:1972:RTW

[AK72] Nikitas A. Alexandridis and Allen Klinger. Real-time Walsh–Hadamard transformation. IEEE Transactions on Computers, C-21(3): 288–292, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008951.

Andrews:1975:HCM

[AK75] M. Andrews and G. A. Korn. Hybrid computer methods for direct functional optimization. IEEE Transactions on Computers, C-24 (10):958–965, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672694.

Akers:1970:RMC

[Ake70] S. B. Akers, Jr. R70-40 module clustering to minimize delay in digital networks. IEEE Transactions on Computers, C-19(10): 987–988, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671406.

Akers:1972:RLA

[Ake72] Sheldon B. Akers. A rectangular logic array. IEEE Transactions on Computers, C-21(8):848–857, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009040.

Akers:1973:UTS

[Ake73] Sheldon B. Akers. Universal test sets for logic networks. IEEE Transactions on Computers, C-22(9):835–839, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009174.

Akers:1974:FDG

[Ake74] S. B. Akers, Jr. Fault diagnosis as a graph coloring problem. IEEE Transactions on Computers, C-23(7):706–713, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672611.

Akers:1976:LSF

[Ake76] S. B. Akers, Jr. A logic system for fault test generation. IEEE Transactions on Computers, C-25(6):620–630, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674662.

Akers:1978:BDD

[Ake78] S. B. Akers. Binary decision diagrams. IEEE Transactions on Computers, C-27(6):509–516, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675141.

Ascher:1971:ISR

[AKM^{+}71] R. N. Ascher, G. M. Koppelman, M. J. Miller, G. Nagy, and G. L. Shelton, Jr. An interactive system for reading unformatted printed text. IEEE Transactions on Computers, C-20 (12):1527–1543, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671759.

Allan:1975:ADA

[AKT75] F. J. Allan, T. Kameda, and S. Toida. An approach to the diagnosability analysis of a system. IEEE Transactions on Computers, C-24(10):1040–1042, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672718.

Asano:1977:WRS

[AKT^{+}77] T. Asano, T. Kitahashi, K. Tanaka, H. Horino, and N. Amano. A wirerouting scheme based on trunkdivision methods. IEEE Transactions on Computers, C-26(8): 764–772, August 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674914.

Arcelli:1971:PPO

[AL71] C. Arcelli and S. Levialdi. Picture processing and overlapping blobs. IEEE Transactions on Computers, C-20(9):1111–1115, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672007.

Aleksander:1971:CUL

[Ale71] I. Aleksander. Comments on “Universal Logic Modules and Their Applications”. IEEE Transactions on Computers, C-20(5):586–587, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671892. See [YT70].

Alltop:1975:CAT

[All75a] W. O. Alltop. A computer algorithm for transposing nonsquare matrices. IEEE Transactions on Computers, C-24(10): 1038–1040, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672717. See [Ekl72] and

comments [Van77].

Allum:1975:LMS

[All75b] J. H. J. Allum. A least mean squares cubic algorithm for online differential of sampled analog signals. IEEE Transactions on Computers, C-24(6):585–590, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672863. See comments

[Wil77b].

Allum:1977:AR

[All77] J. H. J. Allum. Author’s reply. IEEE Transactions on Computers, C-26(4):416, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674853.

Almaini:1978:SMI

[Alm78] A. E. A. Almaini. Sequential machine implementations using universal logic modules. IEEE Transactions on Computers, C-27(10): 951–960, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674975.

Anderson:1973:DTS

[AM73] D. A. Anderson and G. Metze.

Design of totally self-checking check circuits for m-out-of-n codes. IEEE Transactions on Computers, C-22(3):263–269, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672298.

Agarwal:1977:ROF

[AM77] V. K. Agarwal and G. M. Masson. Resolution-oriented fault interrelationships in combinational logic networks. IEEE Transactions on Computers, C-26(11): 1170–1175, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674771.

Agarwal:1979:FFA

[AM79a] V. K. Agarwal and G. M. Masson. A functional form approach to test set coverage in tree networks. IEEE Transactions on Computers, C-28(1):50–52, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675221.

Agarwal:1979:RCP

[AM79b] V. K. Agarwal and G. M. Masson. Recursive coverage projection of test sets. IEEE Transactions on Computers, C-28(11): 865–870, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675268.

Arvillias:1979:TRG

[AM79c] A. C. Arvillias and D. G. Maritsas. Toggle-registers generating in parallel k kth decimations of msequences x^{P} + x^{k} + 1 design tables. IEEE Transactions on Computers, C-28(2):89–101, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675298.

Amari:1972:LPP

[Ama72] S.-I. Amari. Learning patterns and pattern sequences by self-organizing nets of threshold elements. IEEE Transactions on Computers, C-21 (11):1197–1206, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672070.

Amin:1977:AGL

[Ami77] A. T. Amin. An algorithm for grey-level transformations in digitized images. IEEE Transactions on Computers, C-26(11): 1158–1161, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674767.

Ascher:1974:MAH

[AN74] R. N. Ascher and G. Nagy. A means for achieving a high degree of compaction on scan-digitized printed text. IEEE Transactions on Computers, C-23(11): 1174–1179, November 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672419.

Andrews:1970:BTI

[And70] H. C. Andrews. B70-2 transmission of information by orthogonal functions. IEEE Transactions on Computers, C-19(7):656, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671596.

Andrews:1971:MRF

[And71] H. C. Andrews. Multidimensional rotations in feature selection. IEEE Transactions on Computers, C-20(9):1045–1051, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671993.

Anderson:1972:BCG

[And72] T. F. Anderson. B72-2 A compiler generator. IEEE Transactions on Computers, C-21(1):109, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672036.

Anderson:1974:MMR

[And74] G. A. Anderson. Multiple match resolvers: a new design method. IEEE Transactions on Computers, C-23(12):1317–1320, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672449.

Andrews:1976:RM

[And76] P. B. Andrews. Refutations by matings. IEEE Transactions on Computers, C-25(8):801–807, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674698.

Anonymous:1970:IIT

[Ano70a] Anonymous. 1970 index IEEE Transactions on Computers vol. C-19. IEEE Transactions on Computers, C-19(12):1297–b–1297–m, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671463.

Anonymous:1970:ACCa

[Ano70b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (1):85–94, January 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008907.

Anonymous:1970:ACCb

[Ano70c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (2):177–195, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671485.

Anonymous:1970:ACCc

[Ano70d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (3):273–290, March 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671502.

Anonymous:1970:ACCd

[Ano70e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(4):365–385, April 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671522.

Anonymous:1970:ACCe

[Ano70f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(5):467–482, May 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671553.

Anonymous:1970:ACCf

[Ano70g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(6):567–582, June 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671581.

Anonymous:1970:ACCg

[Ano70h] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(7):661–678, July 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671605.

Anonymous:1970:ACCh

[Ano70i] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (8):767–782, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671630.

Anonymous:1970:ACCi

[Ano70j] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(9):861–877, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671660.

Anonymous:1970:ACCj

[Ano70k] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(10):991–1008, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671412.

Anonymous:1970:ACCk

[Ano70l] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (11):1121–1138, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671439.

Anonymous:1970:ACCl

[Ano70m] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (12):1229–1297, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671461.

Anonymous:1970:AI

[Ano70n] Anonymous. Author index. IEEE Transactions on Computers, C-19(1):100, January 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008910.

Anonymous:1970:CGC

[Ano70o] Anonymous. Comment on “Gray Code Counter”. IEEE Transactions on Computers, C-19(7):653–654, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671593.

Anonymous:1970:Ca

[Ano70p] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (1):82–84, January 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008906.

Anonymous:1970:Cb

[Ano70q] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (2):175–176, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Cc

[Ano70r] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (5):460–461, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Cd

[Ano70s] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (6):561–562, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Ce

[Ano70t] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (7):655, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Cf

[Ano70u] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (8):760–761, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Cg

[Ano70v] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (9):852–853, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Ch

[Ano70w] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (10):984–985, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Ci

[Ano70x] Anonymous. Contributors. IEEE Transactions on Computers, C-19(12):1227–1228, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:CPO

[Ano70y] Anonymous. Correction to “Pictorial Output with a Line Printer”. IEEE Transactions on Computers, C-19(9):851, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671649. See [Mac70].

Anonymous:1970:DCI

[Ano70z] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-19(1):94–99, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008908.

Anonymous:1970:EN

[Ano70-27] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-19(5):387, May 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671528.

Anonymous:1970:II

[Ano70-28] Anonymous. Identifier index.

IEEE Transactions on Computers, C-19(1):99–100, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008909.

Anonymous:1970:ITCa

[Ano70-29] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(2):c1–195–c, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCd

[Ano70-30] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(5):c1–482–b, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCe

[Ano70-31] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(6):c1–582–b, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCf

[Ano70-32] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(7):c1–678–b, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCg

[Ano70-33] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(8):c1–782–b, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCi

[Ano70-34] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(10):c1–1008–d, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCj

[Ano70-35] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(11):c1–1138–d, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCk

[Ano70-36] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(12):c1–1297–o, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCh

[Ano70-37] Anonymous. IEEE transactions on computerss. IEEE Transactions on Computers, C-19(9):c1– 877–c, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:IAa

[Ano70-38] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(2):195–b, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671487.

Anonymous:1970:IAb

[Ano70-39] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(3):290–c, March 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671505.

Anonymous:1970:IAc

[Ano70-40] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(4):385–b, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671524.

Anonymous:1970:IAd

[Ano70-41] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(5):482–a, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671554.

Anonymous:1970:IAe

[Ano70-42] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(6):582–a, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671582.

Anonymous:1970:IAf

[Ano70-43] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(7):678–a, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671606.

Anonymous:1970:IAg

[Ano70-44] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(8):782–a, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671631.

Anonymous:1970:IAh

[Ano70-45] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(9):877–b, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671662.

Anonymous:1970:IAi

[Ano70-46] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(10):1008–c, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671415.

Anonymous:1970:IAj

[Ano70-47] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(11):1138–c, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671442.

Anonymous:1970:IAk

[Ano70-48] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(12):1297–n, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671464.

Anonymous:1971:IIT

[Ano71a] Anonymous. 1971 index IEEE Transactions on Computers vol. C-20. IEEE Transactions on Computers, C-20(12):1637–b–1637– f, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671789.

Anonymous:1971:ACCa

[Ano71b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (1):115–132, January 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671686.

Anonymous:1971:ACCb

[Ano71c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (2):243–260, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671819.

Anonymous:1971:ACCc

[Ano71d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (3):364–380, March 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671841.

Anonymous:1971:ACCd

[Ano71e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(4):478–492, April 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671871.

Anonymous:1971:ACCe

[Ano71f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(5):599–616, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671901.

Anonymous:1971:ACCf

[Ano71g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(6):713–726, June 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671928.

Anonymous:1971:ACCg

[Ano71h] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(7):813–830, July 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671945.

Anonymous:1971:ACCh

[Ano71i] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (8):949–964, August 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671976.

Anonymous:1971:ACCi

[Ano71j] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(9):1121–1137, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672010.

Anonymous:1971:ACCj

[Ano71k] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(10):1224–1239, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671708.

Anonymous:1971:ACCk

[Ano71l] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (11):1420–1435, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671743.

Anonymous:1971:ACCl

[Ano71m] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (12):1623–1637, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671787.

Anonymous:1971:Ca

[Ano71n] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (1):112–114, January 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cb

[Ano71o] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (2):240–242, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cc

[Ano71p] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (3):362–363, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cd

[Ano71q] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (4):477, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Ce

[Ano71r] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (6):711–712, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cf

[Ano71s] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (8):945–947, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cg

[Ano71t] Anonymous. Contributors. IEEE Transactions on Computers, C-20(9):1118–1120, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Ch

[Ano71u] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (10):1220, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Ci

[Ano71v] Anonymous. Contributors. IEEE Transactions on Computers, C-20(11):1415–1419, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cj

[Ano71w] Anonymous. Contributors. IEEE Transactions on Computers, C-20(12):1619–1621, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ENa

[Ano71x] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-20(4):381, April 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671845.

Anonymous:1971:ENb

[Ano71y] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-20(10):1139, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671690.

Anonymous:1971:ITCa

[Ano71z] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(1):c1–132–b, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCb

[Ano71-27] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(2):c1–260–b, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCc

[Ano71-28] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(3):c1–380–b, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCd

[Ano71-29] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(4):c1–492–b, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCe

[Ano71-30] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(5):c1–616–b, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCf

[Ano71-31] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(6):c1–726–d, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCh

[Ano71-32] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(8):c1–964–d, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCi

[Ano71-33] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(9):c1–1137–e, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCj

[Ano71-34] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(10):c1–1239–e, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCk

[Ano71-35] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(11):c1–1435–c, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCl

[Ano71-36] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(12):c1–1637–q, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:IAa

[Ano71-37] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(1):132–a, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671687.

Anonymous:1971:IAb

[Ano71-38] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(2):260–a, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671820.

Anonymous:1971:IAc

[Ano71-39] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(3):380–a, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671842.

Anonymous:1971:IAd

[Ano71-40] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(4):492–a, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671872.

Anonymous:1971:IAe

[Ano71-41] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(5):616–a, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671902.

Anonymous:1971:IAf

[Ano71-42] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(6):726–c, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671931.

Anonymous:1971:IAg

[Ano71-43] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(7):830–a, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671946.

Anonymous:1971:IAh

[Ano71-44] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(8):964–c, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671979.

Anonymous:1971:IAi

[Ano71-45] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(9):1137–d, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672014.

Anonymous:1971:IAj

[Ano71-46] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(10):1239–d, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671712.

Anonymous:1971:IAk

[Ano71-47] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(11):1435–b, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671745.

Anonymous:1971:IAl

[Ano71-48] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(12):1637–p, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671792.

Anonymous:1971:NEC

[Ano71-49] Anonymous. New Editor-in-Chief. IEEE Transactions on Computers, C-20(1):1, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671665.

Anonymous:1971:P

[Ano71-50] Anonymous. Preface. IEEE Transactions on Computers, C-20(9):965–966, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671982.

Anonymous:1971:SI

[Ano71-51] Anonymous. Subject index. IEEE Transactions on Computers, C-20(12):1637–f–1637–n, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671790.

Anonymous:1971:TPR

[Ano71-52] Anonymous. Technical paper referees. IEEE Transactions on Computers, C-20(1):2–6, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:IIT

[Ano72a] Anonymous. 1972 index IEEE Transactions on Computers vol. C-21. IEEE Transactions on Computers, C-21(12):1475–b–1475– o, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672125.

Anonymous:1972:ACCa

[Ano72b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (1):111–127, January 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672039.

Anonymous:1972:ACCb

[Ano72c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (2):225–235, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008936.

Anonymous:1972:ACCc

[Ano72d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (3):323–331, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008963.

Anonymous:1972:ACCd

[Ano72e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(4):415–425, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008995.

Anonymous:1972:ACCe

[Ano72f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(5):519–536, May 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672155.

Anonymous:1972:ACCf

[Ano72g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(6):619–627, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009029.

Anonymous:1972:ACCg

[Ano72h] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(7):821–836, July 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672181.

Anonymous:1972:ACCh

[Ano72i] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (8):925–934, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009059.

Anonymous:1972:ACCi

[Ano72j] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(9):1035–1044, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009094.

Anonymous:1972:ACCj

[Ano72k] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(10):1137–1152, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672061.

Anonymous:1972:ACCk

[Ano72l] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (11):1253–1267, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672085.

Anonymous:1972:ACCl

[Ano72m] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (12):1460–1475, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672123.

Anonymous:1972:AII

[Ano72n] Anonymous. Applications are invited for IEEE/Popov Society Exchange Delegates. IEEE Transactions on Computers, C-21(12): 1475–a, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672124.

Anonymous:1972:AIa

[Ano72o] Anonymous. Author index. IEEE Transactions on Computers, C-21(2):240, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008939.

Anonymous:1972:AIb

[Ano72p] Anonymous. Author index. IEEE Transactions on Computers, C-21(3):336, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008966.

Anonymous:1972:AIc

[Ano72q] Anonymous. Author index. IEEE Transactions on Computers, C-21(4):432, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008998.

Anonymous:1972:AId

[Ano72r] Anonymous. Author index. IEEE Transactions on Computers, C-21(6):632, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009032.

Anonymous:1972:AIe

[Ano72s] Anonymous. Author index. IEEE Transactions on Computers, C-21(8):939, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009062.

Anonymous:1972:AIf

[Ano72t] Anonymous. Author index. IEEE Transactions on Computers, C-21 (9):1051, September 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009097.

Anonymous:1972:BRa

[Ano72u] Anonymous. Book review. IEEE Transactions on Computers, C-21(2):224, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008935.

Anonymous:1972:BRb

[Ano72v] Anonymous. Book review. IEEE Transactions on Computers, C-21(3):322, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008962.

Anonymous:1972:BRd

[Ano72w] Anonymous. Book review. IEEE Transactions on Computers, C-21(6):618, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009028.

Anonymous:1972:BRc

[Ano72x] Anonymous. Book reviews. IEEE Transactions on Computers, C-21(4):413–414, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008994.

Anonymous:1972:BRe

[Ano72y] Anonymous. Book reviews. IEEE Transactions on Computers, C-21(8):924, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009058.

Anonymous:1972:BRf

[Ano72z] Anonymous. Book reviews. IEEE Transactions on Computers, C-21(9):1033–1034, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009093.

Anonymous:1972:Ca

[Ano72-27] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (1):105–107, January 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Cb

[Ano72-28] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (2):221–223, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008934.

Anonymous:1972:Cc

[Ano72-29] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (3):321–322, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008961.

Anonymous:1972:Cd

[Ano72-30] Anonymous. Contributors. IEEE Transactions on Computers, C-21(4):412–413, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008993.

Anonymous:1972:Ce

[Ano72-31] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (5):515–516, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Cf

[Ano72-32] Anonymous. Contributors. IEEE Transactions on Computers, C-21(6):617, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009027.

Anonymous:1972:Cg

[Ano72-33] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (7):819–820, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Ch

[Ano72-34] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (8):922–923, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009057.

Anonymous:1972:Ci

[Ano72-35] Anonymous. Contributors. IEEE Transactions on Computers, C-21(9):1032–1033, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009092.

Anonymous:1972:Cj

[Ano72-36] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (10):1132–1133, October 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Ck

[Ano72-37] Anonymous. Contributors. IEEE Transactions on Computers, C-21(11):1248–1249, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Cl

[Ano72-38] Anonymous. Contributors. IEEE Transactions on Computers, C-21(12):1455–1457, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:CBF

[Ano72-39] Anonymous. Correction to “Batch-Fabricated Three-Dimensional Planar Coaxial Interconnections for Microelectronic Systems”. IEEE Transactions on Computers, C-21(8):922, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009056. See comments

[Par71].

Anonymous:1972:CJC

[Ano72-40] Anonymous. Correction to the June 1972 contributors section. IEEE Transactions on Computers, C-21(11):1247, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672079.

Anonymous:1972:DCIa

[Ano72-41] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(2): 236–239, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008937.

Anonymous:1972:DCIb

[Ano72-42] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(3): 331–335, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008964.

Anonymous:1972:DCIc

[Ano72-43] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(4):426–430, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008996.

Anonymous:1972:DCId

[Ano72-44] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(6):628–630, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009030.

Anonymous:1972:DCIe

[Ano72-45] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(8): 935–938, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009060.

Anonymous:1972:DCIf

[Ano72-46] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(9):1045– 1049, September 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009095.

Anonymous:1972:ENa

[Ano72-47] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-21(2):129, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008916.

Anonymous:1972:ENb

[Ano72-48] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-21(6):537, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009005.

Anonymous:1972:ENc

[Ano72-49] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-21(9):941, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009069.

Anonymous:1972:ECR

[Ano72-50] Anonymous. Expanded Computing Reviews now available to IEEE Computer Society members. IEEE Transactions on Computers, C-21(12):1475–p, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672126.

Anonymous:1972:IIa

[Ano72-51] Anonymous. Identifier index. IEEE Transactions on Computers, C-21(2):239–240, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008938.

Anonymous:1972:IIb

[Ano72-52] Anonymous. Identifier index. IEEE Transactions on Computers, C-21(3):335–336, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008965.

Anonymous:1972:IIc

[Ano72-53] Anonymous. Identifier index. IEEE Transactions on Computers, C-21(4):431–432, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008997.

Anonymous:1972:IId

[Ano72-54] Anonymous. Identifier index.

IEEE Transactions on Computers, C-21(6):631, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009031.

Anonymous:1972:IIe

[Ano72-55] Anonymous. Identifier index.

IEEE Transactions on Computers, C-21(8):938–939, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009061.

Anonymous:1972:IIf

[Ano72-56] Anonymous. Identifier index.

IEEE Transactions on Computers, C-21(9):1050, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009096.

Anonymous:1972:ICSa

[Ano72-57] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-21(11):1268, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672086.

Anonymous:1972:ICSb

[Ano72-58] Anonymous. IEEE Computer Society membership & publications.

IEEE Transactions on Computers, C-21(12):1475–q, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672127.

Anonymous:1972:ITCa

[Ano72-59] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(1):c1–127–c, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCb

[Ano72-60] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(5):c1–536–b, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCc

[Ano72-61] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(7):c1–836–b, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCd

[Ano72-62] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(10):c1–1152–b, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCe

[Ano72-63] Anonymous. IEEE Transactions On Computers. IEEE Transactions on Computers, C-21(11):c1–1268–b, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCf

[Ano72-64] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(12):c1–1475–s, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:IAa

[Ano72-65] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(1):127–b, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672041.

Anonymous:1972:IAb

[Ano72-66] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(5):536–a, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672156.

Anonymous:1972:IAc

[Ano72-67] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(7):836–a, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672182.

Anonymous:1972:IAd

[Ano72-68] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(10):1152–a, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672062.

Anonymous:1972:IAe

[Ano72-69] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(11):1268–a, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672087.

Anonymous:1972:IAf

[Ano72-70] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(12):1475–r, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672128.

Anonymous:1972:P

[Ano72-71] Anonymous. Preface. IEEE Transactions on Computers, C-21(7):633–635, July 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672159.

Anonymous:1973:IIT

[Ano73a] Anonymous. 1973 index IEEE Transactions on Computers vol.

C-22. IEEE Transactions on Computers, C-22(12):1144–a–1144– l, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672258.

Anonymous:1973:BRa

[Ano73b] Anonymous. Book reviews. IEEE Transactions on Computers, C-22(6):628, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009120.

Anonymous:1973:BRb

[Ano73c] Anonymous. Book reviews. IEEE Transactions on Computers, C-22(7):704–708, July 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009141.

Anonymous:1973:BRc

[Ano73d] Anonymous. Book reviews. IEEE Transactions on Computers, C-22(9):879–884, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009190.

Anonymous:1973:CEP

[Ano73e] Anonymous. A change in editorial policy for Transactions papers and abstracts. IEEE Transactions on Computers, C-22(1):1, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672185.

Anonymous:1973:CJC

[Ano73f] Anonymous. Correction to a January 1973 contributor entry. IEEE Transactions on Computers, C-22(4):428, April 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672329.

Anonymous:1973:EN

[Ano73g] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-22(11):965, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672224.

Anonymous:1973:ECR

[Ano73h] Anonymous. Expanded computing reviews now available to IEEE Computer Society members. IEEE Transactions on Computers, C-22(2):222–b, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672287.

Anonymous:1973:ICSa

[Ano73i] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-22(2):222–a, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672286.

Anonymous:1973:ICSb

[Ano73j] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-22(3):315–a, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672305.

Anonymous:1973:ICSc

[Ano73k] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-22(4):432, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672334.

Anonymous:1973:ICSd

[Ano73l] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-22(5):548, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672360.

Anonymous:1973:ITCd

[Ano73m] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-22(4):c1–432–b, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1973:ITCe

[Ano73n] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-22(5):c1–548–b, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1973:IAa

[Ano73o] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(1):112–a, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672202.

Anonymous:1973:IAb

[Ano73p] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(2):222–c, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672288.

Anonymous:1973:IAc

[Ano73q] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(3):315–b, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672306.

Anonymous:1973:IAd

[Ano73r] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(4):432–a, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672335.

Anonymous:1973:IAe

[Ano73s] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(5):548–a, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672361.

Anonymous:1973:IAf

[Ano73t] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(10):964–a, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672220.

Anonymous:1973:IAg

[Ano73u] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(11):1044–a, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672237.

Anonymous:1973:IAh

[Ano73v] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(12):1144–m, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672259.

Anonymous:1973:PCI

[Ano73w] Anonymous. Parallel computation: An introduction. IEEE Transactions on Computers, C-22 (8):709–710, August 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009148.

Anonymous:1973:Pa

[Ano73x] Anonymous. Peripherals. IEEE Transactions on Computers, C-22(11):1044–b, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672238.

Anonymous:1973:Pb

[Ano73y] Anonymous. Peripherals. IEEE Transactions on Computers, C-22(12):1144–n, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672260.

Anonymous:1973:RMJ

[Ano73z] Anonymous. Reply by M. James Errico. IEEE Transactions on Computers, C-22(12): 1143–1144, December 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672257.

Anonymous:1974:IIT

[Ano74a] Anonymous. 1974 index IEEE Transactions on Computers vol. C-23. IEEE Transactions on Computers, C-23(12):1330–a–1330– n, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672461.

Anonymous:1974:CGWa

[Ano74b] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(5):560–b, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672585.

Anonymous:1974:CGWb

[Ano74c] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(6):648–b, June 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672602.

Anonymous:1974:CGWc

[Ano74d] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(7):752–b, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672620.

Anonymous:1974:CGWd

[Ano74e] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(8):880–b, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672641.

Anonymous:1974:CGWe

[Ano74f] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(9):992–b, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672669.

Anonymous:1974:ENa

[Ano74g] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-23(4):337, April 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672539.

Anonymous:1974:ENb

[Ano74h] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-23(8):753, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672623.

Anonymous:1974:ENc

[Ano74i] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-23(12):1217, December 1974.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672432.

Anonymous:1974:ICSa

[Ano74j] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(1):112–b, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672390.

Anonymous:1974:ICSc

[Ano74k] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(2):224–b, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672506.

Anonymous:1974:ICSe

[Ano74l] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(3):336–b, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672536.

Anonymous:1974:ICSg

[Ano74m] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(4):448–b, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672556.

Anonymous:1974:ICSn

[Ano74n] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(10):1104–b, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672407.

Anonymous:1974:ICSp

[Ano74o] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(11):1216–b, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672429.

Anonymous:1974:ICSb

[Ano74p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(2): 224–a, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672505.

Anonymous:1974:ICSd

[Ano74q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(3):336–a, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672535.

Anonymous:1974:ICSf

[Ano74r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(4):448–a, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672555.

Anonymous:1974:ICSh

[Ano74s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(5):560–a, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672584.

Anonymous:1974:ICSi

[Ano74t] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(6):648–a, June 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672601.

Anonymous:1974:ICSj

[Ano74u] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(7):752–a, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672619.

Anonymous:1974:ICSk

[Ano74v] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(8): 880–a, August 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672640.

Anonymous:1974:ICSl

[Ano74w] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(9): 992–a, September 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672668.

Anonymous:1974:ICSm

[Ano74x] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(10): 1104–a, October 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672406.

Anonymous:1974:ICSo

[Ano74y] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(11): 1216–a, November 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672428.

Anonymous:1974:ICSq

[Ano74z] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(12): 1330–o, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672462.

Anonymous:1974:IA

[Ano74-27] Anonymous. Information for authors. IEEE Transactions on Computers, C-23(1):112–a, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672389.

Anonymous:1974:PGM

[Ano74-28] Anonymous. People are getting more important. IEEE Transactions on Computers, C-23(12): 1330–p, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672463.

Anonymous:1975:IIT

[Ano75a] Anonymous. 1975 index IEEE Transactions on Computers vol. C-24. IEEE Transactions on Computers, C-24(12):1242–a–1242– l, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672766.

Anonymous:1975:CNY

[Ano75b] Anonymous. Computers the next 5 years. IEEE Transactions on Computers, C-24(12): 1242–m–1242–n, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672767.

Anonymous:1975:DEMa

[Ano75c] Anonymous. Does easier mean better? IEEE Transactions on Computers, C-24(6):680–b, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672878.

Anonymous:1975:DEMb

[Ano75d] Anonymous. Does easier mean better? IEEE Transactions on Computers, C-24(7):760–b, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672897.

Anonymous:1975:DEMc

[Ano75e] Anonymous. Does easier mean better? IEEE Transactions on Computers, C-24(8):856–b, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672919.

Anonymous:1975:DEMd

[Ano75f] Anonymous. Does easier mean better? IEEE Transactions on Computers, C-24(9):952–b, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672941.

Anonymous:1975:ENa

[Ano75g] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-24(3):225, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672795.

Anonymous:1975:ENb

[Ano75h] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-24(9):857, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672922.

Anonymous:1975:ICSd

[Ano75i] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(3):336–b, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672816.

Anonymous:1975:ICSf

[Ano75j] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(4):456–b, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672836.

Anonymous:1975:ICSh

[Ano75k] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(5):584–b, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672860.

Anonymous:1975:ICSn

[Ano75l] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(10):1048–b, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672722.

Anonymous:1975:ICSp

[Ano75m] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(11):1144–b, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672747.

Anonymous:1975:ICSr

[Ano75n] Anonymous. IEEE Computer Society membership & publications.

IEEE Transactions on Computers, C-24(12):1242–p, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672769.

Anonymous:1975:ICSa

[Ano75o] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(1): 112–a, January 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672689.

Anonymous:1975:ICSb

[Ano75p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(2): 224–a, February 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672791.

Anonymous:1975:ICSc

[Ano75q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(3):336–a, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672815.

Anonymous:1975:ICSe

[Ano75r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(4):456–a, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672835.

Anonymous:1975:ICSg

[Ano75s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(5):584–a, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672859.

Anonymous:1975:ICSi

[Ano75t] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(6):680–a, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672877.

Anonymous:1975:ICSj

[Ano75u] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(7):760–a, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672896.

Anonymous:1975:ICSk

[Ano75v] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(8): 856–a, August 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672918.

Anonymous:1975:ICSl

[Ano75w] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(9): 952–a, September 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672940.

Anonymous:1975:ICSm

[Ano75x] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(10): 1048–a, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672721.

Anonymous:1975:ICSo

[Ano75y] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(11): 1144–a, November 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672746.

Anonymous:1975:ICSq

[Ano75z] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(12): 1242–o, December 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672768.

Anonymous:1975:IA

[Ano75-27] Anonymous. Information for authors. IEEE Transactions on Computers, C-24(1):112, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672688.

Anonymous:1975:MRN

[Ano75-28] Anonymous. A method for reducing the number of input variables to synchronous sequential circuits. IEEE Transactions on Computers, C-24(10):1029–1031, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672713.

Anonymous:1975:PGMa

[Ano75-29] Anonymous. People are getting more important. IEEE Transactions on Computers, C-24 (1):112–b, January 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672690.

Anonymous:1975:PGMb

[Ano75-30] Anonymous. People are getting more important. IEEE Transactions on Computers, C-24 (2):224–b, February 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672792.

Anonymous:1976:LR

[Ano76a] Anonymous. 1975 list of reviewers. IEEE Transactions on Computers, C-25(6):668–672, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1976:IIT

[Ano76b] Anonymous. 1976 index IEEE Transactions on Computers vol. C-25. IEEE Transactions on Computers, C-25(12):1361, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674605.

Anonymous:1976:AICa

[Ano76c] Anonymous. Announcing the IEEE Computer Society’s standing subscription plan. IEEE Transactions on Computers, C-25(5):552, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674649.

Anonymous:1976:AICb

[Ano76d] Anonymous. Announcing the IEEE Computer Society’s standing subscription plan. IEEE Transactions on Computers, C-25(8):862, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674705.

Anonymous:1976:CLM

[Ano76e] Anonymous. Computers … by the millions, for the millions. IEEE Transactions on Computers, C-25(8):862, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674708.

Anonymous:1976:EN

[Ano76f] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-25(3):209, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009246.

Anonymous:1976:FI

[Ano76g] Anonymous. Foreword and introduction. IEEE Transactions on Computers, C-25(10): 961–962, October 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674536.

Anonymous:1976:GEI

[Ano76h] Anonymous. Guidelines for Editors of the IEEE Computer Society. IEEE Transactions on Computers, C-25(11):1187–1188, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674580.

Anonymous:1976:ICSc

[Ano76i] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-25(5):552, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674651.

Anonymous:1976:ICSe

[Ano76j] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-25(6):672, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674673.

Anonymous:1976:ICSa

[Ano76k] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(4):456, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674628.

Anonymous:1976:ICSb

[Ano76l] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(5):552, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674650.

Anonymous:1976:ICSd

[Ano76m] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(6):672, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674672.

Anonymous:1976:ICSf

[Ano76n] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(7):768, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674691.

Anonymous:1976:ICSg

[Ano76o] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(8):862, August 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674707.

Anonymous:1976:ICSh

[Ano76p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(9): 960, September 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674724.

Anonymous:1976:ICSi

[Ano76q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(10): 1056, October 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674551.

Anonymous:1976:ICSj

[Ano76r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(11): 1188, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674582.

Anonymous:1976:ICSk

[Ano76s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(12): 1361, December 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674606.

Anonymous:1976:IYC

[Ano76t] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Computers, C-25(8):862, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674706.

Anonymous:1976:MAC

[Ano76u] Anonymous. Management auditing of computer operations. IEEE Transactions on Computers, C-25(4):456, April 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674629.

Anonymous:1976:PPNa

[Ano76v] Anonymous. Papers to be published in the next issue August 1976. IEEE Transactions on Computers, C-25(7):768, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674692.

Anonymous:1976:PPNd

[Ano76w] Anonymous. Papers to be published in the next issue December 1976. IEEE Transactions on Computers, C-25(11):1188, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674583.

Anonymous:1976:PPNe

[Ano76x] Anonymous. Papers to be published in the next issue January 1977. IEEE Transactions on Computers, C-25(12):1361, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674607.

Anonymous:1976:PPNc

[Ano76y] Anonymous. Papers to be published in the next issue November 1976. IEEE Transactions on Computers, C-25(10):1056, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674552.

Anonymous:1976:PPNb

[Ano76z] Anonymous. Papers to be published in the next issue October 1976. IEEE Transactions on Computers, C-25(9):960, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674725.

Anonymous:1976:P

[Ano76-27] Anonymous. Preface. IEEE Transactions on Computers, C-25 (12):1189–1190, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674586.

Anonymous:1976:SSA

[Ano76-28] Anonymous. Special sections announcement. IEEE Transactions on Computers, C-25(11):1188, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674581.

Anonymous:1977:IIT

[Ano77a] Anonymous. 1977 index IEEE Transactions on Computers vol. C-26. IEEE Transactions on Computers, C-26(12):1308, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674800.

Anonymous:1977:ACTa

[Ano77b] Anonymous. Announcing computer technology: Status, limits, alternatives. IEEE Transactions on Computers, C-26(11):1180, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674775.

Anonymous:1977:ACTb

[Ano77c] Anonymous. Announcing computer technology: Status, limits, alternatives. IEEE Transactions on Computers, C-26(12):1308, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674804.

Anonymous:1977:BCP

[Ano77d] Anonymous. Books for computer professionals. IEEE Transactions on Computers, C-26(12): 1308, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674802.

Anonymous:1977:CP

[Ano77e] Anonymous. Call for papers. IEEE Transactions on Computers, C-26 (3):320, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1977:CF

[Ano77f] Anonymous. Compcon 77 Fall.

IEEE Transactions on Computers, C-26(5):512, May 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674873.

Anonymous:1977:CAF

[Ano77g] Anonymous. Computer arithmetic:

Foreword and survey. IEEE Transactions on Computers, C-26(7):609, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674892.

Anonymous:1977:ENc

[Ano77h] Anonymous. Editor’s note. IEEE Transactions on Computers, C-26 (9):837, September 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674931.

Anonymous:1977:ENa

[Ano77i] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-26(1):1, January 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009266.

Anonymous:1977:ENb

[Ano77j] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-26(6):513, June 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674876.

Anonymous:1977:ICSa

[Ano77k] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(3):320, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674837.

Anonymous:1977:ICSb

[Ano77l] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(4):416, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674854.

Anonymous:1977:ICSc

[Ano77m] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(5):512, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674872.

Anonymous:1977:ICSd

[Ano77n] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(6):608, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674888.

Anonymous:1977:ICSe

[Ano77o] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(7):704, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674904.

Anonymous:1977:ICSf

[Ano77p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(8):836, August 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674927.

Anonymous:1977:ICSg

[Ano77q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(9): 932, September 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674943.

Anonymous:1977:ICSh

[Ano77r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(10): 1060, October 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674748.

Anonymous:1977:ICSi

[Ano77s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(11): 1180, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674774.

Anonymous:1977:ICSj

[Ano77t] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(12): 1308, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674803.

Anonymous:1977:IYC

[Ano77u] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Computers, C-26(12):1308, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674801.

Anonymous:1977:MDC

[Ano77v] Anonymous. Maxwell’s displacement current. IEEE Transactions on Computers, C-26(3):318, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674834. See [Cat77].

Anonymous:1977:PPNc

[Ano77w] Anonymous. Papers to be published in the next issue August 1977. IEEE Transactions on Computers, C-26(7):704, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674905.

Anonymous:1977:PPNb

[Ano77x] Anonymous. Papers to be published in the next issue July 1977. IEEE Transactions on Computers, C-26(6):608, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674889.

Anonymous:1977:PPNa

[Ano77y] Anonymous. Papers to be published in the next issue May 1977. IEEE Transactions on Computers, C-26(4):416, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674855.

Anonymous:1977:PPNf

[Ano77z] Anonymous. Papers to be published in the next issue November 1977. IEEE Transactions on Computers, C-26(10):1060, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674749.

Anonymous:1977:PPNe

[Ano77-27] Anonymous. Papers to be published in the next issue October 1977. IEEE Transactions on Computers, C-26(9):932, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674944.

Anonymous:1977:PPNd

[Ano77-28] Anonymous. Papers to be published in the next issue September 1977. IEEE Transactions on Computers, C-26(8):836, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674928.

Anonymous:1978:LR

[Ano78a] Anonymous. 1976–1977 list of referees. IEEE Transactions on Computers, C-27(5):472–476, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See correction [Ano78s].

Anonymous:1978:IIT

[Ano78b] Anonymous. 1978 index IEEE Transactions on Computers vol. C-27. IEEE Transactions on Computers, C-27(12):1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675036.

Anonymous:1978:APAa

[Ano78c] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-27(4):384, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675113.

Anonymous:1978:APAc

[Ano78d] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-27(7):679, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675171.

Anonymous:1978:APAd

[Ano78e] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-27(8):782, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675194.

Anonymous:1978:APAb

[Ano78f] Anonymous. Advance program announcements. IEEE Transactions on Computers, C-27(6):582, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675154.

Anonymous:1978:A

[Ano78g] Anonymous. Announcement. IEEE Transactions on Computers, C-27(7):585, July 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675156.

Anonymous:1978:ANQa

[Ano78h] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27 (9):879, September 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675210.

Anonymous:1978:ANQb

[Ano78i] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27(10):973, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674980.

Anonymous:1978:ANQc

[Ano78j] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27 (11):1089, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675005.

Anonymous:1978:BCP

[Ano78k] Anonymous. Books for computer professionals. IEEE Transactions on Computers, C-27(5):476, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675132.

Anonymous:1978:CPa

[Ano78l] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (1):96, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPb

[Ano78m] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (2):192, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPc

[Ano78n] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (8):782, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPd

[Ano78o] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (9):879, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPe

[Ano78p] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (10):973, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPf

[Ano78q] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (11):1089, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPg

[Ano78r] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (12):1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CLR

[Ano78s] Anonymous. Correction to “1976– 1977 List of Referees”. IEEE Transactions on Computers, C-27 (9):879, September 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675209. See [Ano78a].

Anonymous:1978:CFC

[Ano78t] Anonymous. Correction to front cover of December 1977 issue.

IEEE Transactions on Computers, C-27(3):288, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675095.

Anonymous:1978:EN

[Ano78u] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-27(5):385, May 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675116.

Anonymous:1978:ICSj

[Ano78v] Anonymous. The IEEE Computer Society announces tutorial and selected papers in digital image processing. IEEE Transactions on Computers, C-27(10):973, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674982.

Anonymous:1978:ICSl

[Ano78w] Anonymous. The IEEE Computer Society announces tutorial and selected papers in digital image processing. IEEE Transactions on Computers, C-27(11):1089, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675006.

Anonymous:1978:ICSn

[Ano78x] Anonymous. The IEEE Computer Society announces tutorial and selected papers in digital image processing. IEEE Transactions on Computers, C-27(12):1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675037.

Anonymous:1978:ICSa

[Ano78y] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(1):96, January 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674962.

Anonymous:1978:ICSb

[Ano78z] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(2):192, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675067.

Anonymous:1978:ICSc

[Ano78-27] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(3):288, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675096.

Anonymous:1978:ICSd

[Ano78-28] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(4):384, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675112.

Anonymous:1978:ICSe

[Ano78-29] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(5):476, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675133.

Anonymous:1978:ICSf

[Ano78-30] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(6):582, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675153.

Anonymous:1978:ICSg

[Ano78-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(7):679, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675170.

Anonymous:1978:ICSh

[Ano78-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(8):782, August 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675195.

Anonymous:1978:ICSi

[Ano78-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(9): 879, September 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675211.

Anonymous:1978:ICSk

[Ano78-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(10):973, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674983.

Anonymous:1978:ICSm

[Ano78-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(11): 1089, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675007.

Anonymous:1978:ICSo

[Ano78-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(12): 1227, December 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675039.

Anonymous:1978:ITPa

[Ano78-37] Anonymous. IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27(10):973, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674981.

Anonymous:1978:ITPb

[Ano78-38] Anonymous. IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27(11):1089, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675004.

Anonymous:1978:ITSa

[Ano78-39] Anonymous. IEEE Transactions on Software Engineering. IEEE Transactions on Computers, C-27(3):288, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675097.

Anonymous:1978:ITSb

[Ano78-40] Anonymous. IEEE Transactions On Software Engineering. IEEE Transactions on Computers, C-27(7):679, July 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675169.

Anonymous:1978:ITSc

[Ano78-41] Anonymous. IEEE Transactions On Software Engineering. IEEE Transactions on Computers, C-27(8):782, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675193.

Anonymous:1978:IYCa

[Ano78-42] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Computers, C-27(6):582, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675152.

Anonymous:1978:IYCb

[Ano78-43] Anonymous. if your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-27(12):1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675038.

Anonymous:1978:IA

[Ano78-44] Anonymous. Information for authors. IEEE Transactions on Computers, C-27(5):476, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675134.

Anonymous:1979:IIT

[Ano79a] Anonymous. 1979 index IEEE Transactions on Computers vol. C-28. IEEE Transactions on Computers, C-28(12):947, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675290.

Anonymous:1979:PCF

[Ano79b] Anonymous. 254: Proceedings:

Compcon 79 Fall “Using Microprocessors — Extending our Reach”. IEEE Transactions on Computers, C-28(12):947–o, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675293.

Anonymous:1979:ICPa

[Ano79c] Anonymous. 5th international conference on pattern recognition. IEEE Transactions on Computers, C-28(10):805, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:ICPb

[Ano79d] Anonymous. 5th international conference on pattern recognition. IEEE Transactions on Computers, C-28(11):875, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:APAa

[Ano79e] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(3):281, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675343.

Anonymous:1979:APAb

[Ano79f] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(4):329, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675358.

Anonymous:1979:APAc

[Ano79g] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(5):378, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675374.

Anonymous:1979:APAd

[Ano79h] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(6):458, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675384.

Anonymous:1979:APAe

[Ano79i] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(7):543, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675404.

Anonymous:1979:APAf

[Ano79j] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(8):590, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675419.

Anonymous:1979:ANQa

[Ano79k] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-28(5):378, May 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675371.

Anonymous:1979:ANQb

[Ano79l] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-28 (9):693, September 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675444.

Anonymous:1979:AICa

[Ano79m] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(6):458, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675387.

Anonymous:1979:AICb

[Ano79n] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(7):543, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675406.

Anonymous:1979:AICc

[Ano79o] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(8):590, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675422.

Anonymous:1979:AICd

[Ano79p] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(9):693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675447.

Anonymous:1979:AICe

[Ano79q] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(10):805, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675257.

Anonymous:1979:CPa

[Ano79r] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (2):175, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPb

[Ano79s] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (2):175, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPc

[Ano79t] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (3):281, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPd

[Ano79u] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (3):281, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPe

[Ano79v] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (4):329, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPf

[Ano79w] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (4):329, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPg

[Ano79x] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (6):458, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPh

[Ano79y] Anonymous. call for papers. IEEE Transactions on Computers, C-28 (8):590, August 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPi

[Ano79z] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (9):693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPj

[Ano79-27] Anonymous. Call for papers. IEEE Transactions on Computers, C-28(11):875, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPk

[Ano79-28] Anonymous. Call for papers.

IEEE Transactions on Computers, C-28(11):875, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:ENa

[Ano79-29] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-28(1):1, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675215.

Anonymous:1979:ENb

[Ano79-30] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-28(3):177, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675316.

Anonymous:1979:ENc

[Ano79-31] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-28(4):285, April 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675346.

Anonymous:1979:ENd

[Ano79-32] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-28(8):545, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675409.

Anonymous:1979:ICSa

[Ano79-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(1):86, January 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675233.

Anonymous:1979:ICSb

[Ano79-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(2):175, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675312.

Anonymous:1979:ICSc

[Ano79-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(3):281, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675342.

Anonymous:1979:ICSd

[Ano79-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(4):329, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675359.

Anonymous:1979:ICSe

[Ano79-37] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(5):378, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675373.

Anonymous:1979:ICSf

[Ano79-38] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(6):458, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675386.

Anonymous:1979:ICSg

[Ano79-39] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(7):543, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675405.

Anonymous:1979:ICSh

[Ano79-40] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(8):590, August 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675421.

Anonymous:1979:ICSi

[Ano79-41] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(9): 693, September 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675446.

Anonymous:1979:ICSj

[Ano79-42] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(10):805, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675256.

Anonymous:1979:ICSk

[Ano79-43] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(11): 875, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675274.

Anonymous:1979:ICSl

[Ano79-44] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(12): 947, December 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675294.

Anonymous:1979:ICFa

[Ano79-45] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-28(1):86, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675232.

Anonymous:1979:ICFb

[Ano79-46] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-28(12):947, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675291.

Anonymous:1979:IYCa

[Ano79-47] Anonymous. If your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-28(3):281, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675340.

Anonymous:1979:IYCb

[Ano79-48] Anonymous. If your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-28(4):329, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675357.

Anonymous:1979:IYCc

[Ano79-49] Anonymous. If your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-28(5):378, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675372.

Anonymous:1979:IYCd

[Ano79-50] Anonymous. If your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-28(9):693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675443.

Anonymous:1979:IA

[Ano79-51] Anonymous. Information for authors. IEEE Transactions on Computers, C-28(1):86, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675234.

Anonymous:1979:LR

[Ano79-52] Anonymous. List of referees. IEEE Transactions on Computers, C-28(12):945–947, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:NAI

[Ano79-53] Anonymous. Now available from the IEEE Computer Society. IEEE Transactions on Computers, C-28(12):947–q, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675295.

Anonymous:1979:PCS

[Ano79-54] Anonymous. Proceedings of the Conference on Specifications of Reliable Software. IEEE Transactions on Computers, C-28(12): 947–n, December 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675292.

Ahmed:1974:DCT

[ANR74] N. Ahmed, T. Natarajan, and K. R. Rao. Discrete Cosine Transfom. IEEE Transactions on Computers, C-23(1):90–93, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672377. See comments

[Sha75].

Anzelmo:1971:DSF

[Anz71] F. D. Anzelmo. A data-storage format for information system files. IEEE Transactions on Computers, C-20(1):39–43, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671672.

Atkins:1979:TCC

[AO79] D. E. Atkins and Shau-Chi Ong. Time-component complexity of two approaches to multioperand binary addition. IEEE Transactions on Computers, C-28(12): 918–926, December 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675281.

Aminoff:1974:CRA

[AOR74] C. G. Aminoff, L. Ojala, and E. T. Rautanen. On a class of recursive algorithms for continuous estimation of the mean. IEEE Transactions on Computers, C-23(2):191–194, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672479.

Andrews:1976:OPE

[AP76a] Harry C. Andrews and Claude L. Patterson. Outer product expansions and their uses in digital image processing. IEEE Transactions on Computers, C-25(2): 140–148, February 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009228.

Andrews:1976:DID

[AP76b] Harry C. Andrews and Claude L.

Patterson III. Digital interpolation of discrete images. IEEE Transactions on Computers, C-25 (2):196–202, February 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009235.

Agrawala:1974:MPS

[AR74] A. K. Agrawala and T. G. Rauscher.

Microprogramming: Perspective and status. IEEE Transactions on Computers, C-23(8):817–837, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672631.

Ashjaee:1977:TSC

[AR77] M. J. Ashjaee and S. M. Reddy.

On totally self-checking checkers for separable codes. IEEE Transactions on Computers, C-26 (8):737–744, August 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674911.

Agrawal:1978:MOA

[AR78] D. P. Agrawal and T. R. N.

Rao. On multiple operand addition of signed binary numbers. IEEE Transactions on Computers, C-27(11):1068–1070, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674998.

Arbib:1970:RMT

[Arb70] M. A. Arbib. R70-10 the minimalization of tree automata. IEEE Transactions on Computers, C-19(6):563, June 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671574.

Arbib:1976:AIC

[Arb76] M. A. Arbib. Artificial intelligence: Cooperative computation and man–machine symbiosis. IEEE Transactions on Computers, C-25(12):1346–1352, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674603.

Ari:1979:TLM

[Ari79] M. B. Ari. On transposing large 2^{n} × 2^{n} matrices. IEEE Transactions on Computers, C-28 (1):72–75, January 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675228.

Armstrong:1972:DMS

[Arm72] D. B. Armstrong. A deductive method for simulating faults in logic circuits. IEEE Transactions on Computers, C-21(5):464–471, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672135.

Armstrong:1973:DGG

[Arm73] J. R. Armstrong. Design of a graphic generator for remote terminal application. IEEE Transactions on Computers, C-22(5):464–468, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672342.

Arnold:1970:RTT

[Arn70] T. F. Arnold. R70-21 time and tape complexity of pushdown automaton languages. IEEE Transactions on Computers, C-19(7):658–659, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671601.

Arnold:1973:CCE

[Arn73] T. F. Arnold. The concept of coverage and its effect on the reliability model of a repairable system. IEEE Transactions on Computers, C-22(3):251–254, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672296.

Ahlgren:1971:CRA

[ARS71a] R. C. Ahlgren, H. F. Ryan, and C. W. Swonger. A character recognition application of an iterative procedure for feature selection. IEEE Transactions on Computers, C-20(9):1067–1075, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671998.

Ahmed:1971:FCB

[ARS71b] N. Ahmed, K. R. Rao, and R. B. Schultz. Fast complex BI-FORE transform by matrix partitioning. IEEE Transactions on Computers, C-20(6):707–710, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671926.

Abraham:1974:AAR

[AS74] J. A. Abraham and D. P. Siewiorek. An algorithm for the accurate reliability evaluation of triple modular redundancy networks. IEEE Transactions on Computers, C-23(7):682–692, July 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672609.

Arguello:1972:TFC

[ASS72] R. J. Arguello, H. R. Sellner, and J. A. Stuller. Transfer function compensation of sampled imagery. IEEE Transactions on Computers, C-21(7):812–818, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672179.

Asthana:1978:DCT

[Ast78] A. Asthana. Design and control of a three-stage switch matrix in the presence of fan-out. IEEE Transactions on Computers, C-27(10):886–895, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674967.

Avizienis:1970:UAB

[AT70] Algirdas Avižienis and Chin Tung.

A universal arithmetic building element (ABE) and design methods for arithmetic processors. IEEE Transactions on Computers, C-19 (8):733–745, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671616;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Avizienis.
pdf.

Athani:1975:VAC

[Ath75] V. V. Athani. The 16-vector algorithm for computer controlled digital X–Y plotter. IEEE Transactions on Computers, C-24(8): 831–835, August 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672909.

Atkins:1970:DAU

[Atk70] D. E. Atkins. Design of the arithmetic units of ILLIAC III: Use of redundancy and higher radix methods. IEEE Transactions on Computers, C-19(8):720–733, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671615;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Atkins.
pdf.

Arnold:1970:IRS

[ATN70] Thomas F. Arnold, Chung-Jen Tan, and Monroe M. Newborn. Iteratively realized sequential circuits. IEEE Transactions on Computers, C-19(1):54–66, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008900.

Aho:1974:DMR

[AU74] A. V. Aho and J. D. Ullman. Dynamic memories with rapid random and sequential access. IEEE Transactions on Computers, C-23 (3):272–276, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672515.

Austin:1970:UMP

[Aus70] B. J. Austin. Use of a macro processor in logical design. IEEE Transactions on Computers, C-19 (11):1085–1089, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671429.

Avizienis:1971:AEC

[Avi71] A. Avizienis. Arithmetic error codes: Cost and effectiveness studies for application in digital system design. IEEE Transactions on Computers, C-20 (11):1322–1331, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671727.

Avizienis:1973:AAE

[Avi73] Algirdas Avižienis. Arithmetic algorithms for error-coded operands. IEEE Transactions on Computers, C-22(6):567–572, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009108;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Avizienis.
pdf.

Aviziens:1976:FTS

[Avi76] A. Aviziens. Fault-tolerant systems.

IEEE Transactions on Computers, C-25(12):1304–1312, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674598.

Bruno:1971:TAC

[BA71] J. Bruno and S. M. Altman. A theory of asynchronous control networks. IEEE Transactions on Computers, C-20(6):629–638, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671907.

Babu:1972:EPF

[Bab72] C. Chitti Babu. On the extraction of pattern features from imperfectly identified samples. IEEE Transactions on Computers, C-21(4):410–411, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008992. See comments

[Tou74].

Babu:1974:AR

[Bab74] C. C. Babu. Author’s reply. IEEE Transactions on Computers, C-23(9):989, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672663.

Baer:1974:PPD

[Bae74] J.-L. Baer. On program placement in a directly executable hierarchy of memories. IEEE Transactions on Computers, C-23(8):838–849, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672632.

Baer:1976:MS

[Bae76] J.-L. Baer. Multiprocessing systems. IEEE Transactions on Computers, C-25(12):1271– 1277, December 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674594.

Baker:1975:MER

[Bak75a] P. W. Baker. More efficient radix-2 algorithms for some elementary functions. IEEE Transactions on Computers, C-24(11): 1049–1054, November 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672725.

Baker:1975:PMA

[Bak75b] P. W. Baker. Parallel multiplicative algorithms for some elementary functions. IEEE Transactions on Computers, C-24(3): 322–325, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672808.

Baker:1976:SFB

[Bak76] P. W. Baker. Suggestion for a fast binary sine/cosine generator. IEEE Transactions on Computers, C-25(11):1134–1136, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674566.

Baker:1979:SDE

[Bak79] P. W. Baker. The solution of differential equations on shortword-length computing devices. IEEE Transactions on Computers, C-28(3):205–214, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675321.

Bancilhon:1974:GMS

[Ban74a] F. Bancilhon. A geometric model for stochastic automata. IEEE Transactions on Computers, C-23 (12):1290–1299, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672441.

Banerji:1974:NIM

[Ban74b] D. K. Banerji. A novel implementation method for addition and subtraction in residue number systems. IEEE Transactions on Computers, C-23(1):106–109, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=1672383;   http://
                                        ieeexplore.ieee.org/xpl/tocresult.
                                        jsp?isnumber=35077.

Banerji:1974:URA

[Ban74c] D. K. Banerji. On the use of residue arithmetic for computation. IEEE Transactions on Computers, C-23(12):1315–1317, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672448.

Banerji:1978:CBM

[Ban78] R. B. Banerji. The construction of binary matrices with distinct submatrices. IEEE Transactions on Computers, C-27(2): 162–164, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675050.

Bargainer:1970:RUT

[Bar70] J. D. Bargainer. R70-34 unateness test of a Boolean function and two general synthesis methods using threshold logic. IEEE Transactions on Computers, C-19 (9):858, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671655.

Barnes:1971:TWA

[Bar71] B. H. Barnes. A two-way automaton with fewer states than any equivalent one-way automaton. IEEE Transactions on Computers, C-20(4):474–475, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671866.

Barnard:1974:BPD

[Bar74a] A. C. L. Barnard. B 74-33 peripheral devices. IEEE Transactions on Computers, C-23(5):559, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672582. See comments

[Flo75].

Barnard:1974:BCS

[Bar74b] A. C. L. Barnard. B74-18 computer system organization. IEEE Transactions on Computers, C-23(2):220, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672500.

Barnes:1974:BIM

[Bar74c] E. R. Barnes. B74-15 iterative methods for nonlinear optimization problems. IEEE Transactions on Computers, C-23(2):218, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672497.

Barbacci:1975:CRT

[Bar75a] M. R. Barbacci. A comparison of register transfer languages for describing computers and digital systems. IEEE Transactions on Computers, C-24(2):137–150, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672774.

Barnard:1975:RR

[Bar75b] A. C. L. Barnard. Reviewer’s reply. IEEE Transactions on Computers, C-24(2):224, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672790.

Barak:1977:MAT

[Bar77] A. B. Barak. Multiplicative algorithms for ternary arithmetic using binary logic. IEEE Transactions on Computers, C-26(8):823–826, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674922.

Barnes:1978:CIM

[Bar78] G. E. Barnes. Comments on “The Identification of Maximal Parallelism in Straight-Line Microprograms”. IEEE Transactions on Computers, C-27(3): 286–287, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675090. See [DT76,

Das78, Ada78].

Bashkow:1971:BIC

[Bas71a] T. R. Bashkow. B71-2 introduction to computer organization. IEEE Transactions on Computers, C-20(10):1221, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671704.

Bashkow:1971:BSS

[Bas71b] T. R. Bashkow. B71-6 system structure in data, programs, and computers. IEEE Transactions on Computers, C-20(12):1622, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671786.

Bashkow:1973:CRO

[Bas73] T. R. Bashkow. Comment on review of operating systems survey. IEEE Transactions on Computers, C-22(12):1143, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672256.

Batra:1970:DAU

[Bat70] V. Batra. Design of asynchronous unit delays. IEEE Transactions on Computers, C-19(10): 896–902, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671389.

Batcher:1977:MAM

[Bat77] Kenneth E. Batcher. The multidimensional access memory in STARAN. IEEE Transactions on Computers, C-26(2):

174–177, February 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009297.

Baugh:1971:BNP

[Bau71] C. R. Baugh. Bounds on the number of pseudothreshold functions. IEEE Transactions on Computers, C-20(12):1602–1605, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671774.

Baugh:1972:GRF

[Bau72] C. R. Baugh. Generation of representative functions of the NPN equivalence classes of unate Boolean functions. IEEE Transactions on Computers, C-21(12): 1373–1379, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672102.

Bauer:1977:CFP

[Bau77] Henry R. Bauer. Comments on “On the Floating Point Representation of Complex Numbers”. IEEE Transactions on Computers, C-26(2):191, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009301. See [Yue75c].

Banerji:1972:TAR

[BB72] D. K. Banerji and J. A. Brzozowski. On translation algorithms in residue number systems. IEEE Transactions on Computers, C-21(12):1281–1285, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=1672092;   http://
                                        ieeexplore.ieee.org/xpl/tocresult.
                                        jsp?isnumber=35066.

Butler:1973:SCU

[BB73] J. T. Butler and K. J. Breeding. Some characteristics of universal cell nets. IEEE Transactions on Computers, C-22(10): 897–903, October 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672208.

Bracha-Barak:1974:ACF

[BB74a] A. Bracha-Barak. Application of continued fractions for fast evaluation of certain functions on a digital computer. IEEE Transactions on Computers, C-23(3): 301–309, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672519.

Bracha-Barak:1974:MSP

[BB74b] A. Bracha-Barak. A method for solving polynomial equations by continued fractions. IEEE Transactions on Computers, C-23(10):1093–1097, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672402.

Brubaker:1975:MUL

[BB75] T. A. Brubaker and J. C. Becker.

Multiplication using logarithms implemented with read-only memory. IEEE Transactions on Computers, C-24(8):761–765, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672900.

Ben-Bassat:1978:IFP

[BB78a] M. Ben-Bassat. Irrelevant features in pattern recognition. IEEE Transactions on Computers, C-27 (8):746–749, August 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675182.

Ben-Bassat:1978:MPS

[BB78b] M. Ben-Bassat. Myopic policies in sequential classification. IEEE Transactions on Computers, C-27(2):170–174, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675054.

Bender:1978:AAN

[BB78c] E. A. Bender and J. T. Butler. Asymptotic approximations for the number of fanout-free functions. IEEE Transactions on Computers, C-27(12):1180–1183, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675021.

Bandyopadhyay:1972:CPA

[BBC72a] S. Bandyopadhyay, S. Basu, and A. K. Choudhury. A cellular permuter array. IEEE Transactions on Computers, C-21(10): 1116–1119, October 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672052.

Bandyopadhyay:1972:IAM

[BBC72b] S. Bandyopadhyay, S. Basu, and A. K. Choudhury. An iterative array for multiplication of signed binary numbers. IEEE Transactions on Computers, C-21 (8):921–922, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009055.

Barr:1973:ROD

[BBLT73] R. G. Barr, J. A. Becker, W. P. Lidinsky, and V. V. Tantillo. A research-oriented dynamic microprocessor. IEEE Transactions on Computers, C-22(11):

976–985, November 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672227.

Biermann:1975:SSP

[BBP75] A. W. Biermann, R. I. Baum, and F. E. Petry. Speeding up the synthesis of programs from traces. IEEE Transactions on Computers, C-24(2):122–136, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672773.

Babu:1971:APC

[BC71a] C. C. Babu and Wah-Chun Chan.

An algorithm for pattern classification using eigenvectors. IEEE Transactions on Computers, C-20(5):575–578, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671886.

Bearnson:1971:DML

[BC71b] L. W. Bearnson and C. C. Carroll. On the design of minimum length fault tests for combinational circuits. IEEE Transactions on Computers, C-20(11): 1353–1356, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671730.

Benowitz:1975:AFI

[BCA^{+}75] N. Benowitz, D. F. Calhoun, G. E. Alderson, J. E. Bauer, and C. T. Joeckel. An advanced fault isolation system for digital logic. IEEE Transactions on Computers, C-24(5):489–497, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672844.

Bell:1974:IAC

[BCB74] J. Bell, D. Casasent, and C. G. Bell. An investigation of alternative cache organizations. IEEE Transactions on Computers, C-23(4):346–351, April 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672541.

Bossen:1978:MGE

[BCC78] D. C. Bossen, L. C. Chang, and Chin-Long Chen. Measurement and generation of error correcting codes for package failures. IEEE Transactions on Computers, C-27(3):201–204, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675072.

Bouricius:1973:MBM

[BCH^{+}73] W. G. Bouricius, W. C. Carter, E. P. Hsieh, D. C. Jessep, Jr., and A. B. Wadia. Modeling of a bubble-memory organization with self-checking translators to achieve high reliability. IEEE Transactions on Computers, C-22(3):269–275, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672299.

Browne:1973:ETM

[BCHL73] James C. Browne, K. M. Chandy, John Hogarth, and Chester C.- A. Lee. The effect on throughput of multiprocessing in a multiprogramming environment. IEEE Transactions on Computers, C-22 (8):728–735, August 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009151.

Bouricius:1971:RMF

[BCJ^{+}71] W. G. Bouricius, W. C. Carter, D. C. Jessep, P. R. Schneider, and A. B. Wadia. Reliability modeling for fault-tolerant computers. IEEE Transactions on Computers, C-20(11):1306–1311, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671725.

Banerjee:1979:TPP

[BCKT79] Utrpal Banerjee, Shyh-Ching Chen, David J. Kuck, and Ross A. Towle. Time and parallel processor bounds for FORTRAN-like loops. IEEE Transactions on Computers, C-28(9):660–670, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675434.

Borucki:1975:MUC

[BCL75] W. J. Borucki, D. H. Card, and G. C. Lyle. A method of using cluster analysis to study statistical dependence in multivariate data. IEEE Transactions on Computers, C-24(12):1183–1191, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672755.

Breuer:1976:IMS

[BCS76] Melvin A. Breuer, Shih-Jeh Chang, and Stephen Y. H. Su. Identification of multiple stuck-type faults in combinational networks. IEEE Transactions on Computers, C-25(1):44–54, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009204.

Baugh:1972:ONN

[BCSM72] Charles R. Baugh, C. S. Chandersekaran, Richard S. Swee, and Saburo Muroga. Optimal networks of NOR–OR gates for functions of three variables. IEEE Transactions on Computers, C-21 (2):153–160, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008920.

Biss:1976:SMD

[BCSW76] K. Biss, R. T. Chien, F. Stahl, and S. J. Weissman. Semantic modeling for deductive questionanswering. IEEE Transactions on Computers, C-25(4):358–366, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674618.

Bastien:1971:GTP

[BD71] P. L. Bastien and L. A. Dunn.

Global transformations in pattern recognition of bubble chamber photographs. IEEE Transactions on Computers, C-20(9): 995–1001, September 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671987.

Bennett:1975:MCQ

[BD75a] J. R. Bennett and J. S. Mac Donald. On the measurement of curvature in a quantized environment. IEEE Transactions on Computers, C-24(8):803–820, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672905.

Bezdek:1975:OFP

[BD75b] J. C. Bezdek and J. C. Dunn. Optimal fuzzy partitions: a heuristic for estimating the parameters in a mixture of normal distributions. IEEE Transactions on Computers, C-24(8):835–838, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672910.

Briggs:1977:OSM

[BD77] Faye A. Briggs and Edward S. Davidson. Organization of semiconductor memories for parallelpipelined processors. IEEE Transactions on Computers, C-26(2): 162–169, February 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009295.

Brenner:1970:SPD

[BdB70] A. E. Brenner and P. de Bruyne. A sonic pen: a digital stylus system. IEEE Transactions on Computers, C-19(6):546–548, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671566.

Blumenthal:1977:DNP

[BDR77] Andrew F. Blumenthal, Larry S. Davis, and Azriel Rosenfeld. Detecting natural “plateaus” in onedimensional patterns. IEEE Transactions on Computers, C-26(2): 178–179, February 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009298.

Bovet:1970:DMA

[BE70a] D. P. Bovet and G. Estrin. A dynamic memory allocation algorithm. IEEE Transactions on Computers, C-19(5):403–411, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671531.

Bovet:1970:SMA

[BE70b] D. P. Bovet and G. Estrin. On static memory allocation in computer systems. IEEE Transactions on Computers, C-19(6):492–503, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671559.

Beaudry:1978:PRR

[Bea78] M. D. Beaudry. Performancerelated reliability measures for computing systems. IEEE Transactions on Computers, C-27(6):540–547, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675145.

Beck:1971:SFP

[Bec71] F. Beck. The special features of POLLY: How they are used for automatic scanning. IEEE Transactions on Computers, C-20(9):1002–1006, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671988.

Book:1971:AGE

[BEGO71] R. Book, S. Even, S. Greibach, and G. Ott. Ambiguity in graphs and expressions. IEEE Transactions on Computers, C-20(2): 149–153, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671797.

Bell:1972:DUR

[BEGW72] C. G. Bell, J. L. Eggert, J. Grason, and P. Williams. The description and use of register-transfer modules (RTM’s)(R). IEEE Transactions on Computers, C-21(5):495–500, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672141.

Beizer:1970:TNT

[Bei70] B. Beizer. Towards a new theory of sequential switching networks. IEEE Transactions on Computers, C-19(10):939–956, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671394. See correction

[Bei71].

Beizer:1971:CTN

[Bei71] B. Beizer. Correction to “Towards a New Theory of Sequential Switching Networks”. IEEE Transactions on Computers, C-20(8):943, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671972. See [Bei70].

Belser:1976:CIA

[Bel76] Karl Belser. Comment on “An improved algorithm for the generation of nonparametric curves”. IEEE Transactions on Computers, C-25(1):103, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009217. See [JLH73].

Benyon:1970:RCM

[Ben70] P. R. Benyon. R70-25 construction of multistep integration formulas for simulation purposes. IEEE Transactions on Computers, C-19(8):763–764, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671623.

Bennetts:1971:IMP

[Ben71] R. G. Bennetts. An improved method of prime C-class derivation in the state reduction of sequential networks. IEEE Transactions on Computers, C-20(2): 229–231, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671814.

Benedek:1977:DLB

[Ben77] M. Benedek. Developing large binary to BCD conversion structures. IEEE Transactions on Computers, C-26(7):688–700, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674902.

Bergman:1970:RTG

[Ber70a] S. Bergman. R70-27 tree generating regular systems. IEEE Transactions on Computers, C-19(8):764, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671625.

Berndt:1970:FML

[Ber70b] H. Berndt. Functional microprogramming as a logic design aid. IEEE Transactions on Computers, C-19(10):902–907, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671390.

Berkling:1971:CMB

[Ber71] K. J. Berkling. A computing machine based on tree structures. IEEE Transactions on Computers, C-20(4):404–418, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671849.

Bergland:1972:PIF

[Ber72] G. D. Bergland. A parallel implementation of the Fast Fourier Transform algorithm. IEEE Transactions on Computers, C-21(4):366–370, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008977.

Berges:1977:UPN

[Ber77] George A. Berges. Use of postfix notation for the description of microcellular arrays. IEEE Transactions on Computers, C-26(2): 185–191, February 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009300.

Betancourt:1971:DMT

[Bet71] R. Betancourt. Derivation of minimum test sets for unate logical circuits. IEEE Transactions on Computers, C-20(11): 1264–1269, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671719.

Biermann:1972:SFS

[BF72] A. W. Biermann and J. A. Feldman. On the synthesis of finite-state machines from samples of their behavior. IEEE Transactions on Computers, C-21(6):592–597, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009015.

Birk:1975:AMD

[BF75a] J. E. Birk and D. E. Farmer. An algebraic method for designing multivalued logic circuits using principally binary components. IEEE Transactions on Computers, C-24(11):1101–1104, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672731.

Borgerson:1975:RMG

[BF75b] B. R. Borgerson and R. F. Freitas. A reliability model for gracefully degrading and standby-sparing systems. IEEE Transactions on Computers, C-24(5):517–525, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672848.

Bentley:1978:FAC

[BF78] Jon Louis Bentley and Jerome H. Friedman. Fast algorithms for constructing minimal spanning trees in coordinate spaces. IEEE Transactions on Computers, C-27(2): 97–105, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675043.

Bates:1975:NOP

[BG75] R. H. T. Bates and P. T. Gough.

New outlook on processing radiation received from objects viewed through randomly fluctuating media. IEEE Transactions on Computers, C-24(4):449–456, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672834.

Breitbart:1978:AAE

[BG78] Y. Breitbart and S. Gal. Analysis of algorithms for the evaluation of monotonic Boolean functions. IEEE Transactions on Computers, C-27(11):1083–1087, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675002.

Braun:1979:IAD

[BG79] R. D. Braun and D. D. Givone.

An improved algorithm for deriving checking experiments. IEEE Transactions on Computers, C-28 (2):153–156, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675306.

Barsi:1976:TDD

[BGM76] F. Barsi, F. Grandoni, and P. Maestrini. A theory of diagnosability of digital systems. IEEE Transactions on Computers, C-25(6):585–593, June 1976.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674658.

Bossen:1971:CEA

[BH71a] D. C. Bossen and Se June Hong. Cause-effect analysis for multiple fault detection in combinational networks. IEEE Transactions on Computers, C-20(11): 1252–1257, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671717. See comments

[LB74].

Bredeson:1971:GCP

[BH71b] J. G. Bredeson and P. T. Hulina. Generation of a clock pulse for asynchronous sequential machines to eliminate critical races. IEEE Transactions on Computers, C-20 (2):225–226, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671812.

Bredeson:1971:GPI

[BH71c] J. G. Bredeson and P. T. Hultina. Generation of prime implicants by direct multiplication. IEEE Transactions on Computers, C-20(4):475–476, April 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671867.

Bredeson:1973:SMI

[BH73] J. G. Bredeson and P. T. Hulina.

Synthesis of multiple-input change asynchronous circuits using transitionsensitive flip-flops. IEEE Transactions on Computers, C-22(5): 524–531, May 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672349. See comments

[Bre75].

Bossen:1974:AR

[BH74a] D. C. Bossen and S. J. Hong.

Authors’ reply. IEEE Transactions on Computers, C-23(12): 1326–1327, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672453.

Breuer:1974:PES

[BH74b] M. A. Breuer and R. L. Harrison.

Procedures for eliminating static and dynamic hazards in test generation. IEEE Transactions on Computers, C-23(10):1069–1078, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672400. See comments

[Hla78].

Baum:1976:DCS

[BH76] R. I. Baum and D. K. Hsiao.

Database computers — a step towards data utilities. IEEE Transactions on Computers, C-25 (12):1254–1259, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674592.

Bhandarkar:1975:AMI

[Bha75a] D. P. Bhandarkar. Analysis of memory interference in multiprocessors. IEEE Transactions on Computers, C-24(9):897–908, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672928.

Bhandarkar:1975:PMB

[Bha75b] D. P. Bhandarkar. On the performance of magnetic bubble memories in computer systems. IEEE Transactions on Computers, C-24 (11):1125–1129, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672738.

Bhandarkar:1977:SPI

[Bha77] D. P. Bhandarkar. Some performance issues in multiprocessor system design. IEEE Transactions on Computers, C-26(5):506–511, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674868.

Bhargava:1978:SCR

[Bha78] V. K. Bhargava. Some codes of Rahman and Blake for computer applications. IEEE Transactions on Computers, C-27(8):765–767, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675188.

Banerjee:1979:DDC

[BHK79] J. Banerjee, D. K. Hsiao, and K. Kannan. DBC — a database computer for very large databases. IEEE Transactions on Computers, C-28(6):414–429, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675381.

Bouricius:1971:ADF

[BHP^{+}71] W. G. Bouricius, E. P. Hsieh, G. R.

Putzolu, J. P. Roth, P. R. Schneider, and C. Tan. Algorithms for detection of faults in logic circuits. IEEE Transactions on Computers, C-20(11):1258–1264, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671718.

Beltrami:1972:ARS

[BI72] E. J. Beltrami and J. P. Indusi.

An adaptive random search algorithm for constrained minimization. IEEE Transactions on Computers, C-21(9):1004–1008, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009077.

Biberman:1974:BPI

[Bib74] L. M. Biberman. B74-12 principles of interactive computer graphics. IEEE Transactions on Computers, C-23(2):216, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672494. See comments

[New75a].

Biberman:1975:RR

[Bib75] L. M. Biberman. Reviewer’s reply. IEEE Transactions on Computers, C-24(3):336, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672813.

Biswas:1970:ITS

[Bis70] N. N. Biswas. On identification of totally symmetric Boolean functions. IEEE Transactions on Computers, C-19(7):645–648, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671590. See comments

[Bis73].

Biswas:1971:MBF

[Bis71] N. N. Biswas. Minimization of Boolean functions. IEEE Transactions on Computers, C-20(8): 925–929, August 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671966.

Biswas:1973:CIT

[Bis73] Nripendra N. Biswas. Comments on “Identification of Totally Symmetric Boolean Functions”. IEEE Transactions on Computers, C-22(9):863–864, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009180. See [Bis70].

Biswas:1974:SMI

[Bis74] N. N. Biswas. State minimization of incompletely specified sequential machines. IEEE Transactions on Computers, C-23(1):80–84, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672373.

Bitner:1977:CED

[Bit77] J. R. Bitner. Capacity and efficiency of decision functions. IEEE Transactions on Computers, C-26(11):1147–1151, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674763.

Basu:1975:SPN

[BJ75] D. Basu and T. Jayashree. On a simple postcorrection for nonrestoring division. IEEE Transactions on Computers, C-24(10): 1019–1020, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672709.

Berg:1971:SAO

[BK71a] R. O. Berg and L. L. Kinney. Serial adders with overflow correction. IEEE Transactions on Computers, C-20(6):668–671, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671914.

Budnik:1971:OUP

[BK71b] P. Budnik and D. J. Kuck. The organization and use of parallel memories. IEEE Transactions on Computers, C-20(12): 1566–1569, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671764.

Bassett:1972:IPD

[BK72] James C. Bassett and Charles R. Kime. Improved procedures for determining diagnostic resolution. IEEE Transactions on Computers, C-21(4):385–388, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008980.

Berger:1973:FDF

[BK73] I. Berger and Z. Kohavi. Fault detection in fanout-free combinational networks. IEEE Transactions on Computers, C-22(10): 908–914, October 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672210.

Batni:1976:MLT

[BK76] R. P. Batni and C. R. Kime. A module-level testing approach for combinational networks. IEEE Transactions on Computers, C-25(6):594–604, June 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674659.

Baer:1979:MWC

[BK79] J.-L. Baer and B. Koyama. On the minimization of the width of the control memory of microprogrammed processors. IEEE Transactions on Computers, C-28(4):310–316, April 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675352.

Brent:1973:PEA

[BKM73] R. Brent, D. Kuck, and K. Maruyama.

The parallel evaluation of arithmetic expressions without division. IEEE Transactions on Computers, C-22(5):532–534, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672350.

Butter:1975:THR

[BL75] C. D. Butter and T. C. Lee. Thermoplastic holographic recording of binary patterns in PLZT line composer. IEEE Transactions on Computers, C-24(4):402–406, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672827.

Blankenship:1974:CTC

[Bla74] P. E. Blankenship. Comments on “A Two’s Complement Parallel Array Multiplication Algorithm”. IEEE Transactions on Computers, C-23(12):1327, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672455. See [BW73,

Kro74].

Blankenbaker:1979:CIP

[Bla79] J. V. Blankenbaker. Comments on “Inner Product Computers”. IEEE Transactions on Computers, C-28(12):944, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675288. See [SGR78].

Bongiovanni:1975:DES

[BLZ75] G. Bongiovanni, F. Luccio, and A. Zorat. The discrete equation of the straight line. IEEE Transactions on Computers, C-24 (3):310–313, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672805.

Bowman:1970:MFA

[BM70] R. M. Bowman and E. S. McVey.

A method for the fast approximate solution of large prime implicant charts. IEEE Transactions on Computers, C-19(2):169–173, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671481.

Barsi:1973:ECP

[BM73] F. Barsi and P. Maestrini. Error correcting properties of redundant residue number systems. IEEE Transactions on Computers, C-22(3):307–315, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1672304;   http://
ieeexplore.ieee.org/xpl/tocresult.
jsp?isnumber=35074.

Barsi:1974:EDC

[BM74] F. Barsi and P. Maestrini. Error detection and correction by product codes in residue number systems. IEEE Transactions on Computers, C-23(9):915–924, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1672648;   http://
ieeexplore.ieee.org/xpl/tocresult.
jsp?isnumber=35088.

Bate:1977:TCS

[BM77a] J. A. Bate and J. C. Muzio. Three cell structures for ternary cellular arrays. IEEE Transactions on Computers, C-26(12): 1191–1202, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674780.

Bivins:1977:SAA

[BM77b] Robert L. Bivins and Nicholas C. Metropolis. Significance arithmetic: Application to a partial differential equation. IEEE Transactions on Computers, C-26(7):639– 642, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/12/
                                        35159/01674896.pdf?tp=&isnumber=
                                        35159&arnumber=1674896&punumber=
                                        12;   http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1674896.

Barsi:1978:ACR

[BM78] F. Barsi and P. Maestrini. Arithmetic codes in residue number systems with magnitude index. IEEE Transactions on Computers, C-27(12):1185–1188, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=1675023;   http://
                                        ieeexplore.ieee.org/xpl/RecentIssue.
                                        jsp?punumber=12;   http://ieeexplore.
                                        ieee.org/xpl/tocresult.jsp?isnumber=
                                        35165.

Bohling:1970:ICA

[BO70] Dorothea M. Bohling and Lawrence A.

O’Neill. An interactive computer approach to tolerance analysis. IEEE Transactions on Computers, C-19(1):10–16, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008894.

Bentley:1979:ARC

[BO79] J. L. Bentley and T. A. Ottmann.

Algorithms for reporting and counting geometric intersections. IEEE Transactions on Computers, C-28(9):643–647, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675432.

Boehm:1976:SE

[Boe76] B. W. Boehm. Software engineering. IEEE Transactions on Computers, C-25(12):1226–1241, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674590.

Bohlender:1977:FPC

[Boh77] Gerd Bohlender. Floating-point computation of functions with maximum accuracy. IEEE Transactions on Computers, C-26(7):621–632, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674894.

Bonn:1974:BMC

[Bon74] T. H. Bonn. B74-2 man and the computer. IEEE Transactions on Computers, C-23(1):110, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672385.

Booth:1970:EPS

[Boo70] T. L. Booth. Estimation, prediction, and smoothing in discrete parameter systems. IEEE Transactions on Computers, C-19 (12):1193–1203, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671451.

Born:1972:ITD

[Bor72] R. C. Born. An iterative technique for determining the minimal number of variables for a totally symmetric function with repeated variables. IEEE Transactions on Computers, C-21(10):1129–1131, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672055.

Bose:1979:CMF

[Bos79] B. Bose. Comments on “Multiple Fault Detection in Combinational Network”. IEEE Transactions on Computers, C-28(10):

804–805, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675254. See [DW73b,

GS74].

Boute:1974:DSO

[Bou74a] R. T. Boute. Distinguishing sets for optimal state identification in checking experiments. IEEE Transactions on Computers, C-23(8):874–877, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672636.

Boute:1974:ONO

[Bou74b] R. T. Boute. Optimal and near-optimal checking experiments for output faults in sequential machines. IEEE Transactions on Computers, C-23(11):1207– 1213, November 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672425.

Boyes:1974:BNS

[Boy74] J. D. Boyes. Binary noise sources incorporating modulo-N dividers. IEEE Transactions on Computers, C-23(5):550–552, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672573.

Brown:1972:MPR

[BP72] W. G. S. Brown and E. A. Parrish. A model for pattern recognition systems with binary pattern vectors. IEEE Transactions on Computers, C-21(2):219, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008931.

Bandyopadhyay:1975:CUC

[BPC75] S. Bandyopadhyay, A. Pal, and A. K. Choudhury. Characterization of unate cascade realizability using parameters. IEEE Transactions on Computers, C-24(2): 218–219, February 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672784.

Breeding:1971:BRC

[BR71] K. J. Breeding and R. G. Rich. BCD and radix-2 conversions. IEEE Transactions on Computers, C-20(3):360–361, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671838.

Butler:1974:EMI

[BR74] G. A. Butler and H. B. Ritea. Estimation of mutual information in two-class pattern recognition. IEEE Transactions on Computers, C-23(4):410–420, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672549.

Benjauthrit:1976:GSF

[BR76a] Boonsieng Benjauthrit and Irving S. Reed. Galois switching functions and their applications. IEEE Transactions on Computers, C-25(1):78–86, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009207.

Blevins:1976:ADA

[BR76b] P. R. Blevins and C. V. Ramamoorthy. Aspects of a dynamically adaptive operating system. IEEE Transactions on Computers, C-25(7):713–725, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674680.

Brafmen:1977:IC

[BR77] H. Brafmen and B. J. Reuter.

An incremental computer. IEEE Transactions on Computers, C-26 (11):1072–1081, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674753.

Benjauthrit:1978:FSG

[BR78] B. Benjauthrit and I. S. Reed. On the fundamental structure of Galois switching functions. IEEE Transactions on Computers, C-27 (8):757–762, August 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675185.

Brent:1970:ABN

[Bre70a] R. Brent. On the addition of binary numbers. IEEE Transactions on Computers, C-19(8):758–759, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671620;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Brent.pdf.

Breuer:1970:FPS

[Bre70b] M. A. Breuer. Functional partitioning and simulation of digital circuits. IEEE Transactions on Computers, C-19(11): 1038–1046, November 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671423.

Bredt:1971:APS

[Bre71a] T. H. Bredt. Analysis of parallel systems. IEEE Transactions on Computers, C-20(11): 1403–1407, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671739.

Breeding:1971:SPM

[Bre71b] K. J. Breeding. Some properties of minimal threshold approximations. IEEE Transactions on Computers, C-20(12):1544–1551, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671760.

Brenner:1971:BMP

[Bre71c] M. Brenner. B71-4 mathematical programming in practice. IEEE Transactions on Computers, C-20 (10):1222, October 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671706.

Breuer:1971:RAT

[Bre71d] M. A. Breuer. A random and an algorithmic technique for fault detection test generation for sequential circuits. IEEE Transactions on Computers, C-20(11): 1364–1370, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671733.

Breuer:1972:GFT

[Bre72a] M. A. Breuer. Generation of fault tests for linear logic networks. IEEE Transactions on Computers, C-21(1):79–83, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672026.

Breuer:1972:NTV

[Bre72b] Melvin A. Breuer. A note on three-valued logic simulation. IEEE Transactions on Computers, C-21(4):399–402, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008985.

Breeding:1973:BCO

[Bre73a] K. J. Breeding. B73-5 computeroriented approaches to pattern recognition. IEEE Transactions on Computers, C-22(4):429, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672330.

Brent:1973:PAV

[Bre73b] Richard P. Brent. On the precision attainable with various floatingpoint number systems. IEEE Transactions on Computers, C-22(6):601–607, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=5009113;   http://
                                        www.acsel-lab.com/arithmetic/
                                        arith2/papers/ARITH2_Brent.pdf.

Breuer:1973:TIF

[Bre73c] M. A. Breuer. Testing for intermittent faults in digital circuits. IEEE Transactions on Computers, C-22(3):241–246, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672294.

Breuer:1974:ERD

[Bre74] M. A. Breuer. The effects of races, delays, and delay faults on test generation. IEEE Transactions on Computers, C-23(10): 1078–1092, October 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672401.

Bredeson:1975:CSM

[Bre75] J. Bredeson. Comments on “Synthesis of Multiple-Input Change Asynchronous Machines Using Controlled Excitation and Flip-Flops”. IEEE Transactions on Computers, C-24(11):1142–1143, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672744. See [BH73].

Briley:1973:SNR

[Bri73] B. E. Briley. Some new results on average worst case carry. IEEE Transactions on Computers, C-22(5):459–463, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672341. See comment

[Yue74a].

Brown:1970:CDM

[Bro70a] F. M. Brown. Comment on “The Determination of the Maximum Compatibility Classes”. IEEE Transactions on Computers, C-19(5):459, May 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671541.

Brown:1970:RSB

[Bro70b] F. M. Brown. Reduced solutions of Boolean equations. IEEE Transactions on Computers, C-19 (10):976–981, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671399.

Brown:1971:SPS

[Bro71] F. M. Brown. Single-parameter solutions for flip-flop equations. IEEE Transactions on Computers, C-20(4):452–454, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671856.

Brown:1974:EL

[Bro74] F. M. Brown. Equational logic. IEEE Transactions on Computers, C-23(12):1228–1237, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&

arnumber=1672434. See comments

[Cer76].

Brown:1975:CIP

[Bro75a] F. M. Brown. The constrainedinput problem. IEEE Transactions on Computers, C-24(1): 102–106, January 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672682.

Brown:1975:ERS

[Bro75b] F. M. Brown. Equational realizations of switching functions. IEEE Transactions on Computers, C-24(11):1054–1066, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672726.

Brown:1975:WRS

[Bro75c] F. M. Brown. Weighted realizations of switching functions. IEEE Transactions on Computers, C-24(12):1217–1221, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672759.

Brown:1977:RAS

[Bro77] R. D. Brown. A recursive algorithm for sequency-ordered fast Walsh transforms. IEEE Transactions on Computers, C-26(8):819–822, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674921.

Brzozowski:1970:RSG

[Brz70] J. A. Brzozowski. R70-44 synchronization and general repetitive machines, with applications to ultimate definite automata. IEEE Transactions on Computers, C-19(10):989–990, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671410.

Brzozowski:1971:AFS

[Brz71] J. A. Brzozowski. About feedback and SR flip-flops. IEEE Transactions on Computers, C-20(4):476, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671868.

Barnea:1972:CAF

[BS72] Daniel I. Barnea and Harvey F. Silverman. A class of algorithms for fast digital image registration. IEEE Transactions on Computers, C-21(2):179–186, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008923.

Baskett:1974:BOS

[BS74] F. Baskett and R. Sedgewick. B74-38 operating systems theory. IEEE Transactions on Computers, C-23(8):879–880, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672639.

Bernstein:1975:CAL

[BS75] A. J. Bernstein and P. Siegel.

A computer architecture for level structured systems. IEEE Transactions on Computers, C-24(8): 785–793, August 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672903.

Ballard:1976:LSD

[BS76] D. H. Ballard and J. Sklansky. A ladder-structured decision tree for recognizing tumors in chest radiographs. IEEE Transactions on Computers, C-25(5):503–513, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674638.

Birta:1977:MIA

[BS77] Louis G. Birta and Ratilal R.

Shah. The minimization of inverters in analog computer programming. IEEE Transactions on Computers, C-26(1):2–12, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009267.

Baudet:1978:OSA

[BS78] G. Baudet and D. Stevenson. Optimal sorting algorithms for parallel computers. IEEE Transactions on Computers, C-27(1):84–87, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674957.

Beakley:1970:CNP

[BT70a] G. W. Beakley and F. B. Tuteur. Comments on “A Nonparametric Partitioning Procedure for Pattern Classification”. IEEE Transactions on Computers, C-19(4):362–363, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671520.

Beuscher:1970:CSI

[BT70b] H. J. Beuscher and W. N. Toy. Check schemes for integrated microprogrammed control and data transfer circuitry. IEEE Transactions on Computers, C-19(12): 1153–1159, December 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671447.

Beakley:1972:DFP

[BT72] G. W. Beakley and F. B. Tuteur. Distribution-free pattern verification using statistically equivalent blocks. IEEE Transactions on Computers, C-21(12): 1337–1347, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672098.

Booth:1973:APM

[BT73] T. L. Booth and R. A. Thompson. Applying probability measures to abstract languages. IEEE Transactions on Computers, C-22(5):442–450, May 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672339.

Bednar:1974:ACD

[BT74] G. M. Bednar and J. H. Tracey.

An Asynchronous Circuit Design Language (ACDL). IEEE Transactions on Computers, C-23(9): 971–976, September 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672657.

Bowra:1976:MDM

[BT76] James W. Bowra and H. C.

Torng. The modeling and design of multiple function-unit processors. IEEE Transactions on Computers, C-25(3):210–221, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009247.

Bubenik:1972:WMD

[Bub72] V. Bubenik. Weighting method for the determination of the irredundant set of prime implicants.

IEEE Transactions on Computers, C-21(12):1449–1451, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672116. See comments

[NRS74].

Burhardt:1974:ASO

[Bur74] K. K. Burhardt. An adaptive search optimization algorithm. IEEE Transactions on Computers, C-23(9):890–897, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672645.

Butz:1971:AAH

[But71] A. R. Butz. Alternative algorithm for Hilbert’s space-filling curve. IEEE Transactions on Computers, C-20(4):424–426, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671851.

Butin:1972:CDW

[But72] H. Butin. A compact definition of Walsh functions. IEEE Transactions on Computers, C-21(6):590–592, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009014.

Butler:1975:NFR

[But75] J. T. Butler. On the number of functions realized by cascades and disjunctive networks. IEEE Transactions on Computers, C-24(7):681–690, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672881.

Butler:1976:RCN

[But76] J. T. Butler. Restricted cellular networks. IEEE Transactions on Computers, C-25(11): 1139–1142, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674568.

Butin:1978:OWF

[But78a] H. Butin. On an ordering of Walsh functions. IEEE Transactions on Computers, C-27(1):87–90, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674958.

Butler:1978:TNU

[But78b] J. T. Butler. Tandem networks of universal cells. IEEE Transactions on Computers, C-27(9): 785–799, September 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675199.

Butz:1979:FLD

[But79] A. R. Butz. FFT length in digital filtering. IEEE Transactions on Computers, C-28(8):577–580, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675414.

Buzbee:1973:FPS

[Buz73] Billy L. Buzbee. A fast Poisson solver amenable to parallel computation. IEEE Transactions on Computers, C-22(8):793–796, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009160.

Bennett:1978:CSA

[BV78] J. R. Bennett and W. Vogel. On the characteristics of a simple architecture for finite impulse response digital filtering. IEEE Transactions on Computers, C-27 (12):1197–1202, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675027.

Breitbart:1979:CCC

[BV79] Y. Breitbart and K. Vairavan. The computational complexity of a class of minimization algorithms for switching functions. IEEE Transactions on Computers, C-28(12):941–943, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675286.

Burton:1970:CDF

[BW70] D. P. Burton and M. C. Waters.

Comment on “Delay-Free Asynchronous Circuits with Constrained Line Delays”. IEEE Transactions on Computers, C-19(10):982, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671400.

Baugh:1973:TCP

[BW73] C. R. Baugh and B. A. Wooley. A two’s complement parallel array multiplication algorithm. IEEE Transactions on Computers, C-22(12):1045–1047, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672241. See comments

[Bla74, Kro74].

Baugh:1976:SAD

[BW76a] C. R. Baugh and B. A. Wooley.

Statistical analysis of a differential threshold logic circuit configuration. IEEE Transactions on Computers, C-25(7):745–754, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674683.

Bobrow:1976:MCS

[BW76b] D. G. Bobrow and B. Wegbreit.

A model for control structures for artificial intelligence programming languages. IEEE Transactions on Computers, C-25(4):347–353, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674616.

Brandsma:1977:HST

[BW77] J. R. Brandsma and B. L. A. Waumans. The hardware simulator: a tool for evaluating computer systems. IEEE Transactions on Computers, C-26(1):68–72, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009275.

Brzozowski:1979:TMG

[BY79] J. A. Brzozowski and M. Yoeli. On a ternary model of gate networks. IEEE Transactions on Computers, C-28(3):178–184, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675317.

Byrd:1971:VMC

[Byr71] J. S. Byrd. Variable-mode counting with straight binary counters. IEEE Transactions on Computers, C-20(1):97–98, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671679.

Chen:1974:NID

[CA74] Chiu Kuan Chen and H. C. Andrews. Nonlinear intrinsic dimensionality computations. IEEE Transactions on Computers, C-23 (2):178–184, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672475.

Chow:1977:CAP

[CA77] W. K. Chow and J. K. Aggarwal. Computer analysis of planar curvilinear moving images. IEEE Transactions on Computers, C-26(2):179–185, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009299.

Cairns:1971:FFT

[Cai71] T. W. Cairns. On the Fast Fourier Transform on finite Abelian groups. IEEE Transactions on Computers, C-20(5):569–571, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671884.

Calvert:1970:NPF

[Cal70] T. W. Calvert. Nonorthogonal projections for feature extraction in pattern recognition. IEEE Transactions on Computers, C-19(5):447–452, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671536.

Campbell:1972:BSJ

[Cam72] S. Campbell. B72-4 System/360 Job Control Language. IEEE Transactions on Computers, C-21 (1):108–109, January 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672035.

Cantoni:1971:OCF

[Can71] A. Cantoni. Optimal curve fitting with piecewise linear functions. IEEE Transactions on Computers, C-20(1):59–67, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671675.

Capowski:1976:MTP

[Cap76] J. J. Capowski. The matrix transform processor. IEEE Transactions on Computers, C-25(7):703–712, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674679.

Carl:1971:CAF

[Car71a] J. W. Carl. Comment on “An Algorithm for a Fast Hadamard Matrix Transform of Order Twelve”. IEEE Transactions on Computers, C-20(5):587–588, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671896.

Carl:1971:CSD

[Car71b] J. W. Carl. Comments on “A Simplified Definition of Walsh Functions”. IEEE Transactions on Computers, C-20(12):1617, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671782. See [LM71b,

KS72].

Carter:1973:FTC

[Car73] W. C. Carter. Fault-tolerant computing: An introduction and a viewpoint. IEEE Transactions on Computers, C-22(3): 225–229, March 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672291.

Casasent:1973:HDO

[Cas73] David P. Casasent. A hybrid digital/optical computer system. IEEE Transactions on Computers, C-22(9):852–858, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009178.

Catt:1977:CMD

[Cat77] I. Catt. Correction to “Maxwell’s Displacement Current”. IEEE Transactions on Computers, C-26 (11):1180, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674773. See [Ano77v].

Carter:1972:SDE

[CB72] W. H. Carter and V. D. Brown. A signal-dependent error arising in digitally processed images due to quantization. IEEE Transactions on Computers, C-21(12): 1380–1385, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672103.

Choudhury:1970:AR

[CBD70] A. K. Choudhury, A. K. Basu, and S. C. DeSarkar. Authors’ reply. IEEE Transactions on Computers, C-19(5):459, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671542.

Coffman:1971:PIM

[CBS71] E. G. Coffman, Jr., G. J. Burnett, and R. A. Snowdon. On the performance of interleaved memories with multiple-word bandwidths. IEEE Transactions on Computers, C-20(12):1570–1573, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671765.

Carroll:1972:MTL

[CC72] B. D. Carroll and C. L. Coates, Jr. Minimum two-level threshold gate realizations. IEEE Transactions on Computers, C-21(10): 1086–1098, October 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672048.

Chang:1973:RTD

[CC73] Shi-Kuo Chang and C. K. Chow.

The reconstruction of three-dimensional objects from two orthogonal projections and its application to cardiac cineangiography. IEEE Transactions on Computers, C-22 (1):18–28, January 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672189.

Chakrabarti:1970:CFA

[CCB70] K. K. Chakrabarti, A. K. Choudhury, and M. S. Basu. Complementary function approach to the synthesis of three-level NAND network. IEEE Transactions on Computers, C-19(6):509–514, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671561.

Chang:1974:CPD

[CCES74] H. Y.-P. Chang, S. G. Chappell, C. H. Elmendorf, and L. D. Schmidt. Comparison of parallel and deductive fault simulation methods. IEEE Transactions on Computers, C-23(11): 1132–1138, November 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672413.

Chen:1979:APN

[CCF79] I-Ngo Chen, P. Y. Chen, and Tse-Yun Feng. Associative processing of network flow problems. IEEE Transactions on Computers, C-28(3):184–190, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675318.

Chuang:1973:SMI

[CD73] H. Y. H. Chuang and S. Das. Synthesis of multiple-input change asynchronous machines using controlled excitation and flip-flops. IEEE Transactions on Computers, C-22(12):1103–1109, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672249.

Chuang:1975:AR

[CD75] H. Y. H. Chuang and S. Das. Authors’ reply. IEEE Transactions on Computers, C-24(11): 1143–1144, November 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672745.

Chuang:1978:DFS

[CD78] H. Y. H. Chuang and S. Das.

Design of fail-safe sequential machines using separable codes. IEEE Transactions on Computers, C-27 (3):249–252, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675078.

Carter:1971:SST

[CDJ71] W. C. Carter, K. A. Duke, and D. C. Jessep. A simple selftesting decoder checking circuit. IEEE Transactions on Computers, C-20(11):1413–1414, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671741.

Carter:1973:LTM

[CDJ73] W. C. Carter, K. A. Duke, and D. C. Jessep, Jr. Lookaside techniques for minimum circuit memory translators. IEEE Transactions on Computers, C-22 (3):283–289, March 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672301.

Cordella:1978:ACC

[CDL78] L. P. Cordella, M. J. B. Duff, and S. Levialdi. An analysis of computational cost in image processing: a case study. IEEE Transactions on Computers, C-27(10):904–910, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674969.

Cha:1978:VAT

[CDO78] C. W. Cha, W. E. Donath, and F. Ozguner. 9-V algorithm for test pattern generation of combinational digital circuits. IEEE Transactions on Computers, C-27 (3):193–200, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675071.

Chang:1973:DMS

[CDS73] H. Y.-P. Chang, R. C. Dorr, and D. J. Senese. The design of a microprogrammed self-checking processor of an electronic switching system. IEEE Transactions on Computers, C-22(5):489–500, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672345.

Cheung:1975:MTT

[CE75] C. C. Cheung and R. W. Ehrich. On the minimization of tree-type universal logic circuits. IEEE Transactions on Computers, C-24 (11):1110–1113, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672734.

Cerny:1976:CEL

[Cer76] E. Cerny. Comments on “Equational Logic”. IEEE Transactions on Computers, C-25(1): 102–103, January 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009215. See [Bro74].

Cerny:1978:CFO

[Cer78] E. Cerny. Controllability and fault observability in modular combinational circuits. IEEE Transactions on Computers, C-27(10): 896–903, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674968.

Cook:1970:SDD

[CF70a] R. W. Cook and M. J. Flynn. System design of a dynamic microprocessor. IEEE Transactions on Computers, C-19(3):213–222, March 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671492.

Cooper:1970:AIO

[CF70b] D. B. Cooper and J. H. Freeman. On the asymptotic improvement in the outcome of supervised learning provided by additional nonsupervised learning.

IEEE Transactions on Computers, C-19(11):1055–1063, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671425.

Cook:1973:LNC

[CF73] Robert W. Cook and Michael J. Flynn. Logical network cost and entropy. IEEE Transactions on Computers, C-22(9):823–826, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009171. See comments

[Mas78].

Censier:1978:NSC

[CF78a] L. M. Censier and P. Feautrier. A new solution to coherence problems in multicache systems. IEEE Transactions on Computers, C-27 (12):1112–1118, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675013.

Corsini:1978:USM

[CF78b] P. Corsini and G. Frosini. Uniform shift multiplication algorithms without overflow. IEEE Transactions on Computers, C-27 (3):256–258, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675081.

Corsini:1979:PMG

[CF79] P. Corsini and G. Frosini. Properties of the multidimensional generalized discrete Fourier transform. IEEE Transactions on Computers, C-28(11):819–830, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675262.

Cantor:1974:ORP

[CG74] D. G. Cantor and M. Gerla. Optimal routing in a packet-switched computer network. IEEE Transactions on Computers, C-23(10): 1062–1069, October 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672399.

Carl:1972:AFT

[CH72a] J. W. Carl and C. F. Hall.

The application of filtered transforms to the general classification problem. IEEE Transactions on Computers, C-21(7):785–790, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672175. See comments

[CH73b].

Chien:1972:ECH

[CH72b] R. T. Chien and Se June Hong. Error correction in high-speed arithmetic. IEEE Transactions on Computers, C-21(5):433–438, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672131.

Cappa:1973:AIA

[CH73a] M. Cappa and V. C. Hamacher. An augmented iterative array for high-speed binary division. IEEE Transactions on Computers, C-22 (2):172–175, February 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672273. See comments

[Gar74].

Carl:1973:CAF

[CH73b] J. W. Carl and C. F. Hall. Comments on “The Application of Filtered Transforms to the General Classification Problem”. IEEE Transactions on Computers, C-22(1):111, January 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672198. See [CH72a].

Casasent:1975:ISI

[CH75] D. Casasent and S. Horvitz. Introduction to the special issue on optical computing. IEEE Transactions on Computers, C-24(4):337–339, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672819.

Chang:1970:RDU

[Cha70] H. Y. Chang. R70-41 diagnosis and utilization of faulty universal tree circuits. IEEE Transactions on Computers, C-19(10): 988–989, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671407.

Chang:1971:ARM

[Cha71] Chin Liang Chang. The accelerated relaxation method for linear inequalities. IEEE Transactions on Computers, C-20(2): 222–225, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671811.

Chang:1972:EFP

[Cha72] Shi-Kuo Chang. On the execution of fuzzy programs using finite-state machines. IEEE Transactions on Computers, C-21 (3):241–253, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008946.

Chang:1973:PRP

[Cha73a] Chin-Liang Chang. Pattern recognition by piecewise linear discriminant functions. IEEE Transactions on Computers, C-22(9): 859–862, September 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009179.

Chang:1973:CWO

[Cha73b] Shi-Kuo Chang. The computation of window operations on a parallel organized computer — a case study. IEEE Transactions on Computers, C-22(1):34–40, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672191.

Chakrabarti:1974:DCA

[Cha74a] S. Chakrabarti. Divide-and-correct algorithm for division in a negative base. IEEE Transactions on Computers, C-23(9):981–983, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672659.

Chang:1974:FPN

[Cha74b] Chin-Liang Chang. Finding prototypes for nearest neighbor classifiers. IEEE Transactions on Computers, C-23(11):1179– 1184, November 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672420.

Chang:1974:BDA

[Cha74c] H. Y. Chang. B74-3 design automation of digital systems — volume 1, theory and techniques.

IEEE Transactions on Computers, C-23(1):110–111, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672386.

Chang:1974:PBB

[Cha74d] Shi-Kuo Chang. Parallel balancing of binary search trees. IEEE Transactions on Computers, C-23(4):441–444, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672552.

Chaney:1979:CNS

[Cha79] T. J. Chaney. Comments on “A Note on Synchronizer or Interlock Maloperation”. IEEE Transactions on Computers, C-28(10): 802–804, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675252. See [Wor77,

Wor79].

Chen:1971:TCC

[Che71a] C. H. Chen. Theoretical comparison of a class of feature selection criteria in pattern recognition. IEEE Transactions on Computers, C-20(9):1054–1056, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671995. See comments

[Tou72c].

Chen:1971:BMS

[Che71b] Tien Chi Chen. A binary multiplication scheme based on squaring. IEEE Transactions on Computers, C-20(6):678–680, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671918.

Chen:1972:AR

[Che72] C. H. Chen. Author’s reply. IEEE Transactions on Computers, C-21(6):616, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009025.

Chen:1973:NMI

[Che73] R. M.-M. Chen. New matrix inversion algorithms based on exchange method. IEEE Transactions on Computers, C-22(10): 885–890, October 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672206.

Chen:1974:MLS

[Che74] Chi-Tsong Chen. Minimization of linear sequential machines. IEEE Transactions on Computers, C-23 (1):93–95, January 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672378.

Chen:1978:GPM

[Che78] C. L. Chen. On the generation of permutations in magnetic bubble memories. IEEE Transactions on Computers, C-27 (12):1183–1185, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675022.

Chien:1971:SDM

[Chi71] Yi-Tzuu Chien. A sequential decision model for selecting feature subsets in pattern recognition. IEEE Transactions on Computers, C-20(3):282–290, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671825.

Chinal:1972:SCP

[Chi72] J. P. Chinal. Some comments on postcorrections for nonrestoring division. IEEE Transactions on Computers, C-21(12): 1385–1394, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672104.

Chinal:1977:BFS

[Chi77] Jean P. Chinal. Boolean features of sign-flipping and signed-carry logic in base-2. IEEE Transactions on Computers, C-26(1):82–86, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009278.

Chiba:1978:ILH

[Chi78] T. Chiba. Impact of the LSI on high-speed computer packaging. IEEE Transactions on Computers, C-27(4):319–325, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675104.

Choueka:1974:SA

[Cho74] Y. A. Choueka. Structure automata. IEEE Transactions on Computers, C-23(12):1218– 1227, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672433.

Chow:1976:DCC

[Cho76] C. K. Chow. Determination of cache’s capacity and its matching storage hierarchy. IEEE Transactions on Computers, C-25(2): 157–164, February 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009230.

Chang:1973:MTM

[CHSS73] H. Y. P. Chang, G. W. Heimbigner, D. J. Senese, and T. L. Smith. Maintenance techniques of a microprogrammed self-checking control complex of an electronic switching system. IEEE Transactions on Computers, C-22(5):501–512, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672346.

Chu:1970:BBP

[Chu70] W. W. Chu. Buffer behavior for Poisson arrivals and multiple synchronous constant outputs. IEEE Transactions on Computers, C-19(6):530–534, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671563.

Chu:1971:EBC

[Chu71] J. T. Chu. Error bounds for a contextual recognition procedure. IEEE Transactions on Computers, C-20(10):1203–1207, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671699. See comments

[Tou72b].

Chu:1972:AR

[Chu72] John T. Chu. Author’s reply.

IEEE Transactions on Computers, C-21(9):1027–1028, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009088.

Chu:1974:SNE

[Chu74] J. T. Chu. Some new error bounds and approximations for pattern recognition. IEEE Transactions on Computers, C-23(2): 194–199, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672480.

Chuang:1976:FSA

[Chu76] H. Y. H. Chuang. Fail-safe asynchronous machines with multipleinput changes. IEEE Transactions on Computers, C-25(6):637–642, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674664.

Chu:1979:AHD

[Chu79] Yaohan Chu. Architecture of a hardware data interpreter. IEEE Transactions on Computers, C-28 (2):101–109, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675299.

Cioffi:1978:ASI

[Cio78] G. Cioffi. Autotesting speedindependent sequential circuits. IEEE Transactions on Computers, C-27(1):90–94, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674959.

Chandrasekaran:1974:QCI

[CJ74] B. Chandrasekaran and A. K.

Jain. Quantization complexity and independent measurements. IEEE Transactions on Computers, C-23(1):102–106, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672382.

Carter:1971:LDD

[CJW^{+}71] W. C. Carter, D. C. Jessep, A. B. Wadia, P. R. Schneider, and W. G. Bouricius. Logic design for dynamic and interactive recovery. IEEE Transactions on Computers, C-20(11):1300–1305, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671724.

Chakrabarti:1974:FCT

[CK74] K. Chakrabarti and O. Kolp.

Fan-in constrained tree networks of flexible cells. IEEE Transactions on Computers, C-23(12): 1238–1249, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672435.

Chen:1975:TPP

[CK75] Shyh-Ching Chen and D. J.

Kuck. Time and parallel processor bounds for linear recurrence systems. IEEE Transactions on Computers, C-24(7):701–717, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672884.

Chen:1977:CCS

[CK77] Shyh-Ching Chen and D. J. Kuck. Combinational circuit synthesis with time and component bounds. IEEE Transactions on Computers, C-26(8):712–726, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674909.

Chow:1979:MDL

[CK79] Yuan-Chieh Chow and W. H. Kohler. Models for dynamic load balancing in a heterogeneous multiple processor system. IEEE Transactions on Computers, C-28(5):354–361, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675365.

Chang:1977:EBP

[CKL77] D. Y. Chang, D. J. Kuck, and D. H. Lawrie. On the effective bandwidth of parallel memories. IEEE Transactions on Computers, C-26(5):480–490, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674865.

Cohn:1971:SPW

[CL71] M. Cohn and A. J. Lincoln.

Spectral properties of waveforms generated by smooth sequences. IEEE Transactions on Computers, C-20(3):330–332, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671833.

Cyre:1972:GMC

[CL72] W. R. Cyre and G. J. Lipovski.

On generating multipliers for a cellular Fast Fourier Transform processor. IEEE Transactions on Computers, C-21(1):83–87, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672027.

Chao:1975:HRR

[CL75] Chin Chao and H. H. Loomis, Jr.

High rate realization of finite-state machines. IEEE Transactions on Computers, C-24(7):729–740, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672886.

Cheng:1977:ASO

[CL77] D. Cheng and J. J. Liu. An algorithm for sequency ordering of Hadamard functions. IEEE Transactions on Computers, C-26 (3):308–309, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674828.

Cheng:1979:GOT

[CL79] D. K. Cheng and J. J. Liu. A generalized orthogonal transformation matrix. IEEE Transactions on Computers, C-28(2): 147–150, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675304.

Clegg:1973:USA

[Cle73] F. W. Clegg. Use of SPOOF’s in the analysis of faulty logic networks. IEEE Transactions on Computers, C-22(3):229–234, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672292.

Costes:1978:RAM

[CLL78] A. Costes, C. Landrault, and J. C. Laprie. Reliability and availability models for maintained systems featuring hardware failures and design faults. IEEE Transactions on Computers, C-27(6):548–560, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675146.

Chen:1970:ISL

[CM70] Wei Chen and L. P. McNamee. Iterative solution of large-scale systems by hybrid techniques.

IEEE Transactions on Computers, C-19(10):879–889, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671387.

Chaney:1973:ABS

[CM73a] T. J. Chaney and C. E. Molnar. Anomalous behavior of synchronizer and arbiter circuits. IEEE Transactions on Computers, C-22(4):421–422, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672323.

Chien:1973:PAR

[CM73b] Robert T. Chien and Edward A.

Mark. On performance analysis of retrieval functions. IEEE Transactions on Computers, C-22(9):840–844, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009175.

Cerny:1974:CAS

[CM74] E. Cerny and M. A. Marin. A computer algorithm for the synthesis of memoryless logic circuits. IEEE Transactions on Computers, C-23(5):455–465, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672560.

Carter:1976:IEF

[CM76] W. C. Carter and C. E. Mc-Carthy. Implementation of an experimental fault-tolerant memory system. IEEE Transactions on Computers, C-25(6):557–568, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674655.

Cerny:1977:AUM

[CM77] E. Cerny and M. A. Marin. An approach to unified methodology of combinational switching circuits. IEEE Transactions on Computers, C-26(8):745–756, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674912.

Chang:1979:AR

[CM79a] D. M. Chang and T. H. Mott, Jr. Author’s reply. IEEE Transactions on Computers, C-28 (11):875, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675271.

Current:1979:IPC

[CM79b] K. W. Current and D. A. Mow. Implementing parallel counters with four-valued threshold logic. IEEE Transactions on Computers, C-28(3):200–204, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675320.

Cutler:1979:CCI

[CM79c] R. B. Cutler and S. Muroga.

Comments on “Computing Irredundant Normal Forms from Abbreviated Presence Functions”. IEEE Transactions on Computers, C-28(11):874–875, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675270.

Cutler:1979:CGC

[CM79d] R. B. Cutler and S. Muroga. Comments on “Generalization of Consensus Theory and Application to the Minimization of Boolean Functions”. IEEE Transactions on Computers, C-28(7):542–543, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675403.

Cerny:1979:SMB

[CMS79] E. Cerny, D. Mange, and E. Sanchez.

Synthesis of minimal binary decision trees. IEEE Transactions on Computers, C-28(7):472–482, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675391.

Cody:1973:SDN

[Cod73] William J. Cody, Jr. Static and dynamic numerical characteristics of floating-point arithmetic. IEEE Transactions on Computers, C-22(6):598–601, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009112;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Cody.pdf.

Cohn:1970:AR

[Coh70] M. Cohn. Authors’ reply. IEEE Transactions on Computers, C-19(7):654, July 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671594.

Coles:1977:CHS

[Col77] C. J. Coles. Comments on “Heuristic Synthesis of Microprogrammed Computer Architecture”. IEEE Transactions on Computers, C-26(9):930–932, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674941. See [AAK74].

Cooper:1972:BAI

[Coo72] D. W. Cooper. B72-25 artificial intelligence: The heuristic programming approach. IEEE Transactions on Computers, C-21 (12):1458, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672120.

Cooper:1977:DP

[Coo77] R. G. Cooper. The distributed pipeline. IEEE Transactions on Computers, C-26(11):1123– 1132, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674758.

Corinthios:1971:DCF

[Cor71a] M. J. Corinthios. The design of a class of Fast Fourier Transform computers. IEEE Transactions on Computers, C-20(6):617–623, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671905.

Corinthios:1971:FFT

[Cor71b] M. J. Corinthios. A Fast Fourier Transform for high-speed signal processing. IEEE Transactions on Computers, C-20(8):843–846, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671952.

Corinthios:1977:FTA

[Cor77] Michael J. Corinthios. A fast Z transformation algorithm for system identification. IEEE Transactions on Computers, C-26(1): 55–67, January 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009274.

Covo:1974:AMC

[Cov74] A. A. Covo. Analysis of multiprocessor control organizations with partial program memory replication. IEEE Transactions on Computers, C-23(2):113–120, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672466.

Cox:1972:BFD

[Cox72] J. E. Cox. B72-19 future developments in telecommunications. IEEE Transactions on Computers, C-21(10):1135–1136, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672059.

Coy:1979:DET

[Coy79] W. Coy. On the design of easily testable iterative systems of combinational cells. IEEE Transactions on Computers, C-28(5):367–371, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675368.

Chandy:1972:RRS

[CR72a] K. M. Chandy and C. V. Ramamoorthy. Rollback and recovery strategies for computer programs. IEEE Transactions on Computers, C-21(6):546–556, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009007.

Chandy:1972:DML

[CR72b] K. M. Chandy and R. A. Russell. The design or multipoint linkages in a teleprocessing tree network. IEEE Transactions on Computers, C-21(10):1062–1066, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672045.

Chiang:1972:NAB

[CR72c] A. C. L. Chiang and I. S. Reed.

Notes on the arithmetic BN modulo A codes. IEEE Transactions on Computers, C-21(8):891–894, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009046.

Chien:1972:NDB

[CR72d] Y. T. Chien and R. Ribak. A new data base for syntax-directed pattern analysis and recognition. IEEE Transactions on Computers, C-21(7):790–801, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672176.

Casasent:1978:SFO

[CR78] D. Casasent and P. D. Rapp. System functions for an optical/digital processor. IEEE Transactions on Computers, C-27(8):732–738, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675180.

Chiang:1972:PSP

[CRB72] A. C. L. Chiang, I. S. Reed, and A. V. Banes. Path sensitization, partial Boolean difference, and automated fault diagnosis. IEEE Transactions on Computers, C-21(2):189–195, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008925.

Cassee:1970:TAT

[CS70] P. R. Cassee and M. J. O. Strutt. Is there any advantage of ternary logic as compared with binary? IEEE Transactions on Computers, C-19(6):559, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671569.

Carl:1973:HWT

[CS73] Joseph W. Carl and Richard V. Swartwood. A hybrid Walsh transform computer. IEEE Transactions on Computers, C-22(7):669–672, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009133.

Casasent:1975:ODP

[CS75] D. Casasent and W. M. Sterling. An optical/digital processor: Hardware and applications. IEEE Transactions on Computers, C-24(4):348–358, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672821.

Ciompi:1979:AOD

[CS79] P. Ciompi and L. Simoncini. Analysis and optimal design of selfdiagnosable systems with repair. IEEE Transactions on Computers, C-28(5):362–365, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675366.

Carroll:1974:EAT

[CSJ74] B. D. Carroll, H. G. Shah, and D. M. Jones. An examination of algebraic test generation methods for multiple faults. IEEE Transactions on Computers, C-23(7):743–745, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672617.

Cook:1973:DSC

[CSST73] R. W. Cook, W. H. Sisson, T. F.

Storey, and W. N. Toy. Design of a self-checking microprogram control. IEEE Transactions on Computers, C-22(3):255–262, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672297.

Corinthios:1975:PRF

[CSY75] M. J. Corinthios, K. C. Smith, and J. L. Yen. A parallel radix-4 Fast Fourier Transform computer. IEEE Transactions on Computers, C-24(1):80–92, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672678.

Curtis:1970:SPR

[Cur70] H. Allen Curtis. Systematic procedures for realizing synchronous sequential machines using flipflop memory: Part II. IEEE Transactions on Computers, C-19 (1):66–73, January 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008901.

Curtis:1971:FRC

[Cur71a] H. A. Curtis. The further reduction of CC-tables. IEEE Transactions on Computers, C-20(4):454–456, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671857.

Curtis:1971:RPS

[Cur71b] H. A. Curtis. The realization of polylinear sequential circuits using flip-flop memory. IEEE Transactions on Computers, C-20 (1):87–94, January 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671677.

Curtis:1973:TLS

[Cur73a] H. A. Curtis. Tan-like state assignments for synchronous sequential machines. IEEE Transactions on Computers, C-22(2): 181–187, February 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672275.

Curtis:1973:NTD

[Cur73b] H. Allen Curtis. A new type double-rank sequential machine. IEEE Transactions on Computers, C-22(8):796–803, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009161.

Curtis:1976:SDB

[Cur76] H. A. Curtis. Simplified decomposition of Boolean functions. IEEE Transactions on Computers, C-25(10):1033–1044, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674544.

Curts:1977:ATM

[Cur77] H. A. Curts. Adjacency table method of deriving minimal sums. IEEE Transactions on Computers, C-26(11):1136–1141, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674760.

Curtis:1979:SCM

[Cur79] H. A. Curtis. Short-cut method of deriving nearly optimal arrays of NAND trees. IEEE Transactions on Computers, C-28(7):521–528, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675398.

Cutlip:1970:RSR

[Cut70] W. F. Cutlip. R70-14 some results on cascade decomposition of automata. IEEE Transactions on Computers, C-19(6):565, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671578.

Couranz:1975:TEB

[CW75] G. R. Couranz and D. F. Wann. Theoretical and experimental behavior of synchronizers operating in the metastable region. IEEE Transactions on Computers, C-24(6):604–616, June 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672866.

Conrad:1977:ISL

[CW77] V. Conrad and Y. Wallach. Iterative solution of linear equations on a parallel processor system. IEEE Transactions on Computers, C-26(9):838–847, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674932.

Chandra:1979:MPC

[CW79a] A. K. Chandra and C. K. Wong.

The movement and permutation of columns in magnetic bubble lattice files. IEEE Transactions on Computers, C-28(1):8–15, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675217.

Chen:1979:PMB

[CW79b] I-Ngo Chen and R. Willoner.

An O(n) parallel multiplier with bit-sequential input and output. IEEE Transactions on Computers, C-28(10):721–727, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675239.

Chung:1979:AOI

[CW79c] Kin-Man Chung and C. K. Wong.

Asymptotically optimal interconnection networks from two-state cells. IEEE Transactions on Computers, C-28(7):500–505, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675395.

Cooper:1976:CCA

[CY76] D. B. Cooper and N. Yalabik. On the computational cost of approximating and recognizing noise-perturbed straight lines and quadratic arcs in the plane. IEEE Transactions on Computers, C-25(10):1020–1032, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674543.

Culliney:1979:RSO

[CYNM79] J. N. Culliney, M. H. Young, T. Nakagawa, and S. Muroga. Results of the synthesis of optimal networks of AND and OR gates for four-variable switching functions. IEEE Transactions on Computers, C-28(1):76–85, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675229.

Diaz:1979:UDS

[DAA79] M. Diaz, P. Azema, and J. M. Ayache. Unified design of selfchecking and fail-safe combinational circuits and sequential machines. IEEE Transactions on Computers, C-28(3):276–281, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675338.

Dahlberg:1973:SFR

[Dah73] B. Dahlberg. On symmetric functions with redundant variablesweighted functions. IEEE Transactions on Computers, C-22(5): 450–458, May 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672340.

Danielsson:1970:ICG

[Dan70] P. E. Danielsson. Incremental curve generation. IEEE Transactions on Computers, C-19(9): 783–793, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671634.

Dandapani:1973:DMT

[Dan73] Ramaswami Dandapani. Derivation of minimal test sets for monotonic logic circuits. IEEE Transactions on Computers, C-22(7):657–661, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009131.

Das:1971:FSL

[Das71a] S. K. Das. Feature selection with a linear dependence measure. IEEE Transactions on Computers, C-20(9):1106–1109, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672005. See comments

[TV72].

Das:1971:ASS

[Das71b] S. R. Das. An approach for simplifying switching functions by utilizing the cover table representation. IEEE Transactions on Computers, C-20(3):355–359, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671837.

Das:1971:CNA

[Das71c] S. R. Das. Comments on “A New Algorithm for Generating Prime Implicants”. IEEE Transactions on Computers, C-20(12): 1614–1615, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671778. See [SCL70].

Das:1972:AR

[Das72] Subrata K. Das. Author’s reply. IEEE Transactions on Computers, C-21(4):409, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008990.

Das:1973:NAF

[Das73] S. R. Das. On a new approach for finding all the modified cutsets in an incompatibility graph.

IEEE Transactions on Computers, C-22(2):187–193, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672276.

Dasgupta:1978:CIM

[Das78] S. Dasgupta. Comments on “The Identification of Maximal Parallelism in Straight-Line Microprograms”. IEEE Transactions on Computers, C-27(3): 285–286, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675089. See [DT76,

Bar78, Ada78].

Davis:1970:RSL

[Dav70a] W. A. Davis. R70-42 synthesis of linear sequential machines with unspecified outputs. IEEE Transactions on Computers, C-19(10):989, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671408.

Davis:1970:SMR

[Dav70b] W. A. Davis. Sequential machines realizable with delay elements only. IEEE Transactions on Computers, C-19(4):353–355, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671515.

Davies:1971:PWO

[Dav71] A. C. Davies. Properties of waveforms obtained by nonrecursive digital filtering of pseudorandom binary sequences. IEEE Transactions on Computers, C-20 (3):270–281, March 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671824.

Davies:1972:DGW

[Dav72a] Anthony C. Davies. On the definition and generation of Walsh functions. IEEE Transactions on Computers, C-21(2):187–189, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008924. See comments

[Van73b].

Davis:1972:VAA

[Dav72b] J. M. Davis. Velocity analysis: An application of deterministic estimation to reflection seismology. IEEE Transactions on Computers, C-21(7):730–738, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672168.

Davis:1974:USN

[Dav74] R. L. Davis. Uniform shift networks. ????, C-24(3):317–322, March 1974. CODEN ???? ISSN ????

David:1977:MDA

[Dav77a] R. David. Modular design of asynchronous circuits defined by graphs. IEEE Transactions on Computers, C-26(8):727–737, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674910.

Davis:1977:USA

[Dav77b] L. S. Davis. Understanding shape: Angles and sides. IEEE Transactions on Computers, C-26 (3):236–242, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674812.

David:1978:TSC

[Dav78] R. David. A totally selfchecking 1-out-of-3 checker. IEEE Transactions on Computers, C-27(6):570–572, June 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675149.

Durling:1973:UMR

[DB73] A. E. Durling and T. E. Bullock. A unified method for the reconstruction of sampled data. IEEE Transactions on Computers, C-22(4):388–396, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672319.

DeFiore:1974:QAU

[DB74] C. R. DeFiore and P. B. Berra. A quantitative analysis of the utilization of associative memories in data management. IEEE Transactions on Computers, C-23(2): 121–133, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672467.

David:1976:ARF

[DB76] R. David and G. Blanchet. About random fault detection of combinational networks. IEEE Transactions on Computers, C-25(6):659–664, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674669.

DeSarkar:1971:DIP

[DBC71] S. C. De Sarkar, A. K. Basu, and A. K. Choudhury. On the determination of irredundant prime closed sets. IEEE Transactions on Computers, C-20(8):933–938, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671968.

Das:1973:CMM

[DBC73] S. R. Das, D. K. Banerji, and A. Chattopadhyay. On control memory minimization in microprogrammed digital computers.

IEEE Transactions on Computers, C-22(9):845–848, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009176.

Dudani:1977:AIM

[DBM77] Sahibsingh A. Dudani, Kenneth J.

Breeding, and Robert B. McGhee. Aircraft identification by moment invariants. IEEE Transactions on Computers, C-26(1):39–46, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009272.

Das:1976:CDM

[DDSM76] S. R. Das, C. R. Datta, P. K. Srimani, and K. Mandal. Comments on “Derivation of Minimal Complete Sets of Test-Input Sequences Using Boolean Differences”. IEEE Transactions on Computers, C-25(10):1053–1056, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674548. See [MST^{+}75].

deRennaeSouza:1970:BTA

[de 70] Celso de Renna e Souza. B70-5 theories of abstract automata. IEEE Transactions on Computers, C-19(9):855–857, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671652.

deRennaeSouza:1971:PAM

[de 71] Celso de Renna e Souza. Probabilistic automata with monitored final state sets. IEEE Transactions on Computers, C-20(4):448–452, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671855.

DeWitt:1979:DMO

[De 79] D. J. De Witt. Direct — a multiprocessor organization for supporting relational database management systems. IEEE Transactions on Computers, C-28(6):395–406, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675379.

Dawoud:1973:PDD

[DEA73] D. Sh. Dawoud and N. Z. El-Araby. Parallel digital differential analyzer with arbitrary stored interconnections. IEEE Transactions on Computers, C-22(1):41–46, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672192.

Deekshatulu:1972:BAL

[Dee72] B. L. Deekshatulu. B72-24 adaptation and learning in automatic systems. IEEE Transactions on Computers, C-21(11):1252, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672084.

Denning:1974:BSP

[Den74] P. J. Denning. B74-20 systematic programming: An introduction. IEEE Transactions on Computers, C-23(2):222–223, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672502.

Denman:1976:AAS

[Den76] E. D. Denman. An additional algorithm for a system of coupled algebraic matrix Riccati equations. IEEE Transactions on Computers, C-25(1):91–93, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009210. See comments

[Gou77].

Depeyrot:1970:LSI

[Dep70] M. Depeyrot. Linear system identification using real-time deconvolution. IEEE Transactions on Computers, C-19(12): 1139–1145, December 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671445.

Dertouzos:1973:TBS

[Der73] M. L. Dertouzos. Time bounds on space computations. IEEE Transactions on Computers, C-22 (1):12–17, January 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672188.

Despain:1974:FTC

[Des74] A. M. Despain. Fourier transform computers using CORDIC iterations. IEEE Transactions on Computers, C-23(10):993–1001, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672393.

Despain:1979:VFF

[Des79] A. M. Despain. Very fast Fourier transform algorithms hardware for implementation. IEEE Transactions on Computers, C-28(5):333–341, May 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675363.

Deutsch:1970:POH

[Deu70] E. S. Deutsch. On parallel operations on hexagonal arrays. IEEE Transactions on Computers, C-19(10):982–983, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671401.

Deuser:1971:HMF

[Deu71] L. M. Deuser. A hybrid multispectral feature selection criterion. IEEE Transactions on Computers, C-20(9):1116–1117, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672008.

Devijver:1974:NCB

[Dev74] P. A. Devijver. On a new class of bounds on Bayes risk in multihypothesis pattern recognition. IEEE Transactions on Computers, C-23(1):70–80, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672372. See comments

[Tou75].

Deverell:1975:PIA

[Dev75] J. Deverell. Pipeline iterative arithmetic arrays. IEEE Transactions on Computers, C-24(3): 317–322, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672807.

Devries:1979:CRI

[Dev79] R. C. Devries. Comments on “A Readily Implemented Single-Error-Correcting Unit-Distance Counting Code”. IEEE Transactions on Computers, C-28(3): 253–255, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675328. See [Kau70c].

Doty:1970:DTS

[DF70] K. L. Doty and H. Frank. Discretetime systems with the decomposition property. IEEE Transactions on Computers, C-19(5):412–420, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671532.

Das:1975:FDE

[DF75] P. Das and D. E. Farmer. Faultdetection experiments for paralleldecomposable sequential machines. IEEE Transactions on Computers, C-24(11):1104–1109, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672732.

Deutsch:1978:QSO

[DF78] E. S. Deutsch and J. R. Fram. A quantitative study of the orientation bias of some edge detector schemes. IEEE Transactions on Computers, C-27(3): 205–213, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675073.

Degryse:1972:LT

[DG72] D. Degryse and B. Guerin. A logarithmic transcoder. IEEE Transactions on Computers, C-21 (11):1165–1168, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672066.

Diaz:1974:SRF

[DGC74] M. Diaz, J. C. Geffroy, and M. Courvoisier. On-set realization of fail-safe sequential machines. IEEE Transactions on Computers, C-23(2):133–138, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672468.

Dunworth:1979:ESM

[DH79] A. Dunworth and H. V. Hartog.

An efficient state minimization algorithm for some special classes of incompletely specified sequential machines. IEEE Transactions on Computers, C-28(7):531–535, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675400.

Daniel:1971:DRM

[DI71] J. M. Daniel and J. D. Irwin.

Dynamic resolution of memory access conflicts. IEEE Transactions on Computers, C-20(12): 1583–1586, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671769.

Dias:1975:FMC

[Dia75] F. J. O. Dias. Fault masking in combinational logic circuits. IEEE Transactions on Computers, C-24(5):476–482, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672842.

Dias:1976:TTV

[Dia76] F. J. O. Dias. Truth-table verification of an iterative logic array. IEEE Transactions on Computers, C-25(6):605–613, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674660.

Digby:1973:SMM

[Dig73] David W. Digby. A search memory for many-to-many comparisons. IEEE Transactions on Computers, C-22(8):768–772, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009156.

Dirilten:1972:MMC

[Dir72] Hudai Dirilten. On the mathematical models characterizing faulty four-phase MOS logic arrays. IEEE Transactions on Computers, C-21(3):301–305, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008954.

Das:1972:CCT

[DK72] S. R. Das and N. S. Khabra.

Clause-column table approach for generating all the prime implicants of switching functions. IEEE Transactions on Computers, C-21 (11):1239–1246, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672077.

Dowd:1973:SGH

[DLL^{+}73] P. G. Dowd, D. J. Ladd, G. W. R.

Luderer, P. E. Rosenfeld, and R. L. Salmon. Stare 2 — a graphical hardcopy system. IEEE Transactions on Computers, C-22(4):359–367, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672315.

Ducasse:1975:CRB

[DM75] E. Ducasse and G. Metze. Comments on “The Relationship Between Multivalued Switching Algebra and Boolean Algebra under Different Definitions of Complement”. IEEE Transactions on Computers, C-24(2):220, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672785. See [DM76].

Ducasse:1976:RAR

[DM76] Edgar Ducasse and Gernot Metze.

Rebuttal to the authors’ reply to our comments on “The Relationship Between Multivalued Switching Algebra and Boolean Algebra Under Different Definitions of Complement”. IEEE Transactions on Computers, C-25(3):311–312, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009259. See [Eps73b,

DM75].

DeSousa:1978:SMR

[DM78a] P. T. De Sousa and F. P. Mathur. Sift-out modular redundancy. IEEE Transactions on Computers, C-27(7):624–627, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675162.

Deguchi:1978:TCT

[DM78b] K. Deguchi and I. Morishita. Texture characterization and texturebased image partitioning using twodimensional linear estimation techniques. IEEE Transactions on Computers, C-27(8):739–745, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675181.

Dao:1977:MII

[DMR77] T. T. Dao, E. J. Mccluskey, and L. K. Russell. Multivalued integrated injection logic. IEEE Transactions on Computers, C-26 (12):1233–1241, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674784.

Dirilten:1977:PMU

[DN77] H. Dirilten and T. G. Newman. Pattern matching under affine transformations. IEEE Transactions on Computers, C-26 (3):314–317, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674832.

Dorr:1972:SCC

[Dor72] R. C. Dorr. Self-checking combinational logic binary counters. IEEE Transactions on Computers, C-21(12):1426–1430, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672111.

Doster:1977:CPS

[Dos77] W. Doster. Contextual postprocessing system for cooperation with a multiple-choice characterrecognition system. IEEE Transactions on Computers, C-26(11): 1090–1101, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674755.

Doty:1970:ILD

[Dot70] K. L. Doty. On informationlossless discrete-time systems. IEEE Transactions on Computers, C-19(6):548–551, June 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671567.

Downs:1973:RTA

[Dow73] H. Robert Downs. Real-time algorithms and data management on Illiac IV. IEEE Transactions on Computers, C-22(8):773–777, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009157.

Dandapani:1974:DLN

[DR74] R. Dandapani and S. M. Reddy. On the design of logic networks with redundancy and testability considerations. IEEE Transactions on Computers, C-23(11): 1139–1149, November 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672414.

Davis:1975:DSE

[DR75] L. S. Davis and A. Rosenfeld. Detection of step edges in noisy onedimensional data. IEEE Transactions on Computers, C-24(10): 1006–1010, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672704.

Davis:1977:CSR

[DR77] L. S. Davis and A. Rosenfeld. Curve segmentation by relaxation labeling. IEEE Transactions on Computers, C-26(10):1053–1057, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674746.

DeMori:1975:SPC

[DRS75] R. De Mori, S. Rivoira, and A. Serra. A special-purpose computer for digital signal processing. IEEE Transactions on Computers, C-24(12):1202–1211, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672757.

Drubin:1971:CFF

[Dru71a] M. Drubin. Computation of the Fast Fourier Transform from data stored in external auxiliary memory for any general radix r = 2^{n}, n ≥ 1. IEEE Transactions on Computers, C-20(12):1552–1558, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671761.

Drubin:1971:KPF

[Dru71b] M. Drubin. Kronecker product factorization of the FFT matrix. IEEE Transactions on Computers, C-20(5):590–593, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671899.

Dere:1970:BAP

[DS70] W. Y. Dere and D. J. Sakrison. Berkeley array processor. IEEE Transactions on Computers, C-19(5):444–447, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671535.

Das:1971:DTP

[DS71] S. R. Das and C. L. Sheng. On detecting total or partial symmetry of switching functions. IEEE Transactions on Computers, C-20 (3):352–355, March 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671836.

Das:1972:MTP

[DS72a] S. K. Das and D. F. Stanat. A modified training procedure for linear threshold devices. IEEE Transactions on Computers, C-21(4):396–397, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008983.

DeMori:1972:PSS

[DS72b] R. De Mori and A. Serra. A parallel structure for signed-number multiplication and addition. IEEE Transactions on Computers, C-21 (12):1453–1454, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672118.

Dewdney:1973:TMD

[DS73] A. K. Dewdney and A. L. Szilard. Tours in machines and digraphs. IEEE Transactions on Computers, C-22(7):635–639, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009128.

Delcaro:1974:MTE

[DS74] L. G. Delcaro and G. L. Sicuranza. A method for transposing externally stored matrices. IEEE Transactions on Computers, C-23(9):967–970, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672655.

DeVries:1975:MOO

[DS75] R. C. De Vries and A. Svoboda. Multiple-output optimization with mosaics of Boolean functions. IEEE Transactions on Computers, C-24(8):777–785, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672902.

De-Sarkar:1974:UCR

[DSBC74] S. C. De-Sarkar, S. Bandyopadhyay, and A. K. Choudhury. Unate cascade realization of synchronous sequential machines. IEEE Transactions on Computers, C-23(10): 1008–1019, October 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672395.

Donaldson:1970:UCC

[DT70] R. W. Donaldson and G. T. Toussaint. Use of contextual constraints in recognition of contour-traced handprinted characters. IEEE Transactions on Computers, C-19 (11):1096–1099, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671432.

Dasgupta:1976:IMP

[DT76] S. Dasgupta and J. Tartar. The identification of maximal parallelism in straight-line microprograms. IEEE Transactions on Computers, C-25(10):986–992, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674539. See comments

[Das78].

David:1979:CEL

[DTG79] R. David and R. Tellez-Giron. Comments on “The Error Latency of a Fault in a Sequential Digital Circuit”. IEEE Transactions on Computers, C-28(1):85–86, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675230. See [SM76].

Do-Tu:1978:LAN

[DTI78] Hai Do-Tu and M. Installe. Learning algorithms for nonparametric solution to the minimum error classification problem. IEEE Transactions on Computers, C-27(7):648–659, July 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675165.

Du:1971:CNS

[Du71] Min-Wen Du. Comments on “Nonlinear Sequential Circuits”. IEEE Transactions on Computers, C-20(5):585–586, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671891. See [Kai70].

Du:1972:WFL

[Du72] Min-Wen Du. A way to find a lower bound for the minimal solution of the covering problem. IEEE Transactions on Computers, C-21(3):317–318, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008958.

Dube:1979:PSS

[Dub79] R. P. Dube. Preliminary specification of spline curves. IEEE Transactions on Computers, C-28(4):286–290, April 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675347.

Duin:1976:CSP

[Dui76] R. P. W. Duin. On the choice of smoothing parameters for Parzen estimators of probability density functions. IEEE Transactions on Computers, C-25(11): 1175–1179, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674577.

Duke:1974:BDM

[Duk74] K. A. Duke. B 74-26 discrete mathematical structures and their applications. IEEE Transactions on Computers, C-23(5):555, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672575.

Dunn:1975:GAL

[Dun75] J. C. Dunn. Group averaged linear transforms that detect corners and edges. IEEE Transactions on Computers, C-24(12): 1191–1201, December 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672756.

Dutta:1974:EPH

[Dut74] A. K. Dutta. An experimental procedure for handwritten character recognition. IEEE Transactions on Computers, C-23(5):536–545, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672571.

Druzeta:1978:HRT

[DV78a] A. Druzeta and Z. G. Vranesic.

A higher radix technique for fault detection in many-valued multithreshold networks. IEEE Transactions on Computers, C-27(11): 1070–1073, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674999.

Dubois:1978:DFT

[DV78b] E. Dubois and A. N. Venetsanopoulos. The discrete Fourier transform over finite rings with application to fast convolution. IEEE Transactions on Computers, C-27(7):586–593, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675158.

Druzeta:1974:AME

[DVS74] A. Druzeta, Z. G. Vranesic, and A. S. Sedra. Application of multithreshold elements in the realization of many-valued logic networks. IEEE Transactions on Computers, C-23(11):1194–1198, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672422.

Dollhoff:1972:RSE

[DW72a] Terry L. Dollhoff and Bernard L. Weinberg. A result on set extraction and application to coveringclosure tables. IEEE Transactions on Computers, C-21(6):603–606, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009018.

Drilman:1972:MNN

[DW72b] J. Drilman and P. Weiner. Modular networks and nondeterministic sequential machines. IEEE Transactions on Computers, C-21(10):1124–1129, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672054.

Du:1973:CSS

[DW73a] Min-Wen Du and C. Dennis Weiss. Circuit structure and switching function verification. IEEE Transactions on Computers, C-22(6):618–625, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=

&arnumber=5009116;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Du.pdf.

Du:1973:MFD

[DW73b] Min-Wen Du and C. Dennis Weiss. Multiple fault detection in combinational circuits: Algorithms and computational results. IEEE Transactions on Computers, C-22(3):235–240, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672293. See comments

[GS74, Bos79].

Davies:1978:SMR

[DW78] D. Davies and J. F. Wakerly.

Synchronization and matching in redundant systems. IEEE Transactions on Computers, C-27(6):531–539, June 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675144.

Dinn:1970:CCI

[DWT70] D. F. Dinn, D. A. Winter, and B. G. Trenholm. CINTEL — Computer Interface for Television. IEEE Transactions on Computers, C-19(11):1091–1095, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671431.

Easton:1978:CCS

[Eas78] M. C. Easton. Computation of cold-start miss ratios. IEEE Transactions on Computers, C-27(5):404–408, May 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675119.

Edwards:1975:ARW

[Edw75] C. R. Edwards. The application of the Rademacher–Walsh transform to Boolean function classification and threshold logic synthesis. IEEE Transactions on Computers, C-24(1):48–62, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672675. See comments

[Yue76].

Edwards:1976:AR

[Edw76] C. R. Edwards. Author’s reply. IEEE Transactions on Computers, C-25(7):767, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674689.

Ehrich:1976:RRW

[EF76] R. W. Ehrich and J. P. Foith. Representation of random waveforms by relational trees. IEEE Transactions on Computers, C-25(7):725–736, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674681.

Easton:1979:UBS

[EF79] M. C. Easton and P. A. Franaszek.

Use bit scanning in replacement decisions. IEEE Transactions on Computers, C-28(2): 133–141, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675302.

Erman:1976:SOS

[EFLR76] L. D. Erman, R. D. Fennell, V. R.

Lesser, and D. R. Reddy. System organizations for speech understanding: Implications of network and multiprocessor computer architectures for AI. IEEE Transactions on Computers, C-25(4):414–421, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674623.

El-Ghoroury:1971:RSA

[EGG71] H. N. El-Ghoroury and S. C.

Gupta. Realization of stochastic automata. IEEE Transactions on Computers, C-20(8):889–893, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671959.

El-Ghoroury:1972:ABN

[EGG72] H. N. El-Ghoroury and S. C.

Gupta. Additive Bernoulli noise linear sequential circuits. IEEE Transactions on Computers, C-21(10):1119–1124, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672053.

Ernst:1976:UHO

[EH76] G. W. Ernst and R. J. Hookway. The use of higher order logic in program verification. IEEE Transactions on Computers, C-25(8):844–851, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674703.

Edwards:1978:DSP

[EH78] C. R. Edwards and S. L. Hurst. A digital synthesis procedure under function symmetries and mapping methods. IEEE Transactions on Computers, C-27(11): 985–997, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674988.

Ehrich:1972:NSM

[Ehr72] H.-D. Ehrich. A note on state minimization of a special class of incomplete sequential machines. IEEE Transactions on Computers, C-21(5):500–502, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672142.

Ehrich:1976:RCD

[Ehr76] R. W. Ehrich. A readily computable decision rule with variable dimensionality. IEEE Transactions on Computers, C-25(5):539–542, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674644.

Ehrich:1977:DGE

[Ehr77] R. W. Ehrich. Detection of global edges in textured images. IEEE Transactions on Computers, C-26(6):589–603, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674884.

Etiemble:1977:ITC

[EI77] D. Etiemble and M. Isreal. Implementation of ternary circuits with-binary integrated circuits. IEEE Transactions on Computers, C-26(12):1222–1233, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674783.

Ehrich:1975:ECR

[EK75] R. W. Ehrich and K. J. Koehler. Experiments in the contextual recognition of cursive script. IEEE Transactions on Computers, C-24 (2):182–194, February 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672777.

Eklundh:1972:FCM

[Ekl72] J. O. Eklundh. A fast computer method for matrix transposing. IEEE Transactions on Computers, C-21(7):801–803, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=1672177.  See also

[Ekl73, Sch73, All75a, Ram75, TE76, Ari79].

Eklundh:1973:AR

[Ekl73] J. O. Eklundh. Author’s reply. IEEE Transactions on Computers, C-22(5):543–544, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672354.

Ekstrom:1973:NAI

[Eks73] M. P. Ekstrom. A numerical algorithm for identifying spread functions of shift-invariant imaging systems. IEEE Transactions on Computers, C-22(4):322–328, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672311.

Ekstorm:1974:IIA

[Eks74] M. P. Ekstorm. An iterativeimprovement approach to the numerical solution of vector Toeplitz systems. IEEE Transactions on Computers, C-23(3):320–325, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672522.

Elmasry:1976:LDU

[Elm76] M. I. Elmasry. Logic design using EFL structures. IEEE Transactions on Computers, C-25(9):952–956, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674722.

Elschlager:1974:BDM

[Els74] R. Elschlager. B74-8 design of man-computer dialogues. IEEE Transactions on Computers, C-23 (2):212–213, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672490.

En:1977:NRE

[En77] John En. A new random-error-correction code. IEEE Transactions on Computers, C-26(1):86–88, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009279. See comments

[Paa79].

En:1979:AR

[En79] J. En. Author’s reply. IEEE Transactions on Computers, C-28 (3):257–258, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675330.

Eschenbach:1978:ECR

[EO78] R. F. Eschenbach and B. M. Oliver. An efficient coordinate rotation algorithm. IEEE Transactions on Computers, C-27(12): 1178–1180, December 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675020.

Epstein:1973:EAD

[Eps73a] G. Epstein. An equational axiomatization for the disjoint system of post algebras. IEEE Transactions on Computers, C-22(4):422–423, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672324.

Epstein:1973:CRB

[Eps73b] George Epstein. Comments on “The Relationship Between Multivalued Switching Algebra and Boolean Algebra Under Different Definitions of Complement”. IEEE Transactions on Computers, C-22(9):864, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009181. See [SS72,

DM76].

Ercegovac:1973:REC

[Erc73] Miloš D. Ercegovac. Radix-16 evaluation of certain elementary functions. IEEE Transactions on Computers, C-22(6):561–566, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009107;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Ercegovac.
pdf.

Ercegovac:1977:GHO

[Erc77] Miloš D. Ercegovac. A general hardware-oriented method for evaluation of functions and computations in a digital computer. IEEE Transactions on Computers, C-26(7):667–680, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674900.

Ernbo:1972:AIM

[Ern72] Arne Ernbo. Application of intensity-modulated ink jets to alphanumeric printing devices. IEEE Transactions on Computers, C-21(9):942–947, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009070.

Ernst:1976:DDT

[Ern76] G. W. Ernst. A definition-driven theorem prover. IEEE Transactions on Computers, C-25(4):317–322, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674611.

Ecker:1977:CWC

[ES77] K. Ecker and D. Schutt. Computational work of combinational machines. IEEE Transactions on Computers, C-26(11):1161– 1163, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674768.

Elmasry:1975:TLE

[ET75] M. I. Elmasry and P. M. Thompson. Two-level emitter-function logic structures for logic-in-memory computers. IEEE Transactions on Computers, C-24(3): 250–258, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672799.

Ejiri:1972:PIR

[EUY^{+}72] Masakazu Ejiri, Takeshi Uno, Haruo Yoda, Tatsuo Goto, and Kiyoo Takeyasu. A prototype intelligent robot that assembles objects from plan drawings. IEEE Transactions on Computers, C-21 (2):161–170, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008921.

Elineau:1977:NJK

[EW77] G. Elineau and W. Wiesbeck. A new J–K flip-flop for synchronizers. IEEE Transactions on Computers, C-26(12):1277–1279, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674789.

El-Ziq:1978:CAL

[EZS78] Y. M. El-Ziq and S. Y. H.

Su. Computer-aided logic design of two-level MOS combinational networks with statistical results. IEEE Transactions on Computers, C-27(10):911–923, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674970.

Fino:1976:UMT

[FA76] B. J. Fino and V. R. Algazi.

Unified matrix treatment of the fast Walsh–Hadamard transform. IEEE Transactions on Computers, C-25(11):1142–1146, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674569.

Faiman:1973:BDC

[Fai73] M. Faiman. B73-11 digital circuits and devices. IEEE Transactions on Computers, C-22(5):546–547, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672358.

Farmer:1973:ADF

[Far73] D. E. Farmer. Algorithms for designing fault-detection experiments for sequential machines. IEEE Transactions on Computers, C-22(2):159–167, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672271.

Farhat:1975:NOD

[Far75] N. H. Farhat. Nonlinear optical data processing and filtering: a feasibility study. IEEE Transactions on Computers, C-24(4):443–448, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672833.

Farag:1979:WLR

[Far79] R. F. H. Farag. Word-level recognition of cursive script. IEEE Transactions on Computers, C-28 (2):172–175, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675310.

Fernandez:1973:BNP

[FB73a] Eduardo B. Fernandez and Bertram Bussell. Bounds on the number of processors and time for multiprocessor optimal schedules. IEEE Transactions on Computers, C-22(8):745–751, August 1973.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009153.

Fu:1973:TSS

[FB73b] King-Sun Fu and B. K. Bhargava. Tree systems for syntactic pattern recognition. IEEE Transactions on Computers, C-22 (12):1087–1099, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672247.

French:1974:UWF

[FB74] A. S. French and E. G. Butz.

The use of Walsh functions in the Wiener analysis of nonlinear systems. IEEE Transactions on Computers, C-23(3):225–232, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672509.

Friedman:1975:AFN

[FBS75] J. H. Friedman, F. Baskett, and L. J. Shustek. An algorithm for finding nearest neighbors. IEEE Transactions on Computers, C-24(10):1000–1006, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672703.

Frei:1977:FBD

[FC77] W. Frei and Chung-Ching Chen. Fast boundary detection: a generalization and a new algorithm. IEEE Transactions on Computers, C-26(10):988–998, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674733.

Fu:1972:AR

[FCC72] K. S. Fu, Y. T. Chien, and G. P. Cardillo. Authors’ reply. IEEE Transactions on Computers, C-21(3):320, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008960.

Fisher:1973:BPT

[FD73] L. T. Fisher and D. W. Dearholt. Boundary points of threshold functions. IEEE Transactions on Computers, C-22(12):1132– 1139, December 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672252.

Filene:1974:RIM

[FD74a] R. J. Filene and W. M. Daly. The reliability impact of mission abort strategies on redundant flight computer systems. IEEE Transactions on Computers, C-23(7):739–743, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672616.

Fridrich:1974:MFT

[FD74b] M. Fridrich and W. A. Davis. Minimal fault tests for combinational networks. IEEE Transactions on Computers, C-23(8):850–859, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672633. See comments

[Red77a].

Fram:1975:QEE

[FD75] J. R. Fram and E. S. Deutsch.

On the quantitative evaluation of edge detection schemes and their comparison with human performance. IEEE Transactions on Computers, C-24(6):616–628, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672867.

Freeman:1977:CFA

[FD77a] H. Freeman and L. S. Davis.

A corner-finding algorithm for chain-coded curves. IEEE Transactions on Computers, C-26(3): 297–303, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674825.

Fridrich:1977:MFT

[FD77b] M. Fridrich and W. A. Davis.

Minimal fault tests for redundant combinational networks. IEEE Transactions on Computers, C-26(10):1057–1060, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674747.

Fleischhammer:1979:ABF

[FD79] W. Fleischhammer and O. Dortok. The anomalous behavior of flip-flops in synchronizer circuits. IEEE Transactions on Computers, C-28(3):273–276, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675337.

Fischler:1973:RMP

[FE73] M. A. Fischler and R. A. Elschlager. The representation and matching of pictorial structures. IEEE Transactions on Computers, C-22 (1):67–92, January 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672195.

Fehlauer:1978:DCF

[FE78] J. Fehlauer and B. A. Eisenstein. A declustering criterion for feature extraction in pattern recognition. IEEE Transactions on Computers, C-27(3):261–266, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675083.

Feldman:1979:AOP

[Fel79] M. B. Feldman. An applicationoriented programming language for sequential machine studies. IEEE Transactions on Computers, C-28(8):582–586, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675417.

Feng:1974:DMF

[Fen74] Tse-Yun Feng. Data manipulating functions in parallel processors and their implementations. IEEE Transactions on Computers, C-23(3):309–318, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672520.

Feng:1977:PPP

[Fen77] Tse-Yun Feng. Parallel processors and processing: An introduction. IEEE Transactions on Computers, C-26(2):97–98, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009288.

Ferguson:1977:ANG

[Fer77] J. C. Ferguson. Algorithms in numerical geometry system. IEEE Transactions on Computers, C-26 (12):1290–1292, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674794.

Feustel:1972:BCC

[Feu72] E. A. Feustel. B72-9 compiler construction for digital computers. IEEE Transactions on Computers, C-21(5):517, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672152.

Feustel:1973:ATA

[Feu73] Edward A. Feustel. On the advantages of tagged architecture. IEEE Transactions on Computers, C-22(7):644–656, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009130.

Fung:1975:SSD

[FF75] Lai-Wo Fung and King-Sun Fu. Stochastic syntactic decoding for pattern classification. IEEE Transactions on Computers, C-24(6):662–667, June 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672871.

Foster:1971:CIO

[FG71a] C. C. Foster and R. Gonter. Conditional interpretation of operation codes. IEEE Transactions on Computers, C-20(1):108–111, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671684.

Frosini:1971:PIA

[FG71b] G. Frosini and G. B. Gerace.

Pulse input asynchronous sequential circuits. IEEE Transactions on Computers, C-20(4):436–442, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671853.

Frosini:1971:USS

[FG71c] G. Frosini and G. B. Gerace.

A universal STT state assignment method for pulse input asynchronous sequential circuits. IEEE Transactions on Computers, C-20(8):856–861, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671954.

Frosini:1976:SAP

[FG76] G. Frosini and G. B. Gerace. Synchronous and asynchronous procedures for the synthesis of masterslave circuits. IEEE Transactions on Computers, C-25(11): 1179–1184, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674578.

Foster:1971:MOC

[FGR71] C. C. Foster, R. H. Gonter, and E. M. Riseman. Measures of opcode utilization. IEEE Transactions on Computers, C-20(5):582–584, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671889.

Felstead:1978:CDC

[FHTS78] E. B. Felstead, K. O. Hill, and A. U. Tenne-Sens. Comments on “Direct Coherent Optical Fourier Transforms of Curves”. IEEE Transactions on Computers, C-27(2):190–191, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675063. See [LH75].

Fisher:1974:UPE

[Fis74] L. T. Fisher. Unateness properties of and-exclusive-or logic circuits. IEEE Transactions on Computers, C-23(2):166–172, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672472.

Fishman:1976:POS

[Fis76] D. H. Fishman. A problem-oriented search procedure for theorem proving. IEEE Transactions on Computers, C-25(8):807–815, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674699.

Fukunaga:1970:AKL

[FK70a] K. Fukunaga and W. L. G. Koontz.

Application of the Karhunen–Loève expansion to feature selection and ordering. IEEE Transactions on Computers, C-19(4):311–318, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671511.

Fukunaga:1970:CAG

[FK70b] K. Fukunaga and W. L. G. Koontz.

A criterion and an algorithm for grouping data. IEEE Transactions on Computers, C-19(10): 917–923, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671392.

Fukunaga:1971:ECE

[FK71] K. Fukunaga and D. L. Kessell.

Estimation of classification error. IEEE Transactions on Computers, C-20(12):1521–1527, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671758.

Fujiwara:1974:DDS

[FK74] H. Fujiwara and K. Kinoshita.

Design of diagnosable sequential machines utilizing extra outputs. IEEE Transactions on Computers, C-23(2):138–145, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672469.

Fujiwara:1978:CAP

[FK78a] H. Fujiwara and K. Kinoshita. Connection assignments for probabilistically diagnosable systems. IEEE Transactions on Computers, C-27(3):280–283, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675087.

Fujiwara:1978:CCS

[FK78b] H. Fujiwara and K. Kinoshita. On the computational complexity of system diagnosis. IEEE Transactions on Computers, C-27 (10):881–885, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674966.

Fujiwara:1978:SET

[FK78c] H. Fujiwara and K. Kinoshita. Some existence theorems for probabilistically diagnosable systems. IEEE Transactions on Computers, C-27(4):379–384, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675111.

Frieder:1975:ABC

[FL75] G. Frieder and C. Luk. Algorithms for binary coded balanced and ordinary ternary operations. IEEE Transactions on Computers, C-24(2):212–215, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672781.

Fennell:1977:PAI

[FL77a] Richard D. Fennell and Victor R. Lesser. Parallelism in artificial intelligence problem solving: a case study of Hearsay II. IEEE Transactions on Computers, C-26(2):98–111, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009289.

Finnila:1977:ALA

[FL77b] Charles A. Finnila and Hubert H.

Love. The associative linear array processor. IEEE Transactions on Computers, C-26(2): 112–125, February 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009290.

Fuller:1976:EET

[FLBK76] S. H. Fuller, V. R. Lesser, C. G.

Bell, and C. H. Kaman. The effects-of emerging technology and emulation requirements on microprogramming. IEEE Transactions on Computers, C-25(10): 1000–1009, October 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674541.

Fleisher:1979:ISS

[Fle79] H. Fleisher. Introduction to special section on programmable logic arrays. IEEE Transactions on Computers, C-28(9):593, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675425.

Flores:1972:BCO

[Flo72] I. Flores. B72-26 computer organization and the System/370. IEEE Transactions on Computers, C-21(12):1458–1459, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672121.

Flores:1975:CRP

[Flo75] I. Flores. Comments on the review of “Peripheral Devices”. IEEE Transactions on Computers, C-24(2):223–224, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672789. See [Bar74a].

Fusco:1978:QFM

[FLR78] F. Fusco, S. Levialdi, and L. Ritondale. A quality factor for measuring program efficiency. IEEE Transactions on Computers, C-27 (3):275–280, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675086.

Flynn:1970:DFI

[Fly70] M. J. Flynn. On division by functional iteration. IEEE Transactions on Computers, C-19(8): 702–706, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671612;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Flynn.pdf.

Reprinted in [Swa90].

Flynn:1972:BCS

[Fly72a] M. J. Flynn. B72-3 computer structures, readings and examples. IEEE Transactions on Computers, C-21(1):109–110, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672037.

Flynn:1972:SCO

[Fly72b] Michael J. Flynn. Some computer organizations and their effectiveness. IEEE Transactions on Computers, C-21(9):948–960, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009071.

Friedman:1971:SAO

[FM71] A. D. Friedman and P. R. Menon. Systems of asynchronously operating modules. IEEE Transactions on Computers, C-20(1):100–104, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671681.

Freeman:1972:FTC

[FM72a] Harvey A. Freeman and Gernot Metze. Fault-tolerant computers using “Dotted Logic” redundancy techniques. IEEE Transactions on Computers, C-21(8):867–871, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009042.

Friedman:1972:CDD

[FM72b] A. D. Friedman and P. R. Menon. Comments on “Design of Diagnosable Iterative Arrays”. IEEE Transactions on Computers, C-21(5):511, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672146. See [LY71].

Friedman:1973:RCS

[FM73] A. D. Friedman and P. R. Menon. Restricted checking sequences for sequential machines. IEEE Transactions on Computers, C-22(4):397–399, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672320.

Fantauzzi:1974:MFD

[FM74] G. Fantauzzi and A. Marsella.

Multiple-fault detection and location in fan-out free combinational circuits. IEEE Transactions on Computers, C-23(1):48–55, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672369.

Fuller:1978:MLC

[FM78] S. H. Fuller and P. F. McGehearty.

Minimizing latency in CCD memories. IEEE Transactions on Computers, C-27(3):252–254, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675079.

Fukunage:1975:BBA

[FN75] K. Fukunage and P. M. Narendra. A branch and bound algorithm for computing k-nearest neighbors. IEEE Transactions on Computers, C-24(7):750–753, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672890.

Fujiwara:1975:ETS

[FNSK75] H. Fujiwara, Y. Nagao, T. Sasao, and K. Kinoshita. Easily testable sequential machines with extra inputs. IEEE Transactions on Computers, C-24(8):821–826, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672906.

Fukunaga:1971:AFI

[FO71a] K. Fukunaga and D. R. Olsen. An algorithm for finding intrinsic dimensionality of data. IEEE Transactions on Computers, C-20 (2):176–183, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671801. See comments

[Tru71].

Fukunaga:1971:AR

[FO71b] K. Fukunaga and D. R. Olsen. Authors’ reply. IEEE Transactions on Computers, C-20(12): 1615–1616, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671780.

Fukunaga:1971:TDD

[FO71c] K. Fukunaga and D. R. Olsen. A two-dimensional display for the classification of multivariate data. IEEE Transactions on Computers, C-20(8):917–923, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671964.

Fontao:1971:GMC

[Fon71] R. O. Fontao. A graphical method for checking complete monotonicity. IEEE Transactions on Computers, C-20(4):461–464, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671860.

Fosdick:1973:BAC

[Fos73] L. D. Fosdick. B73-2 approximate calculation of multiple integrals. IEEE Transactions on Computers, C-22(1):112, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672201.

Feng:1975:DPS

[FP75] H.-Y. F. Feng and T. Pavlidis. Decomposition of polygons into simpler components: Feature generation for syntactic pattern recognition. IEEE Transactions on Computers, C-24(6):636–650, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672869.

Flynn:1971:MIV

[FR71] M. J. Flynn and R. F. Rosin. Microprogramming: An introduction and a viewpoint. IEEE Transactions on Computers, C-20(7):727–731, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671934.

Foster:1972:PCE

[FR72] C. C. Foster and E. M. Riseman. Percolation of code to enhance parallel dispatching and execution. IEEE Transactions on Computers, C-21(12):1411–1415, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672108.

Fu:1976:PRI

[FR76] King-Sun Fu and A. Rosenfeld. Pattern recognition and image processing. IEEE Transactions on Computers, C-25(12):1336–1346, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674602.

Frackowiak:1972:SMH

[Fra72] J. Frackowiak. The synthesis of minimal hazardless TANT networks. IEEE Transactions on Computers, C-21(10):1099–1108, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672049.

Franklin:1978:PSO

[Fra78] M. A. Franklin. Parallel solution of ordinary differential equations. IEEE Transactions on Computers, C-27(5):413–420, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675121.

Freeman:1974:BFA

[Fre74a] P. Freeman. B 74-34 Fortran and Algol: a programmed course for students of science and technology. IEEE Transactions on Computers, C-23(5):559–560, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672583.

Freeman:1974:BOS

[Fre74b] P. Freeman. B74-13 operating systems — Infotech state of the art report #14. IEEE Transactions on Computers, C-23(2): 216–217, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672495.

French:1977:CSS

[Fre77] A. S. French. Computer simulation of space-filling molecular models. IEEE Transactions on Computers, C-26(10):1026–1028, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674738.

Freeman:1979:AGD

[Fre79a] H. Freeman. Algorithm for generating a digital straight line on a triangular grid. IEEE Transactions on Computers, C-28(2): 150–152, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675305.

Freeman:1979:AUB

[Fre79b] M. Freeman. Aspects of the upper bounds of finite input-memory and finite output-memory sequential machines. IEEE Transactions on Computers, C-28(3): 249–253, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675327.

Friedman:1970:RAS

[Fri70] A. D. Friedman. R70-17 asynchronous sequential circuits with feedback. IEEE Transactions on Computers, C-19(7):657, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671597.

Friedman:1973:ETI

[Fri73] A. D. Friedman. Easily testable iterative systems. IEEE Transactions on Computers, C-22(12): 1061–1064, December 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672244.

Friedman:1974:DSC

[Fri74] A. D. Friedman. Diagnosis of short-circuit faults in combinational circuits. IEEE Transactions on Computers, C-23(7):746–752, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672618.

Friedman:1977:RPD

[Fri77] J. H. Friedman. A recursive partitioning decision rule for nonparametric classification. IEEE Transactions on Computers, C-26(4):404–408, April 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674849.

Farber:1970:MLB

[FS70] A. S. Farber and E. S. Schlig.

Mathematical “Lower Bounds” and the logic circuit designer. IEEE Transactions on Computers, C-19 (1):80–81, January 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008905.

Foster:1971:CRA

[FS71] C. C. Foster and F. D. Stockton. Counting responders in an associative memory. IEEE Transactions on Computers, C-20 (12):1580–1583, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671768.

Franklin:1974:ART

[FS74] M. A. Franklin and A. Sen. An analytic response time model for singleand dual-density disk systems. IEEE Transactions on Computers, C-23(12):1269–1276, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672439.

Foley:1975:OSD

[FS75] D. H. Foley and J. W. Sammon, Jr. An optimal set of discriminant vectors. IEEE Transactions on Computers, C-24(3):281–289, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672801.

Fukunaga:1978:GCP

[FS78] K. Fukunaga and R. D. Short. Generalized clustering for problem localization. IEEE Transactions on Computers, C-27(2): 176–181, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675056.

Fernandez:1978:ASS

[FSLC78] E. B. Fernandez, R. C. Summers, T. Lang, and C. D. Coleman. Architectural support for system protection and database security. IEEE Transactions on Computers, C-27(8):767–771, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675189.

Friedman:1974:PPA

[FT74] J. H. Friedman and J. W. Tukey.

A projection pursuit algorithm for exploratory data analysis. IEEE Transactions on Computers, C-23(9):881–890, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672644.

Fraley:1976:RAM

[FT76] D. Fraley and J. Talavage. A realization algorithm for multiple preset experiments. IEEE Transactions on Computers, C-25 (11):1130–1132, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674564.

Fung:1979:AMC

[FT79] Kwok-Tung Fung and H. C.

Torng. On the analysis of memory conflicts and bus contentions in a multiple-microprocessor system. IEEE Transactions on Computers, C-28(1):28–37, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675219.

Fu:1974:BFT

[Fu74] K. S. Fu. B 74-30 foundation of the theory of learning systems. IEEE Transactions on Computers, C-23(5):557, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672579.

Fuller:1972:ODS

[Ful72] S. H. Fuller. An optimal drum scheduling algorithm. IEEE Transactions on Computers, C-21(11): 1153–1165, November 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672065.

Frosini:1979:AEF

[FV79] G. Frosini and F. M. Viterbo. An algorithm for evaluating the frequency of a rotating vector. IEEE Transactions on Computers, C-28(8):560–566, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675411.

Friedman:1978:AAP

[FW78] D. P. Friedman and D. S. Wise. Aspects of applicative programming for parallel processing. IEEE Transactions on Computers, C-27(4):289–296, April 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675100.

Gaddess:1970:EDB

[Gad70] Terry G. Gaddess. An errordetecting binary adder: a hardwareshared implementation. IEEE Transactions on Computers, C-19 (1):34–38, January 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008897.

Gagliardi:1972:BFG

[Gag72] U. O. Gagliardi. B72-27 the Fourth Generation. IEEE Transactions on Computers, C-21(12):1459, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672122.

Gaines:1979:MGI

[Gai79] B. R. Gaines. Maryanski’s grammatical inferencer. IEEE Transactions on Computers, C-28(1): 62–64, January 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675224.

Gajewski:1975:PEF

[Gaj75] T. Gajewski. On the periodic equivalents of finite automata. IEEE Transactions on Computers, C-24(10):991–994, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672700.

Galperin:1978:LPC

[Gal78] E. A. Galperin. Loop properties and controllability of linear congruential sequences. IEEE Transactions on Computers, C-27(1):68–76, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674954.

Gardner:1971:FMM

[Gar71] P. L. Gardner. Functional memory and its microprogramming implications. IEEE Transactions on Computers, C-20(7):764–775, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671938.

Garey:1972:RBC

[Gar72a] M. R. Garey. Resident-bubble cellular logic using magnetic domains. IEEE Transactions on Computers, C-21(4):392–396, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008982.

Garey:1972:SBI

[Gar72b] M. R. Garey. Simple binary identification problems. IEEE Transactions on Computers, C-21(6):588–590, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009013.

Gardiner:1974:CAI

[Gar74] A. B. Gardiner. Comments on “An Augmented Iterative Array for High Speed Division”. IEEE Transactions on Computers, C-23 (3):326–327, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672524. See [CH73a].

Garner:1976:SSR

[Gar76] H. L. Garner. A survey of some recent contributions to computer arithmetic. IEEE Transactions on Computers, C-25(12): 1277–1282, December 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674595.

Garner:1978:TCA

[Gar78] H. L. Garner. Theory of computer addition and overflows. IEEE Transactions on Computers, C-27(4):297–301, April 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675101.

Gay:1978:EMS

[Gay78] F. A. Gay. Evaluation of maintenance software in real-time systems. IEEE Transactions on Computers, C-27(6):576–582, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675151.

Grzymala-Busse:1971:DPR

[GB71a] J. W. Grzymala-Busse. On the decomposition of periodic representations of sequential machines. IEEE Transactions on Computers, C-20(8):929–933, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671967.

Grzymala-Busse:1971:PRP

[GB71b] J. W. Grzymala-Busse. Periodic representations and T-partitionable equivalents of sequential machines. IEEE Transactions on Computers, C-20(2):190–198, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671803.

Ghosh:1972:SMT

[GBMC72] S. Ghosh, S. Bandyopadhyay, S. K. Mitra, and A. K. Choudhury. Simple methods for the testing of 2-summability of Boolean functions and isobaricity of threshold functions. IEEE Transactions on Computers, C-21(5):503–507, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672144.

Ghosh:1971:CMN

[GC71a] S. Ghosh and A. K. Choudhury. Cascaded multithreshold networks. IEEE Transactions on Computers, C-20(6):655–662, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671910.

Griswold:1971:LSC

[GC71b] V. M. Griswold and C. C. Carroll. Large-scale circuit interconnection for Boolean function implementation. IEEE Transactions on Computers, C-20(5):572–575, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671885.

Ghosh:1973:PBF

[GC73] S. Ghosh and A. K. Choudhury. Partition of Boolean functions for realization with multithreshold threshold logic elements. IEEE Transactions on Computers, C-22(2):204–215, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672279.

Geadah:1977:NDS

[GC77] Y. A. Geadah and M. J. G.

Corinthios. Natural, dyadic, and sequency order algorithms and processors for the Walsh–Hadamard transform. IEEE Transactions on Computers, C-26(5):435–442, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674860.

Gupta:1974:LTR

[GCC74] A. S. Gupta, D. K. Chattopadhyay, and A. K. Choudhury. Linearity testing and realization of sequential machines. IEEE Transactions on Computers, C-23(5):524–528, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672567.

Gupta:1977:RFT

[GCP^{+}77] A. Sen Gupta, D. K. Chattopadhyay, A. Palit, A. K. Bandyopadhyay, M. S. Basu, and A. K. Choudhury. Realization of faulttolerant and fail-safe sequential machines. IEEE Transactions on Computers, C-26(1):91–96, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009281.

Goke:1972:DRP

[GD72] L. R. Goke and K. L. Doty. Design of a random-pulse computer for classifying binary patterns. IEEE Transactions on Computers, C-21(12):1347–1354, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672099.

Gupta:1973:FSM

[GD73] P. K. Gupta and D. L. Dietmeyer.

Fast state minimization of incompletely specified sequential machines. IEEE Transactions on Computers, C-22(2):215–217, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672280.

Gecsei:1977:INT

[Gec77] J. Gecsei. Interconnection networks from three state cells. IEEE Transactions on Computers, C-26 (8):705–711, August 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674908.

Gelenbe:1971:RMS

[Gel71] S. E. Gelenbe. A realizable model for stochastic sequential machines. IEEE Transactions on Computers, C-20(2):199–204, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671804.

Gelenbe:1973:BTS

[Gel73a] E. Gelenbe. B73-10 time sharing systems. IEEE Transactions on Computers, C-22(5):546, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672357.

Gelenbe:1973:UAE

[Gel73b] Erol Gelenbe. A unified approach to the evaluation of a class of replacement algorithms. IEEE Transactions on Computers, C-22(6):611–618, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=5009115;   http://
                                        www.acsel-lab.com/arithmetic/
                                        arith2/papers/ARITH2_Gelenbe.
                                        pdf.

Gelperin:1976:RBP

[Gel76] D. Gelperin. A resolution-based proof procedure using deletiondirected search. IEEE Transactions on Computers, C-25(4):323–327, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674612.

Geoffrion:1970:RAA

[Geo70] A. M. Geoffrion. R70-4 analysis of algorithms for the zero-one programming problem. IEEE Transactions on Computers, C-19(5):464, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671548.

Gestri:1970:SMS

[Ges70] G. Gestri. Synthesis of multiple sequential machines having different inputs. IEEE Transactions on Computers, C-19(11):

1105–1108, November 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671434.

Geselowitz:1971:UME

[Ges71] D. B. Geselowitz. Use of the multipole expansion to extract significant features of the surface electrocardiogram. IEEE Transactions on Computers, C-20(9):1086–1089, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672000.

Gupta:1978:WSP

[GF78] R. K. Gupta and M. A. Franklin.

Working set and page fault frequency paging algorithms: a performance comparison. IEEE Transactions on Computers, C-27(8): 706–712, August 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675177.

Gabrielian:1973:SSA

[GG73] A. Gabrielian and S. Ginsburg.

Structured storage AFA. IEEE Transactions on Computers, C-22(5):534–537, May 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672351.

Gibson:1975:SCT

[GG75] J. A. Gibson and R. W. Gibbard. Synthesis and comparison of two’s complement parallel multipliers. IEEE Transactions on Computers, C-24(10):1020–1027, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672710.

Gerace:1971:TSP

[GGMM71] G. B. Gerace, L. Gilli, P. Maestrini, and A. R. Meo. TOPI — a special-purpose computer for Boolean analysis and synthesis. IEEE Transactions on Computers, C-20(8):837–842, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671951.

Govindan:1972:CIT

[GH72] G. N. Govindan and R. A. Higgins. Convergence of an identification technique for nonlinear systems. IEEE Transactions on Computers, C-21(11):1216–1219, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672072.

Gaitanis:1978:NDR

[GH78] N. Gaitanis and C. Halatsis. A new double-rank realization of sequential machines. IEEE Transactions on Computers, C-27(12): 1137–1143, December 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675016.

Garey:1977:ASP

[GHJ77] M. R. Garey, F. K. Hwang, and D. S. Johnson. Algorithms for a set partitioning problem arising in the design of multipurpose units. IEEE Transactions on Computers, C-26(4):321–328, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674841.

Gill:1970:SCM

[Gil70a] A. Gill. Single-channel and multichannel finite-state machines. IEEE Transactions on Computers, C-19(11):1073–1078, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671427. See comments

[Ric72].

Gilliland:1970:RSC

[Gil70b] M. C. Gilliland. R70-7 stability controls for the analysis of analog/ digital hybrid loops. IEEE Transactions on Computers, C-19(5):465, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671551.

Gillespie:1974:BPR

[Gil74a] R. C. Gillespie. B74-19 pattern recognition techniques. IEEE Transactions on Computers, C-23 (2):220–222, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672501.

Gillespie:1974:BCO

[Gil74b] R. C. Gillespie. B74-5 computeroriented approaches to pattern recognition. IEEE Transactions on Computers, C-23(1):112, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672388.

Gitman:1970:ADU

[GL70] I. Gitman and M. D. Levine. An algorithm for detecting unimodal fuzzy sets and its application as a clustering technique. IEEE Transactions on Computers, C-19(7):583–593, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671585.

Giloi:1974:FDS

[GL74] W. K. Giloi and H. Liebig. A formalism for description and synthesis of logical algorithms and their hardware implementation. IEEE Transactions on Computers, C-23(9):897–906, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672646.

Glassman:1970:GFF

[Gla70] J. A. Glassman. A generalization of the Fast Fourier Transform. IEEE Transactions on Computers, C-19(2):105–116, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671468.

Gupta:1979:OSC

[GLL79] U. I. Gupta, D. T. Lee, and J. Y.- T. Leung. An optimal solution for the channel-assignment problem. IEEE Transactions on Computers, C-28(11):807–810, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675260.

Gold:1971:FFP

[GLMR71] B. Gold, I. L. Lebow, P. G.

McHugh, and C. M. Rader. The FDP, a fast programmable signal processor. IEEE Transactions on Computers, C-20(1):33–38, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671671.

Givone:1971:MSM

[GLR71] D. D. Givone, M. E. Liebler, and R. P. Roesser. A method of solution for multiple-valued logic expressions. IEEE Transactions on Computers, C-20(4):464–467, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671861.

Glucksman:1971:MCP

[Glu71] H. A. Glucksman. Multicategory classification of patterns represented by high-order vectors of multilevel measurements. IEEE Transactions on Computers, C-20 (12):1593–1598, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671772.

Goldberg:1974:OHS

[GLW74] J. Goldberg, K. N. Levitt, and J. H. Wensley. An organization for a highly survivable memory. IEEE Transactions on Computers, C-23(7):693–705, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672610.

Grasselli:1970:MRO

[GM70] A. Grasselli and U. Montanari. On the minimization of READ-ONLY memories in microprogrammed digital computers. IEEE Transactions on Computers, C-19 (11):1111–1114, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671436.

Gray:1971:LFC

[GM71] F. G. Gray and J. F. Meyer. Locatability of faults in combinational networks. IEEE Transactions on Computers, C-20(11): 1407–1412, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671740.

Gibson:1972:FHC

[GM72a] John A. Gibson and Terence W.

Marks. Fast hybrid computer implementation of the Dynostat algorithm. IEEE Transactions on Computers, C-21(8):872–880, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009043.

Gose:1972:AR

[GM72b] Earl E. Gose and Anthony N. Mucciardi. Authors’ reply. IEEE Transactions on Computers, C-21 (9):1029, September 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009090.

Gray:1976:APF

[GM76] F. G. Gray and J. F. Meyer. Algebraic properties of functions affecting optimum fault-tolerant realizations. IEEE Transactions on Computers, C-25(11):1078– 1088, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674558.

Godbole:1971:AMC

[God71] S. S. Godbole. An algorithm for the machine computation of partial-fractions expansion of functions having multiple poles. IEEE Transactions on Computers, C-20(10):1147–1152, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671692.

Godbole:1973:ECM

[God73] Sadashiva S. Godbole. On efficient computation of matrix chain products. IEEE Transactions on Computers, C-22(9):864–866, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009182.

Golay:1972:SDL

[Gol72] Marcel J. E. Golay. Smoothing of data by least squares procedures and by filtering. IEEE Transactions on Computers, C-21 (3):299–301, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008953.

Golumbic:1976:CM

[Gol76] M. C. Golumbic. Combinatorial merging. IEEE Transactions on Computers, C-25(11): 1164–1167, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674574.

Golay:1978:T

[Gol78] M. J. E. Golay. Topoglyphs.

IEEE Transactions on Computers, C-27(2):164–167, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675051.

Gonenc:1970:MDF

[Gon70] G. Gonenc. A method for the design of fault detection experiments. IEEE Transactions on Computers, C-19(6):551–558, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671568.

Gonzalez:1979:NOS

[Gon79] T. Gonzalez. A note on open shop preemptive schedules. IEEE Transactions on Computers, C-28 (10):782–786, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675246.

Goodman:1970:RRT

[Goo70] E. D. Goodman. R70-18 real-time computation by n-dimensional iterative arrays of finite-state machines. IEEE Transactions on Computers, C-19(7):657–658, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671598.

Good:1971:RBT

[Goo71] I. J. Good. The relationship between two Fast Fourier Transforms. IEEE Transactions on Computers, C-20(3):310–317, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671829.

Gorenstein:1975:BSS

[Gor75] S. Gorenstein. B75-2 systems simulation. IEEE Transactions on Computers, C-24(1):110–111, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672687.

Gouraud:1971:CDC

[Gou71] Henri Gouraud. Continuous shading of curved surfaces. IEEE Transactions on Computers, C-20(6):623–629, June 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671906.

Gourishankar:1977:CAA

[Gou77] V. Gourishankar. Comments on “An Additional Algorithm for a System of Coupled Algebraic Matrix Riccati Equations”. IEEE Transactions on Computers, C-26(3):320, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674836. See [Den76].

Goyal:1976:NAR

[Goy76] L. N. Goyal. A note on Atrubin’s real-time iterative multiplier. IEEE Transactions on Computers, C-25(5):546–548, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674646.

Grewal:1974:BFE

[GP74] K. S. Grewal and J. D. Patterson. A binary feature extraction technique. IEEE Transactions on Computers, C-23(5):545–549, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672572.

Gelenbe:1971:UMR

[GR71a] S. E. Gelenbe and N. Rossi. Uniform modular realizations and linear machines. IEEE Transactions on Computers, C-20(12): 1616–1617, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671781.

Gonzalez:1971:PSP

[GR71b] M. J. Gonzalez, Jr. and C. V. Ramamoorthy. Program suitability for parallel processing. IEEE Transactions on Computers, C-20(6):647–654, June 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671909.

Givone:1972:MLI

[GR72a] D. D. Givone and R. P. Roesser. Multidimensional linear iterative circuits-general properties. IEEE Transactions on Computers, C-21(10):1067–1073, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672046.

Gonzalez:1972:PTE

[GR72b] M. J. Gonzalez, Jr. and C. V. Ramamoorthy. Parallel task execution in a decentralized system. IEEE Transactions on Computers, C-21(12):1310–1322, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672095.

Givone:1973:AR

[GR73a] D. D. Givone and R. P. Roesser. Authors’ reply. IEEE Transactions on Computers, C-22(12):1140, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672254.

Givone:1973:MML

[GR73b] Donald D. Givone and Robert P.

Roesser. Minimization of multidimensional linear iterative circuits. IEEE Transactions on Computers, C-22(7):673–678, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009134.

Grado:1970:REI

[Gra70] G. Grado. R70-5 the effect of IC’s on the design of analog computers. IEEE Transactions on Computers, C-19(5):464–465, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671549.

Gray:1971:LPB

[Gra71] S. B. Gray. Local properties of binary images in two dimensions. IEEE Transactions on Computers, C-20(5):551–561, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671882.

Granlund:1972:FPH

[Gra72] G. H. Granlund. Fourier preprocessing for hand print character recognition. IEEE Transactions on Computers, C-21(2):195–201, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008926.

Gravitz:1973:NMT

[Gra73] M. E. Gravitz. A note on a modified ternary simulator capable of initializing all fault machine memory elements. IEEE Transactions on Computers, C-22 (11):1042–1044, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672236.

Greibach:1975:RCS

[Gre75] S. A. Greibach. A remark on code sets and context-free languages. IEEE Transactions on Computers, C-24(7):741–742, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672887.

Griffith:1973:EDS

[Gri73] A. K. Griffith. Edge detection in simple scenes using a priori information. IEEE Transactions on Computers, C-22(4):371–381, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672317.

Griswold:1974:BEP

[Gri74] R. E. Griswold. B74-44 the elements of programming style. IEEE Transactions on Computers, C-23(12):1329, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672458.

Griffith:1976:EEP

[Gri76] Arnold K. Griffith. Empirical exploration of the performance of the alpha beta tree-searching heuristic. IEEE Transactions on Computers, C-25(1):6–11, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009198.

Griss:1978:UES

[Gri78a] Martin L. Griss. Using an efficient sparse minor expansion algorithm to compute polynomial subresultants and the greatest common denominator. IEEE Transactions on Computers, C-27(10): 945–950, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674974.

Gritsenko:1978:CEC

[Gri78b] V. M. Gritsenko. Comments on “Error-Correcting Codes for Byte-Organized Arithmetic Processors”. IEEE Transactions on Computers, C-27(2):189–190, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675061. See [NR75].

Gault:1972:MFD

[GRR72] J. W. Gault, J. P. Robinson, and S. M. Reddy. Multiple fault detection in combinational networks. IEEE Transactions on Computers, C-21(1):31–36, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672021.

Gault:1974:AR

[GRR74] J. W. Gault, J. P. Robinson, and S. M. Reddy. Authors’ reply. IEEE Transactions on Computers, C-23(3):334, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672531.

Goel:1974:CMF

[GS74] P. Goel and D. P. Siewiorek. Comments on “Multiple Fault Detection in Combinational Networks”. IEEE Transactions on Computers, C-23(3):333–334, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672530. See [DW73b].

Gelenbe:1979:AUS

[GS79] E. Gelenbe and K. Sevcik. Analysis of update synchronization for multiple copy data bases. IEEE Transactions on Computers, C-28(10):737–747, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675241.

Gilbert:1976:PDM

[GSB^{+}76] B. K. Gilbert, M. T. Storma, K. C.

Ballard, L. W. Hobrock, C. E. James, and E. H. Wood. A programmable dynamic memory allocation system for input/output of digital data into standard computer memories at 40 megasamples/s. IEEE Transactions on Computers, C-25(11):1101–1109, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674560.

Gilbert:1976:RTH

[GSJ^{+}76] B. K. Gilbert, M. T. Storma, C. E. James, L. W. Hobrock, E. S. Yang, K. C. Ballard, and E. H. Wood. A real-time hardware system for digital processing of wide-band video images. IEEE Transactions on Computers, C-25(11):1089–1100, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674559.

Gray:1979:DFM

[GST79] F. G. Gray, L. C. C. Shih, and R. A.

Thompson. Diagnosis of faults in modular trees. IEEE Transactions on Computers, C-28(5):342–353, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675364.

Gray:1974:RRC

[GT74] F. G. Gray and R. A. Thompson. Reconfiguration for repair in a class of universal logic modules. IEEE Transactions on Computers, C-23(11):1185–1194, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672421.

Gray:1978:FDB

[GT78a] F. G. Gray and R. A. Thompson. Fault detection in bilateral arrays of combinational cells. IEEE Transactions on Computers, C-27(12):1206–1213, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675029.

Greer:1978:CLD

[GT78b] C. R. Greer and R. A. Thompson. Combinational logic design with decoders. IEEE Transactions on Computers, C-27(9): 869–875, September 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675207.

Guffin:1971:MSC

[Guf71] R. M. Guffin. Microdiagnostics for the Standard Computer MLP-900 processor. IEEE Transactions on Computers, C-20(7):803–808, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671943.

Groginsky:1970:PFF

[GW70] H. L. Groginsky and G. A. Works.

A pipeline Fast Fourier Transform. IEEE Transactions on Computers, C-19(11):1015–1019, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671419.

Greguss:1975:UHF

[GW75] P. Greguss and W. Waigelich. Ultrasonic holographic Fourier spectroscopy via Optical Fourier Transforms. IEEE Transactions on Computers, C-24(4):412–418, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672829.

Gaines:1977:SAA

[GW77] B. R. Gaines and I. H. Witten. Stability and admissibility of adaptive threshold logic convergence. IEEE Transactions on Computers, C-26(1):88–91, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009280.

Garnett:1977:NEB

[GY77] James M. Garnett and Stephen S. Yau. Nonparametric estimation of the Bayes error of feature extractors using ordered nearest neighbor sets. IEEE Transactions on Computers, C-26(1):46–54, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009273.

Hornbuckle:1970:LMA

[HA70] G. D. Hornbuckle and E. I. Ancona. The LX-1 microprocessor and its application to real-time signal processing. IEEE Transactions on Computers, C-19(8):710–720, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671614;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Hornbuckle.
pdf.

Hakimi:1974:CCA

[HA74] S. L. Hakimi and A. T. Amin. Characterization of connection assignment of diagnosable systems. IEEE Transactions on Computers, C-23(1):86–88, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672375.

Huertas:1976:SSA

[HA76] J. L. Huertas and J. I. Acha. Selfsynchronization of asynchronous sequential circuits employing a general clock function. IEEE Transactions on Computers, C-25 (3):297–300, March 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009256.

Hou:1977:LSI

[HA77] Hsieh S. Hou and Harry C. Andrews. Least squares image restoration using spline basis functions. IEEE Transactions on Computers, C-26(9):856–873, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674934.

Habibi:1977:FSW

[Hab77] A. Habibi. Fast suboptimal Wiener filtering of Markov sequences. IEEE Transactions on Computers, C-26(5):443–449, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674861.

Hall:1974:AUD

[Hal74] E. L. Hall. Almost uniform distributions for computer image enhancement. IEEE Transactions on Computers, C-23(2):207–208, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672485.

Hamacher:1971:MCV

[Ham71] V. C. Hamacher. Machine complexity versus interconnection complexity in iterative arrays. IEEE Transactions on Computers, C-20 (3):321–323, March 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671831.

Hampel:1973:MTG

[Ham73] D. Hampel. MultiFunction threshold gates. IEEE Transactions on Computers, C-22(2): 197–203, February 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672278.

Hamacher:1974:BCH

[Ham74] V. C. Hamacher. B 74-28 computer hardware and software: An interdisciplinary introduction. IEEE Transactions on Computers, C-23(5):556, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672577.

Hansalik:1972:LBC

[Han72] W. E. Hansalik. A lower bound on the complexity of arbitrary switching function realizers. IEEE Transactions on Computers, C-21(5):507–510, May 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672145.

Handoko:1975:DTA

[Han75] F. Handoko. A discussion on two algorithms for determining maximum compatibles. IEEE Transactions on Computers, C-24(8):838–840, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672911.

Harrison:1970:BTF

[Har70a] M. A. Harrison. B70-1 truth functions and the problem of their realization by two-terminal graphs. IEEE Transactions on Computers, C-19(5):462, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671544.

Hartmanis:1970:RNC

[Har70b] J. Hartmanis. R70-1 A note on computing time for the recognition of context-free languages by a singletape Turing machine. IEEE Transactions on Computers, C-19(5):463, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671545.

Harr:1971:BAU

[Har71] J. A. Harr. B715 APL users’ guide. IEEE Transactions on Computers, C-20(10):1222–1223, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671707.

Harada:1972:SPN

[Har72] K. Harada. Sequential permutation networks. IEEE Transactions on Computers, C-21(5):472–479, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672136.

Harrison:1973:NCB

[Har73] M. A. Harrison. On the number of classes of binary matrices. IEEE Transactions on Computers, C-22(12):1048–1052, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672242. See comments

[Sag77].

Harrison:1974:BNC

[Har74a] T. J. Harrison. B74-23 numerical control — mathematics and applications. IEEE Transactions on Computers, C-23(3):335, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672532.

Hart:1974:CPA

[Har74b] J. M. Hart. A class of pushdown acceptors for adjunct languages. IEEE Transactions on Computers, C-23(5):449–455, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672559.

Haralick:1976:SEW

[Har76] R. M. Haralick. A storage efficient way to implement the Discrete Cosine Transform. IEEE Transactions on Computers, C-25(7):764–765, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674687.

Hassab:1973:SSC

[Has73] Joseph C. Hassab. A simply solvable class of simultaneous linear equations. IEEE Transactions on Computers, C-22(9): 866–867, September 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009183.

Hassab:1974:CPC

[Has74] J. C. Hassab. The characteristic polynomials for a class of eigenvalue problems. IEEE Transactions on Computers, C-23(9):967, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672654.

Hayes:1971:NMF

[Hay71a] J. P. Hayes. A NAND model for fault diagnosis in combinational logic networks. IEEE Transactions on Computers, C-20 (12):1496–1506, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671755.

Hayes:1971:RBF

[Hay71b] J. P. Hayes. On realizations of Boolean functions requiring a minimal or near-minimal number of tests. IEEE Transactions on Computers, C-20(12): 1506–1513, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671756.

Hayes:1974:MLN

[Hay74] J. P. Hayes. On modifying logic networks to improve their diagnosability. IEEE Transactions on Computers, C-23(1):56–62, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672370.

Hayes:1975:DPS

[Hay75] J. P. Hayes. Detection of pattern-sensitive faults in randomaccess memories. IEEE Transactions on Computers, C-24(2): 150–157, February 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672775.

Hayes:1976:GMF

[Hay76a] J. P. Hayes. A graph model for fault-tolerant computing systems. IEEE Transactions on Computers, C-25(9):875–884, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674712.

Hayes:1976:PIL

[Hay76b] J. P. Hayes. On the properties of irredundant logic networks. IEEE Transactions on Computers, C-25(9):884–892, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674713.

Hayes:1976:TCT

[Hay76c] J. P. Hayes. Transition count testing of combinational logic circuits. IEEE Transactions on Computers, C-25(6):613–620, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674661.

Hayes:1978:GOT

[Hay78a] J. P. Hayes. Generation of optimal transition count tests. IEEE Transactions on Computers, C-27 (1):36–41, January 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674950.

Hayes:1978:PCL

[Hay78b] J. P. Hayes. Path complexity of logic networks. IEEE Transactions on Computers, C-27(5):459–462, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675127.

Hulina:1970:TIS

[HB70] P. T. Hulina and J. G. Bredeson. On Tracey’s internal state assignment method. IEEE Transactions on Computers, C-19(5):458, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671539.

Hsiao:1975:OLS

[HB75a] M. Y. Hsiao and D. C. Bossen. Orthogonal latin square configuration for LSI memory yield and reliability enhancement. IEEE Transactions on Computers, C-24(5):512–516, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672847.

Hunt:1975:SDC

[HB75b] B. R. Hunt and J. R. Breedlove. Scan and display considerations in processing images by digital computer. IEEE Transactions on Computers, C-24(8):848–853, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672915.

Heath:1970:SPC

[HC70] J. R. Heath and C. C. Carroll.

Special-purpose computer organization for double-precision realization of digital filters. IEEE Transactions on Computers, C-19 (12):1146–1152, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671446.

Harlow:1971:IES

[HC71a] C. Harlow and C. L. Coates, Jr.

Inessential errors in sequential machines. IEEE Transactions on Computers, C-20(6):688–690, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671921.

Hsiao:1971:BDF

[HC71b] M. Y. Hsiao and D. K. Chia.

Boolean difference for fault detection in asynchronous sequential machines. IEEE Transactions on Computers, C-20(11): 1356–1361, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671731.

Harlow:1972:FSM

[HC72] Charles A. Harlow and Clarence L.

Coates. Feedback in sequential machine realizations. IEEE Transactions on Computers, C-21(4):371–381, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008978.

Ho:1973:MAR

[HC73] Irving T. Ho and Tien Chi Chen. Multiple addition by residue threshold functions and their representation by array logic. IEEE Transactions on Computers, C-22 (8):762–767, August 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009155.

Hamacher:1974:AR

[HC74] V. C. Hamacher and M. Cappa. Authors’ reply. IEEE Transactions on Computers, C-23(3):327, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672525.

Hemmati:1978:ACA

[HC78] F. Hemmati and D. J. Costello, Jr. An algebraic construction for q-ary shift register sequences. IEEE Transactions on Computers, C-27(12):1192–1195, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675025.

Hackbart:1971:AEF

[HD71] R. R. Hackbart and D. L. Dietmeyer. The avoidance and elimination of function hazards in asynchronous sequential circuits. IEEE Transactions on Computers, C-20(2):184–189, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671802.

Harlow:1973:ARI

[HE73] Charles A. Harlow and Sharon A.

Eisenbeis. The analysis of radiographic images. IEEE Transactions on Computers, C-22(7):678–689, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009135.

Hikita:1979:NMS

[HE79] T. Hikita and H. Enomoto. On the number of multivalued switching functions realizable by cascades. IEEE Transactions on Computers, C-28(5):371–374, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675369.

Hecht:1973:FMF

[Hec73] H. Hecht. Figure of merit for fault-tolerant space computers. IEEE Transactions on Computers, C-22(3):246–251, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672295.

Hedetniemi:1970:RMA

[Hed70] S. Hedetniemi. R70-36 maximin automata. IEEE Transactions on Computers, C-19(9):859, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671657.

Heines:1979:BBC

[Hei79] T. S. Heines. Buffer behavior in computer communication systems. IEEE Transactions on Computers, C-28(8):573–576, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675413.

Hellerman:1972:MCW

[Hel72] L. Hellerman. A measure of computational work. IEEE Transactions on Computers, C-21(5):439–446, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672132. See comments

[Wel74] and discussion [Hel74].

Hellerman:1974:DMC

[Hel74] L. Hellerman. A discussion of “A Measure of Computational Work”. IEEE Transactions on Computers, C-23(2):208–211, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672487. See [Hel72,

Wel74, Mas78].

Henderson:1970:CCF

[Hen70] K. W. Henderson. Comment on “Computation of the Fast Walsh– Fourier Transform”. IEEE Transactions on Computers, C-19(9): 850–851, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671647. See [Lec70a].

Henderson:1971:PIR

[Hen71] K. W. Henderson. Previous investigation of realization of an arbitrary switching function with a network of threshold and parity elements. IEEE Transactions on Computers, C-20(8):941–942, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671970.

Henschen:1976:ITR

[Hen76a] L. J. Henschen. Introduction tutorial on resolution. IEEE Transactions on Computers, C-25(8): 769–772, August 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674695.

Henschen:1976:SRH

[Hen76b] L. J. Henschen. Semantic resolution for Horn sets. IEEE Transactions on Computers, C-25 (8):816–822, August 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674700.

Henderson:1978:SSF

[Hen78] K. W. Henderson. Suggested standard forms for certain real Hadamard matrices. IEEE Transactions on Computers, C-27(5): 469–471, May 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675130.

Hall:1972:OMR

[HEWB72] D. J. Hall, R. M. Endlich, D. E. Wolf, and A. E. Brain. Objective methods for registering landmarks and determining cloud motions from satellite data. IEEE Transactions on Computers, C-21(7):768–776, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672173.

Heydorn:1971:RFE

[Hey71] R. P. Heydorn. Redundancy in feature extraction. IEEE Transactions on Computers, C-20(9): 1051–1054, September 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671994.

Henrichon:1970:AR

[HF70] E. G. Henrichon, Jr. and K. S. Fu.

Authors’ reply. IEEE Transactions on Computers, C-19(4):363–364, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671521.

Hoffman:1971:PRS

[HF71] R. L. Hoffman and K. Fukunaga. Pattern recognition signal processing for mechanical diagnostics signature analysis. IEEE Transactions on Computers, C-20(9):1095–1100, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672002.

Hallin:1972:PAF

[HF72] Thomas G. Hallin and Michael J.

Flynn. Pipelining of arithmetic functions. IEEE Transactions on Computers, C-21(8):880–886, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009044.

Hayes:1974:TPP

[HF74] J. P. Hayes and A. D. Friedman.

Test point placement to simplify fault detection. IEEE Transactions on Computers, C-23(7):727–735, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672614.

Halatsis:1978:INF

[HG78a] C. Halatsis and N. Gaitanis. Irredundant normal forms and minimal dependence sets of a Boolean function. IEEE Transactions on Computers, C-27(11): 1064–1068, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674997.

Halatsis:1978:MCS

[HG78b] C. Halatsis and N. Gaitanis. On the minimization of the control store in microprogrammed computers. IEEE Transactions on Computers, C-27(12):1189–1192, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675024.

Halatsis:1979:PFS

[HG79] C. Halatsis and N. Gaitanis. Positive fail-safe realization of synchronous sequential machines. IEEE Transactions on Computers, C-28(2):167–172, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675309.

Huskey:1976:CCD

[HH76] H. D. Huskey and V. R. Huskey. Chronology of computing devices.

IEEE Transactions on Computers, C-25(12):1190–1199, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674587.

Hill:1977:SSS

[HH77] F. J. Hill and B. Huey. SCIRTSS: a search system for sequential circuit test sequences. IEEE Transactions on Computers, C-26(5):490–502, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674866.

Henle:1976:ATT

[HHJ^{+}76] R. A. Henle, I. T. Ho, W. S.

Johnson, W. D. Pricer, and J. L. Walsh. The application of transistor technology to computers. IEEE Transactions on Computers, C-25(12):1289–1303, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674597.

Hu:1972:ATB

[HI72] Ming-Kuei Hu and A. Iosupovicz.

Analysis of the terminal behavior of some classes of iterative arrays of linear machines. IEEE Transactions on Computers, C-21 (12):1394–1399, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672105.

Hight:1971:MIS

[Hig71] S. L. Hight. Minimal input solutions. IEEE Transactions on Computers, C-20(8):923–925, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671965.

Hight:1973:CDD

[Hig73] S. L. Hight. Complex disjunctive decomposition of incompletely specified Boolean functions. IEEE Transactions on Computers, C-22(1):103–110, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672197.

Higbie:1976:VFP

[Hig76] L. C. Higbie. Vector floating-point data format. IEEE Transactions on Computers, C-25(1):25–32, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009201.

Higbie:1978:OOM

[Hig78] L. C. Higbie. Overlapped operation with microprogramming. IEEE Transactions on Computers, C-27(3):270–275, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675085.

Hill:1979:RBL

[Hil79] J. S. Hill. Revision of the buffer length derivation for a modified ek/ D/1 systems by maritsas and Hartley. IEEE Transactions on Computers, C-28(3):269–273, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675336.

Hutton:1974:SSA

[HK74] D. W. Hutton and W. H. Kautz. A simplified summation array for cellular logic modules. IEEE Transactions on Computers, C-23(2): 203–206, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672483.

Higuchi:1977:SHF

[HK77] T. Higuchi and M. Kameyama.

Static-hazard-free T -gate for ternary memory element and its application to ternary counters. IEEE Transactions on Computers, C-26 (12):1212–1221, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674782. See correction

[HK78].

Higuchi:1978:CSH

[HK78] T. Higuchi and M. Kameyama.

Correction to “Static-Hazard-Free T -Gate for Ternary Memory Element and Its Application to Ternary Counters”. IEEE Transactions on Computers, C-27(12):1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675035. See [HK77].

Hall:1971:SPF

[HKD^{+}71] E. L. Hall, R. P. Kruger, S. J. Dwyer, D. L. Hall, R. W. Mclaren, and G. S. Lodwick. A survey of preprocessing and feature extraction techniques for radiographic images. IEEE Transactions on Computers, C-20(9):1032–1044, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671992.

Hongyuan:1986:CSL

[HL86] Wang Hongyuan and S. C. Lee. Comments on “Sign/Logarithm Arithmetic for FFT Implementation”. IEEE Transactions on Computers, C-35(5):482–484, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1676792. See [SCNS83].

Hlavicka:1970:EHC

[Hla70] J. Hlavicka. Essential hazard correction without the use of delay elements. IEEE Transactions on Computers, C-19(3):232–238, March 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671495.

Hlawiczka:1978:CPE

[Hla78] A. Hlawiczka. Comments on “Procedures for Eliminating Static and Dynamic Hazards in Test Generation”. IEEE Transactions on Computers, C-27(2):191, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675064. See [BH74b].

Hall:1970:GPQ

[HLD70] E. L. Hall, D. D. Lynch, and S. J.

Dwyer. Generation of products and quotients using approximate binary logarithms for digital filtering applications. IEEE Transactions on Computers, C-19(2):97–105, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671467.

Hashimoto:1971:PG

[HN71] A. Hashimoto and K. Noshita.

A property of N-graphs. IEEE Transactions on Computers, C-20 (1):95–97, January 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671678.

Hao:1973:STM

[HN73] H. T. Hao and M. M. Newborn. A study of trigger machines. IEEE Transactions on Computers, C-22(12):1123–1131, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672251.

Harao:1978:SDP

[HN78] M. Harao and S. Noguchi. On some dynamical properties of finite cellular automaton. IEEE Transactions on Computers, C-27(1):42–52, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674951.

Hagler:1975:POD

[HND75] M. O. Hagler, T. G. Newman, and H. Dirilten. A programmable optical-digital scanner-processor for automated two-dimensional data analysis. IEEE Transactions on Computers, C-24(10): 1036–1038, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672716.

Hong:1972:CBF

[HO72] S. J. Hong and D. L. Ostapko. On complementation of Boolean functions. IEEE Transactions on Computers, C-21(9):1022, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009083.

Hakkala:1974:DEA

[HO74] L. Hakkala and L. Ojala. Discretization error analysis in linear DDA connections. IEEE Transactions on Computers, C-23(9):932–941, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672650.

Hartimo:1979:CST

[HO79] I. Hartimo and L. Ojala. Comments on “A Stable, Two-Multiplies-per-Cycle Algorithm for Digital Generation of Sinusoids in Real Time”. IEEE Transactions on Computers, C-28(4):300, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675349. See [Sil77].

Hoagland:1976:MRS

[Hoa76] A. S. Hoagland. Magnetic recording storage. IEEE Transactions on Computers, C-25(12): 1283–1288, December 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674596.

Hobbs:1973:BCT

[Hob73] L. C. Hobbs. B73-6 computing terminals: International computer state of the art report. IEEE Transactions on Computers, C-22(4):429–430, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672331.

Hoevel:1974:IDE

[Hoe74] L. W. Hoevel. “ideal” directly executed languages: An analytical argument for emulation. IEEE Transactions on Computers, C-23(8):759–767, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672625.

Hoel:1976:SVL

[Hoe76] Jeffrey H. Hoel. Some variations of Lee’s algorithm. IEEE Transactions on Computers, C-25(1):19–24, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009200.

Holborow:1972:IBL

[Hol72a] C. E. Holborow. An improved bound on the length of checking experiments for sequential machines with counter cycles. IEEE Transactions on Computers, C-21(6):597–598, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009016.

Holst:1972:STC

[Hol72b] P. A. Holst. Symbolic treatment of certain selector functions. IEEE Transactions on Computers, C-21(5):486–488, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672138.

Holden:1976:TAI

[Hol76] A. D. C. Holden. Trends in artificial intelligence. IEEE Transactions on Computers, C-25(4):313–316, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674610.

Hollaar:1979:DLM

[Hol79] L. A. Hollaar. A design for a list merging network. IEEE Transactions on Computers, C-28(6):406–413, June 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675380.

Hoogendoorn:1977:GMM

[Hoo77] C. H. Hoogendoorn. A general model for memory interference in multiprocessors. IEEE Transactions on Computers, C-26(10):998–1005, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674734.

Hopcroft:1970:RNS

[Hop70] J. E. Hopcroft. R70-2 nested stack automata. IEEE Transactions on Computers, C-19(5):463, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671546.

Hopkins:1971:FTI

[Hop71] A. L. Hopkins, Jr. A faulttolerant information processing concept for space vehicles. IEEE Transactions on Computers, C-20 (11):1394–1403, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671738.

Horelick:1971:CSB

[Hor71a] D. Horelick. Comment on “Serial Binary-to-Decimal and Decimal-to-Binary Conversion”. IEEE Transactions on Computers, C-20(4):476, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671869. See [Rhy70].

Horna:1971:PRT

[Hor71b] O. A. Horna. Pulse reflection in transmission lines. IEEE Transactions on Computers, C-20 (12):1558–1563, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671762.

Horna:1972:NTT

[Hor72] Otakar A. Horna. Nonlinear termination of transmission lines. IEEE Transactions on Computers, C-21(9):1011–1015, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009080.

Hou:1972:AUL

[Hou72] H. S. Hou. Application of uniform loading theory to circuit packaging and memory arrays in high-speed computers. IEEE Transactions on Computers, C-21(5):454–463, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672134.

Howden:1975:MGP

[How75] W. E. Howden. Methodology for the generation of program test data. IEEE Transactions on Computers, C-24(5):554–560, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672852.

Hong:1972:GCM

[HP72] Se June Hong and A. M. Patel. A general class of maximal codes for computer applications. IEEE Transactions on Computers, C-21(12):1322–1331, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672096.

Hight:1973:DMV

[HP73] S. L. Hight and D. P. Petersen.

Dissent in a majority voting system. IEEE Transactions on Computers, C-22(2):168–171, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672272.

Hoehne:1975:DVR

[HP75] H. Hoehne and R. Piloty. Design verification at the register transfer language level. IEEE Transactions on Computers, C-24(9):861–867, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672924.

Hsiao:1977:SAG

[HPP77] M. Y. Hsiao, A. M. Patel, and D. K. Pradhan. Store address generator with on-line fault-detection capability. IEEE Transactions on Computers, C-26(11):1144–1147, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674762.

Hammer:1979:ADR

[HPP79] P. L. Hammer, U. N. Peled, and M. A. Pollatschek. An algorithm to dualize a regular switching function. IEEE Transactions on Computers, C-28(3):238–243, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675324.

House:1970:NRR

[HS70] R. W. House and D. W. Stevens.

A new rule for reducing CC tables. IEEE Transactions on Computers, C-19(11):1108–1111, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671435.

Hopkins:1975:AES

[HS75] A. L. Hopkins, Jr. and T. B. Smith III. The architectural elements of a symmetric fault-tolerant multiprocessor. IEEE Transactions on Computers, C-24(5):498–505, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672845.

Hsiao:1971:GRO

[Hsi71a] D. K. Hsiao. A generalized record organization. IEEE Transactions on Computers, C-20(12): 1490–1495, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671754.

Hsieh:1971:CES

[Hsi71b] E. P. Hsieh. Checking experiments for sequential machines. IEEE Transactions on Computers, C-20(10):1152–1166, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&

                                        arnumber=1671693. See comments

[WY75].

Halatsis:1978:PDS

[HSP78] C. Halatsis, M. Sigala, and G. Philokyprou. Polylinear decomposition of synchronous sequential machines. IEEE Transactions on Computers, C-27(12): 1144–1152, December 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675017.

Hartke:1978:DRD

[HSS78] D. H. Hartke, W. M. Sterling, and J. E. Shemer. Design of a raster display processor for office applications. IEEE Transactions on Computers, C-27(4):337–348, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675106.

Hussain:1972:ROU

[HTD72] A. B. Shahidul Hussain, Godfried T. Toussaint, and Robert W. Donaldson. Results obtained using a simple character recognition procedure on Munson’s handprinted data. IEEE Transactions on Computers, C-21(2):201–205, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008927.

Hu:1972:CSS

[Hu72] S. C. Hu. Cellular synthesis of synchronous sequential machines. IEEE Transactions on Computers, C-21(12):1399–1405, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672106.

Hu:1973:BCI

[Hu73] Sung C. Hu. Boundary conditions for iteratively realized sequential circuits. IEEE Transactions on Computers, C-22(9): 867–868, September 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009184.

Hu:1975:PAD

[Hu75] S. C. Hu. A probabilistic approach of designing more reliable logic gates with asymmetric input faults. IEEE Transactions on Computers, C-24(10):1012–1014, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672706.

Huang:1971:UCA

[Hua71] Jung-Chang Huang. A universal cellular array. IEEE Transactions on Computers, C-20(3): 317–320, March 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671830.

Huber:1976:GES

[Hub76] Franz Huber. On the generation of English sentences. IEEE Transactions on Computers, C-25 (1):90–91, January 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009209.

Huffman:1976:CCP

[Huf76] D. A. Huffman. Curvature and creases: a primer on paper. IEEE Transactions on Computers, C-25(10):1010–1019, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674542.

Hunt:1971:SEU

[Hun71] B. R. Hunt. Spectral effects in the use of Newton–Cotes approximations for computing Discrete Fourier Transforms. IEEE Transactions on Computers, C-20(8):942–943, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671971.

Hunt:1972:MCT

[Hun72] B. R. Hunt. Minimizing the computation time for using the technique of sectioning for digital filtering of pictures. IEEE Transactions on Computers, C-21 (11):1219–1222, November 1972.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672073.

Hunt:1973:ACL

[Hun73] B. R. Hunt. The application of constrained least squares estimation to image restoration by digital computer. IEEE Transactions on Computers, C-22(9):805–812, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009169.

Hunt:1977:BMN

[Hun77] B. R. Hunt. Bayesian methods in nonlinear digital image restoration. IEEE Transactions on Computers, C-26(3):219–229, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674810.

Hurd:1974:EGS

[Hur74] W. J. Hurd. Efficient generation of statistically good pseudonoise by linearly interconnected shift registers. IEEE Transactions on Computers, C-23(2):146–152, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672470.

Hussain:1972:CSS

[Hus72] A. B. Shahidul Hussain. On the correctness of some sequential classification schemes in pattern recognition. IEEE Transactions on Computers, C-21(3):318–320, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008959.

Huskey:1973:PA

[Hus73] H. D. Huskey. The Polish assembler. IEEE Transactions on Computers, C-22(11):1035–1040, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672234.

Hussain:1974:CSP

[Hus74] A. B. S. Hussain. Compound sequential probability ratio test for the classification of statistically dependent patterns. IEEE Transactions on Computers, C-23(4):398–410, April 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672548.

Habibi:1970:FM

[HW70] A. Habibi and P. A. Wintz. Fast multipliers. IEEE Transactions on Computers, C-19(2): 153–157, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671474.

Hulme:1975:PIA

[HW75] B. L. Hulme and R. B. Worrell. A prime implicant algorithm with factoring. IEEE Transactions on Computers, C-24 (11):1129–1131, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672739.

Hannington:1976:FPM

[HW76] G. Hannington and D. G. Whitehead. A floating-point multiplexed DDA system. IEEE Transactions on Computers, C-25(11): 1074–1077, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674557. See comments

[MMB78].

Hannington:1978:AR

[HW78] G. Hannington and D. G. Whitehead. Authors reply. IEEE Transactions on Computers, C-27 (12):1226–1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675034.

Hopper:1979:BRN

[HW79] A. Hopper and J. Wheeler. Binary routing networks. IEEE Transactions on Computers, C-28 (10):699–703, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675237.

Hwa:1972:CEM

[Hwa72] H. R. Hwa. Contradiction equations in a B matrix of vertex weight method and their correspondence with the k-summability property of vertices. IEEE Transactions on Computers, C-21(6):606–610, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009019.

Hwang:1973:PRS

[Hwa73] Kai Hwang. Periodic realization of synchronous sequential machines. IEEE Transactions on Computers, C-22(10):923–927, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672212.

Hwa:1974:MGP

[Hwa74] H. R. Hwa. A method for generating prime implicants of a Boolean expression. IEEE Transactions on Computers, C-23(6):637–641, June 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672596.

Hwang:1979:GMT

[Hwa79] Kai Hwang. Global and modular two’s complement cellular array multipliers. IEEE Transactions on Computers, C-28(4):300–306, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675350.

Irao:1976:RIT

[IA76] V. V. Bapeswara Irao and V. K.

Aatre. Reliability index of teleprocessing tree networks. IEEE Transactions on Computers, C-25 (1):97–98, January 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009213.

Ibaraki:1971:GIM

[Iba71] T. Ibaraki. Gate-interconnection minimization of switching networks using negative gates. IEEE Transactions on Computers, C-20(6):698–706, June 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671924.

Ibaraki:1974:CMF

[Iba74] T. Ibaraki. Comments on “Monotone Functions in Sequential Circuits”. IEEE Transactions on Computers, C-23(9):989–990, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672665. See [Mag73].

Ito:1978:CBC

[IC78] M. R. Ito and R. D. Cameron.

Combined binary code translation and parallel-to-serial conversion using stored logic arrays. IEEE Transactions on Computers, C-27(9):833–841, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675203.

Iosupovicz:1972:CAO

[IH72] A. Iosupovicz and Ming-Kuei Hu. A class of autonomous onedimensional iterative arrays of linear machines. IEEE Transactions on Computers, C-21(10): 1073–1086, October 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672047.

Ibaraki:1971:SNM

[IM71] T. Ibaraki and S. Muroga. Synthesis of networks with a minimum number of negative gates. IEEE Transactions on Computers, C-20(1):49–58, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671674.

Iosupovicz:1975:AUF

[Ios75] A. Iosupovicz. Adaptive universal fault-tolerant logic networks. IEEE Transactions on Computers, C-24(10):1043–1048, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672720.

Iosupovicz:1978:ODB

[Ios78] A. Iosupovicz. Optimal detection of bridge faults and stuck-at faults in two-level logic. IEEE Transactions on Computers, C-27(5):452–455, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675125.

Ikram:1976:STI

[IR76] M. Ikram and D. A. Roy. A simple technique to improve the pi-algorithm for prime implicant determination. IEEE Transactions on Computers, C-25(11): 1184–1187, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674579.

Ibarra:1975:PCF

[IS75] O. H. Ibarra and S. K. Sahni.

Polynomially complete fault detection problems. IEEE Transactions on Computers, C-24(3): 242–249, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672798.

Ingle:1976:RMV

[IS76] Ashok D. Ingle and Daniel P.

Siewiorek. A reliability model for various switch designs in hybrid redundancy. IEEE Transactions on Computers, C-25(2): 115–133, February 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009226.

Ishizuka:1977:MMN

[Ish77a] O. Ishizuka. On multivalued multithreshold networks composed of conventional threshold elements. IEEE Transactions on Computers, C-26(12):1251–1257, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674786.

Ishizuka:1977:SMT

[Ish77b] O. Ishizuka. Synthesis of multithreshold tree networks. IEEE Transactions on Computers, C-26 (12):1294–1297, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674796.

Irving:1976:FFM

[ISN76] Thurman A. Irving, Sajjan G. Shiva, and H. Troy Nagle. Flipflops for multiple-valued logic. IEEE Transactions on Computers, C-25(3):237–246, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009250.

Jain:1974:IRM

[JA74] A. K. Jain and E. Angel. Image restoration, modelling, and reduction of dimensionality. IEEE Transactions on Computers, C-23(5):470–476, May 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672562.

Jump:1978:EPD

[JA78] J. R. Jump and S. R. Ahuja. Effective pipelining of digital systems. IEEE Transactions on Computers, C-27(9):855–865, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675205.

Jacobsohn:1973:CDA

[Jac73] David H. Jacobsohn. A combinatoric division algorithm for fixedinteger divisors. IEEE Transactions on Computers, C-22(6):608–610, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009114;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Jacobsohn.
pdf.

Jain:1977:FKL

[Jai77a] A. K. Jain. A fast Karhunen–Loeve transform for digital restoration of images degraded by white and colored noise. IEEE Transactions on Computers, C-26(6):560–571, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674881.

Jain:1977:OFM

[Jai77b] A. K. Jain. An operator factorization method for restoration of blurred images. IEEE Transactions on Computers, C-26 (11):1061–1071, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674752.

Jain:1977:SMR

[Jai77c] A. K. Jain. A semicausal model for recursive filtering of two-dimensional images. IEEE Transactions on Computers, C-26(4):343–350, April 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674844.

Jayashree:1976:BMU

[JB76a] T. Jayashree and D. Basu. On binary multiplication using the quarter square algorithm. IEEE Transactions on Computers, C-25(9):957–960, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674723.

Jayasri:1976:AOM

[JB76b] T. Jayasri and D. Basu. An approach to organizing microinstructions which minimizes the width of control store words. IEEE Transactions on Computers, C-25(5):514–521, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674639.

Jump:1972:MA

[JF72] J. Robert Jump and Dennis R.

Fritsche. Microprogrammed arrays. IEEE Transactions on Computers, C-21(9):974–984, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009074.

Jakubowski:1977:SDR

[JK77] R. Jakubowski and A. Kasprzak.

A syntactic description and recognition of rotary machine elements. IEEE Transactions on Computers, C-26(10):1039–1043, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674743.

Jordan:1973:IAG

[JLH73] B. W. Jordan, Jr., W. J. Lennon, and B. D. Holm. An improved algorithm for the generation of nonparametric curves. IEEE Transactions on Computers, C-22(12):1052–1060, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672243. See comment

[Bel76].

Janisz:1970:IHS

[JM70] T. Janisz and R. C. Martin. Interconnection of high-speed logic circuits. IEEE Transactions on Computers, C-19(9):831–837, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671642.

Jones:1971:OST

[JM71] S. K. Jones and J. L. Melsa. Optimization and simulation of two classes of nonresetting data reconstructors. IEEE Transactions on Computers, C-20(1):12–18, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671668.

Juliussen:1973:MMC

[JM73] J. E. Juliussen and F. J. Mowle. Multiple microprocessors with common main and control memories. IEEE Transactions on Computers, C-22(11):999–1007, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672229.

Jones:1974:TMS

[JM74] L. H. Jones and R. E. Merwin. Trends in microprogramming: a second reading. IEEE Transactions on Computers, C-23 (8):754–759, August 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672624.

Johnson:1971:MIS

[Joh71] A. M. Johnson. The microdiagnostics for the IBM System/360 Model 30. IEEE Transactions on Computers, C-20(7):798–803, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671942.

Joliat:1976:STP

[Jol76] M. L. Joliat. A simple technique for partial elimination of unit productions from LR(k) parsers. IEEE Transactions on Computers, C-25(7):763–764, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674686.

Jorgensen:1974:BBP

[Jor74a] P. Jorgensen. B74-7 basic programming. IEEE Transactions on Computers, C-23(2):212, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672489.

Jorgensen:1974:BPP

[Jor74b] P. C. Jorgensen. B74-6 PL/I programming for engineering and science. IEEE Transactions on Computers, C-23(2):211–212, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672488.

Joseph:1972:BCT

[Jos72a] E. C. Joseph. B72-18 computers and their uses, 2nd ed. IEEE Transactions on Computers, C-21(10):1135, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672058.

Joseph:1972:BEU

[Jos72b] E. C. Joseph. B72-4 expanding use of computers in the 70’s technology — markets, needs, technology. IEEE Transactions on Computers, C-21(1):110, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672038.

Joseph:1974:BDB

[Jos74] E. C. Joseph. B74-14 data base systems. IEEE Transactions on Computers, C-23(2):217–218, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672496.

Joseph:1978:EDS

[Jos78] J. Joseph. On easily diagnosable sequential machines. IEEE Transactions on Computers, C-27 (2):159–162, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675049.

Jarvis:1973:CUS

[JP73] R. A. Jarvis and E. A. Patrick.

Clustering using a similarity measure based on shared near neighbors. IEEE Transactions on Computers, C-22(11):1025– 1034, November 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672233.

Jagoe:1976:MPC

[JP76] J. Roger Jagoe and Keith A. Paton. Measurement of pneumoconiosis by computer. IEEE Transactions on Computers, C-25(1):95–97, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009212.

Jarvis:1976:NTD

[JR76] J. F. Jarvis and C. S. Roberts. A new technique for displaying continuous tone images on a bilevel display. IEEE Transactions on Communications, 24(8):891–898, August 1976. CODEN IECMBT. ISSN 0090-6778.

Jullien:1978:RNS

[Jul78] G. A. Jullien. Residue number scaling and other operations using ROM arrays. IEEE Transactions on Computers, C-27(4):325–336, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=1675105;   http://
                                        ieeexplore.ieee.org/xpl/RecentIssue.
                                        jsp?punumber=12;   http://ieeexplore.
                                        ieee.org/xpl/tocresult.jsp?isnumber=
                                        35168.

Jump:1970:RUS

[Jum70] J. R. Jump. R70-29 uniform synthesis of sequential circuits. IEEE Transactions on Computers, C-19(8):765, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671627.

Jump:1974:ACA

[Jum74] J. R. Jump. Asynchronous control arrays. IEEE Transactions on Computers, C-23(10):1020–1029, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672396.

Jordan:1973:DSS

[JW73] H. F. Jordan and D. C. M. Wood. On the distribution of sums of successive bits of shift-register sequences. IEEE Transactions on Computers, C-22(4):400–408, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672321.

Joshi:1976:SFM

[JW76] A. K. Joshi and R. M. Weischedel.

Some frills for modal tic-tac-toe: Semantics of predicate complement constructions. IEEE Transactions on Computers, C-25(4):374–389, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674620.

Karlgaard:1977:AR

[KAA77] D. C. Karlgaard and A. M. Abd-Alla. Author’s reply. IEEE Transactions on Computers, C-26 (9):932, September 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674942.

Kain:1970:NSC

[Kai70] R. Y. Kain. Nonlinear sequential circuits. IEEE Transactions on Computers, C-19(3): 249–254, March 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671497. See comments

[Du71].

Kameda:1970:RMT

[Kam70a] T. Kameda. R70-23 multi-tape and multi-head pushdown automata. IEEE Transactions on Computers, C-19(7):659–660, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671603.

Kameda:1970:RRH

[Kam70b] T. Kameda. R70-39 on the relational homomorphisms of automata. IEEE Transactions on Computers, C-19(10):987, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671405.

Kamal:1975:ADI

[Kam75] S. Kamal. An approach to the diagnosis of intermittent faults. IEEE Transactions on Computers, C-24(5):461–467, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672840.

Kambayashi:1979:LDP

[Kam79] Y. Kambayashi. Logic design of programmable logic arrays. IEEE Transactions on Computers, C-28(9):609–617, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675428.

Kandel:1973:CMF

[Kan73a] A. Kandel. Comments on “Minimization of Fuzzy Functions”. IEEE Transactions on Computers, C-22(2):217, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672281. See [SC72a].

Kandel:1973:NHE

[Kan73b] A. Kandel. Note on hazard elimination. IEEE Transactions on Computers, C-22(10):955–956, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672216.

Kandel:1973:MFF

[Kan73c] Abraham Kandel. On minimization of fuzzy functions. IEEE Transactions on Computers, C-22(9):826–832, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009172.

Kang:1975:EUM

[Kan75] A. N. C. Kang. On the efficiency of universal machines. IEEE Transactions on Computers, C-24(10):1010–1012, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672705.

Kandel:1976:DFF

[Kan76] A. Kandel. On the decomposition of fuzzy functions. IEEE Transactions on Computers, C-25 (11):1124–1130, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674563.

Kobylarz:1979:ECF

[KAN79] T. Kobylarz and A. Al-Najjar. An examination of the cost function for programmable logic arrays. IEEE Transactions on Computers, C-28(8):586–590, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675418.

Karplus:1970:RUF

[Kar70] W. J. Karplus. R70-6 use of functional approximation methods in the computer solution of initial value partial differential equation problems. IEEE Transactions on Computers, C-19(5):465, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671550.

Kartashev:1974:TIM

[Kar74] S. P. Kartashev. Theory and implementation of p-multiple sequential machines. IEEE Transactions on Computers, C-23(5):500–523, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672566.

Kartashev:1976:MSR

[Kar76] S. I. Kartashev. A microcomputer with a shift-register memory. IEEE Transactions on Computers, C-25(5):470–484, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674635.

Karpovsky:1977:EDD

[Kar77a] M. G. Karpovsky. Error detection in digital devices and computer programs with the aid of linear recurrent equations over finite commutative groups. IEEE Transactions on Computers, C-26 (3):208–218, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674809.

Karpovsky:1977:FFT

[Kar77b] M. G. Karpovsky. Fast Fourier Transforms on finite non-Abelian groups. IEEE Transactions on Computers, C-26(10):1028– 1030, October 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674739.

Kautz:1970:BSC

[Kau70a] W. H. Kautz. Bypass switching for cellular cascades. IEEE Transactions on Computers, C-19(9):837–839, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671643.

Kautz:1970:NCC

[Kau70b] W. H. Kautz. The necessity of closed circuit loops in minimal combinational circuits. IEEE Transactions on Computers, C-19(2): 162–164, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671477.

Kautz:1970:RIS

[Kau70c] W. H. Kautz. A readily implemented single-error-correcting unitdistance counting code. IEEE Transactions on Computers, C-19 (10):972–975, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671397. See comments

[Dev79].

Kaufman:1973:EFR

[Kau73] A. B. Kaufman. An expandable ferroelectric random access memory. IEEE Transactions on Computers, C-22(2):154–158, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672270.

Kaufman:1974:AOA

[Kau74a] M. T. Kaufman. An almostoptimal algorithm for the assembly line scheduling problem. IEEE Transactions on Computers, C-23(11):1169–1174, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672418.

Kautz:1974:TFW

[Kau74b] W. H. Kautz. Testing for faults in wiring networks. IEEE Transactions on Computers, C-23(4):358–363, April 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672543.

Kohavi:1975:FDC

[KB75] Z. Kohavi and I. Berger. Fault diagnosis in combinational tree networks. IEEE Transactions on Computers, C-24(12):1161– 1167, December 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672752.

Kennedy:1971:HIU

[KC71] J. D. Kennedy and S. J. Clark.

Halftone images using computer graphics. IEEE Transactions on Computers, C-20(5):584–585, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671890.

Kukreja:1973:CSC

[KC73] S. N. Kukreja and I-Ngo Chen. Combinational and sequential cellular structures. IEEE Transactions on Computers, C-22(9): 813–823, September 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009170.

Kelly:1971:FEV

[KCLB71] K. K. Kelly, T. W. Calvert, R. L. Longini, and J. P. Brown. Feature enhancement of vectorcardiograms by linear normalization. IEEE Transactions on Computers, C-20(9):1109–1111, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672006.

Khabra:1973:MPS

[KD73] N. S. Khabra and S. R. Das. Multiform partial symmetry and linearity. IEEE Transactions on Computers, C-22(8):804, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009162.

Kam:1979:SDS

[KD79] John B. Kam and George I. Davida. Structured design of substitutionpermutation encryption networks.

IEEE Transactions on Computers, C-28(10):747–753, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675242.

Kella:1970:SMI

[Kel70] J. Kella. State minimization of incompletely specified sequential machines. IEEE Transactions on Computers, C-19(4):342–348, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671513.

Kella:1971:SMI

[Kel71] J. Kella. Sequential machine identification. IEEE Transactions on Computers, C-20(3): 332–338, March 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671834. See comments

[WZ72].

Keller:1974:TTU

[Kel74] R. M. Keller. Towards a theory of universal speed-independent modules. IEEE Transactions on Computers, C-23(1):21–33, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672366.

Kendall:1974:NAC

[Ken74] W. B. Kendall. A new algorithm for computing correlations. IEEE Transactions on Computers, C-23(1):88–90, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672376.

Kent:1977:HSF

[Ken77a] J. G. Kent. Highlights of a study of floating-point instructions. IEEE Transactions on Computers, C-26(7):660–666, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674899.

Kent:1977:HST

[Ken77b] S. A. Kent. A high-speed threshold gate multiplier. IEEE Transactions on Computers, C-26(12): 1279–1283, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674790.

Kernighan:1973:BSP

[Ker73] B. W. Kernighan. B73-8 systems programming. IEEE Transactions on Computers, C-22(4):430–431, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672333.

Kamal:1972:ACH

[KES72] A. A. Kamal and A. K. El-Sherif.

Analysis and compensation of highspeed electronic analog-computer errors. IEEE Transactions on Computers, C-21(8):886–891, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009045.

Keyes:1977:PUI

[Key77] R. W. Keyes. Physical uncertainty and information. IEEE Transactions on Computers, C-26(10):1017–1025, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674737.

Koontz:1972:NFE

[KF72a] W. L. G. Koontz and K. Fukunaga. A nonlinear feature extraction algorithm using distance transformation. IEEE Transactions on Computers, C-21(1):56–63, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672024.

Koontz:1972:AAN

[KF72b] Warren L. G. Koontz and Keinosuke Fukunaga. Asymptotic analysis of a nonparametric clustering technique. IEEE Transactions on Computers, C-21(9):967–974, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009073.

Koontz:1972:NVS

[KF72c] Warren L. G. Koontz and Keinosuke Fukunaga. A nonparametric valleyseeking technique for cluster analysis. IEEE Transactions on Computers, C-21(2):171–178, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008922.

Kreitzer:1973:VDS

[KF73] N. H. Kreitzer and W. J. Fitzgerald. A video display system for image processing by computer. IEEE Transactions on Computers, C-22(2):128–134, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672266.

Karunanithi:1978:SNT

[KF78] S. Karunanithi and A. D. Friedman. Some new types of logical completeness. IEEE Transactions on Computers, C-27(11):998–1005, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674989.

Karunanithi:1979:ADS

[KF79] S. Karunanithi and A. D. Friedman. Analysis of digital systems using a new measure of system diagnosis. IEEE Transactions on Computers, C-28(2):121–133, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675301.

Kamal:1972:HSM

[KG72] A. A. Kamal and M. A. N. Ghannam. High-speed multiplication systems. IEEE Transactions on Computers, C-21(9):1017–1021, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009082.

Kahveci:1974:SDD

[KH74] A. E. Kahveci and E. L. Hall. Sequency domain design of frequency filters. IEEE Transactions on Computers, C-23(9):976–981, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672658.

Kameyama:1977:SMV

[KH77a] M. Kameyama and T. Higuchi. Synthesis of multiple-valued logic networks based on tree-type universal logic module. IEEE Transactions on Computers, C-26(12): 1297–1302, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674797.

Knaizuk:1977:ATR

[KH77b] J. Knaizuk, Jr. and C. R. P. Hartmann. An algorithm for testing random access memories. IEEE Transactions on Computers, C-26(4):414–416, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674851.

Knaizuk:1977:OAT

[KH77c] J. Knaizuk, Jr. and C. R. P. Hartmann. An optimal algorithm for testing stuck-at faults in random access memories. IEEE Transactions on Computers, C-26 (11):1141–1144, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674761. See comments

[Nai79].

Kime:1970:AMD

[Kim70] C. R. Kime. An analysis model for digital system diagnosis. IEEE Transactions on Computers, C-19 (11):1063–1073, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671426.

Kimbleton:1974:BEM

[Kim74] S. R. Kimbleton. B74-43 evaluation and measurement techniques for digital computer systems. IEEE Transactions on Computers, C-23(12):1328–1329, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672457.

Kime:1975:FTC

[Kim75] C. R. Kime. Fault-tolerant computing: An introduction and a perspective. IEEE Transactions on Computers, C-24(5):457–460, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672839.

Kime:1979:AMD

[Kim79] C. R. Kime. An abstract model for digital system fault diagnosis. IEEE Transactions on Computers, C-28(10):754–767, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675243.

Kinney:1970:DAS

[Kin70] L. L. Kinney. Decomposition of asynchronous sequential switching circuits. IEEE Transactions on Computers, C-19(6):515–529, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671562.

King:1971:PPC

[Kin71a] J. C. King. Proving programs to be correct. IEEE Transactions on Computers, C-20(11):

1331–1336, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671728.

King:1971:DAM

[Kin71b] W. K. King. Design of an associative memory. IEEE Transactions on Computers, C-20(6):671–674, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671915.

Kinney:1971:CSA

[Kin71c] L. L. Kinney. A characterization of some asynchronous sequential networks and state assignments. IEEE Transactions on Computers, C-20(4):426–436, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671852.

Kinney:1972:BIS

[Kin72] L. L. Kinney. B72-11 an introduction to switching system design. IEEE Transactions on Computers, C-21(5):518, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672154.

Kittler:1977:DVM

[Kit77] J. Kittler. On the discriminant vector method of feature selection. IEEE Transactions on Computers, C-26(6):604–606, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674885.

Kittler:1978:CIP

[Kit78] J. Kittler. Classification of incomplete pattern vectors using modified discriminant functions. IEEE Transactions on Computers, C-27(4):367–375, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675109.

Kulkarni:1979:OCA

[KJ79] K. G. Kulkarni and V. Jayakumar. Ordering of connections for automated routing. IEEE Transactions on Computers, C-28(10): 791–794, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675248.

Kjelkerud:1972:CPS

[Kje72] Eskil Kjelkerud. A computer program for the synthesis of switching circuits by decomposition. IEEE Transactions on Computers, C-21(6):568–573, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009009.

Korn:1970:PHC

[KK70] G. A. Korn and H. Kosako. A proposed hybrid-computer method for functional optimization. IEEE Transactions on Computers, C-19 (2):149–153, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671473.

Kohavi:1972:DMF

[KK72] Igal Kohavi and Zvi Kohavi. Detection of multiple faults in combinational logic networks. IEEE Transactions on Computers, C-21(6):556–568, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009008.

Koren:1977:DIF

[KK77a] I. Koren and Z. Kohavi. Diagnosis of intermittent faults in combinational networks. IEEE Transactions on Computers, C-26 (11):1154–1158, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674766.

Koren:1977:SFD

[KK77b] I. Koren and Z. Kohavi. Sequential fault diagnosis in combinational networks. IEEE Transactions on Computers, C-26(4):334–342, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674843.

Kartashev:1978:MNS

[KK78] S. P. Kartashev and S. I. Kartashev. On modular networks satisfying the shift-register rule. IEEE Transactions on Computers, C-27(12):1153–1176, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675018.

Kartashev:1979:MSD

[KK79a] S. I. Kartashev and S. P. Kartashev. A multicomputer system with dynamic architecture. IEEE Transactions on Computers, C-28(10):704–721, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675238.

Koren:1979:PSP

[KK79b] I. Koren and Z. Kohavi. On the properties of sensitized paths. IEEE Transactions on Computers, C-28(3):268–269, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675335.

Klinger:1971:CAC

[KKA71] A. Klinger, A. Kochman, and N. Alexandridis. Computer analysis of chromosome patterns: Feature encoding for flexible decision making. IEEE Transactions on Computers, C-20(9):1014–1022, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671990.

Kinoshita:1973:GDS

[KKK73] E. Kinoshita, H. Kosako, and Y. Kojima. General division in the symmetric residue number system. IEEE Transactions on Computers, C-22(2):134–142, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1672267;   http://
ieeexplore.ieee.org/xpl/tocresult.
jsp?isnumber=35073.

Kinoshita:1974:FPA

[KKK74] Eisuke Kinoshita, Hideo Kosako, and Yoshiaki Kojima. Floatingpoint arithmetic algorithms in the symmetric residue number system. IEEE Transactions on Computers, C-23(1):9–20, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1672365;   http://
ieeexplore.ieee.org/xpl/tocresult.
jsp?isnumber=35077.

Klaschka:1971:MRS

[Kla71] T. F. Klaschka. A method for redundancy scheme performance assessment. IEEE Transactions on Computers, C-20(11):

1371–1376, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671734.

Kang:1977:SRT

[KLCC77] A. N. C. Kang, R. C. T. Lee, Chin-Liang Chang, and Shi-Kuo Chang. Storage reduction through minimal spanning trees and spanning forests. IEEE Transactions on Computers, C-26(5):425–434, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674859.

Kleinrock:1970:STC

[Kle70] L. Kleinrock. Swap-time considerations in time-shared systems. IEEE Transactions on Computers, C-19(6):534–540, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671564.

Kleinrock:1976:CN

[Kle76] L. Kleinrock. On communications and networks. IEEE Transactions on Computers, C-25(12): 1326–1335, December 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674601.

Klir:1971:ULP

[Kli71] G. J. Klir. On universal logic primitives. IEEE Transactions on Computers, C-20(4):467–469, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671862.

Klir:1974:SBP

[Kli74] G. J. Klir. On the solution of Boolean and pseudo-Boolean relations. IEEE Transactions on Computers, C-23(10):1098–1100, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672403.

Kluge:1977:TBT

[Klu77] W. E. Kluge. Traversing binary tree structures with shiftregister memories. IEEE Transactions on Computers, C-26(11): 1112–1122, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674757.

Kashef:1973:APC

[KM73] R. S. Kashef and R. B. McGhee. Augmented parity check codes for encoding of asynchronous sequential machines. IEEE Transactions on Computers, C-22(10): 891–896, October 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672207.

Kashyap:1975:PRP

[KM75a] R. L. Kashyap and M. C. Mittal.

Picture reconstruction from projections. IEEE Transactions on Computers, C-24(9):915–923, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672930.

Ku:1975:BDM

[KM75b] Chia-Tai Ku and G. M. Masson.

The Boolean difference and multiple fault analysis. IEEE Transactions on Computers, C-24(1):62–71, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672676.

Kashyap:1978:RSW

[KM78] R. L. Kashyap and M. C. Mittal. Recognition of spoken words and phrases in multitalker environment using syntactic methods. IEEE Transactions on Computers, C-27(5):442–452, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675124.

Kuck:1972:NOS

[KMC72] D. J. Kuck, Y. Muraoka, and Shyh-Ching Chen. On the number of operations simultaneously executable in FORTRAN-like programs and their resulting speedup.

IEEE Transactions on Computers, C-21(12):1293–1310, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672094.

Korn:1970:RAF

[KN70] I. Korn and A. Nathan. On the representation of arbitrary functions and their generation. IEEE Transactions on Computers, C-19(6):483–486, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671557.

Kim:1972:SSM

[KN72] Joonki Kim and M. M. Newborn. The simplification of sequential machines with input restrictions. IEEE Transactions on Computers, C-21(12):1440–1443, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672114.

Koontz:1975:BBC

[KNF75] W. L. G. Koontz, P. M. Narendra, and K. Fukunaga. A branch and bound clustering algorithm. IEEE Transactions on Computers, C-24(9):908–915, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672929. See comments

[Wet78].

Koontz:1976:GTA

[KNF76] W. L. G. Koontz, P. M. Narendra, and K. Fukunaga. A graphtheoretic approach to nonparametric cluster analysis. IEEE Transactions on Computers, C-25(9):936–944, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674719.

Kanani:1973:NCS

[KO73] Dhirubhai V. Kanani and Kenneth H. O’Keefe. A note on conditional-sum addition for base −2 systems. IEEE Transactions on Computers, C-22(6):626, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009117;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Kanani.
pdf.

Kobayashi:1978:SML

[KO78] H. Kobayashi and H. Ohara. A synthesizing method for large parallel counters with a network of smaller ones. IEEE Transactions on Computers, C-27(8):753–757, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675184.

Koczela:1971:TFT

[Koc71] L. J. Koczela. A three-failure-tolerant computer system. IEEE Transactions on Computers, C-20 (11):1389–1393, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671737.

Kock:1975:RTP

[Koc75] W. E. Kock. A real-time parallel optical processing technique. IEEE Transactions on Computers, C-24(4):407–411, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672828.

Kodandapani:1974:NET

[Kod74] K. L. Kodandapani. A note on easily testable realizations for logic functions. IEEE Transactions on Computers, C-23(3): 332–333, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672528.

Koh:1971:MTT

[Koh71] Kyung Shik Koh. A minimization technique for TANT networks. IEEE Transactions on Computers, C-20(1):105–107, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671683. See comments

[LD72a].

Koh:1972:AR

[Koh72a] Kyung Shik Koh. Author’s reply. IEEE Transactions on Computers, C-21(4):408, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008988.

Kohonen:1972:CMM

[Koh72b] Teuvo Kohonen. Correlation matrix memories. IEEE Transactions on Computers, C-21(4):353–359, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008975.

Kohonen:1974:AAM

[Koh74] T. Kohonen. An adaptive associative memory principle. IEEE Transactions on Computers, C-23(4):444–445, April 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672553.

Kohler:1975:PEC

[Koh75] W. H. Kohler. A preliminary evaluation of the critical path method for scheduling tasks on multiprocessor systems. IEEE Transactions on Computers, C-24 (12):1235–1238, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672764.

Kolp:1972:SMC

[Kol72] O. Kolp. The synthesis of multivalued cellular cascades and the decomposability of group functions. IEEE Transactions on Computers, C-21(5):489–492, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672139.

Kontos:1971:IAC

[Kon71] J. Kontos. An iterative analog computer using pulse-ratio modulation. IEEE Transactions on Computers, C-20(10):1188–1192, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671696.

Koontz:1978:AR

[Koo78] W. L. G. Koontz. Author’s reply. IEEE Transactions on Computers, C-27(2):189, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675060.

Koren:1976:ICN

[Kor76] Yoram Koren. Interpolator for a computer numerical control system. IEEE Transactions on Computers, C-25(1):32–37, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009202.

Koren:1979:ASR

[Kor79] I. Koren. Analysis of the signal reliability measure and an evaluation procedure. IEEE Transactions on Computers, C-28(3): 244–249, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675326.

Kozlay:1971:FEO

[Koz71] D. Kozlay. Feature extraction in an optical character recognition machine. IEEE Transactions on Computers, C-20(9):1063–1067, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671997.

Kontos:1971:SMC

[KP71] J. Kontos and G. Papakonstantinou. A simulation method for computer control systems. IEEE Transactions on Computers, C-20 (1):98–100, January 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671680.

Kamal:1974:IFM

[KP74] S. Kamal and C. V. Page. Intermittent faults: a model and a detection procedure. IEEE Transactions on Computers, C-23(7):713–719, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672612.

Kioustelidis:1975:PLA

[KP75] J. B. Kioustelidis and J. K. Petrou. A piecewise linear approximation of log_{2} x with equal maximum errors in all intervals. IEEE Transactions on Computers, C-24(9):858–861, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672923.

Kuck:1977:ARM

[KPS77] David J. Kuck, Douglass S. Parker Jr., and Ahmed H. Sameh. Analysis of rounding methods in floating-point arithmetic. IEEE Transactions on Computers, C-26(7):643–650, July 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674897.

Karplus:1971:IDC

[KR71a] W. J. Karplus and R. A. Russell. Increasing digital computer efficiency with the aid of error-correcting analog subroutines. IEEE Transactions on Computers, C-20(8):831–837, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671950.

Kleir:1971:OSM

[KR71b] R. L. Kleir and C. V. Ramamoorthy. Optimization strategies for microprograms. IEEE Transactions on Computers, C-20(7):783–794, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671940.

Kohonen:1973:RAD

[KR73] Teuvo Kohonen and Matti Ruohonen. Representation of associated data by matrix operators. IEEE Transactions on Computers, C-22(7):701–702, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009138.

Knuth:1975:AIM

[KR75] D. E. Knuth and G. S. Rao. Activity in an interleaved memory. IEEE Transactions on Computers, C-24(9):943–944, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672937.

Kuhl:1978:MST

[KR78a] J. G. Kuhl and S. M. Reddy. A multicode single transition-time state assignment for asynchronous sequential machines. IEEE Transactions on Computers, C-27(10): 927–934, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674972.

Kuhl:1978:DTS

[KR78b] J. G. Kuhl and S. M. Reddy. On the detection of terminal stuckfaults. IEEE Transactions on Computers, C-27(5):467–469, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675129.

Kraft:1974:BMI

[Kra74] G. D. Kraft. B74-35 minicomputersinternational computer state of the art report. IEEE Transactions on Computers, C-23(6):647–648, June 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672600.

Krishnamurthy:1970:OIS

[Kri70a] E. V. Krishnamurthy. On optimal iterative schemes for high-speed division. IEEE Transactions on Computers, C-19(3):227–231, March 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671494.

Krishnamurthy:1970:RTT

[Kri70b] E. V. Krishnamurthy. On rangetransformation techniques for division. IEEE Transactions on Computers, C-19(2):157–160, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671475.

Krishnamurthy:1971:CTW

[Kri71a] E. V. Krishnamurthy. Complementary two-way algorithms for negative radix conversions. IEEE Transactions on Computers, C-20(5):543–550, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671881.

Krishnamurthy:1971:EIR

[Kri71b] E. V. Krishnamurthy. Economical iterative and range-transformation schemes for division. IEEE Transactions on Computers, C-20(4):470–472, April 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671863.

Kriz:1973:SBO

[Kri73] Thomas A. Kriz. Some binary output sequence properties of deterministic autonomous finitestate machines with probabilistic initialization. IEEE Transactions on Computers, C-22(9): 868–869, September 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009185.

Krishnamurthy:1977:MPU

[Kri77] E. V. Krishnamurthy. Matrix processors using p-adic arithmetic for exact linear computations. IEEE Transactions on Computers, C-26(7):633–639, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674895.

Kroft:1974:CTC

[Kro74] D. Kroft. Comments on “A Two’s Complement Parallel Array Multiplication Algorithm”. IEEE Transactions on Computers, C-23 (12):1327–1328, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672456. See [BW73,

Bla74].

Kruskal:1971:CNM

[Kru71] J. B. Kruskal. Comments on “A Nonlinear Mapping for Data Structure Analysis”. IEEE Transactions on Computers, C-20(12): 1614, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671777. See [Whi72].

Kruse:1973:PPP

[Kru73] B. Kruse. A parallel picture processing machine. IEEE Transactions on Computers, C-22(12): 1075–1087, December 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672246.

Kohavi:1971:DSF

[KS71] Z. Kohavi and D. A. Spires. Designing sets of fault-detection tests for combinational logic circuits. IEEE Transactions on Computers, C-20(12):1463–1469, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671751.

Kitai:1972:CSD

[KS72] R. Kitai and K.-H. Siemens. Comments on “A Simplified Definition of Walsh Functions”. IEEE Transactions on Computers, C-21(5):512, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672148. See [Car71b,

LM71b].

Kogge:1973:PAE

[KS73] Peter M. Kogge and Harold S.

Stone. A parallel algorithm for the efficient solution of a general class of recurrence equations. IEEE Transactions on Computers, C-22(8):786–793, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009159.

Kodandapani:1975:RMC

[KS75] K. L. Kodandapani and R. V. Setlur. Reed–Muller canonical forms in multivalued logic. IEEE Transactions on Computers, C-24(6):628–636, June 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672868.

Keshavan:1977:SET

[KS77a] H. R. Keshavan and M. D. Srinath. Sequential estimation technique for enhancement of noisy images. IEEE Transactions on Computers, C-26(10):971–987, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674732.

Kodandapani:1977:NMR

[KS77b] K. L. Kodandapani and R. V. Setlur. A note on minimal Reed–Muller canonical forms of switching functions. IEEE Transactions on Computers, C-26(3): 310–313, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674830.

Kornerup:1977:UNR

[KS77c] P. Kornerup and B. D. Shriver. A unified numeric representation arithmetic unit and its language support. IEEE Transactions on Computers, C-26(7):651–659, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674898.

Kanefsky:1978:DTA

[KS78a] M. Kanefsky and M. G. Strintzis.

A decision theory approach to picture smoothing. IEEE Transactions on Computers, C-27(1):32–36, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674949.

Kodandapani:1978:CNR

[KS78b] K. L. Kodandapani and S. C. Seth.

On combinational networks with restricted fan-out. IEEE Transactions on Computers, C-27(4):309–318, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675103.

Kodandapani:1978:CAM

[KS78c] K. L. Kodandapani and R. V.

Setlur. A cellular array for multivalued logic functions. IEEE Transactions on Computers, C-27 (11):1055–1059, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674995.

Koren:1979:RAM

[KS79] I. Koren and S. Y. H. Su. Reliability analysis of N-modular redundancy systems with intermittent and permanent faults. IEEE Transactions on Computers, C-28(7):514–520, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675397.

Kamal:1974:GPA

[KSA74] A. K. Kamal, H. Singh, and D. P. Agrawal. A generalized pipeline array. IEEE Transactions on Computers, C-23(5):533–536, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672570.

Kinoshita:1976:MBL

[KSM76] Kozo Kinoshita, Tsutomu Sasao, and Jun Matsuda. On magnetic bubble logic circuits. IEEE Transactions on Computers, C-25 (3):247–253, March 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009251.

Kratz:1974:MAS

[KSW74] G. L. Kratz, W. W. Sproul, and E. T. Walendziewicz. A microprogrammed approach to signal processing. IEEE Transactions on Computers, C-23(8):808–817, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672630.

Kitahashi:1972:OEM

[KT72] Tadahiro Kitahashi and Kohkichi Tanaka. Orthogonal expansion of many-valued logical functions and its application to their realization with a single-threshold element. IEEE Transactions on Computers, C-21(2):211–218, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008930.

Kuck:1970:PHS

[Kuc70] D. J. Kuck. A preprocessing high-speed memory system. IEEE Transactions on Computers, C-19(9):793–802, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671635.

Kulisch:1977:MFC

[Kul77] Ulrich Kulisch. Mathematical foundation of computer arithmetic. IEEE Transactions on Computers, C-26(7):610–621, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674893.

Kulkarni:1978:MAH

[Kul78] A. V. Kulkarni. On the mean accuracy of hierarchical classifiers. IEEE Transactions on Computers, C-27(8):771–776, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675190.

Kunt:1975:CHT

[Kun75] M. Kunt. On computation of the Hadamard transform and the R transform in ordered form. IEEE Transactions on Computers, C-24(11):1120–1121, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672736.

Kunz:1979:EBO

[Kun79] H. O. Kunz. On the equivalence between one-dimensional discrete Walsh–Hadamard and multidimensional discrete Fourier transforms. IEEE Transactions on Computers, C-28(3):267–268, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675334.

Kuo:1974:BSS

[Kuo74] B. C. Kuo. B 74-31 signals, systems, and the computer. IEEE Transactions on Computers, C-23(5):557–558, May 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672580.

Kurtzberg:1974:MCP

[Kur74] J. M. Kurtzberg. On the memory conflict problem in multiprocessor systems. IEEE Transactions on Computers, C-23(3): 286–293, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672517.

Kurtzberg:1973:BPA

[KV73] J. M. Kurtzberg and R. D. Villani. A balanced pipelining approach to multiprocessing on an instruction stream level. IEEE Transactions on Computers, C-22 (2):143–148, February 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672268.

Korn:1976:AHC

[KV76] G. A. Korn and R. Vichnevetsky. Analog/hybrid computation and digital simulation. IEEE Transactions on Computers, C-25 (12):1312–1320, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674599.

Kameda:1970:SMN

[KW70] T. Kameda and P. Weiner. On the state minimization of nondeterministic finite automata. IEEE Transactions on Computers, C-19(7):617–627, July 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671587.

Kambayashi:1970:FMS

[KYO70] Y. Kambayashi, S. Yajima, and I. Ohbayashi. On finite-memory sequential machines. IEEE Transactions on Computers, C-19(3): 254–258, March 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671498.

Lamagna:1979:CMN

[Lam79a] E. A. Lamagna. The complexity of monotone networks for certain bilinear forms, routing problems, sorting, and merging. IEEE Transactions on Computers, C-28(10):773–782, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675245.

Lamport:1979:HMM

[Lam79b] L. Lamport. How to make a multiprocessor computer that correctly executes multiprocess programs. IEEE Transactions on Computers, C-28(9):690–691, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675439.

Lang:1976:IBP

[Lan76] T. Lang. Interconnections between processors and memory modules using the shuffle-exchange network. IEEE Transactions on Computers, C-25(5):496–503, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674637.

Landis:1977:MRR

[Lan77] D. Landis. Multiple-response resolution in associative systems. IEEE Transactions on Computers, C-26(3):230–235, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674811.

Langdon:1978:DCT

[Lan78] G. G. Langdon, Jr. A decomposition chart technique to aid in realizations with multiplexers. IEEE Transactions on Computers, C-27(2):157–159, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675048.

Langdon:1979:DMI

[Lan79] G. G. Langdon, Jr. Database machines: An introduction. IEEE Transactions on Computers, C-28(6):381–383, June 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675377.

Larsen:1971:CAP

[Lar71] R. P. Larsen. Computer-aided preliminary layout design of customized MOS arrays. IEEE Transactions on Computers, C-20(5):512–523, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671878.

Lawler:1970:RGT

[Law70] E. L. Lawler. R70-9 A graphtheoretic model for periodic discrete structures. IEEE Transactions on Computers, C-19(6):563, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671573.

Lawrie:1975:AAD

[Law75] D. H. Lawrie. Access and alignment of data in an array processor. IEEE Transactions on Computers, C-24(12):1145–1155, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672750.

Latino:1974:CCE

[LB74] C. D. Latino and J. G. Bredeson. Comments on “Cause-Effect Analysis for Multiple Fault Detection in Combinational Networks”. IEEE Transactions on Computers, C-23(12):1326, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672452. See [BH71a].

Little:1972:DMA

[LC72] W. D. Little and A. C. Capel.

Digital multiplexing analog signals. IEEE Transactions on Computers, C-21(8):920, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009054.

Lee:1979:ASB

[LC79] Share Young Lee and Hsu Chang.

Associative-search bubble devices for content-addressable memory and array logic. IEEE Transactions on Computers, C-28(9): 627–636, September 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675430.

Lee:1972:CMT

[LD72a] Hsiao P. Lee and Edward S. Davidson. Comments on “A Minimization Technique for TANT Networks”. IEEE Transactions on Computers, C-21(4):407, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008987. See [Koh71].

Lee:1972:TNN

[LD72b] Hsiao-Peng Lee and E. S. Davidson. A transform for NAND network design. IEEE Transactions on Computers, C-21(1):12–20, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672019.

Lee:1974:RTC

[LD74] H.-P. S. Lee and E. S. Davidson. Redundancy testing in combinational networks. IEEE Transactions on Computers, C-23(10): 1029–1047, October 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672397. See comments

[Smi79b].

Lakshminarasimhan:1975:UAF

[LD75] A. L. Lakshminarasimhan and B. V. Dasarathy. A unified approach to feature selection and learning in unsupervised environments. IEEE Transactions on Computers, C-24(9):948–952, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672939.

Lempel:1971:HSG

[LE71] A. Lempel and W. L. Eastman. High speed generation of maximal length sequences. IEEE Transactions on Computers, C-20(2): 227–229, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671813.

Lee:1977:FNS

[LE77] S. C. Lee and A. D. Edgar. The Focus number system. IEEE Transactions on Computers, C-26 (11):1167–1170, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674770. See comments

[LE79, Swa79a].

Lee:1979:AFN

[LE79] S. C. Lee and A. D. Edgar. Addendum to “The Focus Number System”. IEEE Transactions on Computers, C-28(9):693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675442. See [LE77,

Swa79a].

Lechner:1970:CCF

[Lec70a] R. J. Lechner. Comment on “Computation of the Fast Walsh–Fourier Transform”. IEEE Transactions on Computers, C-19(2):174, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671483. See [Hen70].

Lechner:1970:TAL

[Lec70b] R. J. Lechner. A transform approach to logic design. IEEE Transactions on Computers, C-19(7):627–640, July 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671588.

Ledley:1972:AC

[Led72] R. S. Ledley. Analysis of cells. IEEE Transactions on Computers, C-21(7):740–753, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672170.

Leeman:1975:SPC

[Lee75] G. B. Leeman, Jr. Some problems in certifying microprograms. IEEE Transactions on Computers, C-24(5):545–553, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672851.

Lee:1976:VBA

[Lee76] S. C. Lee. Vector Boolean algebra and calculus. IEEE Transactions on Computers, C-25(9): 865–874, September 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674711.

Lee:1978:MFP

[Lee78a] Deok Kyo Lee. Method of folding a piecewise polynomial function in the delta function integral representation. IEEE Transactions on Computers, C-27(12): 1177–1178, December 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675019.

Lee:1978:AMT

[Lee78b] H.-P. S. Lee. An algorithm for minimal TANT network generation. IEEE Transactions on Computers, C-27(12):1202–1206, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675028.

Leinius:1974:BCT

[Lei74] R. P. Leinius. B 74-29 currents in the theory of computing. IEEE Transactions on Computers, C-23(5):556–557, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672578.

Lempel:1970:HBG

[Lem70] A. Lempel. On a homomorphism of the de Bruijn graph and its applications to the design of feedback shift registers. IEEE Transactions on Computers, C-19(12):1204–1209, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671452.

Lenfant:1977:FRS

[Len77] J. Lenfant. Fast random and sequential access to dynamic memories of any size. IEEE Transactions on Computers, C-26(9): 847–855, September 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674933.

Lenfant:1978:PPD

[Len78] J. Lenfant. Parallel permutations of data: a Benes network control algorithm for frequently used permutations. IEEE Transactions on Computers, C-27(7):637–647, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675164.

Lesser:1971:IDE

[Les71] V. R. Lesser. An introduction to the direct emulation of control structures by a parallel microcomputer. IEEE Transactions on Computers, C-20(7):751–764, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671937.

Levitt:1970:RTR

[Lev70] K. N. Levitt. R70-38 the time required for group multiplication. IEEE Transactions on Computers, C-19(9):859–860, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671659.

Lewis:1974:BTP

[Lew74] P. M. Lewis. B74-17 the theory of parsing, translation, and compiling — Volume II: Compiling. IEEE Transactions on Computers, C-23(2):219–220, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672499.

Lew:1977:ACG

[Lew77] A. Lew. On analog computer generation of continuous functions. IEEE Transactions on Computers, C-26(3):193–195, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674807.

Lee:1972:SSA

[LF72] H. C. Lee and King-Sun Fu. A stochastic syntax analysis procedure and its application to pattern classification. IEEE Transactions on Computers, C-21(7):660–666, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672164.

Levy:1974:EFS

[LF74] L. S. Levy and M. Freeman. Every finite-state machine can be simulated (realized) by a synchronous (asynchronous) binary feedback shift-register machine. IEEE Transactions on Computers, C-23(2):174–178, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672474.

Lu:1977:SEC

[LF77] Shin-Yee Lu and King-Sun Fu. Stochastic error-correcting syntax analysis for recognition of noisy patterns. IEEE Transactions on Computers, C-26(12): 1268–1276, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674788.

Lu:1978:ECT

[LF78] Shin-Yee Lu and King-Sun Fu. Error-correcting tree automata for syntactic pattern recognition. IEEE Transactions on Computers, C-27(11):1040–1053, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674993.

Locs:1974:FED

[LG74] G. Locs and J. M. Gary. A Fortran extension for data display. IEEE Transactions on Computers, C-23(12):1257–1263, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672437.

Lanzl:1975:DCO

[LH75] F. L. Lanzl and H. Heitmann. Direct coherent Optical Fourier Transform of curves. IEEE Transactions on Computers, C-24(4):434–442, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672832. See comments

[FHTS78].

Lee:1976:ATA

[LH76] Der Tsai Lee and Se June Hong.

An algorithm for transformation of an arbitrary switching function to a completely symmetric function. IEEE Transactions on Computers, C-25(11):1117–1123, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674562.

Little:1979:ASG

[LH79] W. D. Little and R. Heuft. An area shading graphics display system. IEEE Transactions on Computers, C-28(7):528–531, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675399.

Liu:1974:OOB

[LHSM74] Tso-Kai Liu, K. R. Hohulin, Lih-Er Shiau, and S. Muroga. Optimal one-bit full adders with different types of gates. IEEE Transactions on Computers, C-23(1):63–70, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672371.

Li:1977:STP

[Li77] H. F. Li. Scheduling trees in parallel/pipelined processing environments. IEEE Transactions on Computers, C-26(11):1101– 1112, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674756.

Ligomenides:1977:SSF

[Lig77] P. A. Ligomenides. The skipand-set fast-division algorithm. IEEE Transactions on Computers, C-26(10):1030–1032, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674740.

Lillestrand:1972:TCD

[Lil72] R. L. Lillestrand. Techniques for change detection. IEEE Transactions on Computers, C-21(7):654–659, July 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672163.

Ling:1970:HSC

[Lin70] H. Ling. High-speed computer multiplication using a multiplebit decoding algorithm. IEEE Transactions on Computers, C-19 (8):706–709, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671613;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Ling.pdf.

Liou:1976:SFM

[Lio76] Ming-Lei Liou. Spline fit made easy. IEEE Transactions on Computers, C-25(5):522–527, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674640.

Lipovski:1977:VAM

[Lip77] G. Jack Lipovski. On a varistructured array of microprocessors. IEEE Transactions on Computers, C-26(2):125–138, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009291.

Little:1973:SHC

[Lit73] W. D. Little. Serial hybrid computation and errors in hybrid loops. IEEE Transactions on Computers, C-22(4):367–370, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672316.

Little:1974:AHS

[Lit74] W. D. Little. An algorithm for high-speed digital filters. IEEE Transactions on Computers, C-23(5):466–469, May 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672561.

Liu:1975:SAL

[Liu75] Tso-Kai Liu. Synthesis algorithms for 2-level MOS networks. IEEE Transactions on Computers, C-24(1):72–79, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672677.

Liu:1977:SFF

[Liu77a] T. K. Liu. Synthesis of feed-forward MOS networks with cells of similar complexities. IEEE Transactions on Computers, C-26(8):826–831, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674923.

Liu:1977:SMF

[Liu77b] Tso-Kai Liu. Synthesis of multilevel feed-forward MOS networks. IEEE Transactions on Computers, C-26(6):581–588, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674883.

Liu:1979:IMB

[LJ79] J. W. S. Liu and M. Jino. Intelligent magnetic bubble memories and their applications in data base management systems. IEEE Transactions on Computers, C-28(12):888–906, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675279.

Leonard:1975:MFI

[LK75] M. S. Leonard and K. E. Kilpatrick.

A modified fixed-increment training algorithm. IEEE Transactions on Computers, C-24(6):673–675, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672874.

Lee:1972:MSF

[LL72] Samuel C. Lee and Edward T. Lee.

On multivalued symmetric functions. IEEE Transactions on Computers, C-21(3):312–317, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008957.

Landauro:1975:ODC

[LL75] A. Landauro and J. Lienard.

On overflow detection and correction in digital filters. IEEE Transactions on Computers, C-24 (12):1226–1228, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672761.

Leung:1979:MCR

[LL79] J. Y.-T. Leung and E. K. Lai. On minimum cost recovery from system deadlock. IEEE Transactions on Computers, C-28(9): 671–677, September 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675435.

Loomis:1970:SSHa

[LM70] Herschel H. Loomis and Michael R. McCoy. A scheme for synchronizing high-speed logic: Part I. IEEE Transactions on Computers, C-19(1):39–47, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008898.

Lackey:1971:AR

[LM71a] R. B. Lackey and D. Meltzer. Authors’ reply. IEEE Transactions on Computers, C-20(12):1617, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671783.

Lackey:1971:SDW

[LM71b] R. B. Lackey and D. Meltzer. A simplified definition of Walsh functions. IEEE Transactions on Computers, C-20(2):211–213, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&

arnumber=1671807. See comments

[Car71b, KS72].

Liu:1978:TPE

[LM78] P. S. Liu and F. J. Mowle.

Techniques of program execution with a writable control memory. IEEE Transactions on Computers, C-27(9):816–827, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675201.

Lai:1979:MPB

[LM79] Hung Chi Lai and S. Muroga.

Minimum parallel binary adders with NOR (NAND) gates. IEEE Transactions on Computers, C-28(9):648–659, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675433.

Lange-Nielsen:1972:PSA

[LNL72] T. Lange-Nielsen and G. M. Lance.

A pattern search algorithm for feedback-control system parameter optimization. IEEE Transactions on Computers, C-21(11): 1222–1227, November 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672074.

Lomet:1974:BSE

[Lom74] D. B. Lomet. B74-22 software engineering: International computer state of the art report.

IEEE Transactions on Computers, C-23(2):223–224, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672504.

Loomis:1970:SSHb

[Loo70] H. H. Loomis, Jr. A scheme for synchronizing high-speed logic. Part II. IEEE Transactions on Computers, C-19(2):116–124, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671469.

Losq:1976:HER

[Los76] J. Losq. A highly efficient redundancy scheme: Self-purging redundancy. IEEE Transactions on Computers, C-25(6):569–578, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674656.

Losq:1978:ERC

[Los78] J. Losq. Efficiency of random compact testing. IEEE Transactions on Computers, C-27(6):516–525, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675142.

Loutrel:1970:SHL

[Lou70] P. P. Loutrel. A solution to the hidden-line problem for computerdrawn polyhedra. IEEE Transactions on Computers, C-19(3): 205–213, March 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671491.

Lowe:1973:AIS

[Low73] T. C. Lowe. Analysis of an information system model with transfer penalties. IEEE Transactions on Computers, C-22(5):469–480, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672343.

Luccio:1976:HHB

[LP76] Fabrizio Luccio and Linda Pagli.

On the height of height-balanced trees. IEEE Transactions on Computers, C-25(1):87–90, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009208.

Luccio:1978:RHB

[LP78] F. Luccio and L. Pagli. Rebalancing height balanced trees. IEEE Transactions on Computers, C-27(5):386–396, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675117.

Landman:1971:PVB

[LR71a] B. S. Landman and R. L. Russo.

On a pin versus block relationship for partitions of logic graphs. IEEE Transactions on Computers, C-20(12):1469–1479, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671752.

Laws:1971:CAM

[LR71b] B. A. Laws, Jr. and C. K. Rushforth. A cellular-array multiplier for GF(2^{m}). IEEE Transactions on Computers, C-20(12): 1573–1578, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671766.

Larsen:1972:RCV

[LR72] Ronald W. Larsen and Irving S. Reed. Redundancy by coding versus redundancy by replication for failure-tolerant sequential circuits. IEEE Transactions on Computers, C-21(2):130–137, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008917.

Liu:1975:DMC

[LR75] J. J. Liu and L. D. Rudolph. A direct method of computing the GNAF of an integer. IEEE Transactions on Computers, C-24(10):1042–1043, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672719.

Luczak:1976:DHG

[LR76] E. Luczak and A. Rosenfeld. Distance on a hexagonal grid. IEEE Transactions on Computers, C-25(5):532–533, May 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674642.

Lawson:1971:FCM

[LS71] H. W. Lawson, Jr. and B. K.

Smith. Functional characteristics of a multilingual processor. IEEE Transactions on Computers, C-20(7):732–742, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671935.

Lawrence:1972:RPS

[LS72] J. P. Lawrence and Kenneth Steiglitz. Randomized pattern search. IEEE Transactions on Computers, C-21(4):382–385, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008979.

Lang:1976:SEN

[LS76a] Tomas Lang and Harold S. Stone. A shuffle-exchange network with simplified control. IEEE Transactions on Computers, C-25(1):55–65, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009205.

Loveland:1976:HGT

[LS76b] D. W. Loveland and M. E. Stickel. A hole in goal trees: Some guidance from resolution theory. IEEE Transactions on Computers, C-25(4):335–341, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674614.

Lee:1977:TMS

[LSB77] R. C. T. Lee, J. R. Slagle, and H. Blum. A triangulation method for the sequential mapping of points from N-space to twospace. IEEE Transactions on Computers, C-26(3):288–292, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674822.

Lakshmivarahan:1970:PCU

[LT70] S. Lakshmivarahan and M. A. L. Thathachar. Pattern classification using stochastic approximation techniques. IEEE Transactions on Computers, C-19(7):649–651, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671591.

Larson:1973:PPA

[LT73] Robert E. Larson and Edison Tse. Parallel processing algorithms for the optimal control of nonlinear dynamic systems. IEEE Transactions on Computers, C-22(8):777–786, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009158. See comments

[Tab74].

Lavington:1977:MMC

[LTE77] Simon H. Lavington, Gareth Thomas, and David B. G. Edwards. The MU5 multicomputer communication system. IEEE Transactions on Computers, C-26(1):19–28, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009269.

Lucas:1972:BCB

[Luc72] H. C. Lucas, Jr. B72-20 computers in business — an introduction, 2nd ed. IEEE Transactions on Computers, C-21(10):1136, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672060.

Luderer:1974:BDL

[Lud74] G. W. R. Luderer. B74-16 design of on-line computer systems. IEEE Transactions on Computers, C-23(2):218–219, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672498.

LeVan:1975:DUL

[LV75] T. Le Van and N. Van Houtte. Delayed universal logic modules and sequential machine synthesis. IEEE Transactions on Computers, C-24(8):853–855, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672916. See comments

[Sur78].

Liu:1978:ECC

[LW78] Chao-Kai Liu and Tse Lin Wang. Error-correcting codes in binarycoded-decimal arithmetic. IEEE Transactions on Computers, C-27(11):977–984, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674987.

Landgraff:1971:DDI

[LY71] R. W. Landgraff and S. S. Yau. Design of diagnosable iterative arrays. IEEE Transactions on Computers, C-20(8):867–877, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671956. See comments

[FM72b].

Lynn:1972:NRR

[Lyn72] Donald S. Lynn. New results for Rado’s sigma function for binary Turing Machines. IEEE Transactions on Computers, C-21 (8):894–896, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009047.

Lyons:1970:SRT

[Lyo70] R. E. Lyons. The synthesis of redundant threshold-logic elements. IEEE Transactions on Computers, C-19(5):429–443, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671534.

Marihugh:1971:DGA

[MA71] G. E. Marihugh and R. E. Anderson. The H diagram: a graphical approach to logic design. IEEE Transactions on Computers, C-20(10):1192–1196, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671697.

McKee:1977:CRP

[MA77] J. W. McKee and J. K. Aggarwal. Computer recognition of partial views of curved objects. IEEE Transactions on Computers, C-26(8):790–800, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674917.

Maritsas:1978:PSA

[MAB78] D. G. Maritsas, A. C. Arvillias, and A. C. Bounas. Phase-shift analysis of linear feedback shift register structures generating pseudorandom sequences. IEEE Transactions on Computers, C-27(7):660–669, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675166.

Macleod:1970:POL

[Mac70] I. D. G. Macleod. Pictorial output with a line printer. IEEE Transactions on Computers, C-19 (2):160–162, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671476. See correction

[Ano70y].

Mage:1970:AIC

[Mag70] J. M. Mage. Application of iterative consensus to multiple-output functions. IEEE Transactions on Computers, C-19(4):359, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671517.

Mago:1971:RMA

[Mag71] G. Mago. Realization methods for asynchronous sequential circuits. IEEE Transactions on Computers, C-20(3):290–297, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671826.

Mago:1973:MFS

[Mag73] G. Mago. Monotone functions in sequential circuits. IEEE Transactions on Computers, C-22(10): 928–933, October 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672213. See comments

[Iba74].

Mahl:1972:VSA

[Mah72] Robert Mahl. Visible surface algorithms for quadric patches. IEEE Transactions on Computers, C-21(1):1–4, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672017.

Maheshwari:1974:BCT

[Mah74] S. N. Maheshwari. B74-36 communication, transmission, and transportation networks. IEEE Transactions on Computers, C-23(8):878, August 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672637.

Maison:1971:MSR

[Mai71] F. P. Maison. The MECRA: a self-reconfigurable computer for highly reliable process. IEEE Transactions on Computers, C-20 (11):1382–1388, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671736.

Majithia:1972:CAE

[Maj72a] J. C. Majithia. Cellular array for extraction of squares and square roots of binary numbers. IEEE Transactions on Computers, C-21(9):1023–1024, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009084.

Majithia:1972:CFD

[Maj72b] J. C. Majithia. Comments on “A Fast Digital Computer Method for Recursive Estimation of the Mean”. IEEE Transactions on Computers, C-21(5):511–512, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672147. See [Whi70].

Majithia:1972:STD

[Maj72c] J. C. Majithia. A simple technique for determination of essential multiple output prime implicants. IEEE Transactions on Computers, C-21(9):1024–1026, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009085.

Majithia:1973:CAB

[Maj73] J. C. Majithia. A cellular array for binary to BCD conversion. IEEE Transactions on Computers, C-22(4):423–424, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672325.

Majithia:1976:SCC

[Maj76] J. C. Majithia. Some comments concerning design of pipeline arithmetic arrays. IEEE Transactions on Computers, C-25(11): 1132–1134, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674565.

Makhoul:1971:SAL

[Mak71] J. Makhoul. Speaker adaptation in a limited speech recognition system. IEEE Transactions on Computers, C-20(9):1057–1063, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671996.

Mallach:1972:JMM

[Mal72] E. G. Mallach. Job-mix modeling and system analysis of an aerospace multiprocessor. IEEE Transactions on Computers, C-21(5):446–454, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672133.

Mallach:1973:UIW

[Mal73] E. G. Mallach. On the use of “importance weights” in assessing subsystem reliability. IEEE Transactions on Computers, C-22(4):424–425, April 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672326.

Mamdani:1977:AFL

[Mam77] E. H. Mamdani. Application of fuzzy logic to approximate reasoning using linguistic synthesis. IEEE Transactions on Computers, C-26(12):1182–1191, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674779.

Mandelbaum:1971:PAI

[Man71] D. Mandelbaum. On parallel-acting index registers. IEEE Transactions on Computers, C-20(3):361, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671839.

Mandelbaum:1972:ECS

[Man72a] D. Mandelbaum. On error control in sequential machines. IEEE Transactions on Computers, C-21(5):492–495, May 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672140.

Mandelbaum:1972:ECR

[Man72b] David Mandelbaum. Error correction in residue arithmetic. IEEE Transactions on Computers, C-21(6):538–545, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009006.

Manning:1977:AHI

[Man77] F. B. Manning. An approach to highly integrated, computermaintained cellular arrays. IEEE Transactions on Computers, C-26(6):536–552, June 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674879. See comments

[Agr79c].

Marino:1970:LDS

[Mar70] P. J. Marino. A linear decomposition for sequential machines. IEEE Transactions on Computers, C-19(10):956–963, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671395.

Marinos:1971:DMC

[Mar71] P. N. Marinos. Derivation of minimal complete sets of testinput sequences using Boolean differences. IEEE Transactions on Computers, C-20(1):25–32, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671670.

Marino:1972:NAA

[Mar72] D. Marino. New algorithms for the approximate evaluation in hardware of binary logarithms and elementary functions. IEEE Transactions on Computers, C-21(12): 1416–1421, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672109.

Maritsaa:1973:SPC

[Mar73a] D. G. Maritsaa. On the statistical properties of a class of linear product feedback shiftregister sequences. IEEE Transactions on Computers, C-22(10): 961–962, October 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672218.

Maritsas:1973:AFT

[Mar73b] D. G. Maritsas. The autocorrelation function of the two feedback shift-register pseudorandom source. IEEE Transactions on Computers, C-22(10):962–964, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672219.

Maritsas:1973:HSA

[Mar73c] Dimitris G. Maritsas. A high speed and accuracy digital Gaussian generator of pseudorandom numbers. IEEE Transactions on Computers, C-22(7):629–634, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009127.

Maruyama:1973:PEP

[Mar73d] K. Maruyama. On the parallel evaluation of polynomials. IEEE Transactions on Computers, C-22(1):2–5, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672186.

Marjanovic:1974:MDC

[Mar74] S. Marjanovic. Multiplying D/A converter. IEEE Transactions on Computers, C-23(9): 986–988, September 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672661.

Marinos:1975:AR

[Mar75] P. N. Marinos. Author’s reply.

IEEE Transactions on Computers, C-24(1):109, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672685.

Marinos:1976:AR

[Mar76] P. N. Marinos. Author’s reply. IEEE Transactions on Computers, C-25(10):1056, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674549.

Marino:1977:EAI

[Mar77] L. R. Marino. The effect of asynchronous inputs on sequential network reliability. IEEE Transactions on Computers, C-26 (11):1082–1090, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674754.

Marovac:1978:SMR

[Mar78] N. Marovac. Sieve method for real-time computer-aided extraction of multivalued analog signals in the presence of noise. IEEE Transactions on Computers, C-27(10):935–944, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674973.

Markowsky:1979:DSF

[Mar79] G. Markowsky. Diagnosing single faults in fanout-free combinational circuits. IEEE Transactions on Computers, C-28(11): 863–864, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675266.

Mase:1978:CMC

[Mas78] K. Mase. Comments on “A Measure of Computational Work” and “Logical Network Cost and Entropy”. IEEE Transactions on Computers, C-27(1):94–95, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674960. See [Wel74,

Hel74, CF73].

Matula:1970:FFP

[Mat70] D. W. Matula. A formalization of floating-point numeric base conversion. IEEE Transactions on Computers, C-19(8):681–692, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671610;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Matula.
pdf.

Matheson:1971:PEC

[Mat71a] W. S. Matheson. PCN equivalence class invariants and information quantities. IEEE Transactions on Computers, C-20(6):691–694, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671922.

Matheson:1971:RMU

[Mat71b] W. S. Matheson. Recognition of monotonic and unate cascade realizable functions using an informational model of switching circuits. IEEE Transactions on Computers, C-20(10):1214–1219, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671702.

Mathur:1971:RMA

[Mat71c] F. P. Mathur. On reliability modeling and analysis of ultrareliable fault-tolerant digital systems. IEEE Transactions on Computers, C-20(11):1376–1382, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671735.

Maxwell:1977:CDR

[Max77] P. C. Maxwell. Correct DDA register transfers for trapezoidal integration when solving nonlinear equations of the form ẏ = Ay^{n}. IEEE Transactions on Computers, C-26(11):1151–1153, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674764.

Mazer:1971:TSB

[Maz71] L. Mazer. Topological solution of bilateral switching networks. IEEE Transactions on Computers, C-20(2):234–238, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671816.

McCrea:1975:DDA

[MB75] P. G. McCrea and P. W. Baker.

On Digital Differential Analyzer (DDA) circle generation for computer graphics. IEEE Transactions on Computers, C-24(11): 1109–1110, November 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672733.

Maryanski:1977:IFS

[MB77] F. J. Maryanski and T. L. Booth.

Inference of finite-state probabilistic grammars. IEEE Transactions on Computers, C-26(6):521–536, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674878.

Maryanski:1979:AR

[MB79] F. Maryanski and T. L. Booth. Authors’ reply. IEEE Transactions on Computers, C-28(1):64, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675225.

Miller:1977:STD

[MBM77] A. J. Miller, A. W. Brown, and P. Mars. A simple technique for the determination of delayed maximal length linear binary sequences. IEEE Transactions on Computers, C-26(8):808–811, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674919.

Minnick:1975:CRM

[MBSS75] R. C. Minnick, P. T. Bailey, R. M. Sandfort, and W. L. Semon. Cascade realizations of magnetic bubble logic using a small set of primitives. IEEE Transactions on Computers, C-24(2):215–217, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672782.

Moore:1970:SCA

[MC70] J. B. Moore and K. T. Clark. A simple convergent algorithm for rapid solution of polynomial equations. IEEE Transactions on Computers, C-19(1):79–80, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008904.

McCluskey:1971:FEC

[MC71] E. J. McCluskey and F. W. Clegg. Fault equivalence in combinational logic networks. IEEE Transactions on Computers, C-20 (11):1286–1293, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671722.

Menon:1978:DFS

[MC78] P. R. Menon and S. G. Chappell. Deductive fault simulation with functional blocks. IEEE Transactions on Computers, C-27 (8):689–695, August 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675175.

McCluskey:1979:LDM

[McC79] E. J. McCluskey. Logic design of multivalued I^{2}L logic circuits. IEEE Transactions on Computers, C-28(8):546–559, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675410.

Mcdougall:1972:NIE

[Mcd72] M. H. Mcdougall. A note on the interruption of extended core storage transfers. IEEE Transactions on Computers, C-21(1):87–90, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672028.

McLean:1972:BMD

[McL72] E. R. McLean. B72-23 modem data processors and systems. IEEE Transactions on Computers, C-21(11):1251–1252, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672083.

McCaughern:1972:VSF

[MCRD72] R. W. McCaughern, D. M. Caughey, H. Rombeek, and W. S. Davidson. A video simulation facility. IEEE Transactions on Computers, C-21(7):738–740, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672169.

Mei:1974:BSF

[Mei74] K. C. Y. Mei. Bridging and stuckat faults. IEEE Transactions on Computers, C-23(7):720–727, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672613.

Menger:1970:AR

[Men70a] K. S. Menger. Author’s reply. IEEE Transactions on Computers, C-19(6):560, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671571.

Mengert:1970:SLI

[Men70b] P. H. Mengert. Solution of linear inequalities. IEEE Transactions on Computers, C-19(2): 124–131, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671470.

Meo:1975:ANT

[Meo75] A. R. Meo. Arithmetic networks and their minimization using a new line of elementary units. IEEE Transactions on Computers, C-24(3):258–280, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672800.

Metropolis:1973:ABC

[Met73] Nicholas C. Metropolis. Analyzed binary computing. IEEE Transactions on Computers, C-22(6):573–576, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009109.

Meyer:1971:FTS

[Mey71] J. F. Meyer. Fault tolerant sequential machines. IEEE Transactions on Computers, C-20(10): 1167–1177, October 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671694.

Meyer:1976:CBR

[Mey76] J. F. Meyer. Computationbased reliability analysis. IEEE Transactions on Computers, C-25(6):578–584, June 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674657.

Menon:1971:FDI

[MF71] P. R. Menon and A. D. Friedman. Fault detection in iterative logic arrays. IEEE Transactions on Computers, C-20(5):524–535, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671879.

Merlin:1975:PMD

[MF75] P. M. Merlin and D. J. Farber. A parallel mechanism for detecting curves in pictures. IEEE Transactions on Computers, C-24 (1):96–98, January 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672680.

Malhotra:1976:DEC

[MF76a] Vinod K. Malhotra and Robert D. Fisher. A double error-correction scheme for peripheral systems. IEEE Transactions on Computers, C-25(2):105–115, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009225.

Manning:1976:SCC

[MF76b] Frank B. Manning and Robert R.

Fenichel. Synchronous counters constructed entirely of J–K flipflops. IEEE Transactions on Computers, C-25(3):300–306, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009257.

Moayer:1976:TSA

[MF76c] Bijan Moayer and King-Sun Fu.

A tree system approach for fingerprint pattern recognition. IEEE Transactions on Computers, C-25 (3):262–274, March 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009253.

Montalvo:1976:AIC

[MF76d] Fanya S. Montalvo and Caxton C.

Foster. An algorithm for intercell communication in a tesselated automaton. IEEE Transactions on Computers, C-25(1):99–102, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009214.

Marouf:1978:EDS

[MF78] M. A. Marouf and A. D. Friedman. Efficient design of selfchecking checker for any m-out-of-n code. IEEE Transactions on Computers, C-27(6):482–490, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675138.

Michales:1971:HCS

[MG71a] G. C. Michales and V. Gourishankar. Hybrid computer solution of optimal control problems. IEEE Transactions on Computers, C-20(2):209–211, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671806.

Mucciardi:1971:CST

[MG71b] A. N. Mucciardi and E. E. Gose. A comparison of seven techniques for choosing subsets of pattern recognition properties. IEEE Transactions on Computers, C-20(9):1023–1031, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671991. See comments

[Tou72a].

McDonald:1975:TCQ

[MG75] T. G. McDonald and R. K. Guha. The two’s complement quasi-serial multiplier. IEEE Transactions on Computers, C-24 (12):1233–1235, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672763.

Maritsas:1970:BLE

[MH70a] D. G. Maritsas and M. G. Hartley. Buffer length for Erlang input and constant removal rate. IEEE Transactions on Computers, C-19(9):839–843, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671644.

Maritsas:1970:CSV

[MH70b] D. G. Maritsas and M. G. Hartley. A case study of a versatile generator of repeatable non-Poisson sequences of pseudorandom pulses. IEEE Transactions on Computers, C-19(10):924–938, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671393.

Maritsas:1970:DCG

[MH70c] D. G. Maritsas and M. G. Hartley. Design criteria for a generator of repeatable non-Poisson sequences of pseudorandom pulses. IEEE Transactions on Computers, C-19(9):812–817, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671638.

Maruoka:1973:LNF

[MH73] A. Maruoka and N. Honda. Logical networks of flexible cells.

IEEE Transactions on Computers, C-22(4):347–358, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672314.

Maruoka:1975:RLF

[MH75] A. Maruoka and N. Honda. The range of logical flexibility of tree networks. IEEE Transactions on Computers, C-24(1):9–28, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672673.

Maheshwari:1976:MDS

[MH76] Shachindra N. Maheshwari and S. Louis Hakimi. On models for diagnosable systems and probabilistic fault diagnosis. IEEE Transactions on Computers, C-25 (3):228–236, March 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009249. See corrections

and comments [MH78].

Maheshwari:1978:CCM

[MH78] S. N. Maheshwari and S. L. Hakimi. Corrections and comments on “On Models for Diagnosable Systems and Probabilistic Fault Diagnosis”. IEEE Transactions on Computers, C-27(3):287–288, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675092. See [MH76].

Muroga:1972:DOS

[MI72] Saburo Muroga and Toshihide Ibaraki. Design of optimal switching networks by integer programming. IEEE Transactions on Computers, C-21(6):573–582, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009010.

Michalopouos:1974:BPR

[Mic74] D. A. Michalopouos. B74-46 pattern recognition, learning and thought. IEEE Transactions on Computers, C-23(12):1330, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672460.

Miller:1973:CST

[Mil73] Raymond E. Miller. A comparison of some theoretical models of parallel computation. IEEE Transactions on Computers, C-22 (8):710–717, August 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009149.

Miller:1974:BOS

[Mil74a] E. F. Miller, Jr. B74-25 operating systems: a pragmatic approach. IEEE Transactions on Computers, C-23(3):336, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672534.

Miller:1974:MSR

[Mil74b] E. F. Miller, Jr. A multiple-stream registerless shared-resource processor. IEEE Transactions on Computers, C-23(3):277–285, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672516.

Milgram:1975:CMC

[Mil75] D. L. Milgram. Computer methods for creating photomosaics. IEEE Transactions on Computers, C-24(11):1113–1119, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672735.

Milgram:1977:ATP

[Mil77] D. L. Milgram. Adaptive techniques for photomosaicking. IEEE Transactions on Computers, C-26 (11):1175–1180, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674772.

Minnick:1975:SMB

[Min75] R. C. Minnick. A system of magnetic bubble logic. IEEE Transactions on Computers, C-24(2): 217–218, February 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672783.

Mouftah:1977:DTC

[MJ77] H. T. Mouftah and I. B. Jordan. Design of ternary COS/ MOS memory and sequential circuits. IEEE Transactions on Computers, C-26(3):281–288, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674821. See correction

[MJ78].

Mouftah:1978:CDT

[MJ78] H. T. Mouftah and I. B. Jordan.

Correction to “Design of Ternary COS/MOS Memory and Sequential Circuits”. IEEE Transactions on Computers, C-27(2):192, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675066. See [MJ77].

McMahon:1975:HOC

[MJTW75] D. H. McMahon, G. L. Johnson, S. L. Teeter, and C. G. Whitney. A hybrid optical computer processing technique for fingerprint identification. IEEE Transactions on Computers, C-24(4):358–369, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672822.

Majithia:1971:CAN

[MK71a] J. C. Majithia and R. Kitai. A cellular array for the nonrestoring extraction of square roots. IEEE Transactions on Computers, C-20(12):1617–1618, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671784.

Majithia:1971:IAM

[MK71b] J. C. Majithia and R. Kitai. An iterative array for multiplication of signed binary numbers. IEEE Transactions on Computers, C-20(2):214–216, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671809.

McPherson:1979:TLD

[MK79] J. A. McPherson and C. R. Kime. A two-level diagnostic model for digital systems. IEEE Transactions on Computers, C-28(1):16–27, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675218.

Murakami:1970:SMC

[MKO70] S. Murakami, K. Kinoshita, and H. Ozaki. Sequential machines capable of fault diagnosis. IEEE Transactions on Computers, C-19 (11):1079–1085, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671428.

Muroga:1976:MLN

[ML76] S. Muroga and Hung Chi Lai. Minimization of logic networks under a generalized cost function. IEEE Transactions on Computers, C-25(9):893–907, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674714.

Morreale:1970:CEP

[MM70] E. Morreale and M. Mennucci.

Computer experience on partitioned list algorithms. IEEE Transactions on Computers, C-19 (11):1099–1105, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671433.

Marasa:1973:SSC

[MM73a] John D. Marasa and David W.

Matula. A simulative study of correlated error propagation in various finite-precision arithmetics. IEEE Transactions on Computers, C-22(6):587–597, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009111;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Marasa.
pdf.

Meisel:1973:PAA

[MM73b] W. S. Meisel and D. A. Michalopoulos. A partitioning algorithm with application in pattern classification and the optimization of decision trees. IEEE Transactions on Computers, C-22(1):93–103, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672196.

Mallela:1978:DSI

[MM78a] S. Mallela and G. M. Masson. Diagnosable systems for intermittent faults. IEEE Transactions on Computers, C-27(6):560–566, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675147.

Meyer:1978:EFD

[MM78b] G. G. L. Meyer and G. M. Masson. An efficient fault diagnosis algorithm for symmetric multiple processor architectures. IEEE Transactions on Computers, C-27(11):1059–1063, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674996.

Mitchell:1977:MMM

[MMB77] O. R. Mitchell, C. R. Myers, and W. Boyne. A max-min measure for image texture analysis. IEEE Transactions on Computers, C-26(4):408–414, April 1977.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674850.

McCrea:1978:CFP

[MMB78] P. G. McCrea, P. C. Maxwell, and P. W. Baker. Comments on “A Floating-Point Multiplexed DDA System”. IEEE Transactions on Computers, C-27(12):1226, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675033. See [HW76].

Matsushita:1974:MCC

[MMO74] S. Matsushita and T. Moto-Oka. Magnitude of cross-coupling noise in digital multiwire transmission lines. IEEE Transactions on Computers, C-23(11): 1122–1132, November 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672412.

McGhee:1970:ERD

[MN70] Robert B. McGhee and Ragnar N. Nilsen. The extended resolution digital differential analyzer: a new computing structure for solving differential equations. IEEE Transactions on Computers, C-19(1):1–9, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008893.

Mark:1978:SPS

[MN78] J. W. Mark and S. F. W. Ng. Satellite packet switching with global assignments and batch Poisson arrivals. IEEE Transactions on Computers, C-27(12): 1216–1221, December 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675031.

Mohn:1971:TSF

[Moh71] W. S. Mohn, Jr. Two statistical feature evaluation techniques applied to speaker identification. IEEE Transactions on Computers, C-20(9):979–987, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671985.

Montangero:1974:AOS

[Mon74] C. Montangero. An approach to the optimal specification of read-only memories in microprogrammed digital computers. IEEE Transactions on Computers, C-23(4):375–389, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672546.

Moore:1971:BSS

[Moo71] F. R. Moore. On the bounds for state-set size in the proofs of equivalence between deterministic, nondeterministic, and twoway finite automata. IEEE Transactions on Computers, C-20(10): 1211–1214, October 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671701.

Moore:1976:DSC

[Moo76] R. C. Moore. D-Script: a computational theory of descriptions. IEEE Transactions on Computers, C-25(4):366–373, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674619.

Morishita:1970:AAT

[Mor70a] I. Morishita. Analysis of an adaptive threshold logic unit. IEEE Transactions on Computers, C-19 (12):1181–1192, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671450.

Morreale:1970:CCP

[Mor70b] E. Morreale. Computational complexity of partitioned list algorithms. IEEE Transactions on Computers, C-19(5):421–428, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671533.

Morreale:1970:ROP

[Mor70c] E. Morreale. Recursive operators for prime implicant and irredundant normal form determination. IEEE Transactions on Computers, C-19(6):504–509, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671560.

Morris:1971:TFP

[Mor71] R. Morris. Tapered floating point: a new floating-point representation. IEEE Transactions on Computers, C-20(12):1578–1579, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671767.

Morrin:1974:BWR

[Mor74] T. H. Morrin. A black-white representation of a gray-scale picture. IEEE Transactions on Computers, C-23(2):184–186, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672476.

Morgan:1975:DCU

[Mor75] D. R. Morgan. A/D conversion using geometric feedback AGC. IEEE Transactions on Computers, C-24(11):1074–1078, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672728.

Mor:1976:TVS

[Mor76a] M. Mor. On the three-valued simulation of digital systems. IEEE Transactions on Computers, C-25 (11):1152–1156, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674571.

Morgan:1976:MAT

[Mor76b] C. G. Morgan. Methods for automated theorem proving in nonclassical logics. IEEE Transactions on Computers, C-25(8):852–862, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674704.

Moravec:1979:FIM

[Mor79] H. P. Moravec. Fully interconnecting multiple computers with pipelined sorting nets. IEEE Transactions on Computers, C-28(10): 795–798, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675249.

McCharen:1976:PEA

[MOW76] J. D. McCharen, R. A. Overbeek, and L. A. Wos. Problems and experiments for and with automated theorem-proving programs. IEEE Transactions on Computers, C-25(8):773–782, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674696.

McCluskey:1978:BNP

[MPS78] E. J. McCluskey, K. P. Parker, and J. J. Shedletsky. Boolean network probabilities and network design. IEEE Transactions on Computers, C-27(2):187–189, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675058.

Miczo:1972:RBO

[MR72] A. Miczo and L. D. Rudolph. A relationship between output symbol occurrence rate and observability of autonomous machines. IEEE Transactions on Computers, C-21(8):911–913, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009052.

Mennone:1974:ECL

[MR74] A. Mennone and R. L. Russo. An example computer logic graph and its partitions and mappings. IEEE Transactions on Computers, C-23(11):1198–1204, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672423.

Molberg:1975:NMI

[MR75] J. R. Molberg and D. K. Reynolds. A new method for initial estimates in the numerical solution of Laplace’s equation. IEEE Transactions on Computers, C-24(6):678–680, June 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672876.

Meyer:1976:FTC

[MR76] J. F. Meyer and J.-C. Rault.

Fault-tolerant computing: a introduction. IEEE Transactions on Computers, C-25(6):553–556, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674654.

Mukhopadhyay:1970:MEL

[MS70] A. Mukhopadhyay and G. Schmitz.

Minimization of exclusive or and logical equivalence switching circuits. IEEE Transactions on Computers, C-19(2):132–140, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671471.

Maki:1973:CNS

[MS73a] G. K. Maki and D. H. Sawin III.

Comments on “Next-State Equations of Asynchronous Sequential Machines”. IEEE Transactions on Computers, C-22(2):218, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672282. See [Sau72a].

Mar:1973:FHT

[MS73b] H. Y. L. Mar and C. L. Sheng. Fast Hadamard Transform using the H diagram. IEEE Transactions on Computers, C-22(10): 957–960, October 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672217.

Maki:1974:FTA

[MS74] G. K. Maki and D. H. Sawin III. Fault-tolerant asynchronous sequential machines. IEEE Transactions on Computers, C-23(7):651–657, July 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672606.

Meyer:1975:LDU

[MS75a] J. F. Meyer and R. J. Sundstrom. On-line diagnosis of unrestricted faults. IEEE Transactions on Computers, C-24(5):468–475, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672841.

Mizoguchi:1975:AUL

[MS75b] R. Mizoguchi and M. Shimura. An approach to unsupervised learning classification. IEEE Transactions on Computers, C-24(10): 979–983, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672697.

Mizoguchi:1976:NLT

[MS76] R. Mizoguchi and M. Shimura.

Nonparametric learning without a teacher based on mode estimation. IEEE Transactions on Computers, C-25(11):1109–1117, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674561.

Mehta:1977:SOA

[MS77a] M. A. Mehta and W. B. Smith. Synthesis of optimal ambiguity resolver functions. IEEE Transactions on Computers, C-26(8):782–789, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674916.

Mitchell:1977:TLM

[MS77b] R. L. Mitchell and C. R. Stone.

Table-lookup methods for generating arbitrary random numbers. IEEE Transactions on Computers, C-26(10):1006–1008, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674735.

Mitra:1978:ITD

[MS78] S. K. Mitra and G. K. Sorknes.

On the implementation of a twodimensional FIR filter using a single multiplier. IEEE Transactions on Computers, C-27(8):762–764, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675186.

Maki:1972:ISA

[MSJ72] G. K. Maki, D. H. Sawin III, and B.-R. A. Jeng. Improved state assignment selection tests. IEEE Transactions on Computers, C-21(12):1443–1449, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672115.

Metze:1975:CDM

[MST^{+}75] G. Metze, D. R. Schertz, Kilin To, G. Whitney, C. R. Kime, and J. D. Russell. Comments on “Derivation of Minimal Complete Sets of Test-Input Sequences Using Boolean Differences”. IEEE Transactions on Computers, C-24(1):108, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672684. See [DDSM76].

Maki:1970:SAS

[MT70] G. K. Maki and J. H. Tracey. State assignment selection in asynchronous sequential circuits. IEEE Transactions on Computers, C-19(7):641–644, July 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671589.

Maki:1971:SAP

[MT71] G. K. Maki and J. H. Tracey.

A state assignment procedure for asynchronous sequential circuits. IEEE Transactions on Computers, C-20(6):666–668, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671913.

Marinkovic:1974:AMP

[MT74a] S. B. Marinkovic and Z. Tosic.

Algorithm for minimal polarized polynomial form determination. IEEE Transactions on Computers, C-23(12):1313–1315, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672447.

Mukai:1974:MRF

[MT74b] Y. Mukai and Y. Tohma. A method for the realization of failsafe asynchronous sequential circuits. IEEE Transactions on Computers, C-23(7):736–739, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672615.

Muroga:1970:ETF

[MTB70] S. Muroga, T. Tsuboi, and C. R.

Baugh. Enumeration of threshold functions of eight variables. IEEE Transactions on Computers, C-19(9):818–825, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671639.

Muir:1974:BDB

[Mui74] D. Muir. B74-37 data base management-international computer state of the art report 15. IEEE Transactions on Computers, C-23(8):878–879, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672638.

Mukhopadhyay:1979:HAN

[Muk79] A. Mukhopadhyay. Hardware algorithms for nonnumeric computation. IEEE Transactions on Computers, C-28(6):384–394, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675378.

Murry:1970:GAG

[Mur70] Herschell F. Murry. A general approach for generating natural random variables. IEEE Transactions on Computers, C-19 (12):1210–1213, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671453.

Muth:1976:NVC

[Mut76] P. Muth. A nine-valued circuit model for test generation.

IEEE Transactions on Computers, C-25(6):630–636, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674663.

Morris:1979:PRS

[MVW79] S. B. Morris, A. Valliere, and R. A. Wisniewski. Processes for random and sequential accessing in dynamic memories. IEEE Transactions on Computers, C-28 (3):225–237, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675323.

Mykkeltveit:1975:GCD

[Myk75] J. Mykkeltveit. Generating and counting the double adjacencies in a pure circulating shift register. IEEE Transactions on Computers, C-24(3):299–304, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672803.

Newborn:1972:UMB

[NA72] M. M. Newborn and T. F. Arnold.

Universal modules for bounded signal fan-out synchronous sequential circuits. IEEE Transactions on Computers, C-21(1):63–79, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672025.

Ni:1974:TDD

[NA74] Ming-Duenn Ni and J. K. Aggarwal. Two-dimensional digital filtering and its error analysis. IEEE Transactions on Computers, C-23(9):942–954, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672651.

Ng:1976:CFF

[NA76a] Y. W. Ng and A. Avizienis. Comments on “Fault Folding for Irredundant and Redundant Combinational Circuits”. IEEE Transactions on Computers, C-25(2):207, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009237. See [To73].

Ni:1976:EAT

[NA76b] Ming Duenn Ni and J. K. Aggarwal. Error analysis of twodimensional recursive digital filters employing floating-point arithmetic. IEEE Transactions on Computers, C-25(7):755–759, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674684.

Nadler:1971:ERR

[Nad71] M. Nadler. Error and reject rates in a hierarchical pattern recognizer. IEEE Transactions on Computers, C-20(12):1598–1601, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671773.

Nair:1979:COA

[Nai79] R. Nair. Comments on “An Optimal Algorithm for Testing Stuckat Faults in Random Access Memories”. IEEE Transactions on Computers, C-28(3):258–261, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675331. See [KH77c].

Nakamura:1979:SGM

[Nak79] K. Nakamura. Synthesis of gateminimum multi-output two-level negative gate networks. IEEE Transactions on Computers, C-28 (10):768–772, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675244.

Nambiar:1970:SHP

[Nam70] K. K. Nambiar. Solution to Harrison’s problem. IEEE Transactions on Computers, C-19(2): 166–167, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671479.

Naylor:1971:SSI

[Nay71] W. C. Naylor. Some studies in the interactive design of character recognition systems. IEEE Transactions on Computers, C-20(9):1075–1086, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671999.

Naccarato:1978:RMM

[NC78] D. Naccarato and Y. T. Chien. Recursive methods for matrix inversion in pattern recognition environments. IEEE Transactions on Computers, C-27(9):866–869, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675206.

Newman:1973:NTD

[ND73] Thomas G. Newman and Hudai Dirilten. A nonlinear transformation for digital picture processing. IEEE Transactions on Computers, C-22(9):869–873, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009186.

Necula:1979:AFE

[Nec79] M.-A. Necula. An algorithm for fast evaluation of time functions. IEEE Transactions on Computers, C-28(7):506–513, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675396.

Nelson:1974:SEC

[Nel74] T. J. Nelson. A study of the electrochemical cell as a storage element for the memory access gap. IEEE Transactions on Computers, C-23(12):1277–1290, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672440.

Nevatia:1976:LOB

[Nev76] R. Nevatia. Locating object boundaries in textured environments. IEEE Transactions on Computers, C-25(11):1170–1175, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674576.

Newborn:1970:RFD

[New70] M. Newborn. R70-19 frequency of decomposability among machines with a large number of states. IEEE Transactions on Computers, C-19(7):658, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671599.

Newman:1975:CRP

[New75a] W. M. Newman. Comments on the review of “Principles of Interactive Computer Graphics”. IEEE Transactions on Computers, C-24(3):336, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672812. See [Bib74].

Newton:1975:SPT

[New75b] T. A. Newton. Some parametric techniques in the analog solution of ordinary differential equations. IEEE Transactions on Computers, C-24(1):1–8, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672672.

Newman:1976:TGD

[New76] W. M. Newman. Trends in graphic display design. IEEE Transactions on Computers, C-25 (12):1321–1325, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674600.

Narendra:1977:BBA

[NF77] P. M. Narendra and K. Fukunaga. A branch and bound algorithm for feature subset selection. IEEE Transactions on Computers, C-26(9):917–922, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674939.

Nakamura:1977:OOE

[NFS77] Y. Nakamura, M. Furuya, and S. Sunohara. An optimal orthogonal expansion for classification of patterns. IEEE Transactions on Computers, C-26(12): 1288–1290, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674793.

Nicoud:1971:IAR

[Nic71] Jean-Daniel Nicoud. Iterative arrays for radix conversion. IEEE Transactions on Computers, C-20 (12):1479–1489, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671753.

Nieh:1970:UMS

[Nie70a] T. T. Nieh. On the uniqueness of minimal-state stochastic sequential machines. IEEE Transactions on Computers, C-19(2): 164–166, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671478.

Nievergelt:1970:TRT

[Nie70b] J. Nievergelt. On the time required for timing — the halting problem rephrased. IEEE Transactions on Computers, C-19(5):458–459, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671540.

Nieminen:1975:NMF

[Nie75] J. Nieminen. A new method of formulating a minimum edge set. IEEE Transactions on Computers, C-24(10):1015–1018, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672708.

Nahi:1977:IBE

[NJ77] N. E. Nahi and M. H. Jahanshahi. Image boundary estimation. IEEE Transactions on Computers, C-26(8):772–781, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674915.

Nagaraja:1974:ASL

[NK74] G. Nagaraja and G. Krishna. An algorithm for the solution of linear inequalities. IEEE Transactions on Computers, C-23(4):421–427, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672550.

Noural:1975:UFS

[NK75] F. Noural and R. S. Kashef. A universal four-state cellular computer. IEEE Transactions on Computers, C-24(8):766–776, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672901.

Neuman:1976:AST

[NM76a] F. Neuman and C. F. Martin. The autocorrelation structure of Tausworthe pseudorandom number generators. IEEE Transactions on Computers, C-25(5):460–464, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674633.

Neuman:1976:APC

[NM76b] F. Neuman and R. B. Merrick.

Autocorrelation peaks in congruential pseudorandom number generators. IEEE Transactions on Computers, C-25(5):457–460, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674632.

Nordmann:1977:MAC

[NM77] B. J. Nordmann, Jr. and B. H. Mc-Cormick. Modular asynchronous control design. IEEE Transactions on Computers, C-26(3): 196–207, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674808.

Noe:1973:MNR

[NN73] Jerre D. Noe and Gary J. Nutt.

Macro E-nets for representation of parallel systems. IEEE Transactions on Computers, C-22(8): 718–727, August 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009150.

Noe:1973:RSO

[Noe73] Philip S. Noe. Remarks on the SHRoptimal state assignment procedure. IEEE Transactions on Computers, C-22(9):873–875, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009187.

Noe:1974:MSO

[NR74] P. S. Noe and V. T. Rhyne. A modification to the SHR-optimal state assignment procedure. IEEE Transactions on Computers, C-23 (3):327–329, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672526.

Neumann:1975:ECC

[NR75] P. G. Neumann and T. R. N. Rao. Error-correcting codes for byte-organized arithmetic processors. IEEE Transactions on Computers, C-24(3):226–232, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672796. See comments

[Gri78b].

Nitzan:1976:PIA

[NR76a] D. Nitzan and C. A. Rosen. Programmable industrial automation.

IEEE Transactions on Computers, C-25(12):1259–1270, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674593.

Noe:1976:OSA

[NR76b] P. S. Noe and V. T. Rhyne.

Optimum state assignment for the D flip-flop. IEEE Transactions on Computers, C-25(3): 306–311, March 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009258.

Nagel:1977:CDF

[NR77a] R. N. Nagel and A. Rosenfeld. Computer detection of freehand forgeries. IEEE Transactions on Computers, C-26(9):895–905, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674937.

Ninke:1977:SRB

[NR77b] W. H. Ninke and G. R. Ritchie.

Shift register binary rate multipliers. IEEE Transactions on Computers, C-26(3):276–278, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674819.

Noe:1974:CWM

[NRS74] P. S. Noe, V. T. Rhyne, and S. Suraratrungsi. Comments on “Weighting Method for the Determination of the Irredundant Set of Prime Implicants”. IEEE Transactions on Computers, C-23(6):646, June 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672599. See [Bub72].

Niederjohn:1975:CIE

[NS75] R. J. Niederjohn and P. P. Stick. A computer interface for efficient zero-crossing interval measurement. IEEE Transactions on Computers, C-24(3):329–331, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672810.

Nassimi:1979:BSM

[NS79] D. Nassimi and S. Sahni. Bitonic sort on a mesh-connected parallel computer. IEEE Transactions on Computers, C-28(1):2–7, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675216.

Nutter:1974:ETP

[NSR74] R. S. Nutter, Jr., R. E. Swartwout, and D. C. Rine. Equivalence and transformations for post multivalued algebras. IEEE Transactions on Computers, C-23 (3):294–300, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672518.

Nordquist:1970:NRS

[NT70] W. R. Nordquist and W. N.

Toy. A novel rotate and shift circuit using bidirectional gates. IEEE Transactions on Computers, C-19(9):802–808, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671636.

Nanya:1978:UST

[NT78] T. Nanya and Y. Tohma. On universal single transition time asynchronous state assignments. IEEE Transactions on Computers, C-27(8):781–782, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675192.

Nanya:1979:UMS

[NT79] T. Nanya and Y. Tohma. Universal multicode STT state assignments for asynchronous sequential machines. IEEE Transactions on Computers, C-28(11):811–818, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675261.

Nair:1978:EAT

[NTA78] R. Nair, S. M. Thatte, and J. A. Abraham. Efficient algorithms for testing semiconductor random-access memories. IEEE Transactions on Computers, C-27(6):572–576, June 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675150.

Nakamura:1972:MNG

[NTK72] K. Nakamura, N. Tokura, and T. Kasami. Minimal negative gate networks. IEEE Transactions on Computers, C-21(1):5–11, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672018.

Nussabaumer:1978:FMN

[Nus78] H. J. Nussabaumer. Fast multipliers for number theoretic transforms. IEEE Transactions on Computers, C-27(8):764–765, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675187.

Nutt:1977:MBC

[Nut77] G. J. Nutt. Memory and bus conflict in an array processor. IEEE Transactions on Computers, C-26(6):514–521, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674877.

Niemann:1979:FCA

[NW79] H. Niemann and J. Weiss. A fastconverging algorithm for nonlinear mapping of high-dimensional data to a plane. IEEE Transactions on Computers, C-28(2): 142–147, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675303.

Oberman:1972:FRM

[Obe72] R. M. M. Oberman. A flexible rate multiplier circuit with uniform pulse distribution outputs. IEEE Transactions on Computers, C-21(8):896–899, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009048.

Oberman:1974:NER

[Obe74] R. M. M. Oberman. A new explanation of the reflected binary code. IEEE Transactions on Computers, C-23(6):641–642, June 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672597.

Oberman:1976:JKG

[Obe76] R. M. M. Oberman. The J–K gate.

IEEE Transactions on Computers, C-25(11):1156–1159, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674572.

OCallaghan:1975:ADN

[O’C75] J. F. O’Callaghan. An alternative definition for “neighborhood of a point”. IEEE Transactions on Computers, C-24(11): 1121–1125, November 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672737.

OGorman:1976:FPE

[OC76] Frank O’Gorman and M. B. Clowes. Finding picture edges through collinearity of feature points. IEEE Transactions on Computers, C-25(4):449–456, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674627.

Olsen:1973:RND

[OF73] D. R. Olsen and K. Fukunaga. Representation of nonlinear data surfaces. IEEE Transactions on Computers, C-22(10):915–922, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672211.

Ogus:1974:FTI

[Ogu74] R. C. Ogus. Fault-tolerance of the iterative cell array switch for hybrid redundancy. IEEE Transactions on Computers, C-23(7):667–681, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672608.

Ogus:1975:PCO

[Ogu75] R. C. Ogus. The probability of a correct output from a combinational circuit. IEEE Transactions on Computers, C-24(5):534–544, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672850.

Ostapko:1979:FAT

[OH79] D. L. Ostapko and Se June Hong.

Fault analysis and test generation for programmable logic arrays (PLA’s). IEEE Transactions on Computers, C-28(9): 617–627, September 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675429.

Oja:1979:CPU

[Oja79] E. Oja. On the construction of projectors using products of elementary matrices. IEEE Transactions on Computers, C-28(1):65–66, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675226.

OKeefe:1973:MDS

[O’K73] K. H. O’Keefe. Modularity in design: Skit registers and counters used as system building blocks. IEEE Transactions on Computers, C-22(2):194–197, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672277.

OKeefe:1975:NFB

[O’K75] K. H. O’Keefe. A note on fast base extension for residue number systems with three moduli. IEEE Transactions on Computers, C-24(11):1132–1133, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=1672740;   http://
                                        ieeexplore.ieee.org/xpl/tocresult.
                                        jsp?isnumber=35091.

Olson:1970:NFI

[Ols70] R. R. Olson. Note on feedforward inverses for linear sequential circuits. IEEE Transactions on Computers, C-19(12): 1216–1221, December 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671455.

Opsahl:1972:OLM

[Ops72] G. I. Opsahl. Optimum logic modules. IEEE Transactions on Computers, C-21(1):90–96, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672029.

Orcutt:1976:IPF

[Orc76] S. E. Orcutt. Implementation of permutation functions in Illiac IV-type computers. IEEE Transactions on Computers, C-25(9):929–936, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674718.

Onoe:1976:ATS

[OS76] M. Onoe and M. Saito. Automatic threshold setting for the sequential similarity detection algorithm. IEEE Transactions on Computers, C-25(10):1052–1053, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674547.

Osborne:1977:SLL

[Osb77] M. Osborne. Seniority logic: a logic for a committee machine. IEEE Transactions on Computers, C-26(12):1302–1306, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674798.

Okuda:1976:MCG

[OTK76] Teruo Okuda, Eiichi Tanaka, and Tamotsu Kasai. A method for the correction of garbled words based on the Levenshtein metric. IEEE Transactions on Computers, C-25(2):172–178, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009232.

Osman:1972:UBF

[OW72] Mohamed Y. Osman and C. Dennis Weiss. Universal base functions and modules for realizing arbitrary switching functions. IEEE Transactions on Computers, C-21(9):985–995, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009075.

OKeefe:1973:RBE

[OW73a] Kenneth H. O’Keefe and John L. Wright. Remarks on base extension for modular arithmetic. IEEE Transactions on Computers, C-22(9):833–835, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009173.

Osman:1973:SLR

[OW73b] M. Y. Osman and C. D. Weiss. Shared logic realizations of dynamically self-checked and fault-tolerant logic. IEEE Transactions on Computers, C-22(3):298–306, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672303.

Ostapko:1970:RAS

[OY70] D. L. Ostapko and S. S. Yau. Realization of an arbitrary switching function with a two-level network of threshold and parity elements. IEEE Transactions on Computers, C-19(3):262–269, March 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671500.

Ozawa:1978:CKS

[Oza78] K. Ozawa. Classification of the keyhole shaped tombs by template matching method. IEEE Transactions on Computers, C-27(5):462–467, May 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675128.

Parhami:1978:DSE

[PA78] B. Parhami and A. Avizienis. Detection of storage errors in mass memories using low-cost arithmetic error codes. IEEE Transactions on Computers, C-27(4):302–308, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675102.

Paaske:1979:CNR

[Paa79] E. Paaske. Comments on “A New Random-Error-Correction Code”.

IEEE Transactions on Computers, C-28(3):255–257, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675329. See [En77].

Page:1970:RSL

[Pag70a] C. V. Page. R70-3 on stochastic languages. IEEE Transactions on Computers, C-19(5):463–464, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671547.

Page:1970:SDP

[Pag70b] C. V. Page. The search for a definition of partition pair for stochastic automata. IEEE Transactions on Computers, C-19(12): 1222–1223, December 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671457.

Pager:1972:SNS

[Pag72] D. Pager. Some notes on speeding up certain loops by software, firmware, and hardware means. IEEE Transactions on Computers, C-21(1):97–100, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672031.

Paige:1973:SDF

[Pai73] M. R. Paige. Synthesis of diagnosable FET networks. IEEE Transactions on Computers, C-22(5):513–516, May 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672347.

Paker:1971:BFP

[Pak71] Y. Paker. A binary floating-point resistor. IEEE Transactions on Computers, C-20(1):7–11, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671667.

Papakonstantinou:1972:SMC

[Pap72] G. K. Papakonstantinou. A synthesis method for cutpoint cellular arrays. IEEE Transactions on Computers, C-21(12): 1286–1292, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672093.

Papakonstantinou:1976:CT

[Pap76] G. Papakonstantinou. Cascade transformation. IEEE Transactions on Computers, C-25(1):93–95, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009211.

Papaioannou:1977:OTG

[Pap77a] S. G. Papaioannou. Optimal test generation in combinational networks by pseudo-Boolean programming. IEEE Transactions on Computers, C-26(6):553–560, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674880.

Papakonstantinou:1977:MGP

[Pap77b] G. Papakonstantinou. A method to generate the prime cascades of an arbitrary switching function. IEEE Transactions on Computers, C-26(5):503–506, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674867.

Papachristou:1978:AON

[Pap78] C. A. Papachristou. An algorithm for optimal NAND cascade logic synthesis. IEEE Transactions on Computers, C-27(12): 1099–1111, December 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675012.

Papakonstantinou:1979:MMS

[Pap79] G. Papakonstantinou. Minimization of modulo-2 sum of products. IEEE Transactions on Computers, C-28(2):163–167, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675308.

Parks:1971:BFT

[Par71] H. L. Parks. Batch-fabricated three-dimensional planar coaxial interconnections for microelectronic systems. IEEE Transactions on Computers, C-20(5):504–511, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671877. See correction

[Ano72-39].

Parchmann:1972:NSA

[Par72a] Rainer Parchmann. The number of state assignments for sequential machines. IEEE Transactions on Computers, C-21(6):613–614, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009022.

Parhami:1972:SAP

[Par72b] Behrooz Parhami. Stochastic automata and the problems of reliability in sequential machines. IEEE Transactions on Computers, C-21(4):388–391, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008981.

Parker:1979:CMH

[Par79] D. S. Parker, Jr. Combinatorial merging and Huffman’s algorithm. IEEE Transactions on Computers, C-28(5):365–367, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675367.

Patel:1971:GRL

[Pat71] A. R. Patel. Generation of rightlinear grammars from regular expressions. IEEE Transactions on Computers, C-20(4):472–473, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671864.

Patt:1972:MST

[Pat72] Yale N. Patt. Minimum search tree structures for data partitioned into pages. IEEE Transactions on Computers, C-21(9):961–967, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009072.

Patt:1973:ONO

[Pat73] Y. N. Patt. Optimal and nearoptimal universal logic modules with interconnected external terminals. IEEE Transactions on Computers, C-22(10):903–907, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672209.

Patterson:1976:SSM

[Pat76] D. A. Patterson. Strum: Structured microprogram development system for correct firmware. IEEE Transactions on Computers, C-25 (10):974–985, October 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674538.

Pavlidis:1973:WST

[Pav73] Theodosios Pavlidis. Waveform segmentation through functional approximation. IEEE Transactions on Computers, C-22(7):689–697, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009136.

Pavlidis:1974:BPC

[Pav74] T. Pavlidis. B74-21 patern classification and scene analysis. IEEE Transactions on Computers, C-23(2):223, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672503.

Pavlidis:1975:BNM

[Pav75a] T. Pavlidis. B75-3 numerical methods. IEEE Transactions on Computers, C-24(2):222, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672787.

Pavlidis:1975:OPP

[Pav75b] T. Pavlidis. Optimal piecewise polynomial L^{2} approximation of functions of one and two variables. IEEE Transactions on Computers, C-24(1):98–102, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672681.

Pavlidis:1977:PAN

[Pav77] T. Pavlidis. Polygonal approximations by Newton’s method. IEEE Transactions on Computers, C-26(8):800–807, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674918.

Pawlak:1971:ACN

[Paw71] Z. Pawlak. Another comment on “Negative Radix Conversion”. IEEE Transactions on Computers, C-20(5):587, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671894. See comments

[Paw71, Wad71, Zoh70].

Paz:1970:RFE

[Paz70a] A. Paz. R70-33 fuzzy events realized by finite probabilistic automata. IEEE Transactions on Computers, C-19(9):858, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671654.

Paz:1970:RDR

[Paz70b] A. Paz. R70-45 on decompositions of regular events. IEEE Transactions on Computers, C-19(10):990, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671411.

Paz:1970:RES

[Paz70c] A. Paz. Regular events in stochastic sequential machines. IEEE Transactions on Computers, C-19(5):456–457, May 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671538.

Paz:1971:WDS

[Paz71] A. Paz. Whirl decomposition of stochastic systems. IEEE Transactions on Computers, C-20(10):1208–1211, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671700.

Plantz:1971:AON

[PB71] A. R. Plantz and M. Berman.

Adoption of the octal number system. IEEE Transactions on Computers, C-20(5):593–598, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671900.

Polge:1976:EFF

[PB76a] R. J. Polge and B. K. Bhagavan. Efficient Fast Fourier Transform programs for arbitrary factors with one step loop unscrambling. IEEE Transactions on Computers, C-25(5):534–539, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674643.

Prusinkiewicz:1976:DTE

[PB76b] P. Prusinkiewicz and S. Budkowski. A double track errorcorrection code for magnetic tape. IEEE Transactions on Computers, C-25(6):642–645, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674665.

Polge:1974:FCA

[PBC74] R. J. Polge, B. K. Bhagavan, and J. M. Carswell. Fast computational algorithms for bit reversal. IEEE Transactions on Computers, C-23(1):1–9, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672364.

Preston:1972:DOS

[PC72] K. Preston, Jr. and J. R. Carvalko. On determining optimum simple Golay marking transformations for binary image processing. IEEE Transactions on Computers, C-21(12):1430–1433, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672112.

Prather:1978:RBE

[PC78] R. E. Prather and H. T. Casstevens.

Realization of Boolean expressions by atomic digraphs. IEEE Transactions on Computers, C-27 (8):681–688, August 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675174.

Patrick:1970:DDE

[PCM70] E. A. Patrick, J. P. Costello, and F. C. Monds. Decisiondirected estimation of a twoclass decision boundary. IEEE Transactions on Computers, C-19 (3):197–205, March 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671490.

Pratt:1977:FCT

[PD77] W. K. Pratt and F. Davarian.

Fast computational techniques for pseudoinverse and Wiener image restoration. IEEE Transactions on Computers, C-26(6):571–580, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674882.

Plavsic:1979:SEB

[PD79] V. M. Plavsic and P. E. Danielsson. Sequential evaluation of Boolean functions. IEEE Transactions on Computers, C-28(12): 879–887, December 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675278.

Pearl:1975:ODM

[Pea75] J. Pearl. Optimal dyadic models of time-invariant systems. IEEE Transactions on Computers, C-24(6):598–603, June 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672865.

Pearl:1976:MVE

[Pea76] J. Pearl. Memory versus error characteristics for inexact representations of linear orders. IEEE Transactions on Computers, C-25(9):922–928, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674717.

Pease:1977:IBC

[Pea77] M. C. Pease III. The indirect binary n-cube microprocessor array. IEEE Transactions on Computers, C-26(5):458–473, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674863.

Pechoucek:1976:ART

[Pec76] Miroslav Pechoucek. Anomalous response times of input synchronizers. IEEE Transactions on Computers, C-25(2):133–139, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009227.

Perlman:1970:DSL

[Per70] M. Perlman. The decomposition of the states of a linear feedback shift register into cycles of equal length. IEEE Transactions on Computers, C-19(11):1029–1035, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671421.

Perkins:1978:MBV

[Per78] W. A. Perkins. A model-based vision system for industrial parts. IEEE Transactions on Computers, C-27(2):126–143, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675046.

Pezaris:1971:NBB

[Pez71a] S. D. Pezaris. A 40-ns 17bit by 17-bit array multiplier. IEEE Transactions on Computers, C-20(4):442–447, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671854.

Pezaris:1971:BBA

[Pez71b] S. D. Pezaris. A 40ns 17-bit by 17bit array multiplier. IEEE Transactions on Computers, C-20(4):442– 447 (??), April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pavlidis:1972:STW

[PF72] Theodosios Pavlidis and Geng Seng Fang. A segmentation technique for waveform classification. IEEE Transactions on Computers, C-21(8):901–904, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009050.

Pfleeger:1973:SRI

[Pfl73] C. P. Pfleeger. State reduction in incompletely specified finitestate machines. IEEE Transactions on Computers, C-22(12): 1099–1102, December 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672248.

Pearce:1975:AAM

[PFL75] R. C. Pearce, J. A. Field, and W. D. Little. Asynchronous arbiter module. IEEE Transactions on Computers, C-24(9):931–932, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672932.

Prasad:1975:MFD

[PG75] B. A. Prasad and F. G. Gray.

Multiple fault detection in arrays of combinational cells. IEEE Transactions on Computers, C-24 (8):794–802, August 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672904.

Podraza:1970:EMP

[PGS70] G. V. Podraza, R. S. Gregg, Jr., and J. R. Slager. Efficient MSI partitioning for a digital computer. IEEE Transactions on Computers, C-19(11):1020–1028, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671420.

Polge:1973:NTC

[PH73] R. J. Polge and R. D. Hays. Numerical technique for the convolution of piecewise polynomial functions. IEEE Transactions on Computers, C-22(11):970–975, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672226.

Pavlidis:1974:SPC

[PH74] T. Pavlidis and S. L. Horowitz.

Segmentation of plane curves. IEEE Transactions on Computers, C-23(8):860–870, August 1974.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672634.

Philokyprou:1977:PFT

[PHG77] G. Philokyprou, C. Halatsis, and N. Gaitanis. On the problem of finding d, t, r, and j type partition pairs of a sequential machine. IEEE Transactions on Computers, C-26(3):294–297, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674824.

Picchia:1974:NAR

[Pic74] W. Del Picchia. A numerical algorithm for the resolution of Boolean equations. IEEE Transactions on Computers, C-23(9): 983–986, September 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672660.

Pinter:1971:STF

[Pin71] C. C. Pinter. On simplifying truth functions: a preliminary reduction of coreless formulas. IEEE Transactions on Computers, C-20(8):938–941, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671969.

Plauger:1973:BHL

[Pla73] P. J. Plauger. B73-4 high level languages — international computer state of the art report. IEEE Transactions on Computers, C-22(2):221–222, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672285.

Pless:1977:ESC

[Ple77] Vera S. Pless. Encryption schemes for computer confidentiality. IEEE Transactions on Computers, C-26 (11):1133–1136, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674759.

Plummer:1972:AA

[Plu72] W. W. Plummer. Asynchronous arbiters. IEEE Transactions on Computers, C-21(1):37–42, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672022.

Preparata:1971:DRR

[PM71] F. P. Preparata and D. E. Muller.

On the delay required to realize Boolean functions. IEEE Transactions on Computers, C-20(4):459–461, April 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671859.

Patterson:1974:FSA

[PM74] W. W. Patterson and G. Metze. A fail-safe asynchronous sequential machine. IEEE Transactions on Computers, C-23(4):369–374, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672545.

Parker:1975:ALC

[PM75a] K. P. Parker and E. J. McCluskey. Analysis of logic circuits with faults using input signal probabilities. IEEE Transactions on Computers, C-24(5):573–578, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672857.

Parker:1975:PTG

[PM75b] K. P. Parker and E. J. Mc-Cluskey. Probabilistic treatment of general combinational networks. IEEE Transactions on Computers, C-24(6):668–670, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672872.

Parrish:1976:ICC

[PM76a] E. A. Parrish, Jr. and E. S. McVey. Implications of charge-coupled devices for pattern recognition. IEEE Transactions on Computers, C-25 (11):1146–1152, November 1976.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674570.

Preparata:1976:EPE

[PM76b] F. P. Preparata and D. E.

Muller. Efficient parallel evaluation of Boolean expressions. IEEE Transactions on Computers, C-25(5):548–549, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674647.

Page:1977:PAR

[PM77a] E. W. Page and P. N. Marinos. Programmable array realizations of synchronous sequential machines. IEEE Transactions on Computers, C-26(8):811–818, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674920.

Payne:1977:ACO

[PM77b] H. J. Payne and W. S. Meisel.

An algorithm for constructing optimal binary decision trees. IEEE Transactions on Computers, C-26(9):905–916, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674938.

Parker:1978:SCO

[PM78a] K. P. Parker and E. J. Mc-Cluskey. Sequential circuit output probabilities from regular expressions. IEEE Transactions on Computers, C-27(3):222–231, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675075.

Pearce:1978:ASR

[PM78b] R. C. Pearce and J. C. Majithia. Analysis of a shared resource MIMD computer organization. IEEE Transactions on Computers, C-27(1):64–67, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674953.

Preparata:1977:RDB

[PMB77] F. P. Preparata, D. E. Mulller, and A. B. Barak. Reduction of depth of Boolean networks with a fan-in constraint. IEEE Transactions on Computers, C-26(5):474–479, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674864.

Poe:1974:BAA

[Poe74] A. T. Poe. B74-11 abstract automata. IEEE Transactions on Computers, C-23(2):214–216, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672493.

Pols:1971:RTR

[Pol71] L. C. W. Pols. Real-time recognition of spoken words. IEEE Transactions on Computers, C-20(9):972–978, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671984.

Pradhan:1975:RML

[PP75] D. K. Pradhan and A. M. Patel. Reed–Muller like canonic forms for multivalued functions. IEEE Transactions on Computers, C-24(2):206–210, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672779.

Putzolu:1971:HAT

[PR71] G. R. Putzolu and J. P. Roth.

A heuristic algorithm for the testing of asynchronous circuits. IEEE Transactions on Computers, C-20(6):639–647, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671908.

Pradhan:1972:ECT

[PR72] D. K. Pradhan and S. M. Reddy.

Error-control techniques for logic processors. IEEE Transactions on Computers, C-21(12):1331– 1336, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672097.

Pradhan:1973:FTA

[PR73] Dhiraj K. Pradhan and Sudhakar M. Reddy. Fault-tolerant asynchronous networks. IEEE Transactions on Computers, C-22(7):662–669, July 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009132. See comments

[SM75a].

Paz:1974:LAA

[PR74a] A. Paz and M. Rabinovitz. Linear automata approximation problem. IEEE Transactions on Computers, C-23(3):249–255, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672512.

Pradhan:1974:DTL

[PR74b] D. K. Pradhan and S. M. Reddy. Design of two-level fault-tolerant networks. IEEE Transactions on Computers, C-23(1):41–48, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672368.

Pradhan:1975:AR

[PR75] D. K. Pradhan and S. M. Reddy. Authors’ reply. IEEE Transactions on Computers, C-24(7):758–759, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672893.

Pradhan:1976:TC

[PR76] D. K. Pradhan and S. M. Reddy.

Techniques to construct. IEEE Transactions on Computers, C-25(9):945–949, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674720.

Panda:1978:ISP

[PR78] D. P. Panda and A. Rosenfeld. Image segmentation by pixel classification in (gray level, edge value) space. IEEE Transactions on Computers, C-27(9):875–879, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675208.

Prather:1971:APP

[Pra71a] R. E. Prather. An algebraic proof of the Paull–Unger theorem. IEEE Transactions on Computers, C-20(5):578–580, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671887.

Pratt:1971:AR

[Pra71b] W. K. Pratt. Author’s reply. IEEE Transactions on Computers, C-20(5):588, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671897.

Pratt:1972:GWF

[Pra72] W. K. Pratt. Generalized Wiener filtering computation techniques. IEEE Transactions on Computers, C-21(7):636–641, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672160.

Pradhan:1974:FTC

[Pra74] D. K. Pradhan. Fault-tolerant carry-save adders. IEEE Transactions on Computers, C-23(12): 1320–1322, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672450.

Pradhan:1978:ASA

[Pra78a] D. K. Pradhan. Asynchronous state assignments with unateness properties and fault-secure design. IEEE Transactions on Computers, C-27(5):396–404, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675118.

Pradhan:1978:FTA

[Pra78b] D. K. Pradhan. Fault-tolerant asynchronous networks using read-only memories. IEEE Transactions on Computers, C-27(7):674–679, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675168.

Pradhan:1978:TGS

[Pra78c] D. K. Pradhan. A theory of Galois switching functions. IEEE Transactions on Computers, C-27 (3):239–248, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675077.

Pradhan:1978:UTS

[Pra78d] D. K. Pradhan. Universal test sets for multiple fault detection in AND–EXOR arrays. IEEE Transactions on Computers, C-27 (2):181–187, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675057.

Pardee:1971:GRT

[PRD71] S. Pardee, P. E. Rosenfeld, and P. G. Dowd. G101 — a remote time share terminal with graphic output capabilities. IEEE Transactions on Computers, C-20(8):878–881, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671957.

Preparata:1970:RND

[Pre70] F. P. Preparata. R70-28 A note on definite stochastic sequential machines. IEEE Transactions on Computers, C-19(8):764–765, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671626.

Preparata:1971:DUB

[Pre71a] F. P. Preparata. On the design of universal Boolean functions. IEEE Transactions on Computers, C-20(4):418–423, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671850.

Preston:1971:FEG

[Pre71b] K. Preston, Jr. Feature extraction by Golay hexagonal pattern transforms. IEEE Transactions on Computers, C-20(9):1007–1014, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671989.

Preparata:1972:ULM

[Pre72] Franco P. Preparata. Universal logic modules of a new type. IEEE Transactions on Computers, C-21(6):585–588, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009012.

Preparata:1978:NPS

[Pre78a] F. P. Preparata. New parallelsorting schemes. IEEE Transactions on Computers, C-27(7):669– 673, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675167.

Preston:1978:NAT

[Pre78b] F. S. Preston. A new algorithm for the tangent. IEEE Transactions on Computers, C-27(2):167, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675052.

Patrick:1971:IUP

[PS71] E. A. Patrick and L. Y. L. Shen. Interactive use of problem knowledge for clustering and decision making. IEEE Transactions on Computers, C-20(2):216–222, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671810.

Pavlidis:1978:ACA

[PS78] T. Pavlidis and K. Steiglitz. The automatic counting of asbestos fibers in air samples. IEEE Transactions on Computers, C-27 (3):258–261, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675082.

Pu:1972:CSF

[Pu72] A. T. Pu. On complete systems and finite automata. IEEE Transactions on Computers, C-21(10):1109–1113, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672050.

Pangratz:1979:PRN

[PW79a] H. Pangratz and H. Weinrichter. Pseudo-random number generator based on binary and quinary maximal-length sequences. IEEE Transactions on Computers, C-28(9):637–642, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675431.

Patil:1979:PLA

[PW79b] S. S. Patil and T. A. Welch. A programmable logic approach for VLSI. IEEE Transactions on Computers, C-28(9):594–601, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675426.

Quan:1974:NHS

[Qua74] A. Y. C. Quan. A note on high-speed generation of maximal length sequences. IEEE Transactions on Computers, C-23(2): 201–203, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672482.

Rauscher:1978:DPO

[RA78] T. G. Rauscher and A. K. Agrawala.

Dynamic problem-oriented redefinition of computer architecture via microprogramming. IEEE Transactions on Computers, C-27 (11):1006–1014, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674990.

Rader:1972:DCM

[Rad72] C. M. Rader. Discrete convolutions via Mersenne transforms. IEEE Transactions on Computers, C-21(12):1269–1273, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672090. See correction

[Rad73].

Rader:1973:CDC

[Rad73] Charles M. Rader. Correction to “Discrete Convolutions via Mersenne Transforms”. IEEE Transactions on Computers, C-22(7):703, July 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009140. See [Rad72].

Ramamoorthy:1971:FTC

[Ram71] C. V. Ramamoorthy. Faulttolerant computing: An introduction and an overview. IEEE Transactions on Computers, C-20 (11):1241–1244, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671715.

Ramamoorthy:1973:BIC

[Ram73a] C. V. Ramamoorthy. B73-9 introduction to computer organization and data structures. IEEE Transactions on Computers, C-22(5):545–546, May 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672356.

Ramanujam:1973:DPN

[Ram73b] H. R. Ramanujam. Decomposition of permutation networks. IEEE Transactions on Computers, C-22(7):639–643, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009129.

Ramapriyan:1975:GEA

[Ram75] H. K. Ramapriyan. A generalization of Eklundh’s algorithm for transposing large matrices. IEEE Transactions on Computers, C-24(12):1221–1226, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672760. See [Ekl72].

Ramamoorthy:1976:CSE

[Ram76a] C. V. Ramamoorthy. Computer science and engineering education. IEEE Transactions on Computers, C-25(12):1200–1206, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674588.

Ramapriyan:1976:MAS

[Ram76b] H. K. Ramapriyan. A multilevel approach to sequential detection of pictorial features. IEEE Transactions on Computers, C-25 (1):66–78, January 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009206.

Ramot:1976:NC

[Ram76c] J. Ramot. Nonparametric curves.

IEEE Transactions on Computers, C-25(1):103–104, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009218.

Ramapriyan:1977:ACM

[Ram77a] H. K. Ramapriyan. An algorithm for constrained maximization of the trace of a matrix. IEEE Transactions on Computers, C-26(12):1306–1308, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674799.

Ramapriyan:1977:DHG

[Ram77b] H. K. Ramapriyan. Data handling for the geometric correction of large images. IEEE Transactions on Computers, C-26 (11):1163–1167, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674769.

Rao:1970:BEC

[Rao70] T. R. N. Rao. Biresidue errorcorrecting codes for computer arithmetic. IEEE Transactions on Computers, C-19(5):398–402, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671530.

Rao:1972:ECA

[Rao72] Thammavarapu R. N. Rao. Error correction in adders using systematic subcodes. IEEE Transactions on Computers, C-21(3): 254–259, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008947.

Rao:1978:ARC

[Rao78] T. R. N. Rao. Author’s reply and comments on the arithmetic codes generated by A = (r − 1)p^{2}. IEEE Transactions on Computers, C-27(2):190, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675062.

Ratz:1976:MTT

[Rat76] H. C. Ratz. Matrix transformations for N-tuple analysis of binary patterns. IEEE Transactions on Computers, C-25(5):542–546, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674645.

Rau:1979:IMB

[Rau79a] B. R. Rau. Interleaved memory bandwidth in a model of a multiprocessor computer system. IEEE Transactions on Computers, C-28(9):678–681, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675436.

Rau:1979:PBP

[Rau79b] B. R. Rau. Program behavior and the performance of interleaved memories. IEEE Transactions on Computers, C-28(3): 191–199, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675319.

Ravi:1972:BII

[Rav72] C. V. Ravi. On the bandwidth and interference in interleaved memory systems. IEEE Transactions on Computers, C-21(8):899–901, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009049.

Ricardi:1972:RTE

[RB72] L. J. Ricardi and M. L. Burrows. A recurrence technique for expanding a function in spherical harmonics. IEEE Transactions on Computers, C-21(6):583–585, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009011.

Rao:1975:MIS

[RB75] C. V. S. Rao and N. N. Biswas. Minimization of incompletely specified sequential machines. IEEE Transactions on Computers, C-24 (11):1089–1100, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672730.

Rao:1976:FCC

[RB76] C. V. S. Rao and N. N. Biswas. Further comments on “Closure Partition Method for Minimizing Incomplete Sequential Machines”. IEEE Transactions on Computers, C-25(7):767–768, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674690. See [YB75].

Rao:1978:FCP

[RB78] C. V. Kameswara Rao and K. Balck. Finding the core point in a fingerprint. IEEE Transactions on Computers, C-27(1):77–81, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674955.

Reed:1970:CTF

[RC70] I. S. Reed and A. C. L. Chiang. Coding techniques for failuretolerant counters. IEEE Transactions on Computers, C-19(11): 1035–1038, November 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671422.

Ramamoorthy:1971:SSP

[RC71a] C. V. Ramamoorthy and Lih-Chung Chang. System segmentation for the parallel diagnosis of computers. IEEE Transactions on Computers, C-20(3):261–270, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671823.

Ramamoorthy:1971:EOL

[RC71b] C. V. Ramamoorthy and Yeh-Hao Chin. An efficient organization or large frequency-dependent files for binary searching. IEEE Transactions on Computers, C-20(10):1178–1187, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671695.

Rigas:1971:PAC

[RC71c] H. B. Rigas and D. J. Coombs. Patch: Analog computer patching from a digital simulation language. IEEE Transactions on Computers, C-20(10):1140–1146, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671691.

Ramamoorthy:1972:SMT

[RC72] C. V. Ramamoorthy and L. C. Chang. System modeling and testing procedures for microdiagnostics. IEEE Transactions on Computers, C-21(11):1169–1183, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672067.

Ryan:1974:PMS

[RC74] T. A. Ryan, Jr. and E. G. Coffman, Jr. A problem in multiprogrammed storage allocation. IEEE Transactions on Computers, C-23(11):1116–1122, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672411.

Ramamoorthy:1972:OSS

[RCG72] C. V. Ramamoorthy, K. M. Chandy, and Mario J. Gonzalez. Optimal scheduling strategies in a multiprocessor system. IEEE Transactions on Computers, C-21(2): 137–146, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008918.

Riseman:1971:CWR

[RE71] E. M. Riseman and R. W. Ehrich.

Contextual word recognition using binary digrams. IEEE Transactions on Computers, C-20(4):397–403, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671848.

Redfield:1971:SMP

[Red71a] S. R. Redfield. A study in microprogrammed processors: a medium sized microprogrammed processor. IEEE Transactions on Computers, C-20(7):743–750, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671936.

Redinpo:1971:ITW

[Red71b] G. R. Redinpo. An implementation technique for Walsh functions. IEEE Transactions on Computers, C-20(6):706–707, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671925.

Reddy:1972:DPF

[Red72a] S. M. Reddy. A design procedure for fault-locatable switching circuits. IEEE Transactions on Computers, C-21(12):1421–1426, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672110.

Reddy:1972:ETR

[Red72b] S. M. Reddy. Easily testable realizations for logic functions. IEEE Transactions on Computers, C-21(11):1183–1188, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672068.

Reddy:1973:CTS

[Red73a] S. M. Reddy. Complete test sets for logic functions. IEEE Transactions on Computers, C-22 (11):1016–1020, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672231.

Redinbo:1973:BSC

[Red73b] G. R. Redinbo. Bounds on the sequency content of the circular functions. IEEE Transactions on Computers, C-22(4):426–427, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672327.

Reddy:1974:NSC

[Red74] S. M. Reddy. Note on selfchecking checkers. IEEE Transactions on Computers, C-23(10): 1100–1102, October 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672404.

Reddy:1977:CMF

[Red77a] S. M. Reddy. Comments on “Minimal Fault Tests for Combinational Networks”. IEEE Transactions on Computers, C-26(3): 318–319, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674835. See [FD74b].

Reddy:1977:NTL

[Red77b] S. M. Reddy. A note on testing logic circuits by transition counting. IEEE Transactions on Computers, C-26(3):313–314, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674831.

Reddy:1978:CLC

[Red78a] S. M. Reddy. A class of linear codes for error control in byte-per-card organized digital systems. IEEE Transactions on Computers, C-27(5):455–459, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675126.

Reddy:1978:FTC

[Red78b] S. M. Reddy. Fault-tolerant computing: An introduction. IEEE Transactions on Computers, C-27(6):481–482, June 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675137.

Redinbo:1979:FFA

[Red79] G. R. Redinbo. Finite field arithmetic on an array processor. IEEE Transactions on Computers, C-28(7):461–471, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675390.

Reeves:1970:CTL

[Ree70] R. Reeves. Comment on “A Transform for Logic Networks”. IEEE Transactions on Computers, C-19(6):559–560, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671570.

Reddy:1976:HSU

[REFN76] D. R. Reddy, L. D. Erman, R. D. Fennell, and R. B. Neely. The Hearsay-I speech understanding system: An example of the recognition process. IEEE Transactions on Computers, C-25(4):422–431, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674624.

Regis:1973:MQM

[Reg73] Renaud C. Regis. Multiserver queueing models of multiprocessing systems. IEEE Transactions on Computers, C-22(8):736–745, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009152.

Reiliy:1974:BID

[Rei74a] K. D. Reiliy. B74-42 introduction to discrete structures. IEEE Transactions on Computers, C-23 (11):1215–1216, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672427.

Reilly:1974:BMB

[Rei74b] K. D. Reilly. B74-41 the metaphorical brain, an introduction to cybernetics as artificial intelligence and brain theory. IEEE Transactions on Computers, C-23(10): 1103–1104, October 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672405.

Reiter:1976:SGD

[Rei76] R. Reiter. A semantically guided deductive system for automatic theorem proving. IEEE Transactions on Computers, C-25(4):328–334, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674613.

Rejchrt:1972:SFG

[Rej72] Vladimir J. Rejchrt. Signal flow graph and a Fortran program for Haar–Fourier transform. IEEE Transactions on Computers, C-21(9):1026–1027, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009086.

Remy:1972:BPP

[Rem72] E. H. Remy. B72-10 PL/I programming in technological applications. IEEE Transactions on Computers, C-21(5):517, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672153.

Remler:1974:PRC

[Rem74] M. P. Remler. Pattern recognition by convolution polynomial. IEEE Transactions on Computers, C-23(5):528–530, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672568.

Reusch:1972:NMC

[Reu72] B. Reusch. Note on minimal congruences on transition graphs.

IEEE Transactions on Computers, C-21(1):96–97, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672030.

Reusch:1975:GPI

[Reu75] B. Reusch. Generation of prime implicants from subfunctions and a unifying approach to the covering problem. IEEE Transactions on Computers, C-24(9): 924–930, September 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672931.

Reynolds:1970:GRE

[Rey70] B. G. Reynolds. General repetitive events and machines. IEEE Transactions on Computers, C-19 (2):167–169, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671480.

Riseman:1972:IPP

[RF72] E. M. Riseman and C. C. Foster. The inhibition of potential parallelism by conditional jumps. IEEE Transactions on Computers, C-21(12):1405–1411, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672107.

Reid:1975:CIF

[RF75] R. J. Reid and J. S. Frame. Convergence in iteratively formed correlation matrix memories. IEEE Transactions on Computers, C-24 (8):827–830, August 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672907.

Reddi:1978:RCS

[RF78] S. S. Reddi and E. A. Feustel. A restructurable computer system. IEEE Transactions on Computers, C-27(1):1–20, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674947.

Ramamoorthy:1976:SPP

[RFL76] C. V. Ramamoorthy, T. F. Fox, and H. F. Li. Scheduling parallel processable tasks for a uniprocessor. IEEE Transactions on Computers, C-25(5):485–495, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674636. See corrections

[RL78].

Ramamoorthy:1972:SPI

[RGK72] C. V. Ramamoorthy, James R. Goodman, and K. H. Kim. Some properties of iterative squarerooting methods using high-speed multiplication. IEEE Transactions on Computers, C-21(8):

837–847, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009039.

Rowland:1971:VAD

[RH71] J. R. Rowland and W. M. Holmes.

A variational approach to digital integration. IEEE Transactions on Computers, C-20(8):894–900, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671960.

Rowland:1973:DPP

[RH73] J. R. Rowland and W. M. Holmes.

Dynamic programming for the partitioning of numerical simulations with hybrid computer applications. IEEE Transactions on Computers, C-22(5):516–523, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672348.

Riseman:1974:CPS

[RH74] E. M. Riseman and A. R. Hanson.

A contextual postprocessing system for error correction using binary n-grams. IEEE Transactions on Computers, C-23(5):480–493, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672564.

Ramamoorthy:1975:RAS

[RH75a] C. V. Ramamoorthy and Yih-Wu Han. Reliability analysis of systems with concurrent error detection. IEEE Transactions on Computers, C-24(9):868–878, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672925.

Robinson:1975:ETT

[RH75b] J. P. Robinson and C. W. Hoffner II. Easily tested three-level gate networks for T or more of N symmetric functions. IEEE Transactions on Computers, C-24 (3):331–335, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672811.

Rabbat:1976:CIS

[RHR76a] N. B. Rabbat, S. Q. A. M. A. Hossain, and W. D. Ryan. Corrections to “Improved Simulation of Pulse Propagation Between Digital Circuits”. IEEE Transactions on Computers, C-25(10):1056, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674550. See [RHR76b].

Rabbat:1976:ISP

[RHR76b] N. B. Rabbat, S. Q. A. M. A. Hossain, and W. D. Ryan. Improved simulation of pulse propagation between digital circuits.

IEEE Transactions on Computers, C-25(3):221–228, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009248. See corrections

[RHR76a].

Rhyne:1970:SBD

[Rhy70] V. T. Rhyne. Serial binaryto-decimal and decimal-to-binary conversion. IEEE Transactions on Computers, C-19(9): 808–812, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671637. See comment

[Hor71a].

Rhyne:1971:SPN

[Rhy71] V. T. Rhyne. A simple postcorrection for nonrestoring division. IEEE Transactions on Computers, C-20(2):213–214, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671808.

Ricco:1972:CSC

[Ric72] R. Ricco. Comments on “Single-Channel and Multichannel Finite-State Machines”. IEEE Transactions on Computers, C-21(11): 1247, November 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672078. See [Gil70a].

Rice:1974:BCC

[Ric74a] J. R. Rice. B74-1 complexity of computer computations. IEEE Transactions on Computers, C-23(1):109–110, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672384.

Richardson:1974:RWS

[Ric74b] R. L. Richardson. Reciprocal Walsh series. IEEE Transactions on Computers, C-23(9):964–966, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672653.

Rips:1972:BCA

[Rip72] E. M. Rips. B72-17 computer analysis of circuits. IEEE Transactions on Computers, C-21(10): 1133–1135, October 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672057.

Ristenbatt:1974:BCC

[Ris74] M. P. Ristenbatt. B74-45 computercommunication networks. IEEE Transactions on Computers, C-23 (12):1329–1330, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672459.

Rivest:1977:NFM

[Riv77] R. L. Rivest. The necessity of feedback in minimal monotone combinational circuits. IEEE Transactions on Computers, C-26(6):606–607, June 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674886.

Read:1972:AGT

[RJ72] J. S. Read and S. N. Jayaramamurthy. Automatic generation of texture feature detectors. IEEE Transactions on Computers, C-21(7):803–812, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672178.

Rosenfeld:1973:ADD

[RJ73] Azriel Rosenfeld and Emily Johnston. Angle detection on digital curves. IEEE Transactions on Computers, C-22(9): 875–878, September 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009188.

Russell:1971:SFF

[RK71] J. D. Russell and C. R. Kime.

Structural factors in the fault diagnosis of combinational networks. IEEE Transactions on Computers, C-20(11):1276–1285, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671721.

Russell:1975:SFDa

[RK75a] J. D. Russell and C. R. Kime. System fault diagnosis: Closure and diagnosability with repair. IEEE Transactions on Computers, C-24(11):1078–1089, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672729.

Russell:1975:SFDb

[RK75b] J. D. Russell and C. R. Kime. System fault diagnosis: Masking, exposure, and diagnosability without repair. IEEE Transactions on Computers, C-24(12): 1155–1161, December 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672751.

Rosenfeld:1972:CHL

[RL72] Azriel Rosenfeld and Yung H. Lee. A clustering heuristic for linedrawing analysis. IEEE Transactions on Computers, C-21(8): 904–911, August 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009051.

Ramamoorthy:1978:CSP

[RL78] C. V. Ramamoorthy and H. F.

Li. Corrections to “Scheduling Parallel Processable Tasks for a Uniprocessor”. IEEE Transactions on Computers, C-27(3):288, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675094. See [RFL76].

Ramamoorthy:1971:CDU

[RM71] C. V. Ramamoorthy and W. Mayeda.

Computer diagnosis using the blocking gate approach. IEEE Transactions on Computers, C-20 (11):1294–1299, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671723.

Reynolds:1978:FDC

[RM78] D. A. Reynolds and G. Metze.

Fault detection capabilities of alternating logic. IEEE Transactions on Computers, C-27(12): 1093–1098, December 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675011.

Rhyne:1977:NDS

[RN77] V. T. Rhyne and P. S. Noe. On the number of distinct state assignments for a sequential machine. IEEE Transactions on Computers, C-26(1):73–75, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009276.

Rhyne:1977:NTF

[RNMP77] V. T. Rhyne, P. S. Noe, M. H. Mckinney, and U. W. Pooch. A new technique for the fast minimization of switching functions. IEEE Transactions on Computers, C-26(8):757–764, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674913.

Rao:1975:IDP

[RNR75] K. R. Rao, M. A. Narasimhan, and K. Revuluri. Image data processing by Hadamard–Haar transform. IEEE Transactions on Computers, C-24(9):888–896, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672927.

Robertson:1970:CBM

[Rob70a] J. E. Robertson. The correspondence between methods of digital division and multiplier recoding procedures. IEEE Transactions on Computers, C-19(8):692–701, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671611;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Robertson.
pdf.

Robinson:1970:RPS

[Rob70b] J. P. Robinson. R70-24 A procedure for selecting diagnostic tests. IEEE Transactions on Computers, C-19(7):660, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671604.

Robertson:1979:MBO

[Rob79a] E. L. Robertson. Microcode bit optimization is NP-complete. IEEE Transactions on Computers, C-28(4):316–319, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675353.

Robinson:1979:OGR

[Rob79b] J. P. Robinson. Optimum Golomb rulers. IEEE Transactions on Computers, C-28(12):943–944, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675287.

Rochkind:1970:TFS

[Roc70] M. M. Rochkind. Transferable FORTRAN subroutine for rapid extended sorting. IEEE Transactions on Computers, C-19(3): 270–272, March 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671501.

Roesser:1978:TDM

[Roe78] R. P. Roesser. Two-dimensional microprocessor pipelines for image processing. IEEE Transactions on Computers, C-27(2): 144–156, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675047.

Rom:1978:DRB

[Rom78] R. Rom. Distribution of runs in binary words. IEEE Transactions on Computers, C-27(11): 1087–1089, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675003.

Roochvarg:1973:BTM

[Roo73] A. C. Roochvarg. B73-7 theory of machines and computations. IEEE Transactions on Computers, C-22(4):430, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672332.

Rossol:1971:GHC

[Ros71] L. Rossol. Generation of halftones by computer-controlled microfilm recorder. IEEE Transactions on Computers, C-20(6):662–664, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671911.

Rosen:1974:BCS

[Ros74a] S. Rosen. B74-10 computer semantics, studies of algorithms, processors and languages. IEEE Transactions on Computers, C-23(2):214, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672492.

Rosenfeld:1974:DSL

[Ros74b] A. Rosenfeld. Digital straight line segments. IEEE Transactions on Computers, C-23(12): 1264–1269, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672438.

Rosenfeld:1975:NAD

[Ros75] A. Rosenfeld. A note on automatic detection of texture gradients. IEEE Transactions on Computers, C-24(10):988–991, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672699.

Roth:1970:RDS

[Rot70] J. P. Roth. R70-12 diagnosis of single-gate failures in combinational circuits. IEEE Transactions on Computers, C-19(6):564–565, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671576.

Roth:1977:HV

[Rot77] J. P. Roth. Hardware verification. IEEE Transactions on Computers, C-26(12):1292–1294, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674795.

Roth:1978:PLA

[Rot78] J. P. Roth. Programmed logic array optimization. IEEE Transactions on Computers, C-27(2): 174–176, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675055.

Russo:1971:HPP

[ROW71] R. L. Russo, P. H. Oden, and P. K. Wolff, Sr. A heuristic procedure for the partitioning and mapping of computer logic graphs. IEEE Transactions on Computers, C-20(12):1455–1462, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671750.

Roy:1974:DFE

[Roy74] B. K. Roy. Diagnosis and fault equivalence in combinational circuits. IEEE Transactions on Computers, C-23(9):955–963, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672652.

Ramamoorthy:1973:CTR

[RPL73] C. V. Ramamoorthy, J. H. Park, and H. F. Li. Compilation techniques for recognition of parallel processable tasks in arithmetic expressions. IEEE Transactions on Computers, C-22(11): 986–998, November 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672228.

Robbins:1974:TLE

[RR74] W. E. Robbins and L. D. Rudolph.

On two-level exclusive-or majority networks. IEEE Transactions on Computers, C-23(1):34–41, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672367.

Revuluri:1974:CDS

[RRNA74] K. Revuluri, K. R. Rao, M. A.

Narasimhan, and N. Ahmed. Cyclic and dyadic shifts, Walsh–Hadamard transform, and the H diagram. IEEE Transactions on Computers, C-23(12):1303–1306, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672444.

Reischer:1971:AAS

[RS71a] C. Reischer and D. A. Simovici. Associative algebraic structures in the set of Boolean functions and some applications in automata theory. IEEE Transactions on Computers, C-20(3):298–303, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671827.

Rink:1971:ADC

[RS71b] R. A. Rink and W. Streifer. Application of digital computers to solve analytically a class of second-order non-linear ordinary differential equations. IEEE Transactions on Computers, C-20(8):901–910, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671961.

Reischer:1972:EPA

[RS72a] C. Reischer and D. A. Simovici. On the existence of a periodic analog of a finite connected automaton. IEEE Transactions on Computers, C-21(2):208–211, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008929.

Roy:1972:DMD

[RS72b] P. K. Sinha Roy and C. L. Sheng. A decomposition method of determining maximum compatibles. IEEE Transactions on Computers, C-21(3):309–312, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008956. See comments

[RS73b].

Roy:1973:AR

[RS73a] P. K. Sinha Roy and C. L. Sheng.

Authors’ reply. IEEE Transactions on Computers, C-22(6):627, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009119.

Roy:1973:ARC

[RS73b] P. K. Sinha Roy and C. L.

Sheng. Author’s reply to comments on “Decomposition Method of Determining Maximum Compatibles”. IEEE Transactions on Computers, C-22(6):627, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith2/papers/ARITH2_
Sureshchander.pdf. See [RS72b,

Sur73].

Ramamoorthy:1974:ATC

[RS74] C. V. Ramamoorthy and K. S.

Shankar. Automatic testing for the correctness and equivalence of loopfree microprograms. IEEE Transactions on Computers, C-23 (8):768–782, August 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672626.

Robach:1975:DTM

[RS75] C. Robach and G. Saucier. Diversified test methods for local control units. IEEE Transactions on Computers, C-24(5):562–567, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672854.

Robach:1978:DTC

[RS78] C. Robach and G. Saucier. Dynamic testing of control units. IEEE Transactions on Computers, C-27(7):617–623, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675161.

Rao:1979:ATD

[RSH79] G. S. Rao, H. S. Stone, and T. C. Hu. Assignment of tasks in a distributed processor system with limited memory. IEEE Transactions on Computers, C-28(4):291–299, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675348.

Robach:1976:PTD

[RSL76] C. Robach, G. Saucier, and J. Lebrun. Processor testability and design consequences. IEEE Transactions on Computers, C-25(6):645–652, June 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674666.

Rao:1970:BLR

[RT70] T. R. N. Rao and A. K. Trehan. Binary logic for residue arithmetic using magnitude index. IEEE Transactions on Computers, C-19(8):752–757, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671619;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Rao.pdf.

Rosenfeld:1971:ECD

[RT71] A. Rosenfeld and M. Thurston.

Edge and curve detection for visual scene analysis. IEEE Transactions on Computers, C-20(5):562–569, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671883.

Robertson:1973:SIC

[RT73a] James E. Robertson and Kishor S.

Trivedi. The status of investigations into computer hardware design based on the use of continued fractions. IEEE Transactions on Computers, C-22(6):555–560, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009106;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Robertson.
pdf.

Roome:1973:AMS

[RT73b] W. D. Roome and H. C. Toorng. Algorithms for multiple shift register realizations of sequential machines. IEEE Transactions on Computers, C-22(10):933–943, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672214.

Ramamoorthy:1974:HLL

[RT74] C. V. Ramamoorthy and M. Tsuchiya. A high-level language for horizontal microprogramming. IEEE Transactions on Computers, C-23 (8):791–801, August 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672628.

Reed:1979:NHA

[RT79] I. S. Reed and T. K. Truong. A new hybrid algorithm for computing a fast discrete Fourier transform. IEEE Transactions on Computers, C-28(7):487–492, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675393.

Reed:1977:IPT

[RTKH77] I. S. Reed, T. K. Truong, Y. S. Kwoh, and E. L. Hall. Image processing by transforms over a finite field. IEEE Transactions on Computers, C-26(9):874–881, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674935.

Rosenfeld:1972:ECD

[RTL72] A. Rosenfeld, M. Thurston, and Yung-Han Lee. Edge and curve detection: Further experiments. IEEE Transactions on Computers, C-21(7):677–715, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672166.

Ramamoorthy:1978:DFC

[RTW78] C. V. Ramamoorthy, J. L. Turner, and B. W. Wah. A design of a fast cellular associative memory for ordered retrieval. IEEE Transactions on Computers, C-27(9):800–815, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675200.

Rubin:1974:LPC

[Rub74] F. Rubin. The Lee path connection algorithm. IEEE Transactions on Computers, C-23(9): 907–914, September 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672647. See correction

[Rub76a].

Rubin:1975:IAT

[Rub75a] F. Rubin. An improved algorithm for testing the planarity of a graph. IEEE Transactions on Computers, C-24(2):113–121, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672772.

Rubin:1975:WCB

[Rub75b] F. Rubin. Worst case bounds for maximal compatible subsets. IEEE Transactions on Computers, C-24(8):830–831, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672908.

Rubinfield:1975:PMB

[Rub75c] L. P. Rubinfield. A proof of the modified Booth’s algorithm for multiplication. IEEE Transactions on Computers, C-24(10): 1014–1015, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672707.

Rubin:1976:CLP

[Rub76a] Frank Rubin. Correction to “The Lee Path Connection Algorithm”. IEEE Transactions on Computers, C-25(2):208, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009239. See [Rub74].

Rubin:1976:GNC

[Rub76b] Frank Rubin. Generation of nonparametric curves. IEEE Transactions on Computers, C-25(1):103, January 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009216.

Rubin:1979:DSC

[Rub79] F. Rubin. Decrypting a stream cipher based on J–K flip-flops. IEEE Transactions on Computers, C-28(7):483–487, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675392.

Rudeanu:1974:AAB

[Rud74] S. Rudeanu. An algebraic approach to Boolean equations. IEEE Transactions on Computers, C-23 (2):206–207, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672484.

Rudeanu:1976:SRF

[Rud76] S. Rudeanu. Square roots and functional decompositions of Boolean functions. IEEE Transactions on Computers, C-25(5):528–532, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674641.

Rumbaugh:1977:DFM

[Rum77] James Rumbaugh. A data flow multiprocessor. IEEE Transactions on Computers, C-26(2): 138–146, February 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009292.

Russo:1972:TBL

[Rus72] Roy L. Russo. On the tradeoff between logic performance and circuit-to-pin ratio for LSI. IEEE Transactions on Computers, C-21(2):147–153, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008919.

Rux:1971:BMP

[Rux71] P. T. Rux. B71-3 microprogramming: Principles and practices. IEEE Transactions on Computers, C-20(10):1221–1222, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671705.

Rosenfeld:1973:MAM

[RV73] J. L. Rosenfeld and R. D. Villani. Micromultiprocessing: An approach to multiprocessing at the level of very small tasks. IEEE Transactions on Computers, C-22(2):149–153, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672269.

Rey:1974:SSA

[RV74] C. A. Rey and J. Vaucher. Selfsynchronized asynchronous sequential machines. IEEE Transactions on Computers, C-23(12): 1306–1311, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672445.

Reddy:1974:ETC

[RW74] S. M. Reddy and J. R. Wilson.

Easily testable cellular realizations for the (exactly P)-out-of-n and (p or more)-out-of-n logic functions. IEEE Transactions on Computers, C-23(1):98–100, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672380.

Rosenfeld:1975:IMA

[RW75] A. Rosenfeld and J. S. Weszka.

An improved method of angle detection on digital curves. IEEE Transactions on Computers, C-24(9):940–941, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672935.

Santos:1970:GMS

[SA70] J. Santos and H. Arango. A graphic method for the synthesis of threshold ternary functions. IEEE Transactions on Computers, C-19(10):975–976, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671398.

Satake:1974:FTN

[SA74] S. Satake and T. Aibara. A formula for ternary network cost. IEEE Transactions on Computers, C-23(3):329–332, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672527.

Swartzlander:1975:SLN

[SA75] E. E. Swartzlander, Jr. and A. G. Alexopoulos. The sign/logarithm number system. IEEE Transactions on Computers, C-24(12): 1238–1242, December 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672765. See [HL86].

Schiller:1971:MIG

[SAFV71] W. L. Schiller, R. L. Abraham, R. M. Fox, and A. Van Dam. A microprogrammed intelligent graphics terminal. IEEE Transactions on Computers, C-20(7):775–782, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671939.

Sagalovich:1977:CNC

[Sag77] Y. L. Sagalovich. Comments on “On the Number of Classes of Binary Matrices”. IEEE Transactions on Computers, C-26(5):512, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674871. See [Har73].

Sahin:1974:MBR

[Sah74] K. E. Sahin. Message-based response routing with Selcuk networks. IEEE Transactions on Computers, C-23(12):1250– 1257, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672436.

Sahin:1977:SSC

[Sah77] K. E. Sahin. Segmented Selcuk communication networks and response routing. IEEE Transactions on Computers, C-26(8):831–833, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674924.

Sammon:1970:IPA

[Sam70a] J. W. Sammon, Jr. Interactive pattern analysis and classification. IEEE Transactions on Computers, C-19(7):594–616, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671586.

Sammon:1970:ODP

[Sam70b] J. W. Sammon, Jr. An optimal discriminant plane. IEEE Transactions on Computers, C-19(9):826–829, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671640.

Sandewall:1976:CPC

[San76] E. Sandewall. Conversion of predicate-calculus axioms to corresponding deterministic programs. IEEE Transactions on Computers, C-25(4):342–346, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674615.

Santos:1970:CAS

[SAPR70] J. Santos, H. Arango, M. Pascual, and G. Roing. A cyclic algebra for the synthesis of ternary digital systems. IEEE Transactions on Computers, C-19(7):651–653, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671592.

Sarma:1971:ADC

[Sar71] M. S. Sarma. Application of digital computers to three-dimensional field determination. IEEE Transactions on Computers, C-20(8): 914–917, August 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671963.

Sarwate:1978:SFF

[Sar78] D. V. Sarwate. Semi-fast Fourier transforms over GF(2^{m}). IEEE Transactions on Computers, C-27 (3):283–285, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675088.

Saucier:1972:NSE

[Sau72a] Gabriele Saucier. Next-state equations of asynchronous sequential machines. IEEE Transactions on Computers, C-21(4):397–399, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008984. See comments

[MS73a].

Saucier:1972:SAA

[Sau72b] Gabriele Saucier. State assignment of asynchronous sequential machines using graph techniques. IEEE Transactions on Computers, C-21(3):282–288, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008950.

Sawin:1974:OAS

[Saw74] D. H. Sawin III. Optimization of asynchronous sequential circuit realizations. IEEE Transactions on Computers, C-23(2): 186–188, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672477.

Sawin:1975:DRS

[Saw75] D. H. Sawin III. Design of reliable synchronous sequential circuits. IEEE Transactions on Computers, C-24(5):567–570, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672855.

Sawchuk:1977:RTC

[Saw77] Alexander A. Sawchuk. Realtime correction of intensity nonlinearities in imaging systems. IEEE Transactions on Computers, C-26(1):34–39, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009271.

Schoen:1971:BEI

[SB71] T. A. Schoen and M. R. Belsole, Jr. A Burroughs 220 emulator for the IBM 360/25. IEEE Transactions on Computers, C-20(7):795–798, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671941.

Sidhu:1972:PEA

[SB72] G. S. Sidhu and R. T. Boute.

Property encoding: Application in binary picture encoding and boundary following. IEEE Transactions on Computers, C-21(11): 1206–1216, November 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672071.

Sureshchander:1974:NRS

[SB74a] Sureshchander and P. C. P. Bhatt.

A new representation of sequential machines for determining all incompatible pairs. IEEE Transactions on Computers, C-23(3): 325–326, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672523.

Sureshchander:1974:NUL

[SB74b] Sureshchander and P. C. P. Bhatt.

A note on Q-universal logic modules. IEEE Transactions on Computers, C-23(3):319–320, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672521.

Sarje:1977:ATA

[SB77a] A. K. Sarje and N. N. Biswas. An algorithm for testing 2-asummability of Boolean functions. IEEE Transactions on Computers, C-26(10):1049–1053, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674745.

Snir:1977:DAP

[SB77b] M. Snir and A. B. Barak. A direct approach to the parallel evaluation of rational expressions with a small number of processors. IEEE Transactions on Computers, C-26(10):933–937, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674728.

Sarje:1979:NAA

[SB79] A. K. Sarje and N. N. Biswas. A new approach to 2-asummability testing. IEEE Transactions on Computers, C-28(10):798–801, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675250.

Siy:1972:MFF

[SC72a] P. Siy and C. S. Chen. Minimization of fuzzy functions. IEEE Transactions on Computers, C-21 (1):100–102, January 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672032. See comments

[Kan73a].

Su:1972:NAF

[SC72b] S. Y. H. Su and Yun-Chung Cho. A new approach to the fault location of combinational circuits. IEEE Transactions on Computers, C-21(1):21–30, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672020.

Su:1972:CMM

[SC72c] Stephen Y. H. Su and Peter T.

Cheung. Computer minimization of multivalued switching functions. IEEE Transactions on Computers, C-21(9):995–1003, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009076.

Sung:1974:TAC

[SC74] Chia-Hsiaing Sung and C. L.

Coates, Jr. Tessellation aspect of combinational cellular array testing. IEEE Transactions on Computers, C-23(4):363–369, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672544.

Shliferstein:1977:SPI

[SC77] A. R. Shliferstein and Y. T.

Chien. Some properties of imageprocessing operations on projection sets obtained from digital pictures. IEEE Transactions on Computers, C-26(10):958–970, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674731.

Sklansky:1972:MPP

[SCH72] Jack Sklansky, Robert L. Chazin, and Bruce J. Hansen. Minimumperimeter polygons of digitized silhouettes. IEEE Transactions on Computers, C-21(3): 260–268, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008948.

Schumann:1973:CFC

[Sch73] U. Schumann. Comments on “A Fast Computer Method for Matrix Transposing” and application to the solution of Poisson’s equation. IEEE Transactions on Computers, C-22(5):542–543, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672353. See [Ekl72].

Schertz:1974:FTC

[Sch74a] D. R. Schertz. Fault-tolerant computing: An introduction. IEEE Transactions on Computers, C-23(7):649–650, July 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672605.

Schoeffler:1974:BMD

[Sch74b] J. D. Schoeffler. B74-4 minicomputers in data processing and simulation. IEEE Transactions on Computers, C-23(1):111–112, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672387.

Schkolnick:1975:CPP

[Sch75] M. Schkolnick. On a covering problem for partially specified switching functions. IEEE Transactions on Computers, C-24 (11):1139–1141, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672742.

Schachter:1978:DPC

[Sch78a] B. Schachter. Decomposition of polygons into convex sets. IEEE Transactions on Computers, C-27 (11):1078–1082, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675001.

Schaffner:1978:PDP

[Sch78b] M. R. Schaffner. Processing by data and program blocks. IEEE Transactions on Computers, C-27 (11):1015–1028, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674991.

Schurmann:1978:MWR

[Sch78c] J. Schurmann. A multifont word recognition system for postal address reading. IEEE Transactions on Computers, C-27(8): 721–732, August 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675179.

Schneck:1979:CWU

[Sch79a] P. B. Schneck. Comment on “When to Use Random Testing”. IEEE Transactions on Computers, C-28(8):580–581, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675415. See [Agr78b].

Schwartz:1979:TTD

[Sch79b] M. Schwartz. Throughput and time delay analysis for a common queue configuration in a multiprocessor enviromnent. IEEE Transactions on Computers, C-28(12):939–941, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675285.

Sankar:1973:AAN

[SCK73a] P. V. Sankar, S. Chakrabarti, and E. V. Krishnamurthy. Arithmetic algorithms in a negative base. IEEE Transactions on Computers, C-22(2):120–125, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672264.

Sankar:1973:DDA

[SCK73b] P. V. Sankar, S. Chakrabarti, and E. V. Krishnamurthy. Deterministic division algorithm in a negative base. IEEE Transactions on Computers, C-22(2):125–128, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672265.

Slagle:1970:NAG

[SCL70] J. R. Slagle, Chin-Liang Chang, and R. C. T. Lee. A new algorithm for generating prime implicants. IEEE Transactions on Computers, C-19(4):304–310, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671510. See comments

[Das71c].

Sklansky:1976:PDC

[SCL76] Jack Sklansky, Luigi P. Cordella, and Stefano Levialdi. Parallel detection of concavities in cellular blobs. IEEE Transactions on Computers, C-25(2):187–196, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009234.

Swartzlander:1983:SLA

[SCNS83] E. E. Swartzlander, Jr., D. V. Satish Chandra, H. T. Nagle, Jr., and S. A. Starks. Sign/logarithm arithmetic for FFT implementation. IEEE Transactions on Computers, C-32(6):526–534, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1676274. See comments

[HL86].

Schmidt:1974:ECT

[SD74] D. C. Schmidt and L. E. Druffel. An extension of the clause table approach to multi-output combinational switching networks. IEEE Transactions on Computers, C-23(4):338–346, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672540.

Sethi:1979:IMS

[SD79] A. S. Sethi and N. Deo. Interference in multiprocessor systems with localized memory access probabilities. IEEE Transactions on Computers, C-28(2):157–163, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675307.

Selkow:1974:DKR

[Sel74] S. M. Selkow. Diagnostic keys as a representation for context in pattern recognition. IEEE Transactions on Computers, C-23(9):970–971, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672656.

Servit:1973:HCA

[Ser73] M. Servit. Hazard correction in asynchronous sequential circuits using inertial delay elements. IEEE Transactions on Computers, C-22(11):1041–1042, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672235.

Servit:1975:HCS

[Ser75] M. Servit. Hazard correction in synchronous sequential circuits. IEEE Transactions on Computers, C-24(3):305–310, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672804.

Sethares:1970:CPM

[Set70a] G. C. Sethares. Completely periodic multithreshold functions. IEEE Transactions on Computers, C-19(4):355–358, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671516.

Setlur:1970:MDE

[Set70b] R. V. Setlur. A method to determine the expressive power of a set of connectives. IEEE Transactions on Computers, C-19 (12):1223–1225, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671458.

Sanathanan:1971:HCT

[SF71] C. K. Sanathanan and J. D. Ferguson. Hybrid computation techniques inferred from functional analysis. IEEE Transactions on Computers, C-20(1):19–24, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671669.

Sengbush:1972:DAO

[SF72] R. L. Sengbush and M. R. Foster. Design and application of optimal velocity filters in seismic exploration. IEEE Transactions on Computers, C-21(7):648–654, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672162.

Sadeh:1974:MCS

[SF74] E. Sadeh and M. A. Franklin. Monte Carlo solution of partial differential equations by special purpose digital computer. IEEE Transactions on Computers, C-23(4):389–397, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672547.

Shapira:1977:COP

[SF77] R. Shapira and H. Freeman. A cyclic-order property of bodies with three-face vertices. IEEE Transactions on Computers, C-26(10):1035–1039, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674742.

Shapira:1978:CDB

[SF78] R. Shapira and H. Freeman. Computer description of bodies bounded by quadric surfaces from a set of imperfect projections. IEEE Transactions on Computers, C-27(9):841–854, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675204.

Short:1971:SPD

[SG71] R. A. Short and J. Goldberg. Soviet progress in the design of faulttolerant digital machines. IEEE Transactions on Computers, C-20 (11):1337–1352, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671729.

Salama:1974:CAS

[SG74] A. I. A. Salama and V. Gourishankar. A computational algorithm for solving a system of coupled algebraic matrix Riccati equations. IEEE Transactions on Computers, C-23(1):100–102, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672381.

Swartzlander:1978:IPC

[SGR78] E. E. Swartzlander, Jr., B. K. Gilbert, and I. S. Reed. Inner product computers. IEEE Transactions on Computers, C-27(1):21–31, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674948. See comments

[Bla79].

Sutton:1972:TMA

[SH72] R. N. Sutton and E. L. Hall. Texture measures for automatic classification of pulmonary disease. IEEE Transactions on Computers, C-21(7):667–676, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672165.

Saxton:1977:OFS

[SH77] Terrence L. Saxton and Cheng-Chi Huang. An optimal frame synchronization technique using an associative processor. IEEE Transactions on Computers, C-26(2):

170–174, February 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009296.

Swartzlander:1979:RAS

[SH79] E. E. Swartzlander, Jr. and D. J. Heath. A routing algorithm for signal processing networks. IEEE Transactions on Computers, C-28(8):567–572, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675412.

Shanmugam:1975:CDC

[Sha75] K. S. Shanmugam. Comments on “Discrete Cosine Transform”. IEEE Transactions on Computers, C-24(7):759, July 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672894. See [ANR74].

Shapiro:1978:TLE

[Sha78a] H. D. Shapiro. Theoretical limitations on the efficient use of parallel memories. IEEE Transactions on Computers, C-27(5):421–428, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675122.

Shapiro:1978:TMC

[Sha78b] S. D. Shapiro. Transform method of curve detection for textured image data. IEEE Transactions on Computers, C-27(3):254–255, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675080.

Shelly:1970:REA

[She70a] J. H. Shelly. R70-37 extensions of asynchronous circuits and the delay problem I. good extensions and the delay problem of the first kind. IEEE Transactions on Computers, C-19(9):859, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671658.

Sheng:1970:RRS

[She70b] C. L. Sheng. R70-15 realization of sequential machines with threshold elements. IEEE Transactions on Computers, C-19(6):565–566, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671579.

Shepherdson:1970:RSB

[She70c] J. C. Shepherdson. R70-20 sequential Boolean equations. IEEE Transactions on Computers, C-19(7):658, July 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671600.

Sherlock:1970:SMS

[She70d] J. F. Sherlock. The simulation of a multicomputer system. IEEE Transactions on Computers, C-19(11):1114–1118, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671437.

Shepherd:1971:RSL

[She71] B. J. Shepherd. Right shift for low-cost multiply and divide. IEEE Transactions on Computers, C-20(12):1586–1589, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671770.

Shedletsky:1977:CSI

[She77] J. J. Shedletsky. Comment on the sequential and indeterminate behavior of an end-around-carry adder. IEEE Transactions on Computers, C-26(3):271–272, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674817.

Shedletsky:1978:ECA

[She78a] J. J. Shedletsky. Error correction by alternate-data retry. IEEE Transactions on Computers, C-27(2):106–112, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675044.

Shedletsky:1978:RIN

[She78b] J. J. Shedletsky. A rollback interval for networks with an imperfect self-checking property. IEEE Transactions on Computers, C-27(6):500–508, June 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675140.

Shively:1970:PFI

[Shi70] Richard R. Shively. Preface: First IEEE Conference on Computer Arithmetic, Minneapolis, Minnesota, June 16, 1969. IEEE Transactions on Computers, C-19 (8):679–680, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671609;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_contents.
pdf;   http://www.acsel-lab.
com/arithmetic/arith1/papers/
ARITH1_Shively.pdf.

Shively:1975:BRE

[Shi75] R. R. Shively. Book reviews Editor’s comments. IEEE Transactions on Computers, C-24(3):336, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672814.

Sholl:1974:DTM

[Sho74] H. A. Sholl. Direct transition memory and its application in computer design. IEEE Transactions on Computers, C-23(10): 1048–1061, October 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672398.

Shriver:1971:MNA

[Shr71] B. D. Shriver, Sr. Microprogramming and numerical analysis. IEEE Transactions on Computers, C-20(7):808–812, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671944.

Story:1972:OSA

[SHR72] J. R. Story, H. J. Harrison, and E. A. Reinhard. Optimum state assignment for synchronous sequential circuits. IEEE Transactions on Computers, C-21(12): 1365–1373, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672101.

Si:1978:DTR

[Si78]  Siu-Chong Si.  Dynamic test-

ing of redundant logic networks. IEEE Transactions on Computers, C-27(9):828–832, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675202.

Sickel:1976:STC

[Sic76] S. Sickel. A search technique for clause interconnectivity graphs. IEEE Transactions on Computers, C-25(8):823–835, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674701.

Siewiorek:1974:BDS

[Sie74] D. Siewiorek. B74-40 digital systems: Hardware organization and design. IEEE Transactions on Computers, C-23(9): 991–992, September 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672667.

Siewiorek:1975:RMC

[Sie75] D. P. Siewiorek. Reliability modeling of compensating module failures in majority voted redundancy. IEEE Transactions on Computers, C-24(5):525–533, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672849.

Siegel:1977:ATS

[Sie77] Howard Jay Siegel. Analysis techniques for SIMD machine interconnection networks and the effects of processor address masks.

IEEE Transactions on Computers, C-26(2):153–161, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009294.

Siegel:1979:MSM

[Sie79] H. J. Siegel. A model of SIMD machines and a comparison of various interconnection networks. IEEE Transactions on Computers, C-28(12):907–917, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675280.

Sugino:1976:NLS

[SIF76] K. Sugino, Y. Inagaki, and T. Fukumura. A note on the linear space automata stability problem. IEEE Transactions on Computers, C-25(7):678–683, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674677.

Sikdar:1970:DMM

[Sik70] K. Sikdar. Determination of multipliers mapping an arbitrary integer into a range of certain type. IEEE Transactions on Computers, C-19(12):1221–1222, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671456.

Siler:1973:BIC

[Sil73] K. F. Siler. B73-1 an introduction to Cobol programming. IEEE Transactions on Computers, C-22(1):112, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672200.

Silverston:1977:STM

[Sil77] S. M. Silverston. A stable, two-multiplies-per-cycle algorithm for digital generation of sinusoids in real time. IEEE Transactions on Computers, C-26(11): 1153–1154, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674765. See comments

[HO79].

Silio:1979:ESC

[Sil79] C. B. Silio, Jr. An efficient simplex coverability algorithm in E2 with application to stochastic sequential machines. IEEE Transactions on Computers, C-28(2): 109–120, February 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675300.

Singh:1971:DIA

[Sin71] S. Singh. On delayed-input asynchronous sequential circuits. IEEE Transactions on Computers, C-20(5):500–503, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671876.

Sintonen:1972:RFV

[Sin72] Leo Sintonen. On the realization of functions in N-valued logic. IEEE Transactions on Computers, C-21(6):610–612, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009020.

Sintonen:1977:CMF

[Sin77] L. Sintonen. A clocked multivalued flip-flop. IEEE Transactions on Computers, C-26(3): 292–294, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674823.

Sites:1974:SBD

[Sit74] R. L. Sites. Serial binary division by ten. IEEE Transactions on Computers, C-23(12): 1299–1301, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672442.

Sarkar:1971:EPP

[SK71] B. P. Sarkar and E. V. Krishnamurthy. Economic pseudodivision processes for obtaining square root, logarithm, and arctan.

IEEE Transactions on Computers, C-20(12):1589–1593, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671771.

Sen:1974:RAA

[SK74] S. K. Sen and E. V. Krishnamurthy. Rank-augmented LU-algorithm for computing generalized matrix inverses. IEEE Transactions on Computers, C-23(2):199–201, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672481.

Sastry:1975:PCM

[SK75] K. V. Sastry and R. Y. Kain. On the performance of certain multiprocessor computer organizations. IEEE Transactions on Computers, C-24(11):1066–1074, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672727.

Sameh:1977:CPA

[SK77a] A. H. Sameh and D. J. Kuck. Correction to “A Parallel QR Algorithm for Symmetric Tridiagonal Matrices”. IEEE Transactions on Computers, C-26(8):836, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674926. See [SK77b].

Sameh:1977:PAS

[SK77b] Ahmed H. Sameh and David J.

Kuck. A parallel QR algorithm for symmetric tridiagonal matrices. IEEE Transactions on Computers, C-26(2):147–153, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009293. See correction

[SK77a].

Seth:1977:DFL

[SK77c] Sharad C. Seth and K. L. Kodandapani. Diagnosis of faults in linear tree networks. IEEE Transactions on Computers, C-26(1):29–33, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009270.

Sasao:1978:CRI

[SK78a] T. Sasao and K. Kinoshita. Cascade realization of 3-input 3output conservative logic circuits. IEEE Transactions on Computers, C-27(3):214–221, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675074.

Sasao:1978:RMC

[SK78b] T. Sasao and K. Kinoshita. Realization of minimum circuits with two-input conservative logic elements. IEEE Transactions on Computers, C-27(8):749–752, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675183.

Sasao:1979:CLE

[SK79a] T. Sasao and K. Kinoshita. Conservative logic elements and their universality. IEEE Transactions on Computers, C-28(9): 682–685, September 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675437.

Sasao:1979:NFF

[SK79b] T. Sasao and K. Kinoshita. On the number of fanout-free functions and unate cascade functions. IEEE Transactions on Computers, C-28(1):66–72, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675227.

Stenzel:1977:CHS

[SKG77] W. J. Stenzel, W. J. Kubitz, and G. H. Garcia. A compact high-speed parallel multiplication scheme. IEEE Transactions on Computers, C-26(10):948–957, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674730.

Suenaga:1979:HSA

[SKK79] Y. Suenaga, T. Kamae, and T. Kobayashi. A high-speed algorithm for the generation of straight lines and circular arcs. IEEE Transactions on Computers, C-28(10):728–736, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675240.

Sklansky:1970:RUC

[Skl70] J. Sklansky. R70-16 unate cellular logic. IEEE Transactions on Computers, C-19(6):566, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671580.

Sklansky:1972:MCR

[Skl72] J. Sklansky. Measuring concavity on a rectangular mosaic. IEEE Transactions on Computers, C-21(12):1355–1364, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672100.

Sklansky:1978:HTC

[Skl78] J. Sklansky. On the Hough technique for curve detection. IEEE Transactions on Computers, C-27(10):923–926, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674971.

Su:1978:CPM

[SKM78] S. Y. H. Su, I. Koren, and Y. K. Malaiya. A continuousparameter Markov model and detection procedures for intermittent faults. IEEE Transactions on Computers, C-27(6):567–570, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675148.

Schnurmann:1975:WRT

[SLC75] H. D. Schnurmann, E. Lindbloom, and R. G. Carpenter. The weighted random test-pattern generator. IEEE Transactions on Computers, C-24(7):695–700, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672883.

Slivinski:1970:ETL

[Sli70] T. A. Slivinski. An extension of threshold logic. IEEE Transactions on Computers, C-19(4):319–341, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671512.

Shen:1970:ADD

[SM70a] V. Y.-S. Shen and A. C. McKellar. An algorithm for the disjunctive decomposition of switching functions. IEEE Transactions on Computers, C-19(3):239–248, March 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671496.

Stein:1970:SIO

[SM70b] M. L. Stein and E. J. Mundstock. Sorting implicit outputs in digital simulation. IEEE Transactions on Computers, C-19(9): 844–847, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671645.

Schertz:1971:DMF

[SM71a] D. R. Schertz and G. Metze. On the design of multiple fault diagnosable networks. IEEE Transactions on Computers, C-20(11): 1361–1364, November 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671732.

Schmitt:1971:TAI

[SM71b] N. M. Schmitt and J. L. Melsa.

Two approaches for increasing storage density in modern digital computing systems. IEEE Transactions on Computers, C-20(2): 167–175, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671800.

Stein:1971:SMA

[SM71c] M. L. Stein and W. D. Munro.

Scaling machine arithmetic. IEEE Transactions on Computers, C-20(6):675–678, June 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671917.

Schaefer:1972:EDC

[SM72a] L. F. Schaefer and A. Macovski. Encoding and decoding of color information using two-dimensional spatial filtering. IEEE Transactions on Computers, C-21(7):642–647, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672161.

Schertz:1972:NRF

[SM72b] Donald R. Schertz and Gernot Metze. A new representation for faults in combinational digital circuits. IEEE Transactions on Computers, C-21(8):858–866, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009041.

Siewiorek:1973:ICS

[SM73a] D. P. Siewiorek and E. J. Mc-Cluskey. An iterative cell switch design for hybrid redundancy. IEEE Transactions on Computers, C-22(3):290–297, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672302.

Siewiorek:1973:SCS

[SM73b] D. P. Siewiorek and E. J. Mc-Cluskey. Switch complexity in systems with hybrid redundancy. IEEE Transactions on Computers, C-22(3):276–282, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672300.

Sawin:1974:ASM

[SM74] D. H. Sawin III and G. K.

Maki. Asynchronous sequential machines designed for fault detection. IEEE Transactions on Computers, C-23(3):239–249, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672511. See comments

[Wak75a].

Sawin:1975:CFT

[SM75a] D. H. Sawin III and G. K.

Maki. Comments on “Fault-Tolerant Asynchronous Networks”. IEEE Transactions on Computers, C-24(7):756–758, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672892. See [PR73].

Sawin:1975:FSA

[SM75b] D. H. Sawin III and G. K. Maki.

Fail-safe asynchronous sequential machines. IEEE Transactions on Computers, C-24(6):675–677, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672875.

Schmidt:1975:MRC

[SM75c] D. C. Schmidt and G. Metze. Modular replacement of combinational switching networls. IEEE Transactions on Computers, C-24 (1):29–48, January 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672674.

Shedletsky:1976:ELF

[SM76] J. J. Shedletsky and E. J. Mc-Cluskey. The error latency of a fault in a sequential digital circuit. IEEE Transactions on Computers, C-25(6):655–659, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674668. See comments

[DTG79].

Smith:1978:SFS

[SM78] J. E. Smith and G. Metze. Strongly fault secure logic networks. IEEE Transactions on Computers, C-27(6):491–499, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675139.

Shedletsky:1979:AR

[SM79] J. J. Shedletsky and E. J. Mc-Cluskey. Authors’ reply. IEEE Transactions on Computers, C-28(1):86, January 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675231.

Smith:1971:GSR

[Smi71a] A. R. Smith. General shiftregister sequences of arbitrary cycle length. IEEE Transactions on Computers, C-20(4):456–459, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671858.

Smith:1971:CPD

[Smi71b] D. R. Smith. Complexity of partially defined combinational switching functions. IEEE Transactions on Computers, C-20(2): 204–208, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671805.

Smith:1974:GIS

[Smi74] R. J. Smith II. Generation of internal state assignments for large asynchronous sequential machines. IEEE Transactions on Computers, C-23(9):924–932, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672649.

Smith:1975:DFT

[Smi75] T. B. Smith III. A damageand fault-tolerant input/output network. IEEE Transactions on Computers, C-24(5):505–512, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672846.

Smith:1976:MWS

[Smi76] A. J. Smith. A modified working set paging algorithm. IEEE Transactions on Computers, C-25(9):907–914, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674715.

Smith:1978:ECL

[Smi78] J. E. Smith. On the existence of combinational logic circuits exhibiting multiple redundancy. IEEE Transactions on Computers, C-27(12):1221–1225, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675032.

Smith:1979:AES

[Smi79a] A. J. Smith. An analytic and experimental study of multiple channel controllers. IEEE Transactions on Computers, C-28(1):38–49, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675220.

Smith:1979:CRT

[Smi79b] J. E. Smith. Comments on “Redundancy Testing in Combinational Networks”. IEEE Transactions on Computers, C-28(3): 261–262, March 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675332. See [LD74].

Smith:1979:DFP

[Smi79c] J. E. Smith. Detection of faults in programmable logic arrays. IEEE Transactions on Computers, C-28(11):845–853, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675264.

Smith:1979:NSC

[Smi79d] J. E. Smith. On necessary and sufficient conditions for multiple fault undetectability. IEEE Transactions on Computers, C-28(10): 801–802, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675251.

Smith:1979:USD

[Smi79e] J. E. Smith. Universal system diagnosis algorithms. IEEE Transactions on Computers, C-28(5):374–378, May 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675370.

Shen:1971:FAD

[SMW71] V. Y.-S. Shen, A. C. McKellar, and P. Weiner. A fast algorithm for the disjunctive decomposition of switching functions. IEEE Transactions on Computers, C-20(3):304–309, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671828.

Su:1971:CAS

[SN71] S. Y. H. Su and Chong-Woo Nam. Computer-aided synthesis or multiple-output multilevel NAND networks with fan-in and fan-out constraints. IEEE Transactions on Computers, C-20(12): 1445–1455, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671749.

Sklansky:1972:PMD

[SN72a] J. Sklansky and P. J. Nahin. A parallel mechanism for describing silhouettes. IEEE Transactions on Computers, C-21(11): 1233–1239, November 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672076.

Som:1972:MSP

[SN72b] A. Som and A. K. Nath. On a method of sequential pattern recognition. IEEE Transactions on Computers, C-21(12): 1433–1440, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672113.

Speelpenning:1979:SMP

[SN79] B. Speelpenning and J. Nievergelt. A simple model of processor — resource utilization in networks of communicating modules. IEEE Transactions on Computers, C-28(12):927–929, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675282.

Su:1979:AFI

[SNEL79] S. Y. W. Su, Le Huu Nguyen, A. Emam, and G. J. Lipovski. The architectural features and implementation techniques of the multicell CASSM. IEEE Transactions on Computers, C-28(6):430–445, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675382.

Sakaguchi:1970:NAM

[SNN70] M. Sakaguchi, N. Nishida, and T. Nemoto. A new associative memory system utilizing holography. IEEE Transactions on Computers, C-19(12):1174–1181, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671449.

Schuster:1979:RAP

[SNOS79] S. A. Schuster, H. B. Nguyen, E. A. Ozkarahan, and K. C. Smith. RAP.2 — an associative processor for databases and its applications. IEEE Transactions on Computers, C-28(6):446–458, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675383. Also published

in/as: ASCA, Apr.1978, pp.52–59.

Sutherland:1973:HBS

[SO73] I. E. Sutherland and D. Oestreicher. How big should a printed circuit board be? IEEE Transactions on Computers, C-22(5):537–542, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672352.

Saluja:1979:MRM

[SO79a] K. K. Saluja and E. H. Ong. Minimization of Reed–Muller canonic expansion. IEEE Transactions on Computers, C-28(7):535–537, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675401.

Srihari:1979:EAD

[SO79b] S. N. Srihari and M. K. Ohanesian. An efficient algorithm for determining Hadamard sequency vectors. IEEE Transactions on Computers, C-28(3):243–244, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675325.

Sokal:1972:OCN

[Sok72] Nathan O. Sokal. Optimum choice of noise frequency band and sampling rate for generating random binary digits from clipped white noise. IEEE Transactions on Computers, C-21(6):614–615, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009023.

Smith:1972:ACT

[SP72a] E. A. Smith and D. R. Phillips.

Automated cloud tracking using precisely aligned digital ATS pictures. IEEE Transactions on Computers, C-21(7):715–729, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672167.

Sobolewski:1972:PAAa

[SP72b] John S. Sobolewski and W. H.

Payne. Pseudonoise with arbitrary amplitude distribution — Part I: Theory. IEEE Transactions on Computers, C-21(4):337–345, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008973.

Sobolewski:1972:PAAb

[SP72c] John S. Sobolewski and W. H. Payne. Pseudonoise with arbitrary amplitude distribution — Part II: Hardware implementation. IEEE Transactions on Computers, C-21(4):346–352, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008974.

Steer:1977:DHS

[SP77] D. G. Steer and S. R. Penstone. Digital hardware for sine– cosine functions. IEEE Transactions on Computers, C-26(12): 1283–1286, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674791.

Spira:1971:TNC

[Spi71] P. M. Spira. On the time necessary to compute switching functions. IEEE Transactions on Computers, C-20(1):104–105, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671682.

Spira:1973:CTA

[Spi73] Philip M. Spira. Computation times of arithmetic and Boolean functions in (d, r) circuits. IEEE Transactions on Computers, C-22(6):552–555, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009105;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Spira.pdf.

Smith:1971:ASA

[SR71] J. R. Smith, Jr. and C. H.

Roth, Jr. Analysis and synthesis of asynchronous sequential networks using edge-sensitive flipflops. IEEE Transactions on Computers, C-20(8):847–855, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671953.

Shaham:1972:NDA

[SR72a] Z. Shaham and Z. Riesel. A note on division algorithms based on multiplication. IEEE Transactions on Computers, C-21(5):513–514, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672150.

Sheng:1972:ASM

[SR72b] C. L. Sheng and P. K. Sinha Roy. An approach for the synthesis of multithreshold threshold elements. IEEE Transactions on Computers, C-21(8):913–920, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009053.

Scherba:1973:CTM

[SR73] M. B. Scherba and R. P. Roesser. Computation of the transition matrix of a linear sequential circuit. IEEE Transactions on Computers, C-22(4):427–428, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672328.

Saluja:1974:ETT

[SR74a] K. K. Saluja and S. M. Reddy. Easily testable two-dimensional cellular logic arrays. IEEE Transactions on Computers, C-23(11): 1204–1207, November 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672424.

Saluja:1974:MTL

[SR74b] K. K. Saluja and S. M. Reddy. On minimally testable logic networks. IEEE Transactions on Computers, C-23(5):552–554, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672574.

Saluja:1975:FDT

[SR75] K. K. Saluja and S. M. Reddy.

Fault detecting test sets for Reed– Muller canonic networks. IEEE Transactions on Computers, C-24 (10):995–998, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672701.

Sundaramurthy:1977:SRF

[SR77] M. Sundaramurthy and V. U.

Reddy. Some results in fixed-point Fast Fourier Transform error analysis. IEEE Transactions on Computers, C-26(3):305–308, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674827.

Srinivasan:1971:CECa

[Sri71a] C. V. Srinivasan. Codes for error correction in high-speed memory systems — Part I: Correction of cell defects in integrated memories. IEEE Transactions on Computers, C-20(8):882–888, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671958.

Srinivasan:1971:CECb

[Sri71b] C. V. Srinivasan. Codes for error correction in high-speed memory systems. Part II: Correction of temporary and catastrophic errors. IEEE Transactions on Computers, C-20(12):1514–1520, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671757.

Srini:1975:RFF

[Sri75] V. P. Srini. Realization of fuzzy forms. IEEE Transactions on Computers, C-24(9):941–943, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672936.

Srinivasan:1976:ACI

[Sri76] C. V. Srinivasan. The architecture of coherent information system: a general problem solving system. IEEE Transactions on Computers, C-25(4):390–402, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674621.

Srini:1978:FLS

[Sri78] V. P. Srini. Fault location in a semiconductor random-access memory unit. IEEE Transactions on Computers, C-27(4):349–358, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675107.

Srini:1979:IRM

[Sri79] V. P. Srini. Iterative realization of multivalued logic systems. IEEE Transactions on Computers, C-28(4):306–310, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675351.

Singh:1971:AUD

[SS71] S. Singh and S. G. S. Shiva.

Asynchronous unit delays. IEEE Transactions on Computers, C-20(5):493–499, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671875.

Su:1972:RBM

[SS72] S. Y. H. Su and A. A. Sarris. The relationship between multivalued switching algebra and Boolean algebra under different definitions of complement. IEEE Transactions on Computers, C-21(5):479–485, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672137. See comments

[Eps73b, DM76].

Si:1975:MOS

[SS75a] Siu-Chong Si and A. K. Susskind.

A method for obtaining SPOOF’s. IEEE Transactions on Computers, C-24(5):560–562, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672853.

Su:1975:AR

[SS75b] S. Y. H. Su and A. A. Sarris. Authors’ reply. IEEE Transactions on Computers, C-24(2): 220–221, February 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672786.

Spillman:1977:DSS

[SS77] R. J. Spillman and S. Y. H. Su. Detection of single, stucktype failures in multivalued combinational networks. IEEE Transactions on Computers, C-26(12): 1242–1251, December 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674785.

Sankar:1972:SIM

[SSK72] P. V. Sankar, S. K. Sen, and E. V. Krishnamurthy. Simply invertible matrices. IEEE Transactions on Computers, C-21(5):512–513, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672149.

Suryanarayanan:1972:LCM

[SSR72] K. L. Suryanarayanan, A. C. Soudack, and S. G. Rao. An online computer method for parameter tracking of sampled-data systems. IEEE Transactions on Computers, C-21(3):292–299, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008952.

Szygenda:1976:MDS

[ST76] S. A. Szygenda and E. W.

Thompson. Modeling and digital simulation for design verification and diagnosis. IEEE Transactions on Computers, C-25 (12):1242–1253, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674591.

Sloan:1979:PRR

[ST79] K. R. Sloan, Jr. and S. L. Tanimoto.

Progressive refinement of raster images. IEEE Transactions on Computers, C-28(11):871–874, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675269.

Stabler:1970:MT

[Sta70a] E. P. Stabler. Microprogram transformations. IEEE Transactions on Computers, C-19(10): 908–916, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671391.

Stabler:1970:SDL

[Sta70b] E. P. Stabler. System description languages. IEEE Transactions on Computers, C-19(12):

1160–1173, December 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671448.

Stamopoulos:1974:PAJ

[Sta74] C. D. Stamopoulos. Parallel algorithms for joining two points by a straight-line segment. IEEE Transactions on Computers, C-23(6):642–646, June 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672598.

Stamopoulos:1975:PIP

[Sta75a] C. D. Stamopoulos. Parallel image processing. IEEE Transactions on Computers, C-24(4):424–433, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672831.

Stark:1975:ODC

[Sta75b] H. Stark. An optical-digital computer for parallel processing of images. IEEE Transactions on Computers, C-24(4):340–347, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672820.

Stefanelli:1972:SHS

[Ste72] R. Stefanelli. A suggestion for a high-speed parallel binary divider. IEEE Transactions on Computers, C-21(1):42–55, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672023.

Stellhorn:1977:IFP

[Ste77] W. H. Stellhorn. An inverted file processor for information retrieval. IEEE Transactions on Computers, C-26(12):1258–1267, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674787.

Shoenfelt:1971:TEE

[STG71] J. E. Shoenfelt, C. C. Tappert, and A. J. Goetze. Techniques for efficient encoding of features in pattern recognition. IEEE Transactions on Computers, C-20(9):1104–1106, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672004.

Stiffler:1978:CRA

[Sti78a] J. J. Stiffler. Coding for randomaccess memories. IEEE Transactions on Computers, C-27(6): 526–531, June 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675143.

Stiffler:1978:RFT

[Sti78b] J. J. Stiffler. The reliability of a fault-tolerant configuration having variable coverage. IEEE Transactions on Computers, C-27 (12):1195–1197, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675026.

Stone:1970:BTS

[Sto70a] H. S. Stone. B70-4 theory of scheduling. IEEE Transactions on Computers, C-19(9): 854–855, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671651.

Stone:1970:LMC

[Sto70b] Harold S. Stone. A logic-in-memory computer. IEEE Transactions on Computers, C-19(1):73–78, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008902.

Stone:1970:OHS

[Sto70c] Harold S. Stone. The organization of high-speed memory for parallel block transfer of data. IEEE Transactions on Computers, C-19(1):47–53, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008899.

Storwick:1970:ICT

[Sto70d] R. M. Storwick. Improved construction techniques for (d, k) graphs. IEEE Transactions on Computers, C-19(12):1214–1216, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671454.

Stone:1971:PPP

[Sto71] H. S. Stone. Parallel processing with the perfect shuffle. IEEE Transactions on Computers, C-20(2):153–161, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671798.

Stone:1972:DME

[Sto72] Harold S. Stone. Dynamic memories with enhanced data access. IEEE Transactions on Computers, C-21(4):359–366, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008976.

Stoffel:1974:CDT

[Sto74a] J. C. Stoffel. A classifier design technique for discrete variable pattern recognition problems. IEEE Transactions on Computers, C-23(4):428–441, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672551.

Stoffers:1974:SAD

[Sto74b] K. E. Stoffers. Sequential algorithm for the determination of maximum compatibles. IEEE Transactions on Computers, C-23(1):95–98, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672379.

Stone:1975:DMF

[Sto75] H. S. Stone. Dynamic memories with fast random and sequential access. IEEE Transactions on Computers, C-24(12): 1167–1174, December 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672753.

Stoffers:1977:PSE

[Sto77] K. E. Stoffers. Partitioning of separating edges: a new approach to combinational logic design. IEEE Transactions on Computers, C-26(8):833–836, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674925.

Stucki:1975:SSA

[Stu75] P. Stucki. Solid-state area scanning arrays as interface devices between optical and digital computing systems — a simulation study. IEEE Transactions on Computers, C-24(4):370–380, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672823.

Su:1970:RFL

[Su70] Yueh-Hsung Su. R70-30 fuzzy logic and its application to switching systems. IEEE Transactions on Computers, C-19(8):766, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671628.

Su:1973:BFD

[Su73] S. Y. H. Su. B73-12 fault detection in digital circuits. IEEE Transactions on Computers, C-22(5):547, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672359.

Su:1975:BLL

[Su75] S. Y. H. Su. B75-1 logic and logic design. IEEE Transactions on Computers, C-24(1):110, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672686.

Sung:1976:TSC

[Sun76] Chia-Hsiaing Sung. Testable sequential cellular arrays. IEEE Transactions on Computers, C-25 (1):11–18, January 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009199.

Sundberg:1978:EED

[Sun78] C.-E. W. Sundberg. Erasure and error decoding for semiconductor memories. IEEE Transactions on Computers, C-27(8):696–705, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675176.

Sundberg:1979:PTS

[Sun79] C.-E. Sundberg. Properties of transparent shortened codes for memories with stuck-at faults. IEEE Transactions on Computers, C-28(9):686–690, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675438.

Sureshchander:1970:CSB

[Sur70a] Sureshchander. Comments on “The Synthesis of Binary Sequence Detectors”. IEEE Transactions on Computers, C-19(2): 173–174, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671482.

Sureshchander:1970:RFF

[Sur70b] Sureshchander. RST flip-flop input equations. IEEE Transactions on Computers, C-19(11): 1118–1120, November 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671438.

Sureshchander:1973:CDM

[Sur73] Sureshchander. Comments on “Decomposition Method of Determining Maximum Compatibles”. IEEE Transactions on Computers, C-22(6):627, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=5009118;   http://
www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Sureshchander.

pdf. See [RS72b, RS73b].

Sureshchander:1974:ATA

[Sur74] Sureshchander. An algorithm for testing asummability of Boolean functions. IEEE Transactions on Computers, C-23(2):188–191, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672478.

Sureshchander:1975:MSF

[Sur75] Sureshchander. Minimization of switching functions — a fast technique. IEEE Transactions on Computers, C-24(7):753–756, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672891.

Sureshchander:1978:CDU

[Sur78] Sureshchander. Comments on “Delayed Universal Logic Modules and Sequential Machine Synthesis”. IEEE Transactions on Computers, C-27(2):191, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675065. See [LV75].

Suter:1974:MAA

[Sut74] B. W. Suter. The modular arithmetic of arbitrarily long sequences of digits. IEEE Transactions on Computers, C-23(12): 1301–1303, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672443.

Sheppard:1974:FDB

[SV74] D. A. Sheppard and Z. G. Vranesic. Fault detection of binary sequential machines using R-valued test machines. IEEE Transactions on Computers, C-23(4):352–358, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672542.

Schwartzmann:1975:ADT

[SV75] D. H. Schwartzmann and J. J. Vidal. An algorithm for determining the topological dimensionality of point clusters. IEEE Transactions on Computers, C-24 (12):1175–1182, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672754.

Svoboda:1970:ADC

[Svo70] A. Svoboda. Adder with distributed control. IEEE Transactions on Computers, C-19(8):749–751, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671618;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Svoboda.
pdf.

Svoboda:1973:PPB

[Svo73] Antonin Svoboda. Parallel processing in Boolean algebra. IEEE Transactions on Computers, C-22(9):848–851, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009177.

Sasaki:1970:CCS

[SW70a] A. Sasaki and S. Watanabe. Comment on “Computer Simulation of Pulse Propagation Through a Periodic Loaded Transmission Line”. IEEE Transactions on Computers, C-19(10):983, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671402. See [SW70b].

Sasaki:1970:CSP

[SW70b] Akio Sasaki and S. Watanabe. Computer simulation of pulse propagation through a periodic loaded transmission line. IEEE Transactions on Computers, C-19 (1):25–33, January 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008896. See comment

[SW70a].

Schmookler:1971:HSD

[SW71a] Martin S. Schmookler and Arnold Weinberger. High speed decimal addition. IEEE Transactions on Computers, C-20(8):862–867, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671955.

Smith:1971:ASP

[SW71b] F. W. Smith and M. H. Wright. Automatic ship photo interpretation by the method of moments. IEEE Transactions on Computers, C-20(9):1089–1095, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672001.

Singh:1973:MOA

[SW73] S. Singh and R. Waxman. Multiple operand addition and multiplication. IEEE Transactions on Computers, C-22(2):113–120, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672263.

Smith:1976:MHH

[SW76] Carl H. Smith and Larry D. Wittie. Memory hardware for high speed job selection. IEEE Transactions on Computers, C-25(2): 148–156, February 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009229.

Swamy:1972:GRM

[Swa72a] Sowmitri Swamy. On generalized Reed–Muller expansions. IEEE Transactions on Computers, C-21(9):1008–1009, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009078.

Swartz:1972:BFI

[Swa72b] D. Swartz. B72-22 FORTRAN IV: a modem approach. IEEE Transactions on Computers, C-21 (11):1251, November 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672082.

Swartzlander:1973:PC

[Swa73a] E. E. Swartzlander, Jr. Parallel counters. IEEE Transactions on Computers, C-22(11): 1021–1024, November 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672232.

Swartzlander:1973:QSM

[Swa73b] E. E. Swartzlander, Jr. The quasi-serial multiplier. IEEE Transactions on Computers, C-22(4):317–321, April 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672310.

Swanson:1974:IPM

[Swa74a] R. C. Swanson. Interconnections for parallel memories to unscramble p-ordered vectors. IEEE Transactions on Computers, C-23 (11):1105–1115, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672410.

Swartwout:1974:BDC

[Swa74b] R. E. Swartwout. B74-9 digital computer fundamentals. IEEE Transactions on Computers, C-23 (2):213–214, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672491.

Swartzlander:1979:CFN

[Swa79a] E. E. Swartzlander, Jr. Comment on “The Focus Number System”. IEEE Transactions on Computers, C-28(9):693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675441. See [LE77,

LE79].

Swartzlander:1979:MCS

[Swa79b] E. E. Swartzlander, Jr. Microprogrammed control for specialized processors. IEEE Transactions on Computers, C-28(12): 930–934, December 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675283.

Swartzlander:1990:CAa

[Swa90] Earl E. Swartzlander, Jr. Computer Arithmetic, volume 1 of IEEE Computer Society Press tutorial. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8931-5 (hardcover), 0-8186-5931-9 (microfiche). xiii + 378 pp. LCCN QA76.6 .C633 1990.

Sholl:1975:DAS

[SY75] H. A. Sholl and Shou-Chung Yang.

Design of asynchronous sequential networks using read-only memories. IEEE Transactions on Computers, C-24(2):195–206, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672778.

Tohma:1971:FTS

[TA71] Y. Tohma and S. Aoyagi. Failuretolerant sequential machines with past information. IEEE Transactions on Computers, C-20(4): 392–396, April 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671847.

Tabak:1970:ANP

[Tab70a] D. Tabak. An algorithm for nonlinear process stabilization and control. IEEE Transactions on Computers, C-19(6):487–492, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671558.

Tabak:1970:BCM

[Tab70b] Daniel Tabak. B70-6: Computing Methods in Optimization Problems — 2, L. A. Zadeh, L. W. Neustadt, and A. V. Balakrishnam, Eds. (Academic Press, Inc., N.Y., 1969, 393 pp. IEEE Transactions on Computers, C-19(10):986, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671404.

Tabak:1974:CPP

[Tab74] D. Tabak. Comments on “Parallel Processing Algorithms for the Optimal Control of Nonlinear Dynamic Systems”. IEEE Transactions on Computers, C-23(12):1327, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672454. See [LT73].

Takiyama:1976:RBT

[Tak76] R. Takiyama. A relationship between two types of twolayer threshold-element patternrecognition networks. IEEE Transactions on Computers, C-25(11): 1136–1139, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674567.

Tan:1971:SAA

[Tan71] Chung-Jen Tan. State assignments for asynchronous sequential machines. IEEE Transactions on Computers, C-20(4):382–391, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671846.

Tanimoto:1978:OAC

[Tan78] S. L. Tanimoto. An optimal algorithm for computing Fourier texture descriptors. IEEE Transactions on Computers, C-27(1):81–84, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674956.

Tumbush:1974:SAT

[TB74] G. L. Tumbush and J. E. Brandeberry. A state assignment technique for sequential machines using J–K flip-flops. IEEE Transactions on Computers, C-23(1):85–86, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672374.

Tylaska:1975:IBC

[TB75] T. T. Tylaska and J. D. Bargainer. An improved bound for checking experiments that use simple input-output and characterizing sequences. IEEE Transactions on Computers, C-24(6):670–673, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672873.

Tzidon:1978:PAF

[TBY78] A. Tzidon, I. Berger, and M. Yoeli. A practical approach to fault detection in combinational networks. IEEE Transactions on Computers, C-27(10):968–971, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674978.

Toussaint:1970:ARC

[TD70] G. T. Toussaint and R. W. Donaldson. Algorithms for recognizing contour-traced handprinted characters. IEEE Transactions on Computers, C-19(6):541–546, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671565.

Thayse:1973:BDC

[TD73] A. Thayse and M. Davio. Boolean differential calculus and its application to switching theory. IEEE Transactions on Computers, C-22(4):409–420, April 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672322.

Thayse:1977:LPU

[TD77] A. Thayse and J.-P. Deschamps.

Logic properties of unate discrete and switching functions. IEEE Transactions on Computers, C-26 (12):1202–1212, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674781.

Twogood:1976:EEM

[TE76] R. E. Twogood and M. P. Ekstrom.

An extension of Eklundh’s matrix transposition algorithm and its application in digital image processing. IEEE Transactions on Computers, C-25(9):950–952, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674721. See [Ekl72].

Trivedi:1977:LAD

[TE77] Kishor S. Trivedi and Miloš D. Ercegovac. On-line algorithms for division and multiplication. IEEE Transactions on Computers, C-26(7):681–687, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674901.

Teitelman:1976:CCL

[Tei76] W. Teitelman. Clisp: Conversational Lisp. IEEE Transactions on Computers, C-25(4):354–357, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674617.

Tjaden:1970:DPE

[TF70] G. S. Tjaden and M. J. Flynn. Detection and parallel execution of independent instructions. IEEE Transactions on Computers, C-19(10):889–895, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671388.

Tjaden:1973:RCO

[TF73] Garold S. Tjaden and Michael J. Flynn. Representation of concurrency with ordering matrices. IEEE Transactions on Computers, C-22(8):752–761, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009154.

Tanaka:1978:ECP

[TF78] E. Tanaka and King-Sun Fu. Errorcorrecting parsers for formal languages. IEEE Transactions on Computers, C-27(7):605–616, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675160.

Tam:1972:ASW

[TG72a] Le Dinh Chon Tam and R. Y.

Goulet. On arithmetical shift for Walsh functions. IEEE Transactions on Computers, C-21(12): 1451–1452, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672117.

Tou:1972:RHC

[TG72b] J. T. Tou and R. C. Gonzalez. Recognition of handwritten characters by topological feature extraction and multilevel categorization. IEEE Transactions on Computers, C-21(7):776–785, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672174.

Thomason:1975:SRI

[TG75] M. G. Thomason and R. C. Gonzalez. Syntactic recognition of imperfectly specified patterns. IEEE Transactions on Computers, C-24(1):93–95, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672679.

Tsuchiya:1976:TOH

[TG76] M. Tsuchiya and M. J. Gonzalez. Toward optimization of horizontal microprograms. IEEE Transactions on Computers, C-25(10): 992–999, October 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674540.

Thompson:1978:UMT

[TG78] R. A. Thompson and F. G. Gray. Universal modular trees: a design procedure. IEEE Transactions on Computers, C-27(1):53–63, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674952.

Tront:1979:DMV

[TG79] J. G. Tront and D. D. Givone. A design for multiple-valued logic gates based on MESFET’s. IEEE Transactions on Computers, C-28(11):854–862, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675265.

Trussell:1979:IMM

[TH79] H. J. Trussell and B. R. Hunt.

Improved methods of maximum a posteriori restoration. IEEE Transactions on Computers, C-28 (1):57–62, January 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675223.

Thayse:1978:MJD

[Tha78] A. Thayse. Meet and join derivatives and their use in switching theory. IEEE Transactions on Computers, C-27(8):713–720, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675178.

Therrien:1975:EPP

[The75] C. W. Therrien. Eigenvalue properties of projection operators and their application to the subspace method of feature extraction. IEEE Transactions on Computers, C-24(9):944–948, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672938.

Thomas:1974:OEU

[Tho74] R. T. Thomas. Organization for execution of user microprograms from main memory: Synthesis and analysis. IEEE Transactions on Computers, C-23(8):783–791, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672627.

Thomason:1975:SSD

[Tho75] M. G. Thomason. Stochastic syntax-directed translation schemata for correction of errors in context-free languages. IEEE Transactions on Computers, C-24 (12):1211–1216, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672758.

Thompson:1976:LCU

[Tho76] Richard A. Thompson. Language correction using probabilistic grammars. IEEE Transactions on Computers, C-25(3):275–286, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009254.

Thompson:1977:TBA

[Tho77] W. B. Thompson. Textural boundary analysis. IEEE Transactions on Computers, C-26(3): 272–276, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674818.

Thompson:1978:GCN

[Tho78] C. D. Thompson. Generalized connection networks for parallel processor intercommunication. IEEE Transactions on Computers, C-27(12):1119–1125, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675014.

Takaoka:1972:FSS

[TI72a] T. Takaoka and T. Ibaraki. Nfail-safe sequential machines. IEEE Transactions on Computers, C-21 (11):1189–1196, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672069.

Thorington:1972:ARA

[TI72b] J. M. Thorington, Jr. and J. D.

Irwin. An adaptive replacement algorithm for paged-memory computer systems. IEEE Transactions on Computers, C-21(10): 1053–1061, October 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672044.

Tison:1971:ALS

[Tis71] P. L. Tison. An algebra for logic systems-switching circuits application. IEEE Transactions on Computers, C-20(3):339–351, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671835.

Tseng:1979:IFS

[TJM79] Ben-Dau Tseng, G. A. Jullien, and W. C. Miller. Implementation of FFT structures using the residue number system. IEEE Transactions on Computers, C-28(11):831–845, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1675263;   http://
ieeexplore.ieee.org/xpl/RecentIssue.
jsp?punumber=12;   http://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=
35176.

Tjostheim:1977:RWU

[Tjo77] D. Tjostheim. Recognition of waveforms using autoregressive feature extraction. IEEE Transactions on Computers, C-26(3): 268–270, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674816.

Torode:1974:LMM

[TK74] J. Q. Torode and T. H. Kehl. The logic machine: a modular computer design system. IEEE Transactions on Computers, C-23 (11):1164–1169, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672417.

Tokura:1971:FLN

[TKH71] N. Tokura, T. Kasami, and A. Hashimoto. Failsafe logic nets. IEEE Transactions on Computers, C-20(3):323–330, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671832.

Tacker:1973:DHC

[TL73a] E. C. Tacker and T. D. Linton.

Digital and hybrid computational aspects of the discrete representation theorem of nonlinear estimation. IEEE Transactions on Computers, C-22(1):61–67, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672194.

Tang:1973:DCC

[TL73b] D. T. Tang and C. N. Liu. Distance-2 cyclic chaining of constant-weight codes. IEEE Transactions on Computers, C-22(2):176–180, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672274.

Thurber:1970:SDC

[TM70] K. J. Thurber and J. W. Myrna.

System design of a cellular APL computer. IEEE Transactions on Computers, C-19(4):291–303, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671509.

Takaoka:1971:FSL

[TM71] T. Takaoka and H. Mine. N-fail-safe logical systems. IEEE Transactions on Computers, C-20(5):536–542, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671880.

Tabloski:1976:NET

[TM76] T. F. Tabloski and F. J. Mowle. A numerical expansion technique and its application to minimal multiplexer logic circuits. IEEE Transactions on Computers, C-25(7):684–702, July 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674678.

Tseng:1978:CDC

[TM78a] B. D. Tseng and W. C. Miller. On computing the Discrete Cosine Transform. IEEE Transactions on Computers, C-27(10): 966–968, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674977.

Tsuji:1978:DEM

[TM78b] S. Tsuji and F. Matsumoto. Detection of ellipses by a modified Hough transformation. IEEE Transactions on Computers, C-27 (8):777–781, August 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675191.

To:1973:FFI

[To73] Kilin To. Fault folding for irredundant and redundant combinational circuits. IEEE Transactions on Computers, C-22(11): 1008–1015, November 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672230. See comments

[NA76a].

To:1976:AR

[To76] Kilin To. Author’s reply. IEEE Transactions on Computers, C-25(2):207, February 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009238.

Tohma:1974:DTF

[Toh74] Y. Tohma. Design technique of fail-safe sequential circuits using flip-flops for internal memory. IEEE Transactions on Computers, C-23(11):1149–1154, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672415.

Toida:1976:SDR

[Toi76] S. Toida. System diagnosis and redundant tests. IEEE Transactions on Computers, C-25(11): 1167–1170, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674575.

Tomescu:1970:MMA

[Tom70] Ioan Tomescu. A modified matrix algorithm for determining the complete connection matrix of a switching network. IEEE Transactions on Computers, C-19(1):78–79, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008903.

Tomescu:1972:MMD

[Tom72] I. Tomescu. A matrix method for determining all pairs of compatible states of a sequential machine. IEEE Transactions on Computers, C-21(5):502–503, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672143.

Tomescu:1973:CAS

[Tom73] I. Tomescu. A combinatorial algorithm for solving covering problems. IEEE Transactions on Computers, C-22(2):218–220, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672283.

Tomek:1974:TAP

[Tom74] Ivan Tomek. Two algorithms for piecewise-linear continuous approximation of functions of one variable. IEEE Transactions on Computers, C-23(4):445–448, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672554.

Toma:1975:CLA

[Tom75] C. I. Toma. Cellular logic array for high-speed signed binary number multiplication. IEEE Transactions on Computers, C-24(9):932–935, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672933.

Tohma:1971:RFS

[TOS71] Y. Tohma, Y. Ohyama, and R. Sakai. Realization of failsafe sequential machines by using a k-out-of-n code. IEEE Transactions on Computers, C-20 (11):1270–1275, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671720.

Toussaint:1971:SUB

[Tou71] G. T. Toussaint. Some upper bounds on error probability for multiclass pattern recognition. IEEE Transactions on Computers, C-20(8):943–944, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671973.

Toussaint:1972:CCS

[Tou72a] Godfried T. Toussaint. Comments on “A Comparison of Seven Techniques for Choosing Subsets of Pattern Recognition Properties”. IEEE Transactions on Computers, C-21(9):1028–1029, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009089. See [MG71b].

Toussaint:1972:CEB

[Tou72b] Godfried T. Toussaint. Comments on “Error Bounds for a Contextual Recognition Procedure”. IEEE Transactions on Computers, C-21(9):1027, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009087. See [Chu71].

Toussaint:1972:CTC

[Tou72c] Godfried T. Toussaint. Comments on “Theoretical Comparison of a Class of Feature Selection Criteria in Pattern Recognition”. IEEE Transactions on Computers, C-21(6):615–616, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009024. See [Che71a].

Toussaint:1972:PRC

[Tou72d] Godfried T. Toussaint. Polynomial representation of classifiers with independent discrete-valued features. IEEE Transactions on Computers, C-21(2):205–208, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008928.

Toussaint:1972:SIB

[Tou72e] Godfried T. Toussaint. Some inequalities between distance measures for feature evaluation. IEEE Transactions on Computers, C-21(4):409–410, April 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008991.

Toussaint:1974:CEP

[Tou74] G. T. Toussaint. Comments on “The Extraction of Pattern Features from Imperfectly Identified Samples”. IEEE Transactions on Computers, C-23(9): 988–989, September 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672662. See [Bab72].

Toussaint:1975:CNC

[Tou75] G. T. Toussaint. Comments on “On a New Class of Bounds on Bayes’ Risk in Multihypothesis Pattern Recognition”. IEEE Transactions on Computers, C-24 (8):855–856, August 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672917. See [Dev74].

Toy:1971:MLC

[Toy71a] W. N. Toy. Modular LSI control logic design with error detection. IEEE Transactions on Computers, C-20(2):161–166, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671799.

Toy:1971:NPB

[Toy71b] W. N. Toy. A novel parallel binary counter design with parity prediction and error detection scheme. IEEE Transactions on Computers, C-20(1):44–48, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671673.

Thedchanamoorthy:1972:VMD

[TP72] N. Thedchanamoorthy and J. B. Plant. A versatile multiplying digital-to-analog converter. IEEE Transactions on Computers, C-21(10):1113–1116, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672051.

Thurber:1973:FPP

[TP73] K. J. Thurber and P. C. Patton.

The future of parallel processing. IEEE Transactions on Computers, C-22(12):1140–1143, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672255.

Trivedi:1976:PAA

[Tri76] K. S. Trivedi. Prepaging and applications to array algorithms. IEEE Transactions on Computers, C-25(9):915–921, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674716.

Trivedi:1977:PPA

[Tri77a] K. S. Trivedi. On the paging performance of array algorithms. IEEE Transactions on Computers, C-26(10):938–947, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674729.

Trivedi:1977:UCF

[Tri77b] K. S. Trivedi. On the use of continued fractions for digital computer arithmetic. IEEE Transactions on Computers, C-26(7):700–704, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&

arnumber=1674903. See corrections

[Tri78].

Trivedi:1978:CUC

[Tri78] K. S. Trivedi. Corrections to “On the Use of Continued Fractions for Digital Computer Arithmetic”. IEEE Transactions on Computers, C-27(3):288, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675093. See [Tri77b].

Troxel:1975:SIM

[Tro75] D. E. Troxel. Serial interfaces for minicomputers. IEEE Transactions on Computers, C-24(10): 1027–1028, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672711.

Trunk:1971:CAF

[Tru71] G. V. Trunk. Comments on “An Algorithm for Finding Intrinsic Dimensionality of Data”. IEEE Transactions on Computers, C-20(12):1615, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671779. See [FO71a].

Trunk:1976:SEI

[Tru76] Gerard V. Trunk. Stastical estimation of the intrinsic dimensionality of a noisy signal collection. IEEE Transactions on Computers, C-25(2):165–171, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009231.

Taniguchi:1975:NVF

[TS75] K. Taniguchi and T. Sakai. A new voltage to frequency converter. IEEE Transactions on Computers, C-24(10):1035–1036, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672715.

Toueg:1979:DSD

[TS79] S. Toueg and K. Steiglitz. The design of small-diameter networks by local search. IEEE Transactions on Computers, C-28(7):537–542, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675402.

Tsichritzis:1970:RCA

[Tsi70] D. Tsichritzis. R70-35 classes of automata and transitive closure. IEEE Transactions on Computers, C-19(9):858–859, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671656.

Tung:1970:ACP

[Tun70a] Chin Tung. On the apparent continuity of processing in a paging environment. IEEE Transactions on Computers, C-19(11):

1047–1054, November 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671424.

Tung:1970:SDD

[Tun70b] Chin Tung. Signed-digit division using combinational arithmetic nets. IEEE Transactions on Computers, C-19(8):746–748, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1671617;   http://
www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Tung.pdf.

Turner:1970:RSD

[Tur70] J. Turner. R70-22 subdirect decompositions of transformation graphs. IEEE Transactions on Computers, C-19(7):659, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671602.

Toussaint:1972:CFS

[TV72] Godfried T. Toussaint and Toomas R. Vilmansen. Comments on “Feature Selection with a Linear Dependence Measure”. IEEE Transactions on Computers, C-21(4):408, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008989. See [Das71a].

Turcat:1976:RTC

[TV76] C. Turcat and A. Verdillon. Recursion and testing of combinational circuits. IEEE Transactions on Computers, C-25(6):652–654, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674667.

Turcat:1979:SAT

[TV79] C. Turcat and A. Verdillon. Symmetry, automorphism, and test. IEEE Transactions on Computers, C-28(4):319–325, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675354.

Turn:1976:PSI

[TW76] R. Turn and W. H. Ware. Privacy and security issues in information systems. IEEE Transactions on Computers, C-25(12): 1353–1361, December 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674604.

Torng:1977:OIC

[TW77] H. C. Torng and N. C. Wilhelm. The optimal interconnection of circuit modules in microprocessor and digital system design. IEEE Transactions on Computers, C-26(5):450–457, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674862.

Tang:1978:OTN

[TWB78] D. T. Tang, L. S. Woo, and L. R. Bahl. Optimization of teleprocessing networks with concentrators and multiconnected terminals. IEEE Transactions on Computers, C-27(7):594–604, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675159.

Ulug:1974:UTA

[UB74] M. E. Ulug and B. A. Bowen. A unified theory of the algebraic topological methods for the synthesis of switching systems. IEEE Transactions on Computers, C-23(3):255–267, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672513.

Underwood:1975:VLM

[UC75] S. A. Underwood and C. L. Coates, Jr. Visual learning from multiple views. IEEE Transactions on Computers, C-24(6):651–661, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672870.

Uhr:1972:LRC

[Uhr72] L. Uhr. Layered “Recognition Cone” networks that preprocess, classify, and describe. IEEE Transactions on Computers, C-21(7):758–768, July 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672172.

Uhrig:1973:BDP

[Uhr73] J. Uhrig. B73-3 dynamic programming and partial differential equations. IEEE Transactions on Computers, C-22(2):221, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672284.

Ullmann:1972:TLB

[Ull72] J. R. Ullmann. Transference of learning between recognition classes. IEEE Transactions on Computers, C-21(2):219–220, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5008932.

Ulman:1970:CHT

[Ulm70] L. J. Ulman. Computation of the Hadamard transform and the Rtransform in ordered form. IEEE Transactions on Computers, C-19(4):359–360, April 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671518.

Ulrich:1971:NSS

[Ulr71] J. W. Ulrich. A note on the solution of sequential Boolean equations. IEEE Transactions on Computers, C-20(2):231–234, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671815.

Udupa:1977:NCT

[UM77] K. J. Udupa and I. S. N. Murthy. New concepts for three-dimensional shape analysis. IEEE Transactions on Computers, C-26(10): 1043–1049, October 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674744.

Unger:1971:ASS

[Ung71] S. H. Unger. Asynchronous sequential switching circuits with unrestricted input changes. IEEE Transactions on Computers, C-20 (12):1437–1444, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671748.

Unger:1977:CRA

[Ung77a] B. W. Unger. A computer resource allocation model with some measured and simulation results. IEEE Transactions on Computers, C-26(3):243–259, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674813.

Unger:1977:SSC

[Ung77b] S. H. Unger. Self-synchronizing circuits and nonfundamental mode operation. IEEE Transactions on Computers, C-26(3):278–281, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674820.

Unger:1977:TRI

[Ung77c] S. H. Unger. Tree realizations of iterative circuits. IEEE Transactions on Computers, C-26(4):365–383, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674846.

Unger:1977:GCS

[Ung77d] Stephen H. Unger. The generation of completion signals in iterative combinational circuits. IEEE Transactions on Computers, C-26(1):13–18, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009268.

Usas:1975:TSC

[Usa75] A. M. Usas. A totally selfchecking checker design for the detection of errors in periodic signals. IEEE Transactions on Computers, C-24(5):483–489, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672843.

Vairavan:1971:RCI

[Vai71] K. Vairavan. A remark on the concepts of input-memory and output-memory of sequential machines. IEEE Transactions on Computers, C-20(2):239, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671817.

Vairavan:1972:MIM

[Vai72] K. Vairavan. Minimal inputmemory and output-memory finitestate machines. IEEE Transactions on Computers, C-21(6):598–602, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009017.

VanVoorhis:1972:ILB

[Van72] David C. Van Voorhis. An improved lower bound for sorting networks. IEEE Transactions on Computers, C-21(6):612–613, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009021.

VanLantschoot:1973:DAB

[Van73a] E. J. Van Lantschoot. Double adjacencies between cycles of a circulating shift register.

IEEE Transactions on Computers, C-22(10):944–955, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672215.

VanTill:1973:CDG

[Van73b] J. W. J. Van Till. Comments on “On the Definition and Generation of Walsh Functions”. IEEE Transactions on Computers, C-22(7):702–703, July 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009139. See [Dav72a].

Vandewalle:1975:CPL

[Van75] J. Vandewalle. On the calculation of the piecewise linear approximation to a discrete function. IEEE Transactions on Computers, C-24(8):843–846, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672913.

VanVoorhis:1977:CCA

[Van77] D. C. Van Voorhis. Comments on “A Computer Algorithm for Transposing Nonsquare Matrices”. IEEE Transactions on Computers, C-26(6):607–608, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674887. See [All75a].

Varshney:1979:AMI

[Var79] P. K. Varshney. On analytical modeling of intermittent faults in digital systems. IEEE Transactions on Computers, C-28(10): 786–791, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675247.

Vairavan:1976:CCG

[VD76a] K. Vairavan and R. A. Demillo. On the computational complexity of a generalized scheduling problem. IEEE Transactions on Computers, C-25(11):1067–1073, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674556.

VanHaersmaBuma:1976:CCP

[VD76b] Chris E. Van Haersma Buma and Robert P. W. Duin. Computation of concave piecewise linear discriminant functions using Chebyshev polynomials. IEEE Transactions on Computers, C-25(2): 178–186, February 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009233.

Veelenturf:1978:ISM

[Vee78] L. P. J. Veelenturf. Inference of sequential machines from sample computations. IEEE Transactions on Computers, C-27(2):

167–170, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675053.

Vichnevetsky:1971:MDU

[Vic71] R. Vichnevetsky. The method of decomposition for unstable twopoint boundary value problems. IEEE Transactions on Computers, C-20(8):910–914, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671962.

Vichnevetsky:1974:BCS

[Vic74] R. Vichnevetsky. B74-24 computer simulation of dynamic systems. IEEE Transactions on Computers, C-23(3):335–336, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672533.

Vilfan:1970:IUB

[Vil70] B. Vilfan. An improved upper bound for the finite delay of graphs. IEEE Transactions on Computers, C-19(5):453–456, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671537.

Vilmansen:1972:DDP

[Vil72] Toomas R. Vilmansen. On dependence and discrimination in pattern recognition. IEEE Transactions on Computers, C-21(9): 1029–1031, September 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009091.

Vilfan:1973:APT

[Vil73a] B. Vilfan. Another proof of the two-dimensional Cayley–Hamilton theorem. IEEE Transactions on Computers, C-22(12):1140, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672253.

Vilmansen:1973:FEM

[Vil73b] T. R. Vilmansen. Feature evaluation with measures of probabilistic dependence. IEEE Transactions on Computers, C-22(4):381–388, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672318.

Vilmansen:1974:AMI

[Vil74] T. Vilmansen. On an approximation of mutual information. IEEE Transactions on Computers, C-23(12):1311–1313, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672446.

Vink:1978:MTN

[Vin78] H. A. Vink. Minimal TANT networks of functions with DON’T CARE’S and some complemented input variables. IEEE Transactions on Computers, C-27(11): 1073–1078, November 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675000.

Vranesic:1970:MVA

[VLS70] Z. G. Vranesic, E. S. Lee, and K. C. Smith. A many-valued algebra for switching systems. IEEE Transactions on Computers, C-19(10):964–971, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671396.

VanVoorhis:1978:MSI

[VM78] D. C. Van Voorhis and T. H.

Morrin. Memory systems for image processing. IEEE Transactions on Computers, C-27(2): 113–125, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675045.

VanCampenhout:1978:SMC

[VN78] J. M. Van Campenhout and P. Notredame. A stochastic model for closed-loop preemptive microprocessor I/O organizations. IEEE Transactions on Computers, C-27(12):1126–1136, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675015.

Voith:1974:BDD

[Voi74] R. P. Voith. B 74-32 digital design. IEEE Transactions on Computers, C-23(5):558–559, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672581.

Voith:1975:BCO

[Voi75] R. P. Voith. B75-4 computer organization and programming. IEEE Transactions on Computers, C-24(2):222–223, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672788.

Voith:1977:MUL

[Voi77a] R. P. Voith. Minimum universal logic module sequential circuits with decoders. IEEE Transactions on Computers, C-26(10): 1032–1035, October 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674741.

Voith:1977:UIM

[Voi77b] R. P. Voith. ULM implicants for minimization of universal logic module circuits. IEEE Transactions on Computers, C-26(5):417–424, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674858.

VonPuttkamer:1975:SHB

[Von75] E. Von Puttkamer. A simple hardware buddy system memory allocator. IEEE Transactions on Computers, C-24(10):953–957, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672693.

Vanderbrug:1977:TST

[VR77] G. J. Vanderbrug and A. Rosenfeld. Two-stage template matching. IEEE Transactions on Computers, C-26(4):384–393, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674847.

Vlasenko:1979:UMT

[VR79] V. Vlasenko and K. R. Rao. Unified matrix treatment of discrete transforms. IEEE Transactions on Computers, C-28(12):934–938, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675284.

Vranesic:1977:MVL

[Vra77] Z. G. Vranesic. Multiple-valued logic: An introduction and overview.

IEEE Transactions on Computers, C-26(12):1181–1182, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674778.

Vinea:1971:DCF

[VV71] A. Vinea and V. Vinea. A distance criterion for figural pattern recognition. IEEE Transactions on Computers, C-20(6):680–685, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671919.

VanLantschoot:1974:DWC

[VV74] E. J. Van Lantschoot and J. P. Vandewalle. Design of weighted counters with rational scale using continued fraction expansion. IEEE Transactions on Computers, C-23(3):232–238, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672510.

Vink:1978:RCT

[VVA78] H. A. Vink, B. Van Den Dolder, and J. Al. Reduction of CCtables using multiple implication. IEEE Transactions on Computers, C-27(10):961–966, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674976.

Vranesic:1972:FTS

[VW72] Z. G. Vranesic and K. M. Waliuzzaman. Functional transformation in simplification of multivalued switching functions. IEEE Transactions on Computers, C-21 (1):102–105, January 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672033.

Williams:1973:ETL

[WA73] M. J. Y. Williams and J. B. Angell. Enhancing testability of large-scale integrated circuits via test points and additional logic. IEEE Transactions on Computers, C-22(1):46–60, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672193.

Wadel:1971:CNR

[Wad71] L. B. Wadel. Comment on “Negative Radix Conversion”. IEEE Transactions on Computers, C-20(5):587, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671893. See [Paw71,

Zoh70].

Wagner:1970:RTA

[Wag70] E. G. Wagner. R70-32 tessellation automata. IEEE Transactions on Computers, C-19(9):

857–858, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671653.

Wait:1970:SVT

[Wai70] J. V. Wait. State-variable techniques for digital simulation of bandpass systems. IEEE Transactions on Computers, C-19(9): 830–831, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671641.

Waksman:1970:CI

[Wak70a] A. Waksman. On the complexity of inversions. IEEE Transactions on Computers, C-19(12): 1225–1226, December 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671459.

Waksman:1970:WAI

[Wak70b] A. Waksman. On Winograd’s algorithm for inner products. IEEE Transactions on Computers, C-19(4):360–361, April 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671519.

Waksman:1970:RGF

[Wak70c] A. Waksman. R-70-31 A generalized firing squad problem. IEEE Transactions on Computers, C-19(8):766, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671629.

Wakerly:1974:PSC

[Wak74] J. F. Wakerly. Partially selfchecking circuits and their use in performing logical operations. IEEE Transactions on Computers, C-23(7):658–666, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672607.

Wakerly:1975:CAS

[Wak75a] J. F. Wakerly. Comments on “Asynchronous Sequential Machines Designed for Fault Detection”. IEEE Transactions on Computers, C-24(7):759–760, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672895. See [SM74].

Wakerly:1975:DUM

[Wak75b] J. F. Wakerly. Detection of unidirectional multiple errors using low-cost arithmetic codes. IEEE Transactions on Computers, C-24(2):210–212, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672780.

Wakerly:1975:TFT

[Wak75c] J. F. Wakerly. Transient failures in triple modular redundancy systems with sequential modules. IEEE Transactions on Computers, C-24(5):570–573, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672856.

Wallach:1974:BCA

[Wal74] S. J. Wallach. B 74-39 A course in APL/360 with applications. IEEE Transactions on Computers, C-23(9):990–991, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672666.

Walker:1976:SUT

[Wal76] Donald E. Walker. Speech understanding through syntactic and semantic analysis. IEEE Transactions on Computers, C-25(4):432–439, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674625.

Wang:1971:ANS

[Wan71] K. C. Wang. On autonomous NOR sequential machines. IEEE Transactions on Computers, C-20(5):580–582, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671888.

Wang:1972:BDS

[Wan72] C. P. Wang. B72-21 data structures and management. IEEE Transactions on Computers, C-21 (11):1250, November 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672081.

Wang:1975:AGT

[Wan75a] D. T. Wang. An algorithm for the generation of test sets for combinational logic networks. IEEE Transactions on Computers, C-24(7):742–746, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672888.

Wang:1975:PFC

[Wan75b] D. T. Wang. Properties of faults and criticalities of values under tests for combinational networks. IEEE Transactions on Computers, C-24(7):746–750, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672889.

Wang:1975:CBP

[Wan75c] Y. R. Wang. Characterization of binary patterns and their projections. IEEE Transactions on Computers, C-24(10):1032–1035, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672714.

Wang:1977:CSR

[Wan77] L. Wang. Cross-section reconstruction with a fan-beam scanning geometry. IEEE Transactions on Computers, C-26(3): 264–268, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674815.

Warner:1976:CCT

[WBRG76] Harold Warner, Charles L. Bell, Duane M. Rumbaugh, and Timothy V. Gill. Computer-controlled teaching instrumentation for linguistic studies with the Great Apes. IEEE Transactions on Computers, C-25(1):38–43, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009203.

Wong:1974:PGB

[WC74] C. K. Wong and Shi-Kuo Chang. Parallel generation of binary search trees. IEEE Transactions on Computers, C-23(3):268–271, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672514.

Wong:1976:GPM

[WC76] C. K. Wong and Don Coppersmith. The generation of permutations in magnetic bubble memories. IEEE Transactions on Computers, C-25(3):254–262, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009252.

Wernecke:1977:MEI

[WD77] S. J. Wernecke and L. R. D’Addario.

Maximum entropy image reconstruction. IEEE Transactions on Computers, C-26(4):351–364, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674845.

Weaver:1975:SPT

[Wea75] C. S. Weaver. Some properties of threshold logic unit pattern recognition networks. IEEE Transactions on Computers, C-24 (3):290–298, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672802.

Wegbreit:1972:SEL

[Weg72] Ben Wegbreit. A space-efficient list structure tracing algorithm. IEEE Transactions on Computers, C-21(9):1009–1010, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009079.

Wegner:1976:PLF

[Weg76] P. Wegner. Programming languages — the first 25 years. IEEE Transactions on Computers, C-25(12):1207–1225, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674589.

Wehrfritz:1974:TTL

[Weh74] H. C. Wehrfritz. Techniques for the transformation of logic equations. IEEE Transactions on Computers, C-23(5):477–480, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672563.

Weiss:1972:BLT

[Wei72] C. Dennis Weiss. Bounds on the length of terminal stuck-fault test. IEEE Transactions on Computers, C-21(3):305–309, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008955.

Weiss:1975:NPD

[Wei75] H. Weiss. Nonredundant point distribution for coded aperture imaging with application to threedimensional on-line X-ray information retrieving. IEEE Transactions on Computers, C-24(4):391–394, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672825.

Welch:1974:CMC

[Wel74] T. A. Welch. Comments on “A Measure of Computational Work”. IEEE Transactions on Computers, C-23(2):208, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672486. See [Hel72,

Hel74, Mas78].

Welch:1978:MHC

[Wel78] T. A. Welch. Memory hierarchy configuration analysis. IEEE Transactions on Computers, C-27(5):408–413, May 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675120.

Wesselkamper:1978:DDM

[Wes78] T. C. Wesselkamper. Divided difference methods for Galois switching functions. IEEE Transactions on Computers, C-27(3): 232–238, March 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675076.

Wetjen:1978:CBB

[Wet78] A. Wetjen. Comments on “A Branch and Bound Clustering Algorithm”. IEEE Transactions on Computers, C-27(2):189, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675059. See [KNF75].

Wexelblatt:1971:BPS

[Wex71] R. L. Wexelblatt. B71-1 problem solving by digital computers with PL/I programming. IEEE Transactions on Computers, C-20(8):948, August 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671975.

Warmack:1973:AOS

[WG73] R. E. Warmack and R. C. Gonzalez. An algorithm for the optimal solution of linear inequalities and its application to pattern recognition. IEEE Transactions on Computers, C-22(12):1065–1075, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672245.

Waksman:1974:CRP

[WG74] A. Waksman and M. W. Green. On the consecutive retrieval property in file organization. IEEE Transactions on Computers, C-23 (2):173–174, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672473.

Will:1975:ESC

[WG75] P. M. Will and D. D. Grossman.

An experimental system for computer controlled mechanical assembly. IEEE Transactions on Computers, C-24(9):879–888, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672926.

Wendling:1978:SIW

[WGS78] S. Wendling, G. Gagneux, and G. Stamon. A set of invariants within the power spectrum of unitary transformations. IEEE Transactions on Computers, C-27 (12):1213–1216, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675030.

Wong:1978:SHS

[WH78] R. Y. Wong and E. L. Hall. Sequential hierarchical scene matching. IEEE Transactions on Computers, C-27(4):359–366, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675108.

White:1970:FDC

[Whi70] R. C. White, Jr. A fast digital computer method for recursive estimation of the mean. IEEE Transactions on Computers, C-19(9):847–850, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671646. See comments

[Maj72b].

White:1971:FES

[Whi71a] H. S. White. Finding events in a sea of bubbles. IEEE Transactions on Computers, C-20(9):988–995, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671986.

Whitney:1971:DMN

[Whi71b] A. W. Whitney. A direct method of nonparametric measurement selection. IEEE Transactions on Computers, C-20(9):1100–1103, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672003.

Whitney:1971:AFA

[Whi71c] G. E. Whitney. Algebraic fault analysis for constrained combinational networks. IEEE Transactions on Computers, C-20(2): 141–148, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671796.

White:1972:CNM

[Whi72] I. White. Comment on “A Nonlinear Mapping for Data Structure Analysis”. IEEE Transactions on Computers, C-21(2): 220–221, February 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5008933. See [Kru71].

Wilkes:1970:SEL

[Wil70] M. A. Wilkes. Scroll editing:

An on-line algorithm for manipulating long character strings. IEEE Transactions on Computers, C-19(11):1009–1015, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671418.

Wilkes:1971:SMS

[Wil71] M. V. Wilkes. Slave memories and segmentation. IEEE Transactions on Computers, C-20(6):674–675, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671916.

Williams:1975:UDI

[Wil75] G. H. Williams. Uniform decomposition of incompletely specified sequential machines. IEEE Transactions on Computers, C-24 (8):840–843, August 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672912.

Williams:1977:SIA

[Wil77a] R. P. Williams. Serial integer arithmetic with magnetic bubbles. IEEE Transactions on Computers, C-26(3):260–264, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674814.

Williams:1977:CLM

[Wil77b] T. P. T. Williams. Comments on “A Least Mean Squares Cubic Algorithm for On-Line Differential of Sampled Analog Signals”. IEEE Transactions on Computers, C-26(4):416, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674852. See [All75b].

Wilkens:1978:RSM

[Wil78] E. J. Wilkens. Realizations of sequential machines using random access memory. IEEE Transactions on Computers, C-27(5):429–441, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1675123.

Winder:1970:RMS

[Win70a] R. O. Winder. R70-13 multigate synthesis of general Boolean functions by threshold logic elements. IEEE Transactions on Computers, C-19(6):565, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671577.

Winder:1970:TLA

[Win70b] R. O. Winder. Threshold logic asymptotes. IEEE Transactions on Computers, C-19(4):349–353, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671514.

Winzer:1975:CRC

[Win75] G. Winzer. Character recognition with a coherent optical multichannel correlator. IEEE Transactions on Computers, C-24(4):419–423, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672830.

Winker:1976:EIQ

[Win76] S. K. Winker. An evaluation of an implementation of qualified hyperresolution. IEEE Transactions on Computers, C-25(8):835–843, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674702.

Witsenhausen:1970:HSP

[Wit70] H. S. Witsenhausen. Hybrid solution of partial differential equations. IEEE Transactions on Computers, C-19(9):851, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671648.

Webb:1971:SCF

[WK71] S. R. Webb and J. D. Kennedy. Some comments on the Fast Hadamard Transform of order twelve. IEEE Transactions on Computers, C-20(5):588–590, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671898.

Wasserstorm:1973:ASN

[WK73] E. Wasserstorm and A. Kirszenblat. Analog solutions of nonlinear boundary-value problems by the continuation method. IEEE Transactions on Computers, C-22(11):966–970, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672225.

Weiler:1970:RTO

[WKD70] P. W. Weiler, R. S. Kopp, and R. G. Dorman. A real-time operating system for manned spaceflight. IEEE Transactions on Computers, C-19(5):388–398, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671529.

Wade:1975:NEA

[WKP75] G. Wade, H. Keyani, and Soo-Chang Pei. New experiments and analysis in Bragg-diffraction imaging. IEEE Transactions on Computers, C-24(4):395–401, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672826.

Wang:1974:WPS

[WL74] Tse Lin Wang and Chao-Kai Liu. Weight-preserved single-error-correcting scheme for binary adders. IEEE Transactions on Computers, C-23(10):1002–1007, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672394.

Wong:1975:TDF

[WL75] A. K. C. Wong and T. S. Liu.

Typicality, diversity, and feature pattern of an ensemble. IEEE Transactions on Computers, C-24 (2):158–181, February 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672776.

Wong:1977:DDC

[WL77] Andrew K. C. Wong and T. S.

Liu. A decision-directed clustering algorithm for discrete data. IEEE Transactions on Computers, C-26(1):75–82, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009277.

Wojcik:1971:CBA

[WM71] A. S. Wojcik and G. Metze. On the cost of base N adders. IEEE Transactions on Computers, C-20(10):1196–1203, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671698.

Wickersham:1974:SAS

[WM74] R. S. Wickersham and G. K. Maki. Safe asynchronous sequential circuits. IEEE Transactions on Computers, C-23(5):494–500, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672565.

Wilson:1976:RRS

[WM76] G. A. Wilson and J. Minker. Resolution, refinements, and search strategies: a comparative study. IEEE Transactions on Computers, C-25(8):782–801, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674697.

Waterman:1976:IIT

[WN76] D. A. Waterman and A. Newell. PAS-II: An interactive task-free version of an automatic protocol analysis system. IEEE Transactions on Computers, C-25(4):402–413, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674622.

Weszka:1974:TST

[WNR74] J. S. Weszka, R. N. Nagel, and A. Rosenfeld. A threshold selection technique. IEEE Transactions on Computers, C-23(12): 1322–1326, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672451.

Wood:1979:HDP

[Woo79] R. A. Wood. A high density programmable logic array chip. IEEE Transactions on Computers, C-28(9):602–608, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675427.

Worland:1976:PMN

[Wor76] P. B. Worland. Parallel methods for the numerical solution of ordinary differential equations. IEEE Transactions on Computers, C-25(10):1045–1048, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674545.

Wormald:1977:NSI

[Wor77] E. G. Wormald. A note on synchronizer or interlock maloperation. IEEE Transactions on Computers, C-26(3):317–318, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674833. See comments

[Cha79, Wor79].

Wormald:1979:SJC

[Wor79] E. G. Wormald. Support for T. J. Chaney’s comments on “A Note on Synchronizer or Interlock Maloperation”. IEEE Transactions on Computers, C-28(10):804, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675253. See [Wor77,

Cha79].

Waxman:1977:LCC

[WR77] J. Waxman and J. Rootenberg. Logic circuit for cycle detection in a state diagram. IEEE Transactions on Computers, C-26 (3):303–305, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674826.

Wright:1973:TSS

[Wri73] Thomas J. Wright. A two-space solution to the hidden line problem for plotting functions of two variables. IEEE Transactions on Computers, C-22(1):28–33, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672190.

Williams:1976:SBC

[WS76] Stephen Williams and Ivor R.

Smith. SCR bridge convertor computation using tensor methods. IEEE Transactions on Computers, C-25(1):1–6, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009197.

Walker:1979:RSM

[WSB79] W. K. S. Walker, C.-E. W. Sundberg, and C. J. Black. A reliable spaceborne memory with a single error and erasure correction scheme. IEEE Transactions on Computers, C-28(7):493–500, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675394.

Wulf:1974:BPT

[Wul74] W. A. Wulf. B 74-27 program test methods. IEEE Transactions on Computers, C-23(5):555–556, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672576.

Wong:1973:ACC

[WY73] C. K. Wong and P. C. Yue. The anticipatory control of a cyclically permutable memory. IEEE Transactions on Computers, C-22(5):481–488, May 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672344.

Wakimura:1975:CCE

[WY75] Y. Wakimura and N. Yoshida. Comments on “Checking Experiments for Sequential Machines”. IEEE Transactions on Computers, C-24(11):1141–1142, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672743. See [Hsi71b].

Wyman:1973:LDD

[Wym73] Robert H. Wyman. Loading of decoder drivers. IEEE Transactions on Computers, C-22(9): 878–879, September 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009189.

Wyner:1971:NCC

[Wyn71] A. D. Wyner. Note on circuits and chains of spread k in the n-cube. IEEE Transactions on Computers, C-20(4):474, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671865.

Williams:1972:CSM

[WZ72] George H. Williams and E. G. Zavisca. Comments on “Sequential Machine Identification”. IEEE Transactions on Computers, C-21(6):616, June 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009026. See [Kel71].

Yamada:1970:RRE

[Yam70] H. Yamada. R70-11 representation of events in the von Neumann cellular model. IEEE Transactions on Computers, C-19(6):563–564, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671575.

Yang:1973:CPM

[Yan73] Chao-Chih Yang. Closure partition method for minimizing incomplete sequential machines. IEEE Transactions on Computers, C-22 (12):1109–1122, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672250.

Yang:1974:MDP

[Yan74a] C. C. Yang. On the modeling of demand paging algorithms by finite automata. IEEE Transactions on Computers, C-23(8):870–874, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672635.

Yang:1974:GAC

[Yan74b] Chao-Chih Yang. Generation of all closed partitions on the state set of a sequential machine. IEEE Transactions on Computers, C-23(5):530–533, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672569.

Yang:1975:ETA

[Yan75a] Chao-Chih Yang. On the equivalence of two algorithms for finding all maximal compatibles. IEEE Transactions on Computers, C-24 (10):977–979, October 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672696.

Yang:1975:SPM

[Yan75b] Chao-Chih Yang. Structural preserving morphisms of finite automata and an application to graph isomorphism. IEEE Transactions on Computers, C-24(11): 1133–1139, November 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672741.

Yang:1975:CCP

[YB75] Chao-Chih Yang and M. K. Babinski. Comments on “Closure Partition Method for Minimizing Incomplete Sequential Machines”. IEEE Transactions on Computers, C-24(1):106–108, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672683. See [RB76].

Yau:1976:DMA

[YC76] S. S. Yau and J. Chung. On the design of modulo arithmetic units based on cyclic groups. IEEE Transactions on Computers, C-25(11):1057–1067, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674555.

Yeh:1970:RLF

[Yeh70] R. T. Yeh. R70-43 lattice functions, pair algebras, and finite-state machines. IEEE Transactions on Computers, C-19(10):989, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671409.

Yenersoy:1979:SAM

[Yen79] O. Yenersoy. Synthesis of asynchronous machines using mixedoperation mode. IEEE Transactions on Computers, C-28(4): 325–329, April 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675355.

Yajima:1974:PMP

[YI74] S. Yajima and K. Inagaki. Power minimization problems of logic networks. IEEE Transactions on Computers, C-23(2):153–165, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672471.

Yip:1976:SAZ

[Yip76] P. C. Y. Yip. Some aspects of the zoom transform. IEEE Transactions on Computers, C-25 (3):287–296, March 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009255.

Yelowitz:1976:NRT

[YK76] L. Yelowitz and A. Kandel. New results and techniques in resolution theory. IEEE Transactions on Computers, C-25(7):673–677, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674676.

Yau:1973:ECR

[YL73] S. S.-S. Yau and Yu-Cheng Liu. Error correction in redundant residue number systems. IEEE Transactions on Computers, C-22(1):5–11, January 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=1672187;   http://
                                        ieeexplore.ieee.org/xpl/tocresult.
                                        jsp?isnumber=35069.

Yang:1978:CSC

[YM78] Chao-Chin Yang and C. P. May. A correction and some comments concerning graph isomorphism by finite automata. IEEE Transactions on Computers, C-27(1):95–96, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674961.

Yau:1970:FDR

[YO70] S. S. Yau and M. Orsic. Fault diagnosis and repair of cutpoint cellular arrays. IEEE Transactions on Computers, C-19(3):259–262, March 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671499.

Yoeli:1970:RPA

[Yoe70a] M. Yoeli. R70-26 probabilistic aspects of machine decomposition theory. IEEE Transactions on Computers, C-19(8):764, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671624.

Yoeli:1970:SMC

[Yoe70b] M. Yoeli. The synthesis of multivalued cellular cascades. IEEE Transactions on Computers, C-19 (11):1089–1090, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671430.

Yohe:1973:RFP

[Yoh73] J. Michael Yohe. Roundings in floating-point arithmetic. IEEE Transactions on Computers, C-22(6):577–586, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=
                                        &arnumber=5009110;   http://
                                        www.acsel-lab.com/arithmetic/
                                        arith2/papers/ARITH2_Yohe.pdf.

Young:1970:RAM

[You70] P. C. Young. R70-8 analog methods for on-line system identification using noisy measurements. IEEE Transactions on Computers, C-19(5):465–466, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671552.

Young:1971:RLF

[You71] T. Y. Young. The reliability of linear feature extractors. IEEE Transactions on Computers, C-20(9):967–971, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671983.

Yue:1975:TRM

[YRT75] On-Ching Yue, E. L. Rope, and G. Tricoles. Two reconstruction methods for microwave imaging of buried dielectric anomalies. IEEE Transactions on Computers, C-24(4):381–390, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672824.

Yau:1970:ULM

[YT70] S. S. Yau and C. K. Tang. Universal logic modules and their applications. IEEE Transactions on Computers, C-19(2):141–149, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671472. See comments

[Ale71].

Yau:1971:IRS

[YT71a] S. S. Yau and Y. S. Tang. On identification of redundancy and symmetry of switching functions. IEEE Transactions on Computers, C-20(12):1609–1613, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671776.

Yau:1971:TAS

[YT71b] S. S. Yau and Y. S. Tang. Transformation of an arbitrary switching function to a totally symmetric function. IEEE Transactions on Computers, C-20(12): 1606–1609, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671775.

Yau:1971:EAG

[YT71c] S. S. Yau and Yu-Shan Tang. An efficient algorithm for generating complete test sets for combinational logic circuits. IEEE Transactions on Computers, C-20 (11):1245–1251, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671716.

Yang:1976:ADA

[YT76] Chao-Chih Yang and Marilyn A. Tarpy. An algorithm for deriving all pairs of compatible states by closure classes. IEEE Transactions on Computers, C-25(2): 202–207, February 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009236.

Yachida:1977:VMV

[YT77] M. Yachida and S. Tsuji. A versatile machine vision system for complex industrial parts. IEEE Transactions on Computers, C-26(9):882–894, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674936.

Yuan:1974:MMI

[Yua74] Fu-Min Yuan. Minimal memory inverses of linear sequential circuits. IEEE Transactions on Computers, C-23(11):1155–1163, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672416.

Yuen:1972:UBW

[Yue72] Chung-Kwong Yuen. Upper bounds on Walsh transforms. IEEE Transactions on Computers, C-21 (12):1273–1280, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672091.

Yuen:1973:TF

[Yue73a] C. K. Yuen. On the twiddling factors. IEEE Transactions on Computers, C-22(5):544–545, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672355.

Yuen:1973:PON

[Yue73b] Chung-Kwong Yuen. The printing of octal numerals. IEEE Transactions on Computers, C-22(1):111, January 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672199.

Yuen:1974:CSN

[Yue74a] C. K. Yuen. Comment on “Some New Results on Average Worst Case Carry”. IEEE Transactions on Computers, C-23(3):333, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672529. See [Bri73].

Yuen:1974:EWP

[Yue74b] Chung Kwong Yuen. Evaluation of Walsh power spectrum of nearly white signals. IEEE Transactions on Computers, C-23 (11):1213–1214, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672426.

Yuen:1975:CRW

[Yue75a] C. K. Yuen. Computing robust Walsh–Fourier transform by error product minimization. IEEE Transactions on Computers, C-24 (3):313–317, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672806.

Yuen:1975:NBA

[Yue75b] C. K. Yuen. A note on base −2 arithmetic logic. IEEE Transactions on Computers, C-24(3): 325–329, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672809. See comments

[Agr77].

Yuen:1975:FPR

[Yue75c] C. K. Yuen. On the floating point representation of complex numbers. IEEE Transactions on Computers, C-24(8):846–848, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672914. See comments

[Bau77].

Yuen:1975:TCC

[Yue75d] C. K. Yuen. A theorem on the cyclic correlations of Walsh functions. IEEE Transactions on Computers, C-24(10):1028–1029, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672712.

Yuen:1975:FAW

[Yue75e] Chung-Kwong Yuen. Function approximation by Walsh series. IEEE Transactions on Computers, C-24(6):590–598, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672864.

Yuen:1976:CAR

[Yue76] C.-K. Yuen. Comments on “The Application of the Rademacher– Walsh Transform to Boolean Function Classification and Threshold Logic Synthesis”. IEEE Transactions on Computers, C-25(7):766–767, July 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674688. See [Edw75].

Yuen:1977:ARa

[Yue77a] C. K. Yuen. Author’s reply. IEEE Transactions on Computers, C-26(2):192, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5009302.

Yuen:1977:ARb

[Yue77b] C. K. Yuen. Author’s reply. IEEE Transactions on Computers, C-26(5):511–512, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674870.

Yuen:1977:NRD

[Yue77c] C. K. Yuen. A new representation for decimal numbers. IEEE Transactions on Computers, C-26(12):1286–1288, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674792.

Yuen:1977:LDF

[Yue77d] C. K. Yuen. On Little’s digital filtering algorithm. IEEE Transactions on Computers, C-26(3):309, March 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674829.

Yuen:1977:TRN

[Yue77e] Chung-Kwong Yuen. Testing random number generators by Walsh transform. IEEE Transactions on Computers, C-26(4):329–333, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674842.

Yuen:1978:AGC

[Yue78] C. K. Yuen. Analog-to-gray code conversion. IEEE Transactions on Computers, C-27(10): 971–973, October 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674979.

Yau:1975:MFD

[YY75] S. S. Yau and Shih-Chien Yang.

Multiple fault detection for combinational logic circuits. IEEE Transactions on Computers, C-24 (3):233–242, March 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672797.

Zahn:1971:GTM

[Zah71] C. T. Zahn. Graph-theoretical methods for detecting and describing Gestalt clusters. IEEE Transactions on Computers, C-20 (1):68–86, January 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671676.

Zalcstein:1971:NFC

[Zal71] Y. Zalcstein. A note on fast cyclic convolution. IEEE Transactions on Computers, C-20(6):665–666, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671912.

Zeigler:1971:FHR

[Zei71] B. P. Zeigler. Feedback in homomorphic realizations. IEEE Transactions on Computers, C-20(6):685–688, June 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1671920.

Zucker:1977:ARL

[ZHR77a] S. W. Zucker, R. A. Hummel, and A. Rosenfeld. An application of relaxation labeling to line and curve enhancement. IEEE Transactions on Computers, C-26(4):394–403, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1674848. See correction

[ZHR77b].

Zucker:1977:CAR

[ZHR77b] S. W. Zucker, R. A. Hummel, and A. Rosenfeld. Correction to “An Application of Relaxation Labeling to Line and Curve Enhancement”. IEEE Transactions on Computers, C-26(9):922–929, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674940. See [ZHR77a].

Zisapel:1976:SGM

[Zis76] Y. Zisapel. A simplified general method for static hazard detection. IEEE Transactions on Computers, C-25(11):1160–1164, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674573.

Zisapel:1979:DHC

[ZKK79] Y. Zisapel, M. Krieger, and J. Kella. Detection of hazards in combinational switching circuits. IEEE Transactions on Computers, C-28(1):52–56, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1675222.

Zohar:1970:NRC

[Zoh70] S. Zohar. Negative radix conversion. IEEE Transactions on Computers, C-19(3):222–226, March 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671493. See comments

[Paw71, Wad71].

Zohar:1971:AR

[Zoh71] S. Zohar. Author’s reply. IEEE Transactions on Computers, C-20(5):587, May 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671895.

Zohar:1973:CRD

[Zoh73a] S. Zohar. The counting recursive digital filter. IEEE Transactions on Computers, C-22(4):338–347, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672313.

Zohar:1973:FHF

[Zoh73b] S. Zohar. Fast hardware Fourier transformation through counting. IEEE Transactions on Computers, C-22(5):433–441, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672338. See comments

[Zoh74].

Zohar:1973:NHR

[Zoh73c] S. Zohar. New hardware realizations of nonrecursive digital filters. IEEE Transactions on Computers, C-22(4):328–338, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672312.

Zohar:1973:DCR

[Zoh73d] Shalhav Zohar. A/D conversion for radix (−2). IEEE Transactions on Computers, C-22(7):698–701, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009137.

Zohar:1974:CFH

[Zoh74] S. Zohar. Comments on “Fast Hardware Fourier Transformation Through Counting”. IEEE Transactions on Computers, C-23(9): 989, September 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1672664. See [Zoh73b].

Zohar:1976:RRD

[Zoh76a] S. Zohar. A realization of the RAM digital filter. IEEE Transactions on Computers, C-25(10): 1048–1052, October 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674546.

Zohar:1976:RTR

[Zoh76b] S. Zohar. Rounding and truncation in radix (-2) systems. IEEE Transactions on Computers, C-25(5):464–469, May 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1674634.

Zahn:1972:FDP

[ZR72] Charles T. Zahn and Ralph Z.

Roskies. Fourier descriptors for plane closed curves. IEEE Transactions on Computers, C-21(3): 269–281, March 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=5008949.

Zucker:1975:PST

[ZRD75] S. W. Zucker, A. Rosenfeld, and L. S. Davis. Picture segmentation by texture discrimination. IEEE Transactions on Computers, C-24(12):1228–1233, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672762.

Zobrist:1975:BDF

[ZT75] A. L. Zobrist and W. B. Thompson. Building a distance function for Gestalt grouping. IEEE Transactions on Computers, C-24(7):718–728, July 1975. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

                                        tronic). URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=1672885.